A Complete Bibliography of Publications in the IEEE symposia on Foundations of Computer Science (FOCS)

for 2000–2009

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: http://www.math.utah.edu/~beebe/

01 March 2018 Version 2.07

Title word cross-reference

#SAT [BDP03].

(1 + ϵ) [KSS04]. (2 + ϵ) [Ber09]. (log n)^{Omega(1)} [CKN09]. 0 [Ajt02, KW09]. 1 [Ajt02, KW09]. 1 − 1/e [FV06, FMMM09]. 2 [Aus07, BL04, Cap02, CJLV08, DH09, FJW01, KLR09, KKMO04, KN07b, WN06]. 2 − o(1) [GMPT07]. 3 [AS00, ABS03, BOT02, DRS02, DM02, KS09a, Kho02]. = [BG02]. ^{0} [AIK04, BHN09, MST03, RS08]. NP [Cai01]. _{1} [Raz04]. _{2} [Cai01, Raz04]. _{m} [CGPT06]. b [BDFC00]. ∩ [AR04]. ℓ_{1} [BC03, CS02a, KS09b]. ϵ [MST03]. k [AM02, AN08, AV06, AMR09, HSV09, KSS04, OR00, ORSS06, Sch08, SBI02]. L [ACP08]. L_{1} [IR08, KN07b, NS06, KV05a]. L_{p} [ER09, Kho03, LN06]. N [ACR07, BM09]. N^{1/2} + o(1) [ACR07]. ≠ [Raz04]. O [BM09]. O(1) [GP09]. O(k^{3} log n) [CK09b]. O(log n) [AHK04, Cha06]. O(m^{1.31}) [ST03]. O(n) [FG03]. O(n^{1/(2k−1)}) [BIKR02]. O(n^{2}) [DI00, AHK04]. O(n log n) [FG03, Cha06]. O(n\sqrt{(} log log n)) [HT02]. O(v log n) [She09a]. O ∗ (2^{n}) [Koi06]. O ∗ (n^{4}) [LV03b]. Ω(n^{1/3}) [RY06]. R^{d} [MS08b]. Z_{m} [BGL03].

- [Ajt02]. -approximation [FJW01, KSS04, CK09b, GP09]. -Approximations [She09a]. -biased [MST03]. -clustering [OR00]. -CNFs [ABS03]. -colorability [BOT02]. -colorable [Kho02]. -CSP [Aus07]. -CSPs [Sch08]. -dimensional [Cap02]. -DNF [SBI02]. -Embeddability [KS09b]. -infinity [ACP08]. -lifts [BL04]. -Means [AMR09, ORSS06, AV06, KSS04]. -norm [ER09]. -Query [DH09]. -SAT [AS00, AM02]. -Source [KLR09]. -Sums [CJLV08]. -tag [WN06]. -trees [BDFC00]. -uniform [DRS02, Kho02]. -variable [KKMO04]. -Wise [AN08]. -XORSAT [DM02].

2002 [IEE02].

3CNF [FKO06, Fei07].

41st [IEE00]. 42nd [Ano01, IEE01]. 43rd [IEE02]. 44th [IEE03]. 45th [IEE04]. 46th [Ano05a, Ano05d, Ano05b, Ano05c, IEE05]. 47th [Ano06b, Ano06c, Ano06a, Ano06d, IEE06]. 48th [Ano07b, Ano07c, Ano07a, Ano07d, IEE07]. 49th [IEE08].

50th [IEE09].

Abstract [AK07, CS02b, EY07, Gol02, MS04a, San04]. AC [BHN09, RS08]. access [CFLM02, PY00]. Accidental [Val06]. Ackermann [Pet02]. Acyclic [CMM07b, GMR08, MS08a]. Adaptive [SVV07]. adaptivity [DGV04]. Additive [ASS05, Woo06]. Additively [CDDT09]. Adiabatic [AvDK^{+}04, vDMV01]. adsorbing [MR00]. Advantage [CMM07b]. adversarial [BG03]. Adversary [Rei09]. AdWords [MSVV05]. Affine [BV08, GR05]. Affine-Invariant [BV08]. against [Rei06]. agent [GKTW09]. Agents [DK08, DDDR08, AT01]. aggregate [KDG03]. Agnostic [FGRW09]. Agnostically [KKMS05]. Agreements [ASW06]. Algebraic [BATS09, GP06, Har06]. Algebraic-Geometric [BATS09, GP06]. Algebras [CS04]. Algorithm [AV06, BBN07, BNN06, CK09b, iKMR08, Koi06, Amb04, BLMS^{+}03, BFG01, CP05, CLV03, DPSV02, FJW01, GKPR03, HH00, Jai04, Kle05b, KGR02, KSS04, LV03b, Pak00, VW02]. Algorithmic [ACO08, Cha05, DHK05, DD09, Ind01, Gaf02, HNR05]. Algorithms [ABC09a, AI06, ABHK09, BDP03, BW09, BK06, BH06, CLX08, CHKS06, CS08, CSV07, Chl07, CK08, DR08, GG07, GM07b, Har06, KL06b, LV06, Moi09, NO08, RV06, Sar06, SF08, Val06, AS00, AHK05, BCvD05, BCS06, BCK^{+}03, COR01, CKSU05, CR03, DGRS05, FV06, GK02, Guh00, GRS04b, KV00, KLSS03, KM05, LV03a, PT03, RZ02, Sri01, SS05, Sze04, Tre05, Val04, You01]. All-Pairs [BK06, Ber09, RZ04]. Allocating [CCK09]. allocation [ADK^{+}04, FV06, KK00]. Allocations [ALGG09, CG08, EP06]. Allowing [iK09]. Almost [Cho08, ES07, KL05, Kol01, BS05, DHIP04, GS02, Lan04, SS04]. almost-linear [GS02]. Almost-Natural [Cho08]. Alphabets [GP06]. Always [OSZ03]. among [KPR^{+}09]. Amplification [BS07, IJK06]. Analogue [BZ09]. Analysis [AV06, BEK05, KST09, Mos08, RT09, BLMS^{+}03, DHM04, DS05, KN05, Pet05, Upf03]. AND-OR [ACR07]. Annealing [SVV07, LV03b]. Annual [Ano05a, Ano05d, Ano05b, Ano05c, Ano06b, Ano06c, Ano06a, Ano06d, Ano07b, Ano07c, Ano07a, Ano07d, IEE00, IEE02, IEE03, IEE04, IEE05, IEE06, IEE07, IEE08, IEE09]. Anonymity [IKOS06]. Anonymous [DP07a, DP08a]. Answering [YZ05]. answers [BBK04]. Any [ACR07, Aus07, Hai09, NOV06, RS09a, BS05, KU08, KSS04].

Application [AV06, FGMS06, VV02]. Applications [ABIW09, BARdW08, BG07, DKSS09, FFS06, Fel09, FMS07, GRW06, HN06, IKMT07, Spi07, Ung09, VW07, ALW01, AZ02, BGGL01, BBM01, DN00, ELRS02, FM00, HH00, Ind01, IK00, Sri01, Vet02, WX05]. Approach [BN06, PTW08, BYJKS02, CU03, Str00]. Approximability [CG08, GKTW09, SW00, Von09, ABE^{+}05, PY00]. Approximate [AI06, BK06, Ber09, FMS07, HPV01, KKM^{+}08, KN08, KL06a, MS08a, PTW08, PP09, AHK05, CR04, GHS00, Lau04, RZ04, Tho01]. Approximately [IJK06]. Approximating [BYJKK04, BHN08, CKR01, DGV04, GG07, Nut09, RRSS07, Sku00, AKKV05, Kho03, Kho04a, VVZ02]. Approximation [ABHK09, BCK^{+}03, BKM08, CHKS06, CDT06, CS08, Chl07, COR01, CK09b, DDDR08, FV06, GP09, GM07b, GKPR03, HNO08, HKNO09, KLSS03, KM05, Mar07b, Moi09, NO08, Sar06, SF08, Tre05, Ale03, ASS05, AHK04, BCS06, DHK05, DGRS05, FK00, FJW01, GK02, Guh00, GRS04b, Kle05a, Kle05b, KY01, KGR02, KSS04, OR00, Sri01, SS05]. Approximations [She09a]. arbitrarily [BG03]. Arbitrary [iKMR08]. Arc [BF01, McC01]. Arc-disjoint [BF01]. arcs [CDR00]. Area [BNN06, KOS08]. Argument [Hai09]. Arguments [NOV06]. Arithmetic [AV08, RSY07]. arm [Str00]. Arrangement [AMS07]. arrangements [AAS01, Cha00, Cha03]. array [FG04]. Arrival [BS06]. Arrow [CDDT09, Jai04]. Art [Ano08d, Ano09c]. Ascending [IKMT07]. Assembly [Dot09, AKKR02]. Assignment [DR04, ELRS02, NPS03]. Assumptions [ABX08, KLR09, BCNP04, Mic02]. asymmetric [CGK04, KLSS03]. asymptotic [AM02]. asymptotically [OSZ03]. Atlanta [IEE09]. Auctions [AM08, BN02, DLN08, IKMT07, KL03, LMN03, RS02, Vet02]. Augmentation [BCK^{+}07]. augmented [ADRR04]. Authenticated [GKKO07]. Authentication [BCG^{+}02]. Author [Ano02, Ano03, Ano06e, Ano07e, Ano08a, Ano09a]. automatic [ORS05]. automatizability [ABF^{+}04]. automatizable [AR01b]. auxiliary [GK05]. Available [Ano06h, Ano06i, Ano07h, Ano08f, Ano08g, Ano09f, Ano09g]. Average [BLMS^{+}03, GK07, KL08, Tre08, Ale03, BT03, MR04]. Average-case [Tre08, BT03, MR04]. Award [Ano05j]. Awards [Ano05e]. axioms [IS01].

B [FGMP02]. B-trees [FGMP02]. bad [KVV00, RT00]. Balanced [ABC^{+}09b, EP06, JS02]. balancing [AKU03, MPS02]. Balloon [IKMT07]. Barrier [Roc08, She09a, Bar01, BIKR02, DI00, HSS03]. Barriers [ACO08]. Based [Bon07, BGH07, BPR^{+}08, GM07b, RY06, Roc08, GI01, KK02, KDG03, MR04, SF08, SS05, Sze04]. Basing [ABX08, BPR^{+}08, GMR01]. Bayesian [BDP03, OH08]. BC [IEE02]. Be [BSN08, ACR07, FKN08, MS04a]. Beach [IEE00]. beacons [KSW04]. beat [Mar07a]. Beating [FMMM09, GMR08, KY07]. behavior [BEK05]. Being [PSS08]. benefit [DGV04]. Berkeley [IEE06]. Best [Ano05e, KM00b]. Better [TSU06, FG04]. Betti [BZ09]. between [GKM^{+}00, SVV07, Tho02]. Beyond [GT07, KK05, Ver06, Bar01, FLMM05, PV05]. Bias [BATS09, Moc04]. biased [MST03, Rei06]. Bifamilies [Nut09]. Bilinear [RY06, BL02b]. bimatrix [SvS04]. bin [BLS05, BCS06, KM00b]. Binary [BCMR06, OP06, OH08, DHIP04, NS05]. binomial [AR01a]. bins [Cap02]. bipartite [GKPS02, Kho04b]. bisection [FK00, Kho04b]. Bit [BK09a, MS09b, GRS04a, KZ03]. bit-fixing [GRS04a, KZ03]. Bits [BV07]. Black [DGS09, HP07, Bar01, BLV03]. black-box [Bar01]. Black-White [HP07]. Blackbox [KS09a]. blanket [KKLV00]. Blind [BFK09]. Bodies [KN07b, LV03b]. Body [BNN06]. Boolean [Rei09]. booster [HR00]. Bottom [Cha07]. Bound [BCMR06, ES07, GT07, HHRS07, RSY07, RY06, Rei09, VW07, Wag06, BOT02, CR04, HS05, JRP03, KKLV00, Pet02]. boundary [MSW03, PT00]. Bounded [BK09b, BOBP^{+}02, CFLY08, CS07, DGJ^{+}09, Goe06, GKL03, GP06, iKMR08, LLW08, PR03, BN02, BOT02, BL02b, CW04b, DFSS05, JRP03]. Bounded-concurrent [PR03]. Bounded-Degree [CS07, BOT02]. Bounded-depth [BOBP^{+}02]. Bounds [ABX08, AGM08, BMG07, BLR08, BN06, CG08, DES08, DR08, GG07, GK07, KMT07, Mos08, PTW08, PT06, RRSS07, Sch08, SF08, Woo06, AR01a, ABS03, Aue00, BLV03, BSSV00, BOBP^{+}02, BOGH^{+}03, BL02b, CGPT06, DF01, FHS05, GT00, GKS05, IW03, Kla01, KL03, Kol01, NS05, SBI02, Shi02, Shp01]. Box [DGS09, Bar01, BLV03]. branch [AR02a]. branch-width [AR02a]. branching [New00]. Breaking [BIKR02, HSS03, Roc08, She09a, DI00]. Brief [Bon07]. Broadcast [GKKO07, GRW06, GKS05]. Broadcasting [ALS^{+}08, CR03, CGR00, KP02]. Budget [DLN08]. Budgeted [CG08]. Building [KM00a, PRU01]. Bulk [ACSZ07, CHKS06, GP09, And04]. Buy [ACSZ07, CHKS06, GP09, And04, GKPR03, KGR02]. Buy-at-Bulk [ACSZ07, CHKS06, GP09, And04].

Cache [BDFC00, BBF^{+}03]. Cache-oblivious [BDFC00, BBF^{+}03]. Cake [EP06]. calculus [AR01a]. calibrated [Vov02]. California [IEE00, IEE06].

Cambridge [IEE03]. Can [Baz07, FKN08, KLN^{+}08, Mar07a, ACR07, MS04a]. Canada [IEE02]. capacities [CN02, PTW01]. Cascaded [JW09]. Case [ABX08, KL08, AR01a, Ale03, AV06, AMM00, BLMS^{+}03, BT03, Mic02, MR04, Tre08]. cell [CR04]. cellular [ELRS02]. Certain [Sch08, MPS03]. Chain [FGMS06, MR00, Ran03, Sze04]. chains [CDG^{+}02]. Chair [Ano09d]. channel [AZ02]. Channels [BCMR06, AMTdW00]. Characteristic [KU08]. Characterization [Fel09, AS05, Gaf02, LMN03]. charting [GKS00]. Chasm [AV08]. checking [NR05]. Chinese [GSS00]. Choice [ALGG09]. Choice-Memory [ALGG09]. chromatic [FLS02]. circles [AAS01]. Circuit [BHN09]. Circuits [AV08, Cha07, KS09a, RSY07, CGPT06, CW00]. circular [McC01]. circular-arc [McC01]. Classical [Aar08, GI09, KSdW04]. clique [Kho04b]. Clock [LLW08]. Close [Ber09, BFR^{+}00]. Closed [GGL^{+}08]. closest [AKKV05, Mar05]. closure [DI00, San04]. Clustering [BV08, CGW03, GMMO00, HP01a, KN08, AC05, ADPR00, BBC02, KSS04, OR00]. clusterings [KVV00]. CNFs [ABS03]. Code [BK09a]. Codes [BATS09, BSKR06, GP06, IJK06, KS07, Mos08, Ale02, BSS03, CW04b, FK02, GS02, GSS00, GI01, KL05, Lan04, Lub02, NS05, ORS05, TSZS01]. Coding [BYBJK06, LS07, Sud01]. coefficient [BL02b]. Coin [CK09a, Bar02, Moc04]. coin-flipping [Moc04]. coin-tossing [Bar02]. Collecting [ABHK09]. colliding [Str00]. Collision [KMT07, Shi02]. Collisions [HHRS07]. colorability [BOT02]. colorable [Kho02]. Coloring [Koi06, AMSZ03, DHK05, DRS02, DFHV04, GHS00, Kho02]. colorings [ELRS02, HV03]. colouring [DF01]. colours [GMP04]. Combinatorial [ARV06, BW09, CGK^{+}00, GKTW09, Mei09, CS02b, DR04, FKR^{+}02, LMN03, Str00]. Combinatorics [Tao07]. come [DGRS05]. Commitments [HHRS07, PR05]. Committee [Ano09e]. Committees [Ano05f, Ano06g, Ano08b]. common [Mai00]. Communicating [OHP08]. Communication [BHN09, KW09, RT07, She08, VW07, BYJKS02, BGL03, BN02, JRS02, JRP03, KK02, Kla01]. Compact [Tho01, AR02b, Mic02]. comparisons [FG03]. competitive [BLMS^{+}03, DHIP04, Vet02]. Complete [Fel09, MS09b]. completeness [HK01, HP07]. Complexity [AMR09, BHN09, CM08, CDT06, CD06, CDDT09, DMS07, EY07, GKKO07, GI09, HHRS07, JW09, KU09, KN07a, KW09, RV06, Rei09, She08, SY06, Tre08, Ver06, AKV05, AAS01, ABSRW00, Ale03, Amb03, BDP03, BYJKS02, Bra05, Bur01, BL02b, CSWY01, FG02, Gro03, JRS02, JRP03, Kla01, Mic02, MS03, NR05, PT00, PRS02, RSV00, WN06]. composability [Lin03]. composable [BCNP04, Can01]. Composite [CW09]. Composition [DH09, KU08, BS05, Lin03]. Compressibility [HN06]. Compressible [CKL^{+}09]. Computable [GLM06, Ajt02]. Computation [BOCG^{+}06, BFK09, BCL^{+}06, CGOS07, GT07, KSSV06, MPR06, AvDK^{+}04, BSSV00, Ind00, IK00, IML05, KDG03, Lin03, PR03, Reg02, Sto01, vDMV01]. Computational [AK07, Cha06, Cha05, KLR09, Vad04]. Computations [Yus08]. Compute [BNN06]. Computer [ACR07, Ano01, Ano05a, Ano05d, Ano05b, Ano05c, Ano06b, Ano06c, Ano06a, Ano06d, Ano07b, Ano07c, Ano07a, Ano07d, IEE00, IEE01, IEE02, IEE03, IEE04, IEE05, IEE06, IEE07, IEE08, IEE09, Pap01]. Computing [Aar08, BARdW08, BHKK08, CDT06, DP07a, GPV06, CS04, Jai04]. Concave [KN07a]. Concentration [Kan09].

Concepts [KOS08]. Concise [DLM^{+}07]. Concurrent [BPS06, GM00, PR05, PRS02, BS05, FHS05, PR03]. condensers [TSU06]. condensing [RSW00]. condition [Hay06]. conditions [MSW03]. Conference [Ano07f]. confidence [Aue00, Vov02]. Conflict [ELRS02]. Conflict-free [ELRS02]. confluence [CGN01]. Congestion [ARV06, ACKZ05, FJM04, GY02, KKPS00, Rac02]. Conjecture [DGS09, Ver06, Kho05, KV05a, LN06]. conjectured [Ajt02]. conjectures [NPS03]. Connecting [CPR08]. Connection [SVV07]. connections [ALW01]. Connectivity [CPR08, CK08, CK09b, FGMS06, Nut09, PT07, AZ03, FJW01, Gab00, MPS03]. coNP [AR04]. consequences [Cha03]. Constant [Bar02, MR08, NO08, CDG^{+}02, DFHV04, JS02, KGR02, PR03, RVW00]. constant-degree [RVW00]. constant-factor [KGR02]. Constant-round [Bar02]. Constant-Time [NO08]. Constrained [AMS07]. Constraint [BK09b, CM08, BD03, CM04, Gro03, MS03]. Constraints [IN09, Bul02, CGKN04]. Constructing [BATS09, BL04, HSS03]. constructions [GT00, GI01, IK03]. Contents [Ano05c, Ano08j, Ano09j]. contingency [CDG^{+}02]. Control [GM07b, GY02, KKPS00, ORS05]. Convergence [ABC^{+}09b, MS09a, GMV05]. Convex [BNN06, KN07b, BJ02, LV03b]. convexifying [CDR00]. Coppersmith [NPS03]. Copyright [Ano05b, Ano06a, Ano07a, Ano08c, Ano09b]. core [AGS03]. Coresets [FFS06]. Corporate [Ano05g]. Correcting [PV05, ORS05]. Correlated [GP06]. Correlation [BBC02, Mos08]. Correlations [IKOS09, RT07]. Cost [GKTW09, Nut09, ADK^{+}04, BBF^{+}03, GKPR03, Sku00, MMP00]. cost-sharing [GKPR03]. Costs [CFLY08, GK01]. Counterexample [Raz08]. Counting [BH06, IS01, SVV07, BD03, FG02]. coupling [HV03]. Cover [Ano06b, Ano07b, Ano08d, Ano09c, GMPT07, GV09, Nut09, KKLV00]. Covering [CN02, GGL^{+}08, IN09, KY07, DGRS05, KY01, You01]. Coverings [GV09]. Covert [CGOS07]. Criticality [LNP07]. Cryptographic [HN06, KS06, Can01, GT00, IK03]. Cryptography [AIK04, Bon07, CPS07, DFSS05, DP08b, IKOS06, KLR09, DOPS04, KZ03]. CSP [Aus07, RS09a]. CSPs [KKMO04, Sch08]. CTL [Mai00]. Cubes [KORW08, MS04a, SW01]. curve [KL03, TSU06]. curved [KL06b]. Curves [GLM06, Cha00, Cha03]. Cut [AMS07, CKN09, She09a, KV05a, Lau04, AHK04, KKMO04]. cutting [BOGH^{+}03]. cycles [CDR00]. cyclic [BSS03, Mic02]. Cyclotomic [Che07]. Cylinders [Ezr08].

Data [JW09, Pat08a, ABR00, AZ02, BYJKS02, FM00, GMMO00, Ind00, MN06]. Debreu [CDDT09, Jai04]. decidability [AKKR02]. decidable [CGN01]. decision [GSS00, KV05b, OSSS05]. Decodable [DH09, KS07, GI01]. decodibility [CW04b]. Decoding [BSKR06, FK02, GP06, IJK06, AGS03, Ale02, GSS00, Tre03]. Decomposing [GV09, KLSS03]. Decomposition [DHK05, MR00]. Decompositions [Nut09, Kol01]. Defective [KLR09]. defies [MRS05]. Deformations [BLR08]. Degree [CFLY08, CS07, FGMS06, Goe06, She09b, Amb03, BOT02, DHM04, DF01, DFHV04, JPRZ04, Kle05b, RVW00]. Delaunay [BM09]. deletion [ASS05]. Delsarte [NS05]. demand [DGRS05, KL03]. demand-robust [DGRS05]. Dense [RTTV08, FKO06]. densest [Kho04b]. density [ABS03].

Dependent [GKPS02, AZ02]. Depth [AV08, Cha07, KS09a, AZ03, BOBP^{+}02]. depth-first [AZ03]. Depth-three [Cha07]. Derandomization [Che07]. derandomized [TSU06]. descent [KLMN04]. Design [ACSZ07, CFLY08, CHKS06, CK09b, DD09, MT07, And04, ADK^{+}04, BBHM05, GMM00, GRS04b, IML05, LS05, MMP00]. Designing [MMP01, Rou01]. Detecting [Kle00]. Determinant [Yus08, CS04]. Deterministic [GRS04a, GR05, KZ03, KP02, Sto01, SS04]. Deterministically [AC07]. diagonally [ST03]. diagonally-dominant [ST03]. Diagrams [Cha06, HP01b]. Diameter [KN07b, PRU01]. Dichotomy [CM08, Bul02, BD03]. dictionary [FGMP02]. different [BW03]. Differential [MT07]. digraphs [BF01, Tho01]. Dimension [CS02a, BC03]. dimensional [BLS05, Cap02, Cha02]. Dimensionality [AIP06]. Dimensions [AAL09, AI06, ES07, Ezr08, KORW08, SY06, Kol01, KSS04, VVZ02]. Diophantine [Ale02]. Direct [DG08, IJK06, Ung09, ALW01, CSWY01, KSdW04]. Direct-Product [Ung09]. directed [AZ03, CP05, CKR01, KLSS03, RZ02, YZ05]. directions [Blu03]. discounted [BCK^{+}03]. discounted-reward [BCK^{+}03]. Discrepancy [Cha07, RY08, BL04]. Discrete [KMT07, FOS05, Hal04]. Discretized [DP08a]. Dishonest [GKKO07]. Disjoint [AC07, ACKZ05, BF01, CKS04, Kle05b]. Disjointness [KW09, JRP03]. dispenser [GKR01]. Dispersion [RV06]. Dissection [Yus08, Guh00]. Distance [AK07, ABIW09, BK06, Roc08, SVY09, BYJKK04, CW04b, YZ05, MMP00]. Distance-Based [Roc08]. distances [Mar05, Tho01]. Distinct [RRSS07, IW03]. distinctness [Amb04, Shi02]. distortion [GKL03, Ind01]. Distributed [GPV06].

Distribution [RRSS07]. Distributions [DP08a, Sri01, BFR^{+}00, DHM04, DHKS05, FOS05, Ind00, KM00b, TV00, VW02]. DNF [Baz07, BMOS03, KS08, SBI02]. do [IS01]. domains [FOS05]. dominant [ST03]. down [AIS04]. Dual [BBN07, BN06, DPSV02, PT03]. Dynamic [ABG06, BKP04, Ber09, BJ02, CHK09, CPR08, DHIP04, RZ04, San04, DI00, DI01, RZ02, Upf03]. Dynamics [ABC^{+}09b, HS05, Hay06, KMP01].

Earliest [BS06]. Earth [ABIW09]. Earth-Mover [ABIW09]. Earthmover [NS06]. Easy [KRT08, SS04]. economics [Pap01]. Edge [CKS04, KK04, AMSZ03, ASS05, ACKZ05, DI01, GRS04b, HS01, HS02, MS04a]. edge-deletion [ASS05]. Edge-disjoint [CKS04, ACKZ05]. edges [FR01]. Edit [AK07, BYJKK04]. Effectiveness [ORSS06]. efficiency [AKU03, GT00]. Efficient [ABIW09, BGH07, HSV09, KV00, Mei09, AZ03, BCvD05, CS02b, HK01, IK00, LV03a, Mic02, WX05]. Efficiently [BS06, BYJKK04, GI01]. Eigenvalue [BLR08]. Eigenvalues [KLPT09]. Elections [FKN08]. element [Amb04, Bul02, FJW01, Shi02]. Elements [RRSS07, IW03]. elimination [MS04b]. Embeddability [KS09b, KV05a]. Embedding [iKMR08, KSW04, KLMN04]. Embeddings [CJLV08, CMM07a, CMM06, MS08b, ML08, ABK^{+}05, GKL03, Ind00, Ind01]. Emulators [Woo06]. Encryption [BGH07, BPR^{+}08, GRW06, MS09b, GKM^{+}00]. end [GY02]. end-to-end [GY02]. energy [BKP04]. Entangled [KKM^{+}08, KRT08]. entropies [SU01]. Entropy [HNO08, RVW00, JRS02]. envelopes [KS02]. Envy [CS08]. Envy-free [CS08]. Equations [KN07b, Ale02]. Equilibria [CDT06, CDDT09, DP07a, DP08a, DK08, EY07, BVV05, GMV05, Vet02]. Equilibrium [CD06, MS09a, DPSV02, Jai04, SvS04]. Equivalence [Tho02]. equivalent [AvDK^{+}04, BG02]. Erratum [HS02]. Error [DG08, DH09, iK09, MR08, ORS05, She08, AS05]. Error-correcting [ORS05]. Errorless [BS07]. errors [BBK04, PV05]. Estimating [AK07]. estimation [OSZ03]. Euclidean [BKM08, DES08]. Evacuation [BS06]. Evaluated [ACR07]. even [Kle05b]. even-degree [Kle05b]. Every [OSSS05, Rei09]. evolution [ACL01]. Evolvability [Fel09]. Exact [Dot09, PP09, CM04]. Exchange [ABC^{+}09b]. Exclusion [BH06, Koi06]. Exclusive [GRW06]. existence [HR00]. Existential [GKS00]. Expander [GI01, KPS08, BL04, BF01, Gab00]. Expander-based [GI01]. Expanders [AC07, KPS08, AC02, RVW00]. Expansion [CS07]. expected [AMM00, HT02]. expected-case [AMM00]. Explicit [AC02, GRW06]. Exponential [BC08, BBS08, BHKK08, HP07, MS04a]. Exponentially [SvS04]. exposure [KZ03]. exposure-resilient [KZ03]. Expressions [KS08]. Extended [AK07, EY07, San04]. extending [CW04a]. Extensions [DKSS09]. external [CFLM02]. Extracting [BIW04, IKOS09, RSW00, TV00]. Extractor [KLRZ08]. Extractors [DGW07, DW08, KLR09, RY08, TSZS01, GRS04a, GR05, KZ03, RVW00, SU01, Vad02]. Eyes [GGL^{+}08].

faces [AAS01]. Facilities [FFS06]. Facility [FS08, PTW01, Mey01, Vet02]. factor [FV06, KGR02]. factors [Bur01]. failure [Kle00]. fair [ADK^{+}04]. Fairness [CCK09, KK00]. Fan [Cha07]. Fan-in [Cha07]. Fast [AHK05, Ber09, CGR00, CW00, GK02, KU08, LV06, PT07, CU03, YZ05]. Faster [BK06, BNN06, KM09]. Fat [ES07, PT00]. Fault [BCL^{+}06, GPV06]. Fault-Tolerant [BCL^{+}06, GPV06]. favorite [Blu03]. feature [CGK^{+}00]. feedback [BLMS^{+}03]. Few [iK09, BIW04]. fewer [GMP04]. Fiat [GK03]. Fibonacci [CLX08]. Fields [GKS07, BW03, GR05, JPRZ04, KR04]. FIFO [BG03]. final [Pet05]. find [Gab00]. Finding [AC07, HHRS07, SvS04]. Finite [GKS07, RT07, KLMN04, NPS03, Wat00]. first [AZ03]. fit [KM00b]. Fitting [AC05, HPV01]. Fixed [DK08, EY07, GT07]. Fixed-Point [GT07]. fixing [GRS04a, KZ03]. Flipping [CK09a, Moc04]. Flow [BNN06, GK07, She09a, Sku00]. Flow-time [GK07]. Flows [BS06, BLR08]. FOCS [IEE02, IEE03, IEE04, IEE05, IEE06, IEE07]. Fool [Baz07]. Fools [DGJ^{+}09]. Forbidden [Lev02]. Forest [BKM08, OP06, GK02]. Forest-Like [OP06]. Foreword [Ano05h, Ano06f, Ano08e]. form [Sto01]. Formation [ASW06, KU09]. Formula [ACR07]. Formulas [Baz07, Fei07, RY08, ANP03, FKO06]. Forward [Ano07g]. Foundations [Ano01, Ano05a, Ano05d, Ano05b, Ano05c, Ano06b, Ano06c, Ano06a, Ano06d, Ano07b, Ano07c, Ano07a, Ano07d, IEE00, IEE01, IEE02, IEE03, IEE04, IEE05, IEE06, IEE07, IEE08, IEE09, VV02]. Four [AV08, Kol01]. Fourier [CW00, HH00, KN05, MRS05]. fractals [GKL03]. Fractional [KPR^{+}09, KY07, GK02]. fragment [Mai00]. Framework [HSV09]. Free [BK09a, CS08, ELRS02]. Frege [BOBP^{+}02]. Frequency [BHN08, ELRS02]. Frobenius [Sto01]. frontier [GKS00]. frugality [KK05]. Full [GPV06, HSS03]. Full-Information [GPV06]. full-text [HSS03]. Fully [Ber09, DI00, DI01]. Function [IN09, NOV06, Rei09, New00]. Functions [BARdW08, FMV07, GKTW09, LV06, Mos08, She08, Bra05, GMR01, LV03a, Mic02, MOO05, VV02, WX05]. Further [PT06].

Game [Pap01, BS05, SvS04]. Games

[ARV06, BHH^{+}08, CMM06, DP07a, DP08a, KKM^{+}08, KRT08, MS09a, AKV05, BVV05, FJM04, Kho05, KV05a, Tre05, RS09b]. Gap [CKN09, BL04, JS02, KV05a, CG08]. Gaps [KS09b, RS09b, ABL02, BOGH^{+}03, GMPT07, KO06]. Gates [CLX08, CGPT06, IS01]. Gaussian [KOS08, MR04, Mos08, MS04b]. General [Lin03, Rei09, HS05, KR04, KY01, Rac02]. Generalization [OP06, Wag06]. Generalized [Mic02, FJM04, MSVV05]. Generation [KM09]. generator [SU01]. generators [ABSRW00, Ind00, MST03]. generic [GT00]. Genus [iKMR08]. Geometric [ABC09a, BATS09, GP06, KOS08, Mar07b, PTW08, RV06, Wag06, ELRS02, Ind01, OR00]. geometries [GKL03]. Geometry [Cha06, CPR08, Cha05, LV03a]. Georgia [IEE09]. girth [DF01]. Given [FGMS06]. Glauber [KMP01]. Global [BL02a, CMM07a, LLW08, KM00a]. go [Bar01]. Goemans [LN06]. Good [OH08, OSZ03, KVV00]. Goods [CCK09, DK08]. Gossip [KDG03, KK02]. Gossip-based [KDG03, KK02]. gossiping [CGR00]. Graph [ACL06, AK02, HKNO09, Koi06, Moi09, Spi07, AS05, AR02b, DHK05, GT01, Guh00, Kho04b, KRR^{+}00, RVW00, Tor00]. Graphs [AN08, BBS08, CJLV08, CS07, FGMS06, FLS02, GM07a, iKMR08, KLPT09, SVY09, ACL01, ALW01, Alo01, AR02b, AZ03, BL04, BOT02, CKS04, CP05, CLV03, DHM04, DF01, DFHV04, FR01, Gab00, GKPS02, GMP04, GKS00, HS05, KMP01, Kle05b, McC01, McS01, RZ02, RZ04, YZ05]. Greedy [ML08, CP05]. Grothendieck [CW04a]. ground [CGN01]. Group [CKSU05, PT03, RY06, CU03, LN06, MRS05]. Group-theoretic [CKSU05, CU03]. groups [ALW01, BCvD05, Wat00]. Guarantees [Moi09, ABK^{+}05]. guises [Vad02]. Guruswami [PV05].

Halfspaces [DGJ^{+}09, FGKP06, FGRW09, GR06, KS06, She09b, KKMS05, KO02]. Hamiltonian [GI09]. Hard [FGRW09, GMR08, KKM^{+}08, MS08a, AGS03, CN02, DR01, PTW01, Rou01]. hard-core [AGS03]. Hardness [AKKV05, AK07, ACP08, And04, ACKZ05, BS07, CLX08, GKS07, GHS00, GR06, IJK06, KN07a, Kho02, Kho03, Kho04a, KS08, KS06, PSS08, CN04, DRS02, KO06, RS02, Tor00]. harmony [Cap02]. Hashing [AI06, BG07]. heaps [Pet05]. Heat [BNN06]. heavy [DHKS05]. heavy-tailed [DHKS05]. Heisenberg [LN06]. Helly [Hal04]. Helly-type [Hal04]. heterogeneous [FJM04, KK04]. Heuristics [BS07]. Hidden [CSV07, ABK^{+}02b, BCvD05]. Hiding [HHRS07]. Hierarchical [AC05, GMM00]. Hierarchies [Chl07]. Hierarchy [BZ09, FS04, GMPT07]. High [AI06, KORW08, She09b, AZ02, Kho03, VVZ02]. Higher [AAL09, KLPT09, PT06]. Hirsch [Ver06]. Histograms [BO07]. History [BG07]. History-Independent [BG07]. Holographic [CLX08, Val04]. Homology [BCSE^{+}07]. homomorphism [Gro03]. Honest [BOCG^{+}06]. hub [AFKM01]. huge [FLS02, GGN03]. hull [BJ02]. hyperbolic [KMP01]. Hypercontractive [BARdW08]. Hypergraph [FMS07, DRS02, GHS00, HNR05]. Hypergraphs [BC08, KL06a, Kho02].

I/O [AZ03]. I/O-efficient [AZ03]. ICP [AV06]. ideal [IML05]. identities [CS04]. Identity [BGH07, BPR^{+}08, KS09a, BFF^{+}01]. IEEE [Ano01, IEE01, IEE02, IEE03, IEE04, IEE05, IEE06, IEE07, IEE08, IEE09, Ano05a, Ano05d, Ano05b, Ano05c, Ano06b, Ano06c, Ano06a, Ano06d, Ano07b, Ano07c, Ano07a, Ano07d]. II [Cha03]. iii [Ano08l, Ano09l]. im [DOPS04]. Impact [ARV06]. Imperfect [CPS07, DOPS04]. implementation [GGN03]. implementations [FHS05]. Implicit [FGMP02, AR02b]. implying [Hay06]. Impossibility [BPR^{+}08, BC03, GMR01, GK05, KK02]. Improved [ABHK09, ABG06, BCS06, BN06, CG08, DS05, GP06, RZ02, Sar06, HH00]. Improvements [NO08]. Improving [FV06]. in-approximability [SW00]. in-place [FG03]. Inapproximability [AMS07, Aus07, GMR08, MS08b, KKMO04]. incentive [KR05]. Inclusion [BH06, Koi06]. Inclusion-Exclusion [BH06, Koi06]. incomplete [KM00a]. Increasing [GG07]. incrementally [MMP01]. Independence [Baz07, DGJ^{+}09, BFF^{+}01]. Independent [AN08, BG07, Moi09, BIW04, GRS04a]. Index [Ano03, BYBJK06, LS07, Ano02, Ano06e, Ano07e, Ano08a, Ano09a]. indices [HSS03]. Indistinguishable [MPR06]. induced [AR02b]. induced-universal [AR02b]. Inequality [BARdW08, Kan09, Ung09, Cha03, CW04a]. inference [BDP03]. Inferring [BCSE^{+}07]. infinite [AKKR02]. infinity [ACP08]. influences [MOO05]. influential [OSSS05]. Information [ALS^{+}08, Ano09h, BYBJK06, GPV06, GM07b, RY06, BYJKS02, BIKR02, BL02a, CGW03, KDG03, Lev02, Raz05]. information-theoretic [BIKR02]. Informational [CSWY01]. Input [MPR06, GK05]. Input-Indistinguishable [MPR06]. Instability [BG03]. Instance [ABC09a]. Instance-Optimal [ABC09a]. Instances [HN06]. Integer [Che07, HT02, KY01]. Integrality [CKN09, GMPT07, KS09b, RS09b, ABL02, BOGH^{+}03, CGK04, KV05a]. Integration [LV06]. Interaction [MS09a, JRS02]. Interactive [HHRS07, Hai09, JUW09, KR06, OHP08]. Internet [MPS03]. Intersection [KW09, She09b]. Intersections [KS06, KO02]. interval [COR01, CGKN04]. intractable [CM04]. introduction [Pap01]. Intrusion [DP07b]. Intrusion-Resilient [DP07b]. Invariance [MOO05]. Invariant [BV08, GI09]. Invariants [CJLV08]. inverse [Pet02, Rud06, San04]. inverse-Ackermann [Pet02]. Ising [MSW03]. Island [IEE07]. Isomorhism [BC08]. isomorphism [AK02, Tor00]. Isotropic [BV08]. issues [GKR01]. Italy [IEE04]. iterative [FJW01].

Job [MS08a, COR01]. jobs [CGKN04]. Joins [AGM08]. Jumping [VW07]. juntas [FKR^{+}02].

Kakeya [DW08, DKSS09]. Katona [OW09]. Kernel [KN08]. Kernelization [BFL^{+}09]. Kesten [BCMR06]. key [GKM^{+}00]. KKL [OW09]. knapsack [DGV04]. knapsacks [Mic02]. knowing [ABL02, KL03]. Knowledge [BPS06, KR06, NOV06, BGGL01, BLV03, Gol02, KM00a, MRK03, PRS02, Vad04, Wat02]. Knuth [Ano05i]. Kolmogorov [VV02]. Kruskal [OW09].

labeled [FLMM05]. labeling [CN04]. Large [Sar06, Alo01, GR05]. Lasserre [Sch08]. latency [CGRT03]. Lattice [AR04, AR03, GMP04, Reg02]. lattices [Ajt02, Kho04a, Mic02]. law [Ajt02]. Laws [LNP07]. LDCs [BARdW08]. Leakage [DP08b]. Leakage-Resilient [DP08b]. Learn [KLN^{+}08]. Learnability [ABF^{+}04]. Learner [OH08]. Learning [ABK^{+}02b, ABX08, BBK04, BMOS03, FOS05, FGKP06, FGRW09, Fel09, GR06, KST09, KS08, KO02, KS06, KOS08, Aue00, BBHM05, Blu03, BG02, DHKS05, KKMS05, VW02]. lemma [SBI02, Tre03]. Lemmas [BW09]. Length [BSN08, GG07, Roc08, GS02]. Level [Sch08, BLMS^{+}03, Sri01]. level-sets [Sri01]. levels [Cha00, Cha03]. lexicographic [Hal04]. lifts [BL04]. Light [DES08]. Like [OP06, FK02]. Limited [CS08]. Limits [BCL^{+}06, DLN08, Wat02]. Line [AGIK07, GY02, MSVV05, Vov02]. Linear [Ale02, ABS03, AMS07, Ber09, CW09, FHS05, KS07, iKMR08, iK09, KM00b, KN07b, KY07, LS07, McC01, Sch08, ABL02, BSSV00, Cha02, FR01, FK02, GS02, HT02, HP01b, Jai04, KL05, Kle05a, KSS04, LS05, NS05, ST03]. Linear-time [McC01, Kle05a]. Linearization [CJLV08, HPV01]. Linial [LN06]. List [BSKR06, GP06, IJK06, Tre03, AGS03, CW04b]. List-Decoding [IJK06, Tre03]. Lloyd [ORSS06]. Lloyd-Type [ORSS06]. Load [MPS02, AKU03]. Local [ACL06, ABC^{+}09b, BCSE^{+}07, CMM07a, FGMS06, GT07, HKNO09, KS09b, LLW08, MS09a, NO08, SY06, AKU03, BL02a]. Locally [BSS03, DG08, GS02, KS07, KL05]. Location [ABG06, Cha06, FS08, Pat06, AMM00, Mey01, PTW01, Vet02]. Log [KMT07, JS02]. log-Sobolev [JS02]. logarithmic [PRS02]. Logconcave [LV03a, LV06]. logic [GKS00]. Logics [IK03]. LOGSNP [KR06]. Long [BK09a, Mos08, BEK05]. long-run [BEK05]. Longest [GG07]. Look [Bon07]. loops [dVL02]. lose [AKV05]. lossless [TSU06]. Lossy [CJLV08]. Lovász [GMPT07]. Low [ABN08, BC08, Cha02, DES08, DG08, DH09, BG03, GKL03, Ind01, JPRZ04, MOO05, PRU01]. low-degree [JPRZ04]. low-diameter [PRU01]. Low-dimensional [Cha02]. low-distortion [GKL03, Ind01]. Low-Error [DH09]. Lower [AR01a, ABX08, BLV03, BMG07, BL02b, CG08, CGPT06, DES08, DR08, GG07, GK07, GT00, GKS05, HHRS07, Kla01, PTW08, PT06, RRSS07, RSY07, RY06, Sch08, Shp01, SF08, VW07, Woo06, BSSV00, BOT02, BOBP^{+}02, CR04, DF01, FHS05, HS05, IW03, JRP03, Pet02, SBI02, Shi02]. Lower-Bounds [ABX08]. LP [GRS04b]. LT [Lub02]. LTL [Mai00].

Machine [Blu03, CGKN04, BBHM05]. machines [KM05, Vov02, WN06]. Machtey [Ano05j]. Majority [BOCG^{+}06, GKKO07]. Makespan [EOW08]. Malleable [BPS06, PR05]. man [Bar02]. manage [BKP04]. Manipulated [FKN08]. many [AAS01, SvS04, Vad02]. maps [BL02b]. Market [DPSV02, DK08, Jai04]. Markets [CDDT09]. Markov [Ran03, CDG^{+}02, FGMS06, MR00, Sze04]. Markovian [GM07b, HV03]. Mass [RV06]. Massachusetts [IEE03]. massive [ACL01]. Match [PTW08]. Matching [FMMM09, Har06, PP09, ABK^{+}02b, MSVV05]. matchings [MS04b]. mathematical [Pap01]. Matrices [Sar06]. Matrix [BARdW08, Yus08, CU03, CKSU05, Rud06, San04, Shp01, WX05, YZ05]. Matrix-Valued [BARdW08, WX05]. Matroid [Har06]. matroids [JS02]. Matthews [KKLV00]. Max [KL06a, Lau04, KKMO04]. MAX-CUT [KKMO04]. max-Steiner-tree-packing [Lau04]. MAXCUTGAIN [KO06]. Maximal [RY08]. Maximal-Partition [RY08]. Maximization [CG08, CHK09, Von09, CS08]. Maximize [CCK09]. Maximizing [CW04a, FMV07]. Maximum [CMM07b, GMR08, MS04b, ANP03, DF01]. May [BSN08, DGRS05]. Means [AMR09, ORSS06, AV06, KSS04]. Measured [KLMN04]. measurement [BCvD05]. measures [KK00, MR04]. Mechanism [BBHM05, DD09, MT07, IML05, LS05]. mechanisms [AT01, KK05, KK02, PT03]. median [MP00]. membership [BBK04, RSV00]. Memory [ALGG09, CFLM02, MPS02, NR05]. mental [BS05]. Mergers [DW08, DKSS09]. Message [Ano09d, JUW09, CSWY01]. messages [BCG^{+}02]. Meta [BFL^{+}09]. Method [Aar08, AIP06, AV06, DKSS09, Ver06, AHK05, DS05, HNR05, KLMN04, MR00]. Methods [ORSS06]. Metric [ABK^{+}05, CFLY08, CMM07a, MS08b, ML08, BBM01, CN04, MMP00]. Metrical [BLR08, BBM01]. metrics [AC05, KV05a, KLMN04, LN06]. middle [Bar02]. Min [BCK^{+}07, KL06a, Kho04b, Lau04, SU01, Sku00]. min-bisection [Kho04b]. min-cost [Sku00]. min-entropies [SU01]. Min-Max [KL06a]. min-Steiner-cut [Lau04]. Min-Sum [BCK^{+}07]. minded [CS08]. Minimization [IN09, KN07a, CGKN04]. Minimizing [FS08, GK07, KS08, Rac02]. Minimum [EOW08, Goe06, Nut09, CGRT03, FK00, Pet02]. minor [DHK05]. Mixed [RY08, You01]. Mixed-Sources [RY08]. Mixing [BBS08, LNP07, Ran03, CDG^{+}02, GMP04, HS05, Hay06, MSW03]. mixtures [DHKS05, FOS05, VW02]. Mobile [FS08]. MOD [CGPT06]. Model [PP09, ADRR04, Bar02, BG03, DFSS05, MSW03, VV02]. Models [CKL^{+}09, GM07a, BG02, KRR^{+}00]. Moderately [BC08]. Modular [KU08]. Moduli [CW09]. Modulo [KN07b]. Moments [BHN08, Kan09]. Monomials [FGRW09]. Monotone [FMV07, OW09, BW03]. motion [HP01a, Str00]. Movement [FS08]. Mover [ABIW09]. moves [FG03]. MPLS [GKR01]. Muller [TSZS01]. Multi [CGOS07, DLN08, GKTW09, OHP08, VW07, BLMS^{+}03, Lin03, SS05]. Multi-agent [GKTW09]. multi-level [BLMS^{+}03]. Multi-Party [CGOS07, VW07, Lin03]. multi-stage [SS05]. Multi-unit [DLN08].

Multicommodity [Moi09, She09a, FJM04, GKPR03, KK04, KGR02]. Multicommodity-Type [Moi09]. multicuts [CKR01]. multidimensional [BCS06]. multigraphs [KLSS03]. Multilinear [Raz04, RSY07, RY08]. Multilinear-NC [Raz04]. Multinomial [DP08a]. Multiobjective [RT09]. Multiparty [BHN09, BOCG^{+}06]. Multiple [BS06, BHN08, KM05]. multiplication [CU03, CKSU05, YZ05]. multiplicative [AHK05]. Multiplicities [DKSS09]. Multivariate [GKS07, Bur01]. my [Blu03]. myopic [AS00].

Nash [BVV05, CDT06, CD06, DP08a, EY07, SvS04, Vet02]. Natural [Cho08, AS05, BFG01]. NC [AIK04, MST03, Raz04]. Near [AI06, IR08, KMT07, PTW08, PT06, SVV07, FR01, HP01b, LS05]. Near-Neighbor [PTW08, PT06]. Near-Optimal [AI06, IR08, SVV07, LS05]. Nearest [AI06, ACP08, CR04]. Nearly [ABN08, AMM00, Rei09]. negative [FR01, KV05a]. negatively [KL06b]. Neighbor [AI06, ACP08, PTW08, PT06, AC02]. neighbour [CR04]. Nested [Guh00, Yus08]. net [BS05]. Nets [OW09]. Network [ASW06, ACSZ07, CFLY08, CHKS06, CK09b, KLRZ08, KU09, And04, ADK^{+}04, GMM00, GRS04b, Kle00, MMP00]. Networks [ABC^{+}09b, CPR08, DR08, FGMS06, GPV06, KSSV06, AFGZ01, CGR00, CR03, ELRS02, FJM04, KK04, KR05, KP02, MMP01, PRU01, Rac02, Rou01]. Nevada [IEE01]. nine [GRS04b]. No [FG04, SW00]. no-wait [SW00]. Noise [GR06, KPS08, MOO05, Mos08, Rei06]. Noisy [DR08, FGKP06, OH08, GKS05]. Non [BCK^{+}07, BPS06, CHK09, CHKS06, DGS09, FMV07, KR06, LS07, AR01a, ABE^{+}05, BLV03, DS02, FKO06, HV03, PR05, Str00]. non-approximability [ABE^{+}05]. non-binomial [AR01a]. Non-Black-Box [DGS09, BLV03]. non-colliding [Str00]. Non-Interactive [KR06]. Non-Linear [LS07]. Non-Malleable [BPS06, PR05]. non-Markovian [HV03]. Non-Monotone [FMV07]. Non-Preemptive [BCK^{+}07]. non-satisfiability [FKO06]. Non-Uniform [CHKS06, CHK09]. Nonembeddability [KN05]. Nonlinear [CSV07]. nonuniform [PTW01]. Norm [IR08, Rud06, CS02a, ER09]. Norms [JW09, Kho03]. November [IEE00, IEE02]. NP [AR03, AR04, BT03, HN06, NOV06]. Number [DK08, CDG^{+}02, PR03]. Numbers [BZ09, FLS02].

O-efficient [AZ03]. obfuscation [GK05]. objects [FHS05, GGN03]. Oblivious [ER09, GP09, BDFC00, BBF^{+}03, GM00, GKM^{+}00]. observation [BL02a]. obtaining [GRS04a]. October [IEE01, IEE03, IEE04, IEE05, IEE06, IEE07, IEE08, IEE09]. offs [PY00]. Often [FKN08]. Old [DW08]. On-line [GY02, Vov02, MSVV05]. One [BK09a, NOV06, VW07, AS05, AT01, DS02, Mic02]. one-parameter [AT01]. one-sided [AS05]. one-time [DS02]. One-Way [NOV06, VW07, Mic02]. Online [AC07, BN06, EOW08, FMMM09, Mey01, Aue00, DHIP04, KL03, MP00, NR05, Pet02]. Only [BOCG^{+}06]. open [Blu03]. Opportunistic [FM00]. Optimal [AS00, ABC09a, AMS07, AI06, AIS04, BK09a, BSN08, CK09a, IR08, KKMO04, KMT07, OH08, dVL02, AMM00, BCvD05, CR04, FLMM05, FG04, KSdW04, LS05, OSZ03, PY00, RS02, SVV07]. Optimality [AIP06, Mar07b, CFLM02, DHIP04, MOO05]. Optimization [KKPS00, LV06, RT09, SS04, SS05]. Optimum [LS07]. Oracles [BK06, SVY09, Tho01]. order [AM02, GKS00]. Ordering [GMR08]. Orders [OP06]. Organization [Ano07f, FG04]. Organizing [Ano06g, Ano09e]. orientable [dVL02]. Orientations [KL06a, SW01]. orienteering [BCK^{+}03]. Orthogonal [AAL09, ABR00, KL05]. Other [Cha06, EY07, Koi06, BG02, Gro03, KKMO04]. our [GGL^{+}08]. Outcomes [ABC^{+}09b]. Outperforming [LS07]. overlays [KS02].

P [AR01b]. P2P [PRU01]. PAC [BG02]. packet [AFGZ01]. Packing [BN06, Cap02, KY07, BLS05, BCS06, KM00b, Lau04, You01]. pad [DS02]. PAExact [BG02]. Page [Ano05d, Ano08c, Ano08i, Ano08k, Ano08l, Ano09b, Ano09l, Ano09k]. PageRank [ACL06]. Paging [BBN07]. pairing [Pet05]. Pairings [Bon07, BGH07]. Pairs [BK06, Ber09, DI01, RZ04]. Paper [Ano05e]. paradigm [Can01, GK03]. Parallel [BHH^{+}08, Hai09, Raz08, CW00, You01]. Parameter [DDDR08, AT01]. Parameterized [DMS07, FG02]. Parisi [NPS03]. Parities [FGKP06]. parking [Mey05]. Partial [GM07b, OP06, PTW08]. Partial-Information [GM07b]. Partition [RY08, KS02]. Partitioning [ACL06, BLR08, FMS07, Koi06, McS01]. Partitions [BH06, HKNO09, MN06]. Party [CGOS07, VW07, Lin03, PR03]. Paths [AC07, BK06, Ber09, CGRT03, GT07, ACKZ05, BF01, CKS04, DI01, FR01, HS01, HS02, Kle05b, RZ04]. Pattern [PP09]. pay [DGRS05]. PCA [BV08]. PCP [DR04, MR08, Raz05]. PCP-theorem [DR04]. PCPs [DH09, GS02, HK01, Mei09]. Pebbling [HP07]. Peer [FGMS06, KSSV06]. Peer-to-Peer [FGMS06, KSSV06]. Peering [ASW06]. Space [HP07]. Write [BHN08]. Pennsylvania [IEE05, IEE08]. perfect [CLV03, FH00, HK01, Lan04]. Performance [Upf03]. permit [Mey05]. Permutations [BPR^{+}08]. persistence [ELZ00]. PEZ [GKR01]. Phase [ACO08]. Philadelphia [IEE08]. Phylogeny [Roc08, AC05]. pigeonhole [BOBP^{+}02]. Pittsburgh [IEE05]. place [FG03]. placement [GMM00]. Planar [ABG06, DR01, FR01, GV09, Mar07b, Pat06, AZ03, AMM00, BJ02, CKS04, Kle05a, Kle05b, Str00, Tho01, NS06]. Planarity [iK09]. planes [BOGH^{+}03]. Planning [PT07, Str00]. Plans [AGM08]. Play [CMM06, BS05]. Player [CD06, AKV05]. Point [ABG06, Cha06, GT07, Pat06, AMM00, VVZ02]. Pointer [VW07]. Points [EY07, GLM06]. Pollard [KMT07]. polygonal [CDR00]. Polylogarithmic [Baz07, FK00]. Polynomial [Aar08, Amb03, AMR09, BZ09, BHKK08, BKM08, DK08, DGW07, KS09a, OR00, Roc08, Ajt02, AR01a, BS05, CS04, CGN01, CLV03, FS04, Jai04, Pak00, PV05]. Polynomial-Time [BKM08]. polynomially [IS01]. Polynomials [BSKR06, BV07, GKS07, KL08, Ale02, BGL03, Bur01, IK00, JPRZ04, KR04]. Polytopes [Ver06]. Popping [IKMT07]. portfolios [KV00]. possibility [DOPS04]. posted [KL03]. posted-price [KL03]. Postselection [Rei06]. potentially [CM04]. Power [AGIK07, ABK^{+}02a, Cha07, DD09, EOW08, LNP07, AIS04, BW03, Hal04, Wat02]. power-down [AIS04]. powerful [vDMV01]. pre [AKKV05]. pre-processing [AKKV05]. Precedence [AMS07]. predicates [AGS03, GMR01]. prediction [BEK05]. Preemptive [BCK^{+}07]. Preprocessing [KR06]. Pretty [OH08]. price [ADK^{+}04, KL03]. prices [HS01]. Pricing [CHK09, HS02, KK04]. Primal [BBN07, BN06, DPSV02, PT03]. Primal-Dual [BBN07, BN06, PT03]. primal-dual-type [DPSV02]. prime [JPRZ04]. primitive [ADRR04].

Primitives [BMG07]. principles [BOBP^{+}02]. priority [Tho02]. Privacy [JRS02, MT07]. Private [AMTdW00, Lan04, RY06, BIKR02]. Privately [KLN^{+}08]. Prize [Ano05i, ABHK09]. Prize-Collecting [ABHK09]. Probabilistic [Ung09, BEK05, FS04]. Probability [Kan09, OSZ03]. probe [CR04]. Problem [GV09, ORSS06, RRSS07, AR03, AKKV05, ACKZ05, BCvD05, BD03, CSWY01, COR01, CM04, DGV04, FJW01, FGMP02, GKS05, GKPR03, IW03, Kho03, Kho04a, Kle05b, KGR02, Mar05, MP00, Mey05, NPS03, Pet02, Sku00]. Problems [BK09b, BS06, CM08, CS08, FS08, GKTW09, GI09, Har06, HSV09, KPR^{+}09, Koi06, Moi09, Nut09, PT06, Von09, AR04, ASS05, BCS06, BBM01, Blu03, BT03, CGK^{+}00, COR01, CN02, DGRS05, FV06, FG02, Fri02, GK02, Gro03, GMM00, KKPS00, KV05a, MS03, Reg02, Shi02]. procedures [BOGH^{+}03, KV05b]. Proceedings [Ano06h, Ano06i, Ano07h, Ano08f, Ano08g, Ano09f, Ano09g, IEE00, IEE01, IEE02, IEE03, IEE04, IEE07, IEE08, IEE09]. processes [Upf03]. processing [AKKV05]. Product [DG08, IJK06, Ung09, ALW01, BCvD05, FOS05, KSdW04, Pak00, RVW00, Shp01]. Profit [CS08]. Profit-maximization [CS08]. Program [Ano09d, ABL02]. Programming [Chl07, AHK05, Cha02, FK02, LS05, NS05]. Programs [KY07, Rei09, ABE^{+}05, BW03, BEK05, CW04a, CS02b, KY01, New00]. Projections [Sar06]. Proof [DMS07, ABSRW00, DR04, PT03]. Proofs [BSN08, Cho08, JUW09, KR06, NPS03, OHP08, Wat00]. properties [Ajt02, AS05, GT01, MPS03, Wat00]. property [CS02b, FKR^{+}02]. propositional [ABSRW00]. Protection [ACSZ07]. protocol [AKU03]. Protocols [HHRS07, KLRZ08, KK02, BCNP04, BGL03, Can01]. Prover [OHP08]. Provers [KRT08, OHP08]. Providence [IEE07]. Proving [AGS03, ABL02]. proximity [MN06]. pseudo [SU01]. pseudo-random [SU01]. Pseudorandom [ABSRW00, BV07, RTTV08, Ind00]. PSPACE [HP07, JUW09]. PSPACE-completeness [HP07]. PTAS [Kho04b]. public [GKM^{+}00]. Publisher [Ano09h].

quadratic [ABE^{+}05, CW04a]. qualitative [CGW03]. Quantum [AA03, Aar08, AGIK07, Amb04, ACR07, BARdW08, BOCG^{+}06, BFK09, BCL^{+}06, CK09a, CSV07, GI09, JUW09, KSdW04, Moc04, OH08, OHP08, Raz05, Reg02, RT07, Rei09, Shi02, SY06, Sze04, AR03, AvDK^{+}04, AMTdW00, Amb03, BCvD05, BCG^{+}02, CW00, DFSS05, HH00, JRS02, JRP03, Kla01, RSV00, Wat00, Wat02, vDMV01]. quantum-storage [DFSS05]. Quasi [KN07a]. Quasi-Concave [KN07a]. queries [BBK04, YZ05]. Query [AGM08, Ber09, DH09, Fel09, HK01, KR05, MR08, Rei09, SY06, Amb03]. queues [Tho02]. queuing [BG03].

radii [VVZ02]. radio [CGR00, CR03, KP02]. radius [PV05]. Ramsey [BBM01, MN06]. Ramsey-type [BBM01]. Random [ACL01, Ajt02, AN08, BBS08, CM08, CMM07b, GM07a, GMR08, KS07, KM09, MS04a, Sar06, Ver06, AS00, AM02, ANP03, ABS03, ABK^{+}02a, AKU03, Bar02, BVV05, BFF^{+}01, BMOS03, CM04, DHM04, FKO06, Fri02, GGN03, Lan04, McS01, MS03, MPS04, NPS03, Rud06, SU01]. randomised [CR04]. Randomization [DD09]. Randomized [BBN07, Dot09, GT07, KSSV00, RV06, AMSZ03, BSSV00]. Randomizing [IK00]. Randomly [DF01, DFHV04, HV03]. Randomness [KLR09, Tao07, Vad02, BIW04, DOPS04, FH00, RSW00, TV00, WX05]. randomness-efficient [WX05]. Range [AAL09, DG08, ABR00]. Rank [BC08, BOGH^{+}03, DGW07, RS08, Yus08]. rapid [Hay06]. Rapidly [CDG^{+}02]. rates [BG03]. ratio [CGK04]. Rational [IML05]. Rationalizing [KU09]. reachability [RZ02, Tho01]. Read [BHN08]. Read/ Write [BHN08]. Real [BZ09, Bra05, DI01, FHS05]. real-world [FHS05]. realizing [Bar02]. reasoning [IK03]. recognition [McC01]. recognizing [CLV03]. Reconstructing [GKS07]. Reconstruction [BCMR06, GM07a, KPS08, Roc08]. Recovery [IR08]. recursion [BEK05]. recursive [CP05]. recycler [FH00]. Redondo [IEE00]. Reducibility [KPR^{+}09]. Reduction [AIP06, BC03, CS02a]. Reductions [CLX08, KL08, BT03, MR04]. Reed [Ale02, BSKR06, CW04b, TSZS01]. Reference [CPS07]. Refuting [Fei07]. regarding [GT01]. regions [AA03, ELRS02]. regret [KL03]. regular [KLSS03]. Regularity [BW09, HNR05]. related [BBM01, COR01, GK02]. Relational [AGM08]. relationship [GKM^{+}00]. relative [JRS02]. Relaxations [Chl07, RS09b]. relaxed [ABK^{+}05, BCNP04]. remainder [GSS00]. rent [GKPR03, KGR02]. rent-or-buy [GKPR03, KGR02]. Reordering [EOW08]. repeated [RSW00]. Repetition [Hai09, Raz08]. Repetitions [BHH^{+}08]. replacement [HP01b, Pak00]. Reporting [AAL09]. representation [IK00]. Representations [DLM^{+}07, AR02b]. Requirement [Roc08]. Resettability [DGS09]. Resettably [BGGL01]. Resettably-sound [BGGL01]. Resilient [DP07b, DP08b, KZ03]. Resolution [AR01b, BSN08, CM08, HP07, DR01, MS03, SBI02]. Resolving [DGS09]. Resource [BCK^{+}07, KK00]. restrictions [SBI02]. Results [AMS07, Cha06, FGKP06, Kan09, ML08, BDP03, Blu03, FGMP02, KK02, Kho02, KKMO04, KY01, SW00]. Retrieval [HSV09, RY06, BIKR02]. Revenue [CHK09]. Reviewers [Ano05k, Ano06j, Ano07i, Ano08h, Ano09i]. reward [BCK^{+}03]. Rewards [GM07b]. rewrite [CGN01]. Rho [KMT07]. Rhode [IEE07]. ribbons [AKKR02]. Rich [PT06]. robot [Str00]. robust [DGRS05]. Rome [IEE04]. Rooted [KL06a]. Rooted-Orientations [KL06a]. Roster [Ano08i]. Roughly [BCMR06]. Round [GKKO07, HHRS07, RS09a, Bar02, IK00, JRP03, PRS02]. round-complexity [PRS02]. round-efficient [IK00]. Rounding [BHH^{+}08, KORW08, LV06, FJW01, GKPS02, GRS04b]. rounds [PR03]. Routing [BN06, ER09, AFGZ01, GKR01, RT00, Vet02]. rows [CDG^{+}02]. Ruling [Kho04b]. rumor [KSSV00]. run [BEK05].

S [Cai01]. Safraless [KV05b]. salesman [Fri02]. samplable [TV00]. Sampled [BCSE^{+}07]. sampler [WX05]. samplers [TSU06]. Sampling [DR08, LV06, MR00, SVV07, SF08, SS05, CDG^{+}02, FH00, HV03, LV03a, MRS05]. Sampling-based [SF08, SS05]. SAT [AS00, AM02]. Satisfaction [BK09b, CM08, BD03, CM04, Gro03, MS03]. Satisfiability [AR02a, ANP03, CM04, FKO06]. saves [GRS04b]. Scalable [KSSV06]. scaling [BKP04]. Scheduling [AMS07, AZ02, BCK^{+}07, EOW08, AMSZ03, AFGZ01, COR01, CGKN04, KM05, SW00]. Scheme [BKM08, Kle05a]. Schemes [DDDR08, Mar07b, OR00]. Schrijver [GMPT07]. Science [IEE00, IEE01, IEE02, IEE03, IEE04, IEE05, IEE06, IEE07, IEE08, IEE09, Ano01, Ano05a, Ano05d, Ano05b, Ano05c, Ano06b, Ano06c, Ano06a, Ano06d, Ano07b, Ano07c, Ano07a, Ano07d]. scientist [Pap01]. SDP [CKN09, KO06, KS09b, RS09b]. SDPs [GMPT07]. Search [AM08, GT07, OP06, OH08, PTW08, SY06, AA03, AFKM01, AZ03, DHIP04]. Searching [OP06, ABR00, BBF^{+}03, CR04, FG04]. second [GKS00]. second-order [GKS00]. Secret [DP07b]. Secure [BOCG^{+}06, KSSV06, IK00, IML05, Lin03, PR03]. security [Can01, GK03]. seed [GRS04a]. seen [Gro03]. selection [CGK^{+}00, COR01, GK01, VV02]. Self [Dot09, AKKR02]. Self-Assembly [Dot09, AKKR02]. selfish [FJM04, KK04, RT00, Rou01]. Semi [ALW01, MPS04]. Semi-direct [ALW01]. semi-random [MPS04]. Semidefinite [Chl07, AHK05]. semidirect [BCvD05]. Sensor [GV09]. Separable [CDDT09]. Separating [BW03]. Separation [BSN08]. Sequence [Roc08]. Sequential [You01, BEK05]. Service [ASW06]. Set [BH06, GRW06, GGL^{+}08, KW09, BCNP04, Bul02, JRP03, RSV00]. Set-Disjointness [KW09]. set-up [BCNP04]. Sets [BATS09, DW08, DKSS09, KW09, RTTV08, KSW04, MRK03, Sri01, VVZ02]. Settling [CD06, CDDT09]. severely [BN02]. shadow [DS05]. Shallow [DES08]. Shallow-Low-Light [DES08]. Shamir [GK03]. shape [HPV01]. Shapes [Dot09]. shared [Bar02]. Sharing [DP07b, GKPR03]. Sharp [Aus07]. shop [SW00]. Shops [MS08a]. Short [BSN08]. Shortest [Ber09, DI01, FR01, HS01, HS02, Kho03, Kho04a, RZ04]. Shuffling [MPS04]. Side [ALS^{+}08, BYBJK06, Gro03]. sided [AS05]. Sign [RS08]. Sign-Rank [RS08]. Signatures [BMG07]. Simple [SU01, Cha03, ELRS02, Gaf02, GKPR03, Hay06, KSS04]. Simpler [iKMR08].

Simplex [KY07, Ver06, DS05].

simplification [ELZ00]. simulate [IS01]. Simulated [LV03b, SVV07]. Simulating [RT07]. Simulation [DGS09, Bar01, BS05]. Simultaneous [DGS09, BGL03, CSWY01]. Single [CS08, CK08, DDDR08, GP09, HS05, Hay06, Sku00]. Single-minded [CS08]. Single-Parameter [DDDR08]. single-site [HS05, Hay06]. Single-Source [CK08]. Sink [GMV05, SW01]. site [HS05, Hay06]. Size [ACR07, AGM08, BHN09, Moi09, RRSS07, RSY07, HP01b]. Sketches [ABIW09]. Sketching [HNO08]. Skew [LLW08]. skewed [DHM04, KM00b]. Sliding [BO07]. Small [AR02b, BK06, BATS09, KW09, ABS03, KSW04, Mar05, New00, SBI02, WN06]. Small-Bias [BATS09]. Smooth [BO07]. Smoothed [AV06, AMR09, CDT06, Fei07, KST09, KN07a, RT09, Ver06, BLMS^{+}03, DS05]. Sobolev [JS02]. societies [Vet02]. Soft [GSS00, Ale02]. Soft-decision [GSS00]. Solomon [Ale02, BSKR06, CW04b]. solvability [Gaf02]. Solving [BS06, ST03]. Some [ML08]. Sorkin [NPS03]. sort [BM09]. Sorting [GK01, ADRR04, FG03, FG04, HT02, Tho02]. sound [BGGL01]. Source [AFGZ01, CK08, GP09, KLR09, Sku00]. Sources [BS06, DGW07, RY08, BIW04, GRS04a, GR05, KZ03, PY00]. Space [BSN08, BGH07, HSV09, JW09, BSSV00, HT02, HSS03, KSdW04]. Space-Efficient [BGH07, HSV09]. Spaces [BCSE^{+}07, ML08, BBM01, KL06b]. Spacious [BSN08]. Span [Rei09, BW03]. Spanners [DES08, Woo06]. Spanning [ABN08, Goe06, KM09, Pet02]. Sparse [Che07, IR08, KS07, SVY09, ST03]. Sparsest [AMS07, CKN09, She09a, AHK04]. Sparsification [Yus08]. spatial [AA03, GMP04]. Spectral [BLR08, DHM04, JS02, McS01, Spi07, BL04, KVV00, VW02]. speed [AZ02, BKP04, Sze04]. speed-up [Sze04]. Speedups [HP07]. Spherical [KORW08]. Spider [Nut09]. Spider-Cover [Nut09]. spin [Hay06]. Sponsored [AM08]. Sponsors [Ano05g]. SPP [AK02]. spreading [KSSV00]. Stability [AKU03, KPR^{+}09, ADK^{+}04, MOO05]. Stable [Ind00, BFG01]. stage [SS05]. staircase [MR00]. standard [AvDK^{+}04]. states [JRS02]. Static [CFLM02]. Statistical [Fel09, NOV06, Wat02]. Statistically [HHRS07]. Statistically-Hiding [HHRS07]. statistics [BYJKS02]. stealing [BFG01]. Steiner [ABHK09, BKM08, GK02, KM00a, KL06a, Lau04]. steps [SvS04]. Stigum [BCMR06]. Stochastic [FMMM09, GM07b, KRR^{+}00, SS04, DGV04, GRS04b, SS05]. storage [DFSS05]. Straightening [CDR00]. Strategic [ASW06]. Strategies [CHK09, AIS04]. Strategy [DGS09, PT03]. Stratified [BCSE^{+}07]. Stream [JW09, BYJKS02, Ind00]. Streaming [GG07, HNO08, PP09, ADRR04]. Streams [BHN08, GMMO00]. Stretch [ABN08, BK06]. Strict [BOCG^{+}06]. String [CPS07, HSV09, Bar02, CFLM02]. strings [ABK^{+}02a]. Strong [CK09a, GMP04, RS09b, RRSS07, Raz08, AZ03, KSdW04, MRS05]. Strongly [BG07]. Structure [ARV06, Tao07, VV02]. structured [GK01]. Structures [CSV07, Har06, ABR00, FM00, MN06, Pat08a]. Structuring [FLMM05]. study [Vad04]. style [Pet02]. Sub [MR08]. Sub-Constant [MR08]. Subgraph [CMM07b, GMR08, Kho04b]. subgraphs [Alo01]. subgroup [BCvD05]. Sublinear [KPS08]. Sublogarithmic [Pat06]. Submodular [CG08, FMV07, GKTW09, IN09, SF08, Von09]. Subsequence [GG07]. Subsets [RTTV08]. Subspace [BSKR06]. substring [Mar05]. Succinct [KR06, Wat00, Lan04]. Succincter [Pat08b]. succinctness [FLMM05]. Sudan [PV05]. Sum [BCK^{+}07, CSWY01]. Sums [CJLV08]. Sunspots [CPS07]. Super [BSSV00, BS05]. Super-linear [BSSV00]. super-polynomial [BS05]. Supply [CS08]. Support [RRSS07]. Surface [BNN06, iKMR08, KOS08, dVL02]. survey [Sud01]. Survivable [CK09b]. Switch [AMSZ03, FGMS06]. Switching [SBI02]. Symmetric [BMG07, BGL03, BCMR06, She08, Fri02, MRS05, ST03]. Symmetry [Von09]. Symposium [Ano01, Ano05a, Ano05d, Ano05b, Ano05c, Ano06b, Ano06c, Ano06a, Ano06d, Ano07b, Ano07c, Ano07a, Ano07d, IEE00, IEE01, IEE02, IEE03, IEE04, IEE05, IEE06, IEE07, IEE08, IEE09]. Synchronization [LLW08]. Syntactically [RSY07]. synthesis [AFKM01]. system [dVL02]. Systems [AGIK07, CW09, GRW06, BBM01, CGN01, Hay06, ST03, WN06].

Table [Ano05c, Ano08j, Ano09j]. tables [CDG^{+}02]. tag [WN06]. tailed [DHKS05]. tale [BLS05]. task [BBM01]. tautologies [AR02a, DR01]. technique [FH00, KS02]. Techniques [Cha05]. temperature [BKP04]. term [CGN01]. Test [BK09a]. Testable [KS07, AS05, BSS03, GS02, KL05]. testers [CS02b, DR04]. Testing [ADPR00, Alo01, BFR^{+}00, BFF^{+}01, Che07, CS07, DLM^{+}07, DG08, FKR^{+}02, HKNO09, JPRZ04, KR04, KS09a, New00, BOT02, GT01]. Tetrahedra [ES07]. text [HSS03]. Their [FFS06, Ajt02, DN00, Vad02]. Theorem [BZ09, CM08, Hai09, Wag06, BBM01, Bul02, BD03, CFLM02, DR04, JRS02, Lau04, Raz05]. Theorems [KL06a, Ung09, FS04, GT01, Hal04, KN05, KSdW04]. theoretic [BIKR02, CU03, CKSU05]. theoretical [Pap01]. Theory [AM08, HNO08, Spi07, DHK05, Pap01, Sud01]. Three [AAL09, ES07, Ezr08, GT01, SY06, Bul02, Cha07]. three-element [Bul02]. Threshold [BHN09, She09b, Ung09, AM02, CM04, DM02, Rei06]. thresholds [KO02]. Tight [ABN08, BCMR06, DES08, ES07, GT07, HHRS07, IW03, KY01, Mos08, Rei09, Kol01]. Tiling [GI09]. Time [BC08, Ber09, BBS08, BHKK08, BKM08, BM09, Cha06, DK08, HP07, iKMR08, iK09, LNP07, NO08, Pat06, Ajt02, ACR07, AZ02, AHK04, BSSV00, CGN01, DS02, FR01, FS04, GK07, GRS04b, HT02, HSS03, Jai04, KKLV00, KSdW04, Kle05a, KP02, KSS04, MSW03, McC01, OR00, PV05, ST03, WN06]. time-and-space [HSS03]. time-space [BSSV00, KSdW04]. time-varying [AZ02]. Time/Space [HP07]. tiny [FLS02]. Title [Ano05d, Ano06c, Ano07c, Ano08k, Ano08l, Ano09k, Ano09l]. TOC [Ano06d, Ano07d]. Toda [BZ09]. Tolerance [KPS08, ACK^{+}00]. Tolerant [BCL^{+}06, GPV06]. Tolls [FJM04]. Tools [Cha05]. Top [HSV09]. Top- [HSV09]. Topological [CJLV08, ELZ00]. topology [CR03, KP02, MPS03]. tossing [Bar02]. tours [CGRT03]. tractability [GKS00]. tractable [AR01b]. trade [PY00]. trade-offs [PY00]. Tradeoff [ALGG09, BSSV00]. Tradeoffs [CMM07a, KSdW04]. traffic [Vet02]. Transdichotomous [Cha06]. transfer [GM00, GKM^{+}00]. transform [CW00, HH00]. Transitions [ACO08]. transitive [DI00, San04]. Translationally [GI09]. transpositions [MPS04]. Trapdoor [BPR^{+}08, GMR01]. Traveling [GKR01]. travelling [Fri02]. Tree [iKMR08, AC05, DHIP04, Lau04, OSSS05, Pet02, ABHK09]. Tree-Width [iKMR08]. Trees [ABN08, DES08, Goe06, KM09, OP06, BDFC00, CGRT03, FLMM05, FGMP02, KM00a, KMP01, MSW03]. treewidth [Mar07a]. triangles [PT00]. Triangulation [KSW04]. Triangulations [BM09]. Truthful [AT01, DDDR08, LS05, PSS08, KK05, LMN03]. Tseitin [AR02a]. TSP [ABHK09, BCK^{+}03, CGK04, KLSS03, Kle05a]. turbo [FK02]. turbo-like [FK02]. Turing [OSZ03, WN06]. tutorial [Gol02, Sud01]. Tutte [BHKK08]. Two [CD06, JUW09, MR08, She09b, SY06, AKV05, BLS05, MMP00, PR03]. Two-Message [JUW09]. two-party [PR03]. Two-Player [CD06, AKV05]. Type [Moi09, ORSS06, BBM01, DPSV02, Hal04, KV05a]. types [HR00]. Typical [Kan09].

UGC [KO06]. UGC-hardness [KO06]. Unbounded [She08]. Unbounded-Error [She08]. unconditional [Vad04]. Uncrossable [Nut09]. undirected [ACKZ05, RZ04]. Uniform [CHKS06, IJK06, CHK09, DRS02, Gaf02, Kho02]. Union [ES07, Ezr08, PT00]. Unique [BHH^{+}08, CMM06, KRT08, SW01, AC02, Kho05, KV05a, Tre05, RS09b]. unique-neighbor [AC02]. unit [DLN08]. Universal [BFK09, AR02b, KV00, Lin03, WN06]. Universality [ACK^{+}00, DS02]. Universally [BCNP04, Can01]. unknown [CR03, KP02]. unless [AR01b]. unsplittable [Sku00]. Update [Ber09, AHK05]. Updates [PT07]. Upper [GK07, Wag06, ABS03, Aue00, Kol01]. USA [IEE01, IEE05, IEE08, IEE09]. Use [CPS07]. user [AZ02]. user-dependent [AZ02]. users [FJM04, KK04, Rou01]. Using [Aue00, CMM06, Chl07, DH09, Gab00, Kan09, AGS03, ACL06, AHK05, BIW04, DR08, KSW04, MR00, Tre03, YZ05]. Utilities [CDDT09, Jai04].

Value [BHN08, KL03]. Valued

[BARdW08, WX05]. Vancouver [IEE02]. variable [KKMO04, OSSS05]. variables [BFF^{+}01]. varying [AZ02]. VCG [KK05]. vector [AKKV05, FLS02, Kho03, Kho04a]. Vectors [ACL06]. Vegas [IEE01]. verification [Pet02]. Verifiers [Mei09]. version [HNR05]. Vertex [BHKK08, CK08, CK09b, GMPT07, DS05, Gab00]. Vertex-Connectivity [CK09b]. Vertex-Exponential [BHKK08]. vertical [Kol01]. Vertices [iK09]. Via [BN06, AFKM01, AMSZ03, BBHM05, BS05, BLR08, DPSV02, FK02, GKPR03, HPV01, HNO08, KN05, KOS08, Koi06, LS05, MT07, MS04b, NO08, Nut09, PT03, RSW00, San04, Sar06, Yus08]. Vickrey [HS02, HS01]. violations [Cha02]. volume [LV03b]. Voronoi [Cha06, HP01b]. vs [Ale03, Amb03, BL04].

W [AR01b]. wait [SW00]. walk [ABS03, Amb04]. Walks [Ver06, BMOS03, CP05, MR00]. waste [KM00b]. waves [RVW00]. Way [NOV06, VW07, Mic02]. weak [Moc04]. weaker [BOBP^{+}02]. Web [AFKM01, CKL^{+}09, KRR^{+}00, PY00]. weight [FR01]. Weighted [BBN07, FFS06, Kle05a]. weights [AHK05, DI01]. Welfare [CG08]. Well [OH08, Vov02]. well-calibrated [Vov02]. White [HP07]. Width [BK09b, iKMR08, AR02a, New00]. win [AKV05]. win-lose [AKV05]. Windows [BO07]. Wireless [DR08, AZ02]. Wise [AN08]. Without [BGH07, ABL02]. Witnesses [FKO06]. work [BFG01]. work-stealing [BFG01]. world [FHS05]. Worst [ABX08, AV06, KL08, MR04, BT03, Mic02]. Worst-Case [ABX08, AV06, MR04, BT03, Mic02]. worth [HS01, HS02].

XOR [Tre03]. XORSAT [DM02].

zag [ALW01, RVW00]. Zaps [DN00]. Zero [BPS06, Che07, Gol02, KR06, MRK03, NOV06, BGGL01, BLV03, PRS02, Vad04, Wat02]. Zero-Knowledge [KR06, NOV06, Gol02, MRK03, BGGL01, Wat02]. zig [ALW01, RVW00]. zig-zag [ALW01, RVW00]. ZPP [Cai01].

References

Aaronson:2003:QSS

[AA03] S. Aaronson and A. Ambainis. Quantum search of spatial regions. In IEEE [IEE03], pages 200–209. CODEN ASF-PDV. ISBN 0-7695-2040-5. ISSN 0272-5428. LCCN QA76 .S979 2003. URL http:/

                                                   /ieeexplore.ieee.org/iel5/
                                                   8767/27770/01238194.pdf?isnumber=
                                                   27770&prod=CNF&arnumber=1238194&
                                                   arSt=+200&ared=+209&arAuthor=
                                                   Aaronson%2C+S.%3B+Ambainis%
                                                   2C+A.;   http://ieeexplore.
                                                   ieee.org/xpls/abs_all.jsp?
                                                   isnumber=27770&arnumber=1238194&
                                                   count=66&index=21.  IEEE

Computer Society Order Num-

ber PR02040.

Afshani:2009:ORR

[AAL09] P. Afshani, L. Arge, and K. D. Larsen. Orthogonal range reporting in three and higher dimensions. In IEEE [IEE09], pages 149–158. ISBN 0-7695-3850-9. ISSN 0272-5428. LCCN QA76 .S95 2009. IEEE Computer Society order number P3850.

Aaronson:2008:PMQ

[Aar08] S. Aaronson. The polynomial method in quantum and classical computing. In IEEE [IEE08], page 3. ISBN 0-7695-3436-

8. ISSN 0272-5428. LCCN QA76 .S95 2008. URL http://

ieeexplore.ieee.org/servlet/
opac?punumber=4690923. IEEE

Computer Society order number P3436.

Agarwal:2001:CMF

[AAS01] P. K. Agarwal, B. Aronov, and M. Sharir. On the complexity of many faces in arrangements of circles. In IEEE [IEE01], pages 74–83. CODEN ASFPDV. ISBN 0-7695-1390-5, 0-7695-1391-3 (case), 0-7695-1392-1 (microfiche). ISSN 0272-5428. LCCN ????

Afshani:2009:IOG

[ABC09a] P. Afshani, J. Barbay, and T. M.

Chan. Instance-optimal geometric algorithms. In IEEE [IEE09], pages 129–138. ISBN 0-7695-3850-9. ISSN 0272-5428. LCCN QA76 .S95 2009. IEEE Computer Society order number P3850.

Azar:2009:CLD

[ABC^{+}09b] Y. Azar, B. Birnbaum, L. E.

Celis, N. R. Devanur, and Y. Peres. Convergence of local dynamics to balanced outcomes in exchange networks. In IEEE [IEE09], pages 293–302. ISBN 0-7695-3850-9. ISSN 0272-5428. LCCN QA76 .S95 2009. IEEE Computer Society order number P3850.

Arora:2005:NAQ

[ABE^{+}05] S. Arora, E. Berger, Hazan Elad, G. Kindler, and M. Safra. On non-approximability for quadratic programs. In IEEE [IEE05], pages 206–215. ISBN 0-7695-2468-0, 0-7695-2362-5. ISSN 0272-5428. LCCN QA76 .S979 2005. IEEE Computer Society order number P2468.

Alekhnovich:2004:LA

[ABF^{+}04] M. Alekhnovich, M. Braverman, V. Feldman, A. R. Klivans, and T. Pitassi. Learnability and automatizability. In IEEE [IEE04], pages 621–630. CODEN ASFPDV. ISBN 0-7695-2228-9. ISSN 0272-5428. LCCN QA276. URL http:/

/ieeexplore.ieee.org/iel5/
9430/29918/01366282.pdf?isnumber=
29918&prod=CNF&arnumber=1366282&
arSt=+621&ared=+630&arAuthor=
Alekhnovich%2C+M.%3B+Braverman%
2C+M.%3B+Feldman%2C+V.%3B+
Klivans%2C+A.R.%3B+Pitassi%
2C+T.;   http://ieeexplore.
ieee.org/xpls/abs_all.jsp?
isnumber=29918&arnumber=1366282&
count=64&index=63.  IEEE

Computer Society Order Number P2228.

Arge:2006:IDP

[ABG06] Lars Arge, Gerth Stolting Brodal, and Loukas Georgiadis. Improved dynamic planar point location. In IEEE [IEE06], pages 305–314. ISBN 0-7695-2720-5, 0-7695-2362-5. ISSN 0272-5428. LCCN QA76 .S974 2006. URL http://

ieeexplore.ieee.org/servlet/
opac?punumber=4031329. IEEE

Computer Society Order Number P2720.

Archer:2009:IAA

[ABHK09] A. Archer, M. H. Bateni, M. T.

Hajiaghayi, and H. Karloff. Improved approximation algorithms for Prize-Collecting Steiner Tree and TSP. In IEEE [IEE09], pages 427–436. ISBN 0-7695-3850-9. ISSN 0272-5428. LCCN QA76 .S95 2009. IEEE Computer Society order number P3850.

Andoni:2009:ESE

[ABIW09] A. Andoni, Khanh Do Ba, P. Indyk, and D. Woodruff. Efficient sketches for earth-mover distance, with applications. In IEEE [IEE09], pages 324–330. ISBN 0-7695-3850-9. ISSN 0272-5428. LCCN QA76 .S95 2009. IEEE Computer Society order number P3850.

Allender:2002:PRS

[ABK^{+}02a] E. Allender, H. Buhrman, M. Koucky, D. van Melkebeek, and D. Ronneburger. Power from random strings. In IEEE [IEE02], pages 669–678. CO-DEN ASFPDV. ISBN 0-7695-1822-2. ISSN 0272-5428. LCCN QA267. URL http:/

/ieeexplore.ieee.org/iel5/
8411/26517/01181992.pdf?isnumber=
26517&prod=CNF&arnumber=1181992&
arSt=+669&ared=+678&arAuthor=
Allender%2C+E.%3B+Buhrman%
2C+H.%3B+Koucky%2C+M.%3B+
van+Melkebeek%2C+D.%3B+Ronneburger%
2C+D.;   http://ieeexplore.
ieee.org/xpls/abs_all.jsp?
isnumber=26517&arnumber=1181992&
count=82&index=66.  IEEE

Computer Society Order Num-

ber PR01822.

Alon:2002:LHM

[ABK^{+}02b] N. Alon, R. Beigel, S. Kasif, S. Rudich, and B. Sudakov. Learning a hidden matching. In IEEE [IEE02], pages 197–206. CODEN ASFPDV. ISBN 0-7695-1822-2. ISSN 0272-5428. LCCN QA267. URL http:/

                                        /ieeexplore.ieee.org/iel5/
                                        8411/26517/01181943.pdf?isnumber=
                                        26517&prod=CNF&arnumber=1181943&
                                        arSt=+197&ared=+206&arAuthor=
                                        Alon%2C+N.%3B+Beigel%2C+R.
                                        %3B+Kasif%2C+S.%3B+Rudich%
                                        2C+S.%3B+Sudakov%2C+B.;
                                        http://ieeexplore.ieee.org/
                                        xpls/abs_all.jsp?isnumber=
                                        26517&arnumber=1181943&count=
                                        82&index=21.  IEEE Com-

puter Society Order Number

PR01822.

Abraham:2005:MER

[ABK^{+}05] I. Abraham, Y. Bartal, J. Kleinberg, T.-H. H. Chan, O. Neiman, Kedar Dhamdhere, A. Slivkins, and Anupam Gupta. Metric embeddings with relaxed guarantees. In IEEE [IEE05], pages 83–100. ISBN 0-7695-2468-0, 0-7695-2362-5. ISSN 0272-5428. LCCN QA76 .S979 2005. IEEE Computer Society order number P2468.

Arora:2002:PIG

[ABL02] S. Arora, B. Bollobas, and L. Lovasz. Proving integrality gaps without knowing the linear program. In IEEE [IEE02], pages 313–322. CO-DEN ASFPDV. ISBN 0-7695-1822-2. ISSN 0272-5428. LCCN QA267. URL http:/

/ieeexplore.ieee.org/iel5/
8411/26517/01181954.pdf?isnumber=
26517&prod=CNF&arnumber=1181954&
arSt=+313&ared=+322&arAuthor=
Arora%2C+S.%3B+Bollobas%2C+
B.%3B+Lovasz%2C+L.;   http:
//ieeexplore.ieee.org/xpls/
abs_all.jsp?isnumber=26517&
arnumber=1181954&count=82&

index=32. IEEE Computer Society Order Number PR01822.

Abraham:2008:NTL

[ABN08] I. Abraham, Y. Bartal, and O. Neiman. Nearly tight low stretch spanning trees. In IEEE [IEE08], pages 781–

790. ISBN 0-7695-3436-8. ISSN 0272-5428. LCCN QA76 .S95 2008. URL http://

ieeexplore.ieee.org/servlet/
opac?punumber=4690923. IEEE

Computer Society order number P3436.

Alstrup:2000:NDS

[ABR00] S. Alstrup, G. Stolting Brodal, and T. Rauhe. New data structures for orthogonal range searching. In IEEE [IEE00], pages 198–207. CO-DEN ASFPDV. ISBN 0-7695-0850-2, 0-7695-0851-0 (case), 0-7695-0852-9 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 2000. IEEE Computer Society Order Number PR00850.

Alekhnovich:2003:LUB

[ABS03] M. Alekhnovich and E. Ben-Sasson. Linear upper bounds for random walk on small density random 3-CNFs. In IEEE [IEE03], pages 352–361. CO-DEN ASFPDV. ISBN 0-7695-2040-5. ISSN 0272-5428. LCCN QA76 .S979 2003. URL http:/

/ieeexplore.ieee.org/iel5/
8767/27770/01238209.pdf?isnumber=
27770&prod=CNF&arnumber=1238209&
arSt=+352&ared=+361&arAuthor=
Alekhnovich%2C+M.%3B+Ben-
Sasson%2C+E.;   http://ieeexplore.
ieee.org/xpls/abs_all.jsp?
isnumber=27770&arnumber=1238209&
count=66&index=36.  IEEE

Computer Society Order Number PR02040.

Alekhnovich:2000:PGP

[ABSRW00] M. Alekhnovich, E. Ben-Sasson, A. A. Razborov, and A. Wigderson. Pseudorandom generators in propositional proof complexity. In IEEE [IEE00], pages 43–53. CODEN ASF-PDV. ISBN 0-7695-0850-2, 0-7695-0851-0 (case), 0-7695-0852-9 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 2000. IEEE Computer Society Order Number PR00850.

Applebaum:2008:BLB

[ABX08] B. Applebaum, B. Barak, and D. Xiao. On basing lower-bounds for learning on worst-case assumptions. In IEEE [IEE08], pages 211–

220. ISBN 0-7695-3436-8. ISSN 0272-5428. LCCN QA76 .S95 2008. URL http://

ieeexplore.ieee.org/servlet/
opac?punumber=4690923. IEEE

Computer Society order number P3436.

Alon:2002:EUN

[AC02] N. Alon and M. Capalbo.

Explicit unique-neighbor expanders. In IEEE [IEE02], pages 73–79. CODEN ASFPDV. ISBN 0-7695-1822-2. ISSN 0272-5428. LCCN QA267. URL

http://ieeexplore.ieee.org/
iel5/8411/26517/01181884.
pdf?isnumber=26517&prod=CNF&
arnumber=1181884&arSt=+73&
ared=+79&arAuthor=Alon%2C+
N.%3B+Capalbo%2C+M.;   http:
//ieeexplore.ieee.org/xpls/
abs_all.jsp?isnumber=26517&
arnumber=1181884&count=82&

index=9. IEEE Computer Society Order Number PR01822.

Ailon:2005:FTM

[AC05] N. Ailon and M. Charikar. Fitting tree metrics: Hierarchical clustering and phylogeny. In IEEE [IEE05], pages 73– 82. ISBN 0-7695-2468-0, 0-7695-2362-5. ISSN 0272-5428. LCCN QA76 .S979 2005. IEEE Computer Society order number P2468.

Alon:2007:FDP

[AC07] N. Alon and M. Capalbo.

Finding disjoint paths in expanders deterministically and online. In IEEE [IEE07], pages 518–524. ISBN 0-7695-3010-9. ISSN 0272-5428. LCCN QA76 .S974 2007. URL http://

ieeexplore.ieee.org/servlet/
opac?punumber=4389466. IEEE

Computer Society order number P3010.

Alon:2000:UT

[ACK^{+}00] N. Alon, M. Capalbo, Y. Kohayakawa, V. Rodl, A. Rucinski, and E. Szemeredi. Universality and tolerance. In IEEE [IEE00], pages 14–21. CODEN ASFPDV. ISBN 0-7695-0850-2, 0-7695-0851-0 (case), 0-7695-0852-9 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 2000. IEEE Computer Society Order Number PR00850.

Andrews:2005:HUE

[ACKZ05] M. Andrews, J. Chuzhoy, Sanjeev Khanna, and L. Zhang. Hardness of the undirected edgedisjoint paths problem with congestion. In IEEE [IEE05], pages 226–241. ISBN 0-7695-2468-0, 0-7695-2362-5. ISSN 0272-5428. LCCN QA76 .S979 2005. IEEE Computer Society order number P2468.

Aiello:2001:REM

[ACL01] W. Aiello, Fan Chung, and Linyuan Lu. Random evolution in massive graphs. In IEEE [IEE01], pages 510–519. CO-DEN ASFPDV. ISBN 0-7695-1390-5, 0-7695-1391-3 (case), 0-7695-1392-1 (microfiche). ISSN 0272-5428. LCCN ????

Andersen:2006:LGP

[ACL06] Reid Andersen, Fan Chung, and Kevin Lang. Local graph partitioning using pagerank vectors. In IEEE [IEE06], pages 475–486.

ISBN 0-7695-2720-5, 0-7695-2362-5. ISSN 0272-5428. LCCN QA76 .S974 2006. URL http://

ieeexplore.ieee.org/servlet/
opac?punumber=4031329. IEEE

Computer Society Order Number P2720.

Achlioptas:2008:ABP

[ACO08] D. Achlioptas and A. Coja-Oghlan. Algorithmic barriers from phase transitions. In IEEE [IEE08], pages 793– 802. ISBN 0-7695-3436-8. ISSN 0272-5428. LCCN QA76 .S95 2008. URL http://

ieeexplore.ieee.org/servlet/
opac?punumber=4690923. IEEE

Computer Society order number P3436.

Andoni:2008:HNN

[ACP08] A. Andoni, D. Croitoru, and M. Patrascu. Hardness of nearest neighbor under L-infinity. In IEEE [IEE08], pages 424– 433. ISBN 0-7695-3436-8. ISSN 0272-5428. LCCN QA76 .S95 2008. URL http://

ieeexplore.ieee.org/servlet/
opac?punumber=4690923. IEEE

Computer Society order number P3436.

Ambainis:2007:FSC

[ACR07] A. Ambainis, A. M. Childs, and B. W. Reichardt. Any AND-OR formula of size N can be evaluated in time N^{1/2} + o(1) on a quantum computer. In IEEE [IEE07], pages 363– 372. ISBN 0-7695-3010-9. ISSN 0272-5428. LCCN QA76 .S974 2007. URL http://

                                        ieeexplore.ieee.org/servlet/
                                        opac?punumber=4389466. IEEE

Computer Society order number

P3010.

Antonakopoulos:2007:BBN

[ACSZ07] S. Antonakopoulos, C. Chekuri, B. Shepherd, and L. Zhang. Buy-at-bulk network design with protection. In IEEE [IEE07], pages 634–644. ISBN 0-7695-3010-9. ISSN 0272-5428. LCCN QA76 .S974 2007. URL

                                        http://ieeexplore.ieee.org/
                                        servlet/opac?punumber=4389466.

IEEE Computer Society order

number P3010.

Anshelevich:2004:PSN

[ADK^{+}04] E. Anshelevich, A. Dasgupta, J. Kleinberg, E. Tardos, T. Wexler, and T. Roughgarden. The price of stability for network design with fair cost allocation. In IEEE [IEE04], pages 295–304. CODEN ASFPDV. ISBN 0-7695-2228-9. ISSN 0272-5428. LCCN QA276. URL http:/

                                        /ieeexplore.ieee.org/iel5/
                                        9430/29918/01366249.pdf?isnumber=
                                        29918&prod=CNF&arnumber=1366249&
                                        arSt=+295&ared=+304&arAuthor=
                                        Anshelevich%2C+E.%3B+Dasgupta%
                                        2C+A.%3B+Kleinberg%2C+J.%
                                        3B+Tardos%2C+E.%3B+Wexler%
                                        2C+T.%3B+Roughgarden%2C+T.
                                        ;   http://ieeexplore.ieee.
                                        org/xpls/abs_all.jsp?isnumber=
                                        29918&arnumber=1366249&count=

64&index=30. IEEE Computer Society Order Number P2228.

Alon:2000:TC

[ADPR00] N. Alon, S. Dar, M. Parnas, and D. Ron. Testing of clustering. In IEEE [IEE00], pages 240–250. CODEN ASF-PDV. ISBN 0-7695-0850-2, 0-7695-0851-0 (case), 0-7695-0852-9 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 2000. IEEE Computer Society Order Number PR00850.

Aggarwal:2004:SMA

[ADRR04] G. Aggarwal, M. Datar, S. Rajagopalan, and M. Ruhl. On the streaming model augmented with a sorting primitive. In IEEE [IEE04], pages 540–549. CODEN ASFPDV. ISBN 0-7695-2228-9. ISSN 0272-5428. LCCN QA276. URL http:/

/ieeexplore.ieee.org/iel5/
9430/29918/01366274.pdf?isnumber=
29918&prod=CNF&arnumber=1366274&
arSt=+540&ared=+549&arAuthor=
Aggarwal%2C+G.%3B+Datar%2C+
M.%3B+Rajagopalan%2C+S.%3B+
Ruhl%2C+M.;   http://ieeexplore.
ieee.org/xpls/abs_all.jsp?
isnumber=29918&arnumber=1366274&
count=64&index=55.  IEEE

Computer Society Order Number P2228.

Andrews:2001:SRS

[AFGZ01] M. Andrews, A. Fernandez, A. Goel, and L. Zhang. Source routing and scheduling in packet networks. In IEEE [IEE01], pages 168–177. CODEN ASF-PDV. ISBN 0-7695-1390-5, 0-7695-1391-3 (case), 0-7695-1392-1 (microfiche). ISSN 0272-5428. LCCN ????

Achlioptas:2001:WSH

[AFKM01] D. Achlioptas, A. Fiat, A. R. Karlin, and F. McSherry. Web search via hub synthesis. In IEEE [IEE01], pages 500–509. CODEN ASFPDV. ISBN 0-7695-1390-5, 0-7695-1391-3 (case), 0-7695-1392-1 (microfiche). ISSN 0272-5428. LCCN ????

Aharonov:2007:PQS

[AGIK07] D. Aharonov, D. Gottesman, S. Irani, and J. Kempe. The power of quantum systems on a line. In IEEE [IEE07], pages 373–383. ISBN 0-7695-3010-9. ISSN 0272-5428. LCCN QA76 .S974 2007. URL http://

                                        ieeexplore.ieee.org/servlet/
                                        opac?punumber=4389466. IEEE

Computer Society order number P3010.

Atserias:2008:SBQ

[AGM08] A. Atserias, M. Grohe, and D. Marx. Size bounds and query plans for relational joins. In IEEE [IEE08], pages 739– 748. ISBN 0-7695-3436-8. ISSN 0272-5428. LCCN QA76 .S95 2008. URL http://

                                        ieeexplore.ieee.org/servlet/
                                        opac?punumber=4690923. IEEE

Computer Society order number P3436.

Akavia:2003:PHC

[AGS03] A. Akavia, S. Goldwasser, and S. Safra. Proving hardcore predicates using list decoding. In IEEE [IEE03], pages 146–157. CODEN ASF-PDV. ISBN 0-7695-2040-5. ISSN 0272-5428. LCCN QA76 .S979 2003. URL http:/

/ieeexplore.ieee.org/iel5/
8767/27770/01238189.pdf?isnumber=
27770&prod=CNF&arnumber=1238189&
arSt=+146&ared=+157&arAuthor=
Akavia%2C+A.%3B+Goldwasser%
2C+S.%3B+Safra%2C+S.;   http:
//ieeexplore.ieee.org/xpls/
abs_all.jsp?isnumber=27770&
arnumber=1238189&count=66&

index=16. IEEE Computer Society Order Number PR02040.

Arora:2004:ASC

[AHK04] S. Arora, E. Hazan, and S. Kale.

O(\sqrt{log n}) approximation to SPARSEST CUT in Õ(n^{2}) time. In IEEE [IEE04], pages 238–

247. CODEN ASFPDV. ISBN 0-7695-2228-9. ISSN 0272-5428. LCCN QA276. URL http:/

/ieeexplore.ieee.org/iel5/
9430/29918/01366243.pdf?isnumber=
29918&prod=CNF&arnumber=1366243&
arSt=+238&ared=+247&arAuthor=
Arora%2C+S.%3B+Hazan%2C+E.
%3B+Kale%2C+S.;   http://
ieeexplore.ieee.org/xpls/
abs_all.jsp?isnumber=29918&
arnumber=1366243&count=64&

index=24. IEEE Computer Society Order Number P2228.

Arora:2005:FAA

[AHK05] S. Arora, E. Hazan, and S. Kale.

Fast algorithms for approximate semidefinite programming using the multiplicative weights update method. In IEEE [IEE05], pages 339–348. ISBN 0-7695-2468-0, 0-7695-2362-5. ISSN 0272-5428. LCCN QA76 .S979 2005. IEEE Computer Society order number P2468.

Andoni:2006:NOH

[AI06] Alexandr Andoni and Piotr Indyk. Near-optimal hashing algorithms for approximate nearest neighbor in high dimensions. In IEEE [IEE06], pages 459–468. ISBN 0-7695-2720-5, 0-7695-2362-5. ISSN 0272-5428. LCCN QA76 .S974 2006. URL http://

                                        ieeexplore.ieee.org/servlet/
                                        opac?punumber=4031329. IEEE

Computer Society Order Num-

ber P2720.

Applebaum:2004:CNS

[AIK04] B. Applebaum, Y. Ishai, and E. Kushilevitz. Cryptography in NC^{0}. In IEEE [IEE04], pages 166–175. CO-DEN ASFPDV. ISBN 0-7695-2228-9. ISSN 0272-5428. LCCN QA276. URL http:/

                                        /ieeexplore.ieee.org/iel5/
                                        9430/29918/01366236.pdf?isnumber=
                                        29918&prod=CNF&arnumber=1366236&
                                        arSt=+166&ared=+175&arAuthor=
                                        Applebaum%2C+B.%3B+Ishai%
                                        2C+Y.%3B+Kushilevitz%2C+E.
                                        ;   http://ieeexplore.ieee.
                                        org/xpls/abs_all.jsp?isnumber=
                                        29918&arnumber=1366236&count=

64&index=17. IEEE Computer Society Order Number P2228.

Andoni:2006:ODR

[AIP06] Alexandr Andoni, Piotr Indyk, and Mihai Patrascu. On the optimality of the dimensionality reduction method. In IEEE [IEE06], pages 449–458. ISBN 0-7695-2720-5, 0-7695-2362-5. ISSN 0272-5428. LCCN QA76 .S974 2006. URL http://

ieeexplore.ieee.org/servlet/
opac?punumber=4031329. IEEE

Computer Society Order Number P2720.

Augustine:2004:OPS

[AIS04] J. Augustine, S. Irani, and C. Swamy. Optimal powerdown strategies. In IEEE [IEE04], pages 530–539. CO-DEN ASFPDV. ISBN 0-7695-2228-9. ISSN 0272-5428. LCCN QA276. URL http:/

/ieeexplore.ieee.org/iel5/
9430/29918/01366273.pdf?isnumber=
29918&prod=CNF&arnumber=1366273&
arSt=+530&ared=+539&arAuthor=
Augustine%2C+J.%3B+Irani%
2C+S.%3B+Swamy%2C+C.;   http:
//ieeexplore.ieee.org/xpls/
abs_all.jsp?isnumber=29918&
arnumber=1366273&count=64&

index=54. IEEE Computer Society Order Number P2228.

Ajtai:2002:RLC

[Ajt02] M. Ajtai. Random lattices and a conjectured 0-1 law about their polynomial time computable properties. In IEEE [IEE02], pages 733–742. CO-DEN ASFPDV. ISBN 0-7695-1822-2. ISSN 0272-5428. LCCN QA267. URL http:/

/ieeexplore.ieee.org/iel5/
8411/26517/01181998.pdf?isnumber=
26517&prod=CNF&arnumber=1181998&
arSt=+733&ared=+742&arAuthor=
Ajtai%2C+M.;   http://ieeexplore.
ieee.org/xpls/abs_all.jsp?
isnumber=26517&arnumber=1181998&

                                        count=82&index=72.  IEEE

Computer Society Order Num-

ber PR01822.

Arvind:2002:GIS

[AK02] V. Arvind and P. P. Kurur. Graph isomorphism is in SPP. In IEEE [IEE02], pages 743–

750. CODEN ASFPDV. ISBN 0-7695-1822-2. ISSN 0272-5428. LCCN QA267. URL http:/

                                        /ieeexplore.ieee.org/iel5/
                                        8411/26517/01181999.pdf?isnumber=
                                        26517&prod=CNF&arnumber=1181999&
                                        arSt=+743&ared=+750&arAuthor=
                                        Arvind%2C+V.%3B+Kurur%2C+
                                        P.P.;   http://ieeexplore.
                                        ieee.org/xpls/abs_all.jsp?
                                        isnumber=26517&arnumber=1181999&
                                        count=82&index=73.  IEEE

Computer Society Order Num-

ber PR01822.

Andon:2007:CHE

[AK07] A. Andon and R. Krauthgamer. The computational hardness of estimating edit distance [extended abstract]. In IEEE [IEE07], pages 724–734. ISBN 0-7695-3010-9. ISSN 0272-5428. LCCN QA76 .S974 2007. URL

                                        http://ieeexplore.ieee.org/
                                        servlet/opac?punumber=4389466.

IEEE Computer Society order

number P3010.

Adleman:2002:DSA

[AKKR02] L. Adleman, J. Kari, L. Kari, and D. Reishus. On the decidability of self-assembly of infinite ribbons. In IEEE [IEE02], pages 530–537. CO-DEN ASFPDV. ISBN 0-7695-1822-2. ISSN 0272-5428.

LCCN QA267. URL http:/

/ieeexplore.ieee.org/iel5/
8411/26517/01181977.pdf?isnumber=
26517&prod=CNF&arnumber=1181977&
arSt=+530&ared=+537&arAuthor=
Adleman%2C+L.%3B+Kari%2C+
J.%3B+Kari%2C+L.%3B+Reishus%
2C+D.;   http://ieeexplore.
ieee.org/xpls/abs_all.jsp?
isnumber=26517&arnumber=1181977&
count=82&index=53.  IEEE

Computer Society Order Number PR01822.

Alekhnovich:2005:HAC

[AKKV05] M. Alekhnovich, S. A. Khot, G. Kindler, and N. K. Vishnoi. Hardness of approximating the closest vector problem with preprocessing. In IEEE [IEE05], pages 216–225. ISBN 0-7695-2468-0, 0-7695-2362-5. ISSN 0272-5428. LCCN QA76 .S979 2005. IEEE Computer Society order number P2468.

Anagnostopoulos:2003:SER

[AKU03] A. Anagnostopoulos, A. Kirsch, and E. Upfal. Stability and efficiency of a random local load balancing protocol. In IEEE [IEE03], pages 472–481. CO-DEN ASFPDV. ISBN 0-7695-2040-5. ISSN 0272-5428. LCCN QA76 .S979 2003. URL http:/

/ieeexplore.ieee.org/iel5/
8767/27770/01238220.pdf?isnumber=
27770&prod=CNF&arnumber=1238220&
arSt=+472&ared=+481&arAuthor=
Anagnostopoulos%2C+A.%3B+
Kirsch%2C+A.%3B+Upfal%2C+
E.;   http://ieeexplore.ieee.
org/xpls/abs_all.jsp?isnumber=
27770&arnumber=1238220&count=

                                        66&index=47.  IEEE Com-

puter Society Order Number

PR02040.

Abbott:2005:CTP

[AKV05] T. Abbott, D. Kane, and P. Valiant. On the complexity of two-player win-lose games. In IEEE [IEE05], pages 113–

122. ISBN 0-7695-2468-0, 0-7695-2362-5. ISSN 0272-5428. LCCN QA76 .S979 2005. IEEE Computer Society order number P2468.

Alekhnovich:2002:LDE

[Ale02] M. Alekhnovich. Linear Diophantine equations over polynomials and soft decoding of Reed–Solomon codes. In IEEE [IEE02], pages 439–448. CO-DEN ASFPDV. ISBN 0-7695-1822-2. ISSN 0272-5428. LCCN QA267. URL http:/

                                        /ieeexplore.ieee.org/iel5/
                                        8411/26517/01181968.pdf?isnumber=
                                        26517&prod=CNF&arnumber=1181968&
                                        arSt=+439&ared=+448&arAuthor=
                                        Alekhnovich%2C+M.;   http:
                                        //ieeexplore.ieee.org/xpls/
                                        abs_all.jsp?isnumber=26517&
                                        arnumber=1181968&count=82&

index=44. IEEE Computer Society Order Number PR01822.

Alekhnovich:2003:MAC

[Ale03] M. Alekhnovich. More on average case vs approximation complexity. In IEEE [IEE03], pages 298–307. CODEN ASF-PDV. ISBN 0-7695-2040-5. ISSN 0272-5428. LCCN QA76 .S979 2003. URL http:/

                                        /ieeexplore.ieee.org/iel5/

8767/27770/01238204.pdf?isnumber=
27770&prod=CNF&arnumber=1238204&
arSt=+298&ared=+307&arAuthor=
Alekhnovich%2C+M.;   http:
//ieeexplore.ieee.org/xpls/
abs_all.jsp?isnumber=27770&
arnumber=1238204&count=66&

index=31. IEEE Computer Society Order Number PR02040.

Alon:2009:CMT

[ALGG09] N. Alon, E. Lubetzky, and O. Gurel-Gurevich. Choicememory tradeoff in allocations. In IEEE [IEE09], pages 230–238. ISBN 0-7695-3850-9. ISSN 0272-5428. LCCN QA76 .S95 2009. IEEE Computer Society order number P3850.

Alon:2001:TSL

[Alo01] N. Alon. Testing subgraphs in large graphs. In IEEE [IEE01], pages 434–439. CO-DEN ASFPDV. ISBN 0-7695-1390-5, 0-7695-1391-3 (case), 0-7695-1392-1 (microfiche). ISSN 0272-5428. LCCN ????

Alon:2008:BSI

[ALS^{+}08] N. Alon, E. Lubetzky, U. Stav, A. Weinstein, and A. Hassidim. Broadcasting with side information. In IEEE [IEE08], pages 823–832. ISBN 0-7695-3436-

8. ISSN 0272-5428. LCCN QA76 .S95 2008. URL http://

ieeexplore.ieee.org/servlet/
opac?punumber=4690923. IEEE

Computer Society order number P3436.

Alon:2001:SDP

[ALW01] N. Alon, A. Lubotzky, and A. Wigderson. Semi-direct product in groups and zig-zag product in graphs: connections and applications. In IEEE [IEE01], pages 630–637. CO-DEN ASFPDV. ISBN 0-7695-1390-5, 0-7695-1391-3 (case), 0-7695-1392-1 (microfiche). ISSN 0272-5428. LCCN ????

Achlioptas:2002:AOR

[AM02] D. Achlioptas and C. Moore. The asymptotic order of the random k-SAT threshold. In IEEE [IEE02], pages 779–788. CODEN ASFPDV. ISBN 0-7695-1822-2. ISSN 0272-5428. LCCN QA267. URL http:/

/ieeexplore.ieee.org/iel5/
8411/26517/01182003.pdf?isnumber=
26517&prod=CNF&arnumber=1182003&
arSt=+779&ared=+788&arAuthor=
Achlioptas%2C+D.%3B+Moore%
2C+C.;   http://ieeexplore.
ieee.org/xpls/abs_all.jsp?
isnumber=26517&arnumber=1182003&
count=82&index=77.  IEEE

Computer Society Order Number PR01822.

Aggarwal:2008:TSS

[AM08] G. Aggarwal and S. Muthukrishnan. Theory of sponsored search auctions. In IEEE [IEE08], page 7. ISBN 0-7695-3436-

8. ISSN 0272-5428. LCCN QA76 .S95 2008. URL http://

ieeexplore.ieee.org/servlet/
opac?punumber=4690923. IEEE

Computer Society order number P3436.

Ambainis:2003:PDV

[Amb03] A. Ambainis. Polynomial degree vs. quantum query complexity. In IEEE [IEE03], pages 230–239. CODEN ASF-PDV. ISBN 0-7695-2040-5. ISSN 0272-5428. LCCN QA76 .S979 2003. URL http:/

/ieeexplore.ieee.org/iel5/
8767/27770/01238197.pdf?isnumber=
27770&prod=CNF&arnumber=1238197&
arSt=+230&ared=+239&arAuthor=
Ambainis%2C+A.;   http://
ieeexplore.ieee.org/xpls/
abs_all.jsp?isnumber=27770&
arnumber=1238197&count=66&

index=24. IEEE Computer Society Order Number PR02040.

Ambainis:2004:QWA

[Amb04] A. Ambainis. Quantum walk algorithm for element distinctness. In IEEE [IEE04], pages 22–31. CODEN ASFPDV. ISBN 0-7695-2228-9. ISSN 0272-5428. LCCN QA276. URL

http://ieeexplore.ieee.org/
iel5/9430/29918/01366221.
pdf?isnumber=29918&prod=CNF&
arnumber=1366221&arSt=+22&
ared=+31&arAuthor=Ambainis%
2C+A.;   http://ieeexplore.
ieee.org/xpls/abs_all.jsp?
isnumber=29918&arnumber=1366221&
count=64&index=2. IEEE Com-

puter Society Order Number P2228.

Arya:2000:NOE

[AMM00] S. Arya, T. Malamatos, and D. M. Mount. Nearly optimal expected-case planar point location. In IEEE [IEE00], pages 208–218. CODEN ASF-PDV. ISBN 0-7695-0850-2, 0-7695-0851-0 (case), 0-7695-0852-9 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 2000. IEEE Computer Society Order Number PR00850.

Arthur:2009:MPS

[AMR09] D. Arthur, B. Manthey, and H. Roglin. k-means has polynomial smoothed complexity. In IEEE [IEE09], pages 405–414. ISBN 0-7695-3850-9. ISSN 0272-5428. LCCN QA76 .S95 2009. IEEE Computer Society order number P3850.

Ambuhl:2007:IRS

[AMS07] C. Ambuhl, M. Mastrolilli, and O. Svensson. Inapproximability results for sparsest cut, optimal linear arrangement, and precedence constrained scheduling. In IEEE [IEE07], pages 329–337. ISBN 0-7695-3010-

9. ISSN 0272-5428. LCCN QA76 .S974 2007. URL http://

                                        ieeexplore.ieee.org/servlet/
                                        opac?punumber=4389466. IEEE

Computer Society order number

P3010.

Aggarwal:2003:SSR

[AMSZ03] G. Aggarwal, R. Motwani, D. Shah, and An Zhu. Switch scheduling via randomized edge coloring. In IEEE [IEE03], pages 502–512. CODEN ASF-PDV. ISBN 0-7695-2040-5. ISSN 0272-5428. LCCN QA76 .S979 2003. URL http:/

                                        /ieeexplore.ieee.org/iel5/
                                        8767/27770/01238223.pdf?isnumber=

27770&prod=CNF&arnumber=1238223&
arSt=+502&ared=+512&arAuthor=
Aggarwal%2C+G.%3B+Motwani%
2C+R.%3B+Shah%2C+D.%3B+An+
Zhu;   http://ieeexplore.
ieee.org/xpls/abs_all.jsp?
isnumber=27770&arnumber=1238223&
count=66&index=50.  IEEE

Computer Society Order Number PR02040.

Ambainis:2000:PQC

[AMTdW00] A. Ambainis, M. Mosca, A. Tapp, and R. de Wolf. Private quantum channels. In IEEE [IEE00], pages 547–553. CODEN ASF-PDV. ISBN 0-7695-0850-2, 0-7695-0851-0 (case), 0-7695-0852-9 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 2000. IEEE Computer Society Order Number PR00850.

Alon:2008:WIR

[AN08] N. Alon and A. Nussboim.

k-wise independent random graphs. In IEEE [IEE08], pages 813–822. ISBN 0-7695-3436-

8. ISSN 0272-5428. LCCN QA76 .S95 2008. URL http://

ieeexplore.ieee.org/servlet/
opac?punumber=4690923. IEEE

Computer Society order number P3436.

Andrews:2004:HBB

[And04] M. Andrews. Hardness of buyat-bulk network design. In IEEE [IEE04], pages 115–124. CODEN ASFPDV. ISBN 0-7695-2228-9. ISSN 0272-5428. LCCN QA276. URL http:/

/ieeexplore.ieee.org/iel5/
9430/29918/01366231.pdf?isnumber=

                                        29918&prod=CNF&arnumber=1366231&
                                        arSt=+115&ared=+124&arAuthor=
                                        Andrews%2C+M.;   http://ieeexplore.
                                        ieee.org/xpls/abs_all.jsp?
                                        isnumber=29918&arnumber=1366231&
                                        count=64&index=12.  IEEE

Computer Society Order Num-

ber P2228.

Anonymous:2001:ISF

[Ano01] Anonymous. 42nd IEEE Symposium on Foundations of Computer Science. In IEEE [IEE01], pages i–viii. CODEN ASF-PDV. ISBN 0-7695-1390-5, 0-7695-1391-3 (case), 0-7695-1392-1 (microfiche). ISSN 0272-5428. LCCN ????

Anonymous:2002:AI

[Ano02] Anonymous. Author index. In IEEE [IEE02], pages 811–813. CODEN ASFPDV. ISBN 0-7695-1822-2. ISSN 0272-5428. LCCN QA267. URL http:/

                                        /ieeexplore.ieee.org/iel5/
                                        8411/26517/01182007.pdf?isnumber=
                                        26517&prod=CNF&arnumber=1182007&
                                        arSt=+811&ared=+813&arAuthor=
                                        ;   http://ieeexplore.ieee.
                                        org/xpls/abs_all.jsp?isnumber=
                                        26517&arnumber=1182007&count=
                                        82&index=81.  IEEE Com-

puter Society Order Number

PR01822.

Anonymous:2003:AI

[Ano03] Anonymous. Author index. In IEEE [IEE03], pages 660–

661. CODEN ASFPDV. ISBN 0-7695-2040-5. ISSN 0272-5428. LCCN QA76 .S979 2003. URL http:/

                                        /ieeexplore.ieee.org/iel5/

8767/27770/01238238.pdf?isnumber=
27770&prod=CNF&arnumber=1238238&
arSt=+660&ared=+661&arAuthor=
;   http://ieeexplore.ieee.
org/xpls/abs_all.jsp?isnumber=
27770&arnumber=1238238&count=
66&index=65.  IEEE Com-

puter Society Order Number PR02040.

Anonymous:2005:AISa

[Ano05a] Anonymous. 46th annual IEEE Symposium on Foundations of Computer Science. In IEEE [IEE05], page ?? ISBN 0-7695-2468-0, 0-7695-2362-5. ISSN 0272-5428. LCCN QA76 .S979 2005. IEEE Computer Society order number P2468.

Anonymous:2005:AISc

[Ano05b] Anonymous. 46th annual IEEE Symposium on Foundations of Computer Science — copyright. In IEEE [IEE05], page iv. ISBN 0-7695-2468-0, 0-7695-2362-5. ISSN 0272-5428. LCCN QA76 .S979 2005. IEEE Computer Society order number P2468.

Anonymous:2005:AISd

[Ano05c] Anonymous. 46th annual IEEE Symposium on Foundations of Computer Science — table of contents. In IEEE [IEE05], pages v–ix. ISBN 0-7695-2468-0, 0-7695-2362-5. ISSN 0272-5428. LCCN QA76 .S979 2005. IEEE Computer Society order number P2468.

Anonymous:2005:AISb

[Ano05d] Anonymous. 46th annual IEEE Symposium on Foundations of Computer Science — title page. In IEEE [IEE05], pages i– iii. ISBN 0-7695-2468-0, 0-7695-2362-5. ISSN 0272-5428. LCCN QA76 .S979 2005. IEEE Computer Society order number P2468.

Anonymous:2005:BPA

[Ano05e] Anonymous. Best paper awards. In IEEE [IEE05], page xi. ISBN 0-7695-2468-0, 0-7695-2362-5. ISSN 0272-5428. LCCN QA76 .S979 2005. IEEE Computer Society order number P2468.

Anonymous:2005:C

[Ano05f] Anonymous. Committees. In IEEE [IEE05], page xi. ISBN 0-7695-2468-0, 0-7695-2362-5. ISSN 0272-5428. LCCN QA76 .S979 2005. IEEE Computer Society order number P2468.

Anonymous:2005:CS

[Ano05g] Anonymous. Corporate sponsors. In IEEE [IEE05], page xiii. ISBN 0-7695-2468-0, 0-7695-2362-5. ISSN 0272-5428. LCCN QA76 .S979 2005. IEEE Computer Society order number P2468.

Anonymous:2005:F

[Ano05h] Anonymous. Foreword. In IEEE [IEE05], page x. ISBN 0-7695-2468-0, 0-7695-2362-5. ISSN 0272-5428. LCCN QA76 .S979 2005. IEEE Computer Society order number P2468.

Anonymous:2005:KP

[Ano05i] Anonymous. Knuth prize. In IEEE [IEE05], page xi. ISBN 0-7695-2468-0, 0-7695-2362-5. ISSN 0272-5428. LCCN QA76 .S979 2005. IEEE Computer Society order number P2468.

Anonymous:2005:MA

[Ano05j] Anonymous. Machtey award. In IEEE [IEE05], page xi. ISBN 0-7695-2468-0, 0-7695-2362-5. ISSN 0272-5428. LCCN QA76 .S979 2005. IEEE Computer Society order number P2468.

Anonymous:2005:R

[Ano05k] Anonymous. Reviewers. In IEEE [IEE05], page xii. ISBN 0-7695-2468-0, 0-7695-2362-5. ISSN 0272-5428. LCCN QA76 .S979 2005. IEEE Computer Society order number P2468.

Anonymous:2006:AISc

[Ano06a] Anonymous. 47th annual IEEE Symposium on Foundations of Computer Science — copyright. In IEEE [IEE06], page iv. ISBN 0-7695-2720-5, 0-7695-2362-5. ISSN 0272-5428. LCCN QA76 .S974 2006. URL http://

ieeexplore.ieee.org/servlet/
opac?punumber=4031329. IEEE

Computer Society Order Number P2720.

Anonymous:2006:AISa

[Ano06b] Anonymous. 47th annual IEEE Symposium on Foundations of Computer Science — cover. In IEEE [IEE06], page c1. ISBN 0-7695-2720-5, 0-7695-2362-5. ISSN 0272-5428. LCCN QA76 .S974 2006. URL http://

ieeexplore.ieee.org/servlet/
opac?punumber=4031329. IEEE

Computer Society Order Number P2720.

Anonymous:2006:AISb

[Ano06c] Anonymous. 47th annual IEEE Symposium on Foundations of Computer Science — title. In IEEE [IEE06], pages i–iii. ISBN 0-7695-2720-5, 0-7695-2362-5. ISSN 0272-5428. LCCN QA76 .S974 2006. URL http://

                                        ieeexplore.ieee.org/servlet/
                                        opac?punumber=4031329. IEEE

Computer Society Order Number P2720.

Anonymous:2006:AISd

[Ano06d] Anonymous. 47th annual IEEE Symposium on Foundations of Computer Science — TOC. In IEEE [IEE06], pages v–x. ISBN 0-7695-2720-5, 0-7695-2362-5. ISSN 0272-5428. LCCN QA76 .S974 2006. URL http://

                                        ieeexplore.ieee.org/servlet/
                                        opac?punumber=4031329. IEEE

Computer Society Order Number P2720.

Anonymous:2006:AI

[Ano06e] Anonymous. Author index. In IEEE [IEE06], pages 749–750. ISBN 0-7695-2720-5, 0-7695-2362-5. ISSN 0272-5428. LCCN QA76 .S974 2006. URL http://

                                        ieeexplore.ieee.org/servlet/
                                        opac?punumber=4031329. IEEE

Computer Society Order Number P2720.

Anonymous:2006:F

[Ano06f] Anonymous. Foreword. In IEEE [IEE06], page xi. ISBN 0-7695-2720-5, 0-7695-2362-5.

ISSN 0272-5428. LCCN QA76 .S974 2006. URL http://

ieeexplore.ieee.org/servlet/
opac?punumber=4031329. IEEE

Computer Society Order Number P2720.

Anonymous:2006:OC

[Ano06g] Anonymous. Organizing committees. In IEEE [IEE06], page xii. ISBN 0-7695-2720-5, 0-7695-2362-5. ISSN 0272-5428. LCCN QA76 .S974 2006. URL http://

ieeexplore.ieee.org/servlet/
opac?punumber=4031329. IEEE

Computer Society Order Number P2720.

Anonymous:2006:PAa

[Ano06h] Anonymous. Proceedings available. In IEEE [IEE06], page ?? ISBN 0-7695-2720-5, 0-7695-2362-5. ISSN 0272-5428. LCCN QA76 .S974 2006. URL http://

ieeexplore.ieee.org/servlet/
opac?punumber=4031329. IEEE

Computer Society Order Number P2720.

Anonymous:2006:PAb

[Ano06i] Anonymous. Proceedings available. In IEEE [IEE06], page ?? ISBN 0-7695-2720-5, 0-7695-2362-5. ISSN 0272-5428. LCCN QA76 .S974 2006. URL http://

ieeexplore.ieee.org/servlet/
opac?punumber=4031329. IEEE

Computer Society Order Number P2720.

Anonymous:2006:R

[Ano06j] Anonymous. Reviewers. In IEEE [IEE06], pages xiii–xiv.

ISBN 0-7695-2720-5, 0-7695-2362-5. ISSN 0272-5428. LCCN QA76 .S974 2006. URL http://

ieeexplore.ieee.org/servlet/
opac?punumber=4031329. IEEE

Computer Society Order Number P2720.

Anonymous:2007:AISc

[Ano07a]  Anonymous.  48th annual

IEEE Symposium on Foundations of Computer Science — copyright. In IEEE [IEE07], page iv. ISBN 0-7695-3010-9. ISSN 0272-5428. LCCN QA76 .S974 2007. URL http://

ieeexplore.ieee.org/servlet/
opac?punumber=4389466. IEEE

Computer Society order number P3010.

Anonymous:2007:AISa

[Ano07b]  Anonymous.  48th annual

IEEE Symposium on Foundations of Computer Science — cover. In IEEE [IEE07], page c1. ISBN 0-7695-3010-9. ISSN 0272-5428. LCCN QA76 .S974 2007. URL http://

ieeexplore.ieee.org/servlet/
opac?punumber=4389466. IEEE

Computer Society order number P3010.

Anonymous:2007:AISb

[Ano07c]  Anonymous.  48th annual

IEEE Symposium on Foundations of Computer Science — title. In IEEE [IEE07], pages i–iii. ISBN 0-7695-3010-9. ISSN 0272-5428. LCCN QA76 .S974 2007. URL http://

ieeexplore.ieee.org/servlet/
opac?punumber=4389466. IEEE

Computer Society order number P3010.

Anonymous:2007:AISd

[Ano07d]  Anonymous.  48th annual

IEEE Symposium on Foundations of Computer Science — TOC. In IEEE [IEE07], pages v–x. ISBN 0-7695-3010-9. ISSN 0272-5428. LCCN QA76 .S974 2007. URL http://

ieeexplore.ieee.org/servlet/
opac?punumber=4389466. IEEE

Computer Society order number P3010.

Anonymous:2007:AI

[Ano07e] Anonymous. Author index.

In IEEE [IEE07], pages 735– 736. ISBN 0-7695-3010-9. ISSN 0272-5428. LCCN QA76 .S974 2007. URL http://

ieeexplore.ieee.org/servlet/
opac?punumber=4389466. IEEE

Computer Society order number P3010.

Anonymous:2007:CO

[Ano07f] Anonymous. Conference organization. In IEEE [IEE07], page xii. ISBN 0-7695-3010-9. ISSN 0272-5428. LCCN QA76 .S974 2007. URL http://

ieeexplore.ieee.org/servlet/
opac?punumber=4389466. IEEE

Computer Society order number P3010.

Anonymous:2007:F

[Ano07g] Anonymous. Forward. In IEEE [IEE07], page xi. ISBN 0-7695-3010-9. ISSN 0272-5428. LCCN QA76 .S974 2007. URL http://

ieeexplore.ieee.org/servlet/

                                        opac?punumber=4389466. IEEE

Computer Society order number P3010.

Anonymous:2007:PA

[Ano07h] Anonymous. Proceedings available. In IEEE [IEE07], page ?? ISBN 0-7695-3010-9. ISSN 0272-5428. LCCN QA76 .S974 2007. URL http://

                                        ieeexplore.ieee.org/servlet/
                                        opac?punumber=4389466. IEEE

Computer Society order number P3010.

Anonymous:2007:R

[Ano07i] Anonymous. Reviewers. In IEEE [IEE07], page xiii. ISBN 0-7695-3010-9. ISSN 0272-5428. LCCN QA76 .S974 2007. URL

                                        http://ieeexplore.ieee.org/
                                        servlet/opac?punumber=4389466.

IEEE Computer Society order number P3010.

Anonymous:2008:AI

[Ano08a] Anonymous. Author index. In IEEE [IEE08], pages 833– 835. ISBN 0-7695-3436-8. ISSN 0272-5428. LCCN QA76 .S95 2008. URL http://

                                        ieeexplore.ieee.org/servlet/
                                        opac?punumber=4690923. IEEE

Computer Society order number P3436.

Anonymous:2008:C

[Ano08b]  Anonymous.  Committees.

In IEEE [IEE08], page xiii. ISBN 0-7695-3436-8. ISSN 0272-5428. LCCN QA76 .S95 2008. URL http://

                                        ieeexplore.ieee.org/servlet/
                                        opac?punumber=4690923. IEEE

Computer Society order number P3436.

Anonymous:2008:CP

[Ano08c] Anonymous. Copyright page.

In IEEE [IEE08], page iv. ISBN 0-7695-3436-8. ISSN 0272-5428. LCCN QA76 .S95 2008. URL http://

ieeexplore.ieee.org/servlet/
opac?punumber=4690923. IEEE

Computer Society order number P3436.

Anonymous:2008:CA

[Ano08d] Anonymous. Cover art. In IEEE [IEE08], page C1. ISBN 0-7695-3436-8. ISSN 0272-5428. LCCN QA76 .S95 2008. URL http://

ieeexplore.ieee.org/servlet/
opac?punumber=4690923. IEEE

Computer Society order number P3436.

Anonymous:2008:F

[Ano08e] Anonymous. Foreword. In IEEE [IEE08], page xii. ISBN 0-7695-3436-8. ISSN 0272-5428. LCCN QA76 .S95 2008. URL http://

ieeexplore.ieee.org/servlet/
opac?punumber=4690923. IEEE

Computer Society order number P3436.

Anonymous:2008:PAa

[Ano08f] Anonymous. Proceedings available. In IEEE [IEE08], page ?? ISBN 0-7695-3436-8. ISSN 0272-5428. LCCN QA76 .S95 2008. URL http://

ieeexplore.ieee.org/servlet/
opac?punumber=4690923. IEEE

Computer Society order number P3436.

Anonymous:2008:PAb

[Ano08g] Anonymous. Proceedings available. In IEEE [IEE08], page ?? ISBN 0-7695-3436-8. ISSN 0272-5428. LCCN QA76 .S95 2008. URL http://

                                        ieeexplore.ieee.org/servlet/
                                        opac?punumber=4690923. IEEE

Computer Society order number P3436.

Anonymous:2008:R

[Ano08h] Anonymous. Reviewers. In IEEE [IEE08], pages xiv–xvi. ISBN 0-7695-3436-8. ISSN 0272-5428. LCCN QA76 .S95 2008. URL http://

                                        ieeexplore.ieee.org/servlet/
                                        opac?punumber=4690923. IEEE

Computer Society order number P3436.

Anonymous:2008:RP

[Ano08i]  Anonymous.  Roster page.

In IEEE [IEE08], page 836. ISBN 0-7695-3436-8. ISSN 0272-5428. LCCN QA76 .S95 2008. URL http://

                                        ieeexplore.ieee.org/servlet/
                                        opac?punumber=4690923. IEEE

Computer Society order number P3436.

Anonymous:2008:TC

[Ano08j] Anonymous. Table of contents. In IEEE [IEE08], pages v–xi. ISBN 0-7695-3436-8. ISSN 0272-5428. LCCN QA76 .S95 2008. URL http://

                                        ieeexplore.ieee.org/servlet/
                                        opac?punumber=4690923. IEEE

Computer Society order number P3436.

Anonymous:2008:TP

[Ano08k]  Anonymous.  Title page i.

In IEEE [IEE08], page i. ISBN 0-7695-3436-8. ISSN 0272-5428. LCCN QA76 .S95 2008. URL http://

ieeexplore.ieee.org/servlet/
opac?punumber=4690923. IEEE

Computer Society order number P3436.

Anonymous:2008:TPI

[Ano08l] Anonymous. Title page iii.

In IEEE [IEE08], page iii. ISBN 0-7695-3436-8. ISSN 0272-5428. LCCN QA76 .S95 2008. URL http://

ieeexplore.ieee.org/servlet/
opac?punumber=4690923. IEEE

Computer Society order number P3436.

Anonymous:2009:AI

[Ano09a] Anonymous. Author index. In IEEE [IEE09], pages 775–777. ISBN 0-7695-3850-9. ISSN 0272-5428. LCCN QA76 .S95 2009. IEEE Computer Society order number P3850.

Anonymous:2009:CP

[Ano09b] Anonymous. Copyright page. In IEEE [IEE09], page iv. ISBN 0-7695-3850-9. ISSN 0272-5428. LCCN QA76 .S95 2009. IEEE Computer Society order number P3850.

Anonymous:2009:CA

[Ano09c] Anonymous. Cover art. In IEEE [IEE09], page C1. ISBN 0-7695-3850-9. ISSN 0272-5428. LCCN QA76 .S95 2009. IEEE Computer Society order number P3850.

Anonymous:2009:MPC

[Ano09d] Anonymous. Message from program chair. In IEEE [IEE09], page xi. ISBN 0-7695-3850-9. ISSN 0272-5428. LCCN QA76 .S95 2009. IEEE Computer Society order number P3850.

Anonymous:2009:OC

[Ano09e] Anonymous. Organizing committee. In IEEE [IEE09], page xii. ISBN 0-7695-3850-9. ISSN 0272-5428. LCCN QA76 .S95 2009. IEEE Computer Society order number P3850.

Anonymous:2009:PAa

[Ano09f] Anonymous. Proceedings available. In IEEE [IEE09], page ?? ISBN 0-7695-3850-9. ISSN 0272-5428. LCCN QA76 .S95 2009. IEEE Computer Society order number P3850.

Anonymous:2009:PAb

[Ano09g] Anonymous. Proceedings available. In IEEE [IEE09], page ?? ISBN 0-7695-3850-9. ISSN 0272-5428. LCCN QA76 .S95 2009. IEEE Computer Society order number P3850.

Anonymous:2009:PI

[Ano09h] Anonymous. Publisher’s information. In IEEE [IEE09], page 778. ISBN 0-7695-3850-9. ISSN 0272-5428. LCCN QA76 .S95 2009. IEEE Computer Society order number P3850.

Anonymous:2009:R

[Ano09i] Anonymous. Reviewers. In IEEE [IEE09], pages xiii–xiv. ISBN 0-7695-3850-9. ISSN 0272-5428. LCCN QA76 .S95 2009. IEEE Computer Society order number P3850.

Anonymous:2009:TC

[Ano09j] Anonymous. Table of contents.

In IEEE [IEE09], pages v–x. ISBN 0-7695-3850-9. ISSN 0272-5428. LCCN QA76 .S95 2009. IEEE Computer Society order number P3850.

Anonymous:2009:TP

[Ano09k] Anonymous. Title page i. In IEEE [IEE09], page i. ISBN 0-7695-3850-9. ISSN 0272-5428. LCCN QA76 .S95 2009. IEEE Computer Society order number P3850.

Anonymous:2009:TPI

[Ano09l] Anonymous. Title page iii. In IEEE [IEE09], page iii. ISBN 0-7695-3850-9. ISSN 0272-5428. LCCN QA76 .S95 2009. IEEE Computer Society order number P3850.

Achlioptas:2003:MSR

[ANP03] D. Achlioptas, A. Naor, and Y. Peres. On the maximum satisfiability of random formulas. In IEEE [IEE03], pages 362–370. CODEN ASF-PDV. ISBN 0-7695-2040-5. ISSN 0272-5428. LCCN QA76 .S979 2003. URL http:/

/ieeexplore.ieee.org/iel5/
8767/27770/01238210.pdf?isnumber=

27770&prod=CNF&arnumber=1238210&
arSt=+362&ared=+370&arAuthor=
Achlioptas%2C+D.%3B+Naor%
2C+A.%3B+Peres%2C+Y.;   http:
//ieeexplore.ieee.org/xpls/
abs_all.jsp?isnumber=27770&
arnumber=1238210&count=66&

index=37. IEEE Computer Society Order Number PR02040.

Alekhnovich:2001:LBP

[AR01a] M. Alekhnovich and A. A. Razboro. Lower bounds for polynomial calculus: nonbinomial case. In IEEE [IEE01], pages 190–199. CODEN ASF-PDV. ISBN 0-7695-1390-5, 0-7695-1391-3 (case), 0-7695-1392-1 (microfiche). ISSN 0272-5428. LCCN ????

Alekhnovich:2001:RAU

[AR01b] M. Alekhnovich and A. A. Razboro. Resolution is not automatizable unless W[P] is tractable. In IEEE [IEE01], pages 210–219. CODEN ASF-PDV. ISBN 0-7695-1390-5, 0-7695-1391-3 (case), 0-7695-1392-1 (microfiche). ISSN 0272-5428. LCCN ????

Alekhnovich:2002:SBW

[AR02a] M. Alekhnovich and A. A. Razborov. Satisfiability, branchwidth and Tseitin tautologies. In IEEE [IEE02], pages 593–

603. CODEN ASFPDV. ISBN 0-7695-1822-2. ISSN 0272-5428. LCCN QA267. URL http:/

/ieeexplore.ieee.org/iel5/
8411/26517/01181983.pdf?isnumber=
26517&prod=CNF&arnumber=1181983&
arSt=+593&ared=+603&arAuthor=

Alekhnovich%2C+M.%3B+Razborov%
2C+A.A.;   http://ieeexplore.
ieee.org/xpls/abs_all.jsp?
isnumber=26517&arnumber=1181983&
count=82&index=59.  IEEE

Computer Society Order Number PR01822.

Alstrup:2002:SIU

[AR02b] S. Alstrup and T. Rauhe.

Small induced-universal graphs and compact implicit graph representations. In IEEE [IEE02], pages 53–62. CO-DEN ASFPDV. ISBN 0-7695-1822-2. ISSN 0272-5428. LCCN QA267. URL http:/

/ieeexplore.ieee.org/iel5/
8411/26517/01181882.pdf?isnumber=
26517&prod=CNF&arnumber=1181882&
arSt=+53&ared=+62&arAuthor=
Alstrup%2C+S.%3B+Rauhe%2C+
T.;   http://ieeexplore.ieee.
org/xpls/abs_all.jsp?isnumber=
26517&arnumber=1181882&count=
82&index=7.  IEEE Com-

puter Society Order Number PR01822.

Aharonov:2003:LPQ

[AR03] D. Aharonov and O. Regev.

A lattice problem in quantum NP. In IEEE [IEE03], pages 210–219. CODEN ASF-PDV. ISBN 0-7695-2040-5. ISSN 0272-5428. LCCN QA76 .S979 2003. URL http:/

/ieeexplore.ieee.org/iel5/
8767/27770/01238195.pdf?isnumber=
27770&prod=CNF&arnumber=1238195&
arSt=+210&ared=+219&arAuthor=
Aharonov%2C+D.%3B+Regev%2C+
O.;   http://ieeexplore.ieee.
org/xpls/abs_all.jsp?isnumber=

                                        27770&arnumber=1238195&count=
                                        66&index=22.  IEEE Com-

puter Society Order Number

PR02040.

Aharonov:2004:LPN

[AR04] D. Aharonov and O. Regev. Lattice problems in NP ∩ coNP. In IEEE [IEE04], pages 362–371. CODEN ASFPDV. ISBN 0-7695-2228-9. ISSN 0272-5428. LCCN QA276. URL http:/

                                        /ieeexplore.ieee.org/iel5/
                                        9430/29918/01366256.pdf?isnumber=
                                        29918&prod=CNF&arnumber=1366256&
                                        arSt=+362&ared=+371&arAuthor=
                                        Aharonov%2C+D.%3B+Regev%2C+
                                        O.;   http://ieeexplore.ieee.
                                        org/xpls/abs_all.jsp?isnumber=
                                        29918&arnumber=1366256&count=

64&index=37. IEEE Computer Society Order Number P2228.

Ackermann:2006:ICS

[ARV06] Heiner Ackermann, Heiko Roglin, and Berthold Vocking. On the impact of combinatorial structure on congestion games. In IEEE [IEE06], pages 613–622. ISBN 0-7695-2720-5, 0-7695-2362-5. ISSN 0272-5428. LCCN QA76 .S974 2006. URL http://

                                        ieeexplore.ieee.org/servlet/
                                        opac?punumber=4031329. IEEE

Computer Society Order Num-

ber P2720.

Achhoptas:2000:OMA

[AS00] D. Achhoptas and G. B. Sorkin. Optimal myopic algorithms for random 3-SAT. In IEEE [IEE00], pages 590–600. CO-DEN ASFPDV. ISBN 0-7695-0850-2, 0-7695-0851-0 (case), 0-7695-0852-9 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 2000. IEEE Computer Society Order Number PR00850.

Alon:2005:CNG

[AS05] N. Alon and A. Shapira. A characterization of the (natural) graph properties testable with one-sided error. In IEEE [IEE05], pages 429–438. ISBN 0-7695-2468-0, 0-7695-2362-5. ISSN 0272-5428. LCCN QA76 .S979 2005. IEEE Computer Society order number P2468.

Alon:2005:AAE

[ASS05] N. Alon, A. Shapira, and B. Sudakov. Additive approximation for edge-deletion problems. In IEEE [IEE05], pages 419– 428. ISBN 0-7695-2468-0, 0-7695-2362-5. ISSN 0272-5428. LCCN QA76 .S979 2005. IEEE Computer Society order number P2468.

Anshelevich:2006:SNF

[ASW06] Elliot Anshelevich, Bruce Shepherd, and Gordon Wilfong. Strategic network formation through peering and service agreements. In IEEE [IEE06], pages 77–86. ISBN 0-7695-2720-5, 0-7695-2362-5. ISSN 0272-5428. LCCN QA76 .S974 2006. URL http://

ieeexplore.ieee.org/servlet/
opac?punumber=4031329. IEEE

Computer Society Order Number P2720.

Archer:2001:TMO

[AT01] A. Archer and E. Tardos. Truthful mechanisms for oneparameter agents. In IEEE [IEE01], pages 482–491. CO-DEN ASFPDV. ISBN 0-7695-1390-5, 0-7695-1391-3 (case), 0-7695-1392-1 (microfiche). ISSN 0272-5428. LCCN ????

Auer:2000:UUC

[Aue00] P. Auer. Using upper confidence bounds for online learning. In IEEE [IEE00], pages 270–279. CODEN ASFPDV. ISBN 0-7695-0850-2, 0-7695-0851-0 (case), 0-7695-0852-9 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 2000. IEEE Computer Society Order Number PR00850.

Austrin:2007:TSI

[Aus07] P. Austrin. Towards sharp inapproximability for any 2-CSP. In IEEE [IEE07], pages 307–317. ISBN 0-7695-3010-9. ISSN 0272-5428. LCCN QA76 .S974 2007. URL http://

                                        ieeexplore.ieee.org/servlet/
                                        opac?punumber=4389466. IEEE

Computer Society order number P3010.

Arthur:2006:WCS

[AV06] David Arthur and Sergei Vassilvitskii. Worst-case and smoothed analysis of the ICP algorithm, with an application to the k-means method. In IEEE [IEE06], pages 153–164. ISBN 0-7695-2720-5, 0-7695-2362-5. ISSN 0272-5428. LCCN QA76 .S974 2006. URL http://

ieeexplore.ieee.org/servlet/
opac?punumber=4031329. IEEE

Computer Society Order Number P2720.

Agrawal:2008:ACC

[AV08] M. Agrawal and V. Vinay.

Arithmetic circuits: a chasm at depth four. In IEEE [IEE08], pages 67–75. ISBN 0-7695-3436-

8. ISSN 0272-5428. LCCN QA76 .S95 2008. URL http://

ieeexplore.ieee.org/servlet/
opac?punumber=4690923. IEEE

Computer Society order number P3436.

Aharonov:2004:AQC

[AvDK^{+}04] D. Aharonov, W. van Dam, J. Kempe, Z. Landau, S. Lloyd, and O. Regev. Adiabatic quantum computation is equivalent to standard quantum computation. In IEEE [IEE04], pages 42–

51. CODEN ASFPDV. ISBN 0-7695-2228-9. ISSN 0272-5428. LCCN QA276. URL http:/

/ieeexplore.ieee.org/iel5/
9430/29918/01366223.pdf?isnumber=
29918&prod=CNF&arnumber=1366223&
arSt=+42&ared=+51&arAuthor=
Aharonov%2C+D.%3B+van+Dam%
2C+W.%3B+Kempe%2C+J.%3B+Landau%
2C+Z.%3B+Lloyd%2C+S.%3B+Regev%
2C+O.;   http://ieeexplore.
ieee.org/xpls/abs_all.jsp?
isnumber=29918&arnumber=1366223&
count=64&index=4. IEEE Com-

puter Society Order Number P2228.

Andrews:2002:STV

[AZ02] M. Andrews and L. Zhang.

Scheduling over a time-varying user-dependent channel with applications to high speed wireless data. In IEEE [IEE02], pages 293–302. CO-DEN ASFPDV. ISBN 0-7695-1822-2. ISSN 0272-5428. LCCN QA267. URL http:/

/ieeexplore.ieee.org/iel5/
8411/26517/01181952.pdf?isnumber=
26517&prod=CNF&arnumber=1181952&
arSt=+293&ared=+302&arAuthor=
Andrews%2C+M.%3B+Zhang%2C+
L.;   http://ieeexplore.ieee.
org/xpls/abs_all.jsp?isnumber=
26517&arnumber=1181952&count=
82&index=30.  IEEE Com-

puter Society Order Number PR01822.

Arge:2003:ESC

[AZ03] L. Arge and N. Zeh. I/Oefficient strong connectivity and depth-first search for directed planar graphs. In IEEE [IEE03], pages 261–270. CODEN ASF-PDV. ISBN 0-7695-2040-5. ISSN 0272-5428. LCCN QA76 .S979 2003. URL http:/

/ieeexplore.ieee.org/iel5/
8767/27770/01238200.pdf?isnumber=
27770&prod=CNF&arnumber=1238200&
arSt=+261&ared=+270&arAuthor=
Arge%2C+L.%3B+Zeh%2C+N.;
http://ieeexplore.ieee.org/
xpls/abs_all.jsp?isnumber=
27770&arnumber=1238200&count=
66&index=27.  IEEE Com-

puter Society Order Number PR02040.

Barak:2001:HGB

[Bar01] B. Barak. How to go beyond the black-box simulation barrier. In IEEE [IEE01], pages 106–115. CODEN ASF-PDV. ISBN 0-7695-1390-5, 0-7695-1391-3 (case), 0-7695-1392-1 (microfiche). ISSN 0272-5428. LCCN ????

Barak:2002:CRC

[Bar02] B. Barak. Constant-round cointossing with a man in the middle or realizing the shared random string model. In IEEE [IEE02], pages 345–355. CO-DEN ASFPDV. ISBN 0-7695-1822-2. ISSN 0272-5428. LCCN QA267. URL http:/

/ieeexplore.ieee.org/iel5/
8411/26517/01181957.pdf?isnumber=
26517&prod=CNF&arnumber=1181957&
arSt=+345&ared=+355&arAuthor=
Barak%2C+B.;   http://ieeexplore.
ieee.org/xpls/abs_all.jsp?
isnumber=26517&arnumber=1181957&
count=82&index=35.  IEEE

Computer Society Order Number PR01822.

Ben-Aroya:2008:HIM

[BARdW08] A. Ben-Aroya, O. Regev, and R. de Wolf. A hypercontractive inequality for matrixvalued functions with applications to quantum computing and LDCs. In IEEE [IEE08], pages 477–486. ISBN 0-7695-3436-8. ISSN 0272-5428. LCCN QA76 .S95 2008. URL http://

ieeexplore.ieee.org/servlet/
opac?punumber=4690923. IEEE

Computer Society order number P3436.

Ben-Aroya:2009:CSB

[BATS09] A. Ben-Aroya and A. Ta-Shma.

Constructing small-bias sets from algebraic-geometric codes. In IEEE [IEE09], pages 191–197. ISBN 0-7695-3850-9. ISSN 0272-5428. LCCN QA76 .S95 2009. IEEE Computer Society order number P3850.

Bazzi:2007:PIC

[Baz07] L. M. J. Bazzi. Polylogarithmic independence can fool DNF formulas. In IEEE [IEE07], pages 63–73. ISBN 0-7695-3010-

9. ISSN 0272-5428. LCCN QA76 .S974 2007. URL http://

                                        ieeexplore.ieee.org/servlet/
                                        opac?punumber=4389466. IEEE

Computer Society order number

P3010.

Bansal:2002:CC

[BBC02] N. Bansal, A. Blum, and S. Chawla. Correlation clustering. In IEEE [IEE02], pages 238–247. CODEN ASFPDV. ISBN 0-7695-1822-2. ISSN 0272-5428. LCCN QA267. URL http:/

                                        /ieeexplore.ieee.org/iel5/
                                        8411/26517/01181947.pdf?isnumber=
                                        26517&prod=CNF&arnumber=1181947&
                                        arSt=+238&ared=+247&arAuthor=
                                        Bansal%2C+N.%3B+Blum%2C+A.
                                        %3B+Chawla%2C+S.;   http:/
                                        /ieeexplore.ieee.org/xpls/
                                        abs_all.jsp?isnumber=26517&
                                        arnumber=1181947&count=82&

index=25. IEEE Computer Society Order Number PR01822.

Bender:2003:CCO

[BBF^{+}03] M. A. Bender, G. S. Brodal, R. Fagerberg, D. Ge, Simai He, Haodung Hu, J. Iacono, and A. Lopez-Ortiz. The cost of cache-oblivious searching. In IEEE [IEE03], pages 271–282. CODEN ASF-PDV. ISBN 0-7695-2040-5. ISSN 0272-5428. LCCN QA76 .S979 2003. URL http:/

/ieeexplore.ieee.org/iel5/
8767/27770/01238201.pdf?isnumber=
27770&prod=CNF&arnumber=1238201&
arSt=+271&ared=+282&arAuthor=
Bender%2C+M.A.%3B+Brodal%
2C+G.S.%3B+Fagerberg%2C+R.
%3B+Ge%2C+D.%3B+Simai+He%
3B+Haodung+Hu%3B+Iacono%2C+
J.%3B+Lopez-Ortiz%2C+A.;
http://ieeexplore.ieee.org/
xpls/abs_all.jsp?isnumber=
27770&arnumber=1238201&count=
66&index=28.  IEEE Com-

puter Society Order Number PR02040.

Balcan:2005:MDM

[BBHM05] M.-F. Balcan, A. Blum, J. D.

Hartline, and Y. Mansour. Mechanism design via machine learning. In IEEE [IEE05], pages 605–614. ISBN 0-7695-2468-0, 0-7695-2362-5. ISSN 0272-5428. LCCN QA76 .S979 2005. IEEE Computer Society order number P2468.

Bisht:2004:LEA

[BBK04] L. Bisht, N. H. Bshouty, and L. Khoury. Learning with errors in answers to membership queries. In IEEE [IEE04], pages 611–620. CO-DEN ASFPDV. ISBN 0-7695-2228-9. ISSN 0272-5428. LCCN QA276. URL http:/

                                        /ieeexplore.ieee.org/iel5/
                                        9430/29918/01366281.pdf?isnumber=
                                        29918&prod=CNF&arnumber=1366281&
                                        arSt=+611&ared=+620&arAuthor=
                                        Bisht%2C+L.%3B+Bshouty%2C+
                                        N.H.%3B+Khoury%2C+L.;   http:
                                        //ieeexplore.ieee.org/xpls/
                                        abs_all.jsp?isnumber=29918&
                                        arnumber=1366281&count=64&

index=62. IEEE Computer Society Order Number P2228.

Bartal:2001:RTT

[BBM01] Y. Bartal, B. Bollobas, and M. Mendel. A Ramsey-type theorem for metric spaces and its applications for metrical task systems and related problems. In IEEE [IEE01], pages 396–405. CODEN ASFPDV. ISBN 0-7695-1390-5, 0-7695-1391-3 (case), 0-7695-1392-1 (microfiche). ISSN 0272-5428. LCCN ????

Bansal:2007:PDR

[BBN07] N. Bansal, N. Buchbinder, and J. Naor. A primal-dual randomized algorithm for weighted paging. In IEEE [IEE07], pages 507–517. ISBN 0-7695-3010-

9. ISSN 0272-5428. LCCN QA76 .S974 2007. URL http://

                                        ieeexplore.ieee.org/servlet/
                                        opac?punumber=4389466. IEEE

Computer Society order number

P3010.

Bhamidi:2008:MTE

[BBS08] S. Bhamidi, G. Bresler, and A. Sly. Mixing time of exponential random graphs. In IEEE [IEE08], pages 803–

812. ISBN 0-7695-3436-8.

ISSN 0272-5428. LCCN QA76 .S95 2008. URL http://

ieeexplore.ieee.org/servlet/
opac?punumber=4690923. IEEE

Computer Society order number P3436.

Brinkman:2003:IDR

[BC03] B. Brinkman and M. Charikar.

On the impossibility of dimension reduction in ℓ_{1}. In IEEE [IEE03], pages 514–523. CO-DEN ASFPDV. ISBN 0-7695-2040-5. ISSN 0272-5428. LCCN QA76 .S979 2003. URL http:/

/ieeexplore.ieee.org/iel5/
8767/27770/01238224.pdf?isnumber=
27770&prod=CNF&arnumber=1238224&
arSt=+514&ared=+523&arAuthor=
Brinkman%2C+B.%3B+Charikar%
2C+M.;   http://ieeexplore.
ieee.org/xpls/abs_all.jsp?
isnumber=27770&arnumber=1238224&
count=66&index=51.  IEEE

Computer Society Order Number PR02040.

Babai:2008:IHL

[BC08] L. Babai and P. Codenotti. Isomorhism of hypergraphs of low rank in moderately exponential time. In IEEE [IEE08], pages 667–676. ISBN 0-7695-3436-

8. ISSN 0272-5428. LCCN QA76 .S95 2008. URL http://

ieeexplore.ieee.org/servlet/
opac?punumber=4690923. IEEE

Computer Society order number P3436.

Barnum:2002:AQM

[BCG^{+}02] H. Barnum, C. Crepeau, D. Gottesman, A. Smith, and A. Tapp. Authentication of quantum messages. In IEEE [IEE02], pages 449–458. CO-DEN ASFPDV. ISBN 0-7695-1822-2. ISSN 0272-5428. LCCN QA267. URL http:/

/ieeexplore.ieee.org/iel5/
8411/26517/01181969.pdf?isnumber=
26517&prod=CNF&arnumber=1181969&
arSt=+449&ared=+458&arAuthor=
Barnum%2C+H.%3B+Crepeau%2C+
C.%3B+Gottesman%2C+D.%3B+
Smith%2C+A.%3B+Tapp%2C+A.;
 http://ieeexplore.ieee.org/
xpls/abs_all.jsp?isnumber=
26517&arnumber=1181969&count=
82&index=45.  IEEE Com-

puter Society Order Number PR01822.

Blum:2003:AAO

[BCK^{+}03] A. Blum, S. Chawla, D. R. Karger, T. Lane, A. Meyerson, and M. Minkoff. Approximation algorithms for orienteering and discounted-reward TSP. In IEEE [IEE03], pages 46–55. CO-DEN ASFPDV. ISBN 0-7695-2040-5. ISSN 0272-5428. LCCN QA76 .S979 2003. URL http:/

/ieeexplore.ieee.org/iel5/
8767/27770/01238180.pdf?isnumber=
27770&prod=CNF&arnumber=1238180&
arSt=+46&ared=+55&arAuthor=
Blum%2C+A.%3B+Chawla%2C+S.
%3B+Karger%2C+D.R.%3B+Lane%
2C+T.%3B+Meyerson%2C+A.%3B+
Minkoff%2C+M.;   http://ieeexplore.
ieee.org/xpls/abs_all.jsp?
isnumber=27770&arnumber=1238180&
count=66&index=7. IEEE Com-

puter Society Order Number PR02040.

Bansal:2007:NPM

[BCK^{+}07] N. Bansal, Ho-Leung Chan, R. Khandekar, K. Pruhs, B. Schicber, and C. Stein. Nonpreemptive min-sum scheduling with resource augmentation. In IEEE [IEE07], pages 614–624. ISBN 0-7695-3010-9. ISSN 0272-5428. LCCN QA76 .S974 2007. URL http://

ieeexplore.ieee.org/servlet/
opac?punumber=4389466. IEEE

Computer Society order number P3010.

Buhrman:2006:NLF

[BCL^{+}06] Harry Buhrman, Richard Cleve, Monique Laurent, Noah Linden, Alexander Schrijver, and Falk Unger. New limits on faulttolerant quantum computation. In IEEE [IEE06], pages 411–419. ISBN 0-7695-2720-5, 0-7695-2362-5. ISSN 0272-5428. LCCN QA76 .S974 2006. URL http://

ieeexplore.ieee.org/servlet/
opac?punumber=4031329. IEEE

Computer Society Order Number P2720.

Borgs:2006:KSR

[BCMR06] Christian Borgs, Jennifer Chayes, Elchanan Mossel, and Sebastien Roch. The Kesten– Stigum reconstruction bound is tight for roughly symmetric binary channels. In IEEE [IEE06], pages 518–530. ISBN 0-7695-2720-5, 0-7695-2362-5. ISSN 0272-5428. LCCN QA76 .S974 2006. URL http://

ieeexplore.ieee.org/servlet/
opac?punumber=4031329. IEEE

Computer Society Order Num-

ber P2720.

Barak:2004:UCP

[BCNP04] B. Barak, R. Canetti, J. B. Nielsen, and R. Pass. Universally composable protocols with relaxed set-up assumptions. In IEEE [IEE04], pages 186–195. CODEN ASFPDV. ISBN 0-7695-2228-9. ISSN 0272-5428. LCCN QA276. URL http:/

                                        /ieeexplore.ieee.org/iel5/
                                        9430/29918/01366238.pdf?isnumber=
                                        29918&prod=CNF&arnumber=1366238&
                                        arSt=+186&ared=+195&arAuthor=
                                        Barak%2C+B.%3B+Canetti%2C+
                                        R.%3B+Nielsen%2C+J.B.%3B+
                                        Pass%2C+R.;   http://ieeexplore.
                                        ieee.org/xpls/abs_all.jsp?
                                        isnumber=29918&arnumber=1366238&
                                        count=64&index=19.  IEEE

Computer Society Order Num-

ber P2228.

Bansal:2006:IAA

[BCS06] Nikhil Bansal, Alberto Caprara, and Maxim Sviridenko. Improved approximation algorithms for multidimensional bin packing problems. In IEEE [IEE06], pages 697–708. ISBN 0-7695-2720-5, 0-7695-2362-5. ISSN 0272-5428. LCCN QA76 .S974 2006. URL http://

                                        ieeexplore.ieee.org/servlet/
                                        opac?punumber=4031329. IEEE

Computer Society Order Num-

ber P2720.

Bendich:2007:ILH

[BCSE^{+}07] P. Bendich, D. Cohen-Steiner, H. Edelsbrunner, J. Harer, and D. Morozov. Inferring local homology from sampled stratified spaces. In IEEE [IEE07], pages 536–546. ISBN 0-7695-3010-9. ISSN 0272-5428. LCCN QA76 .S974 2007. URL http://

ieeexplore.ieee.org/servlet/
opac?punumber=4389466. IEEE

Computer Society order number P3010.

Bacon:2005:OME

[BCvD05] D. Bacon, A. M. Childs, and W. van Dam. From optimal measurement to efficient quantum algorithms for the hidden subgroup problem over semidirect product groups. In IEEE [IEE05], pages 469–478. ISBN 0-7695-2468-0, 0-7695-2362-5. ISSN 0272-5428. LCCN QA76 .S979 2005. IEEE Computer Society order number P2468.

Bulatov:2003:TDT

[BD03] A. A. Bulatov and V. Dalmau.

Towards a dichotomy theorem for the counting constraint satisfaction problem. In IEEE [IEE03], pages 562–571. CO-DEN ASFPDV. ISBN 0-7695-2040-5. ISSN 0272-5428. LCCN QA76 .S979 2003. URL http:/

/ieeexplore.ieee.org/iel5/
8767/27770/01238229.pdf?isnumber=
27770&prod=CNF&arnumber=1238229&
arSt=+562&ared=+571&arAuthor=
Bulatov%2C+A.A.%3B+Dalmau%
2C+V.;   http://ieeexplore.
ieee.org/xpls/abs_all.jsp?
isnumber=27770&arnumber=1238229&
count=66&index=56.  IEEE

Computer Society Order Number PR02040.

Bender:2000:COT

[BDFC00] M. A. Bender, E. D. Demaine, and M. Farach-Colton. Cacheoblivious b-trees. In IEEE [IEE00], pages 399–409. CO-DEN ASFPDV. ISBN 0-7695-0850-2, 0-7695-0851-0 (case), 0-7695-0852-9 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 2000. IEEE Computer Society Order Number PR00850.

Bacchus:2003:ACR

[BDP03] F. Bacchus, S. Dalmao, and T. Pitassi. Algorithms and complexity results for #sat and Bayesian inference. In IEEE [IEE03], pages 340–351. CO-DEN ASFPDV. ISBN 0-7695-2040-5. ISSN 0272-5428. LCCN QA76 .S979 2003. URL http:/

                                        /ieeexplore.ieee.org/iel5/
                                        8767/27770/01238208.pdf?isnumber=
                                        27770&prod=CNF&arnumber=1238208&
                                        arSt=+340&ared=+351&arAuthor=
                                        Bacchus%2C+F.%3B+Dalmao%2C+
                                        S.%3B+Pitassi%2C+T.;   http:
                                        //ieeexplore.ieee.org/xpls/
                                        abs_all.jsp?isnumber=27770&
                                        arnumber=1238208&count=66&

index=35. IEEE Computer Society Order Number PR02040.

Brazdil:2005:APL

[BEK05] T. Brazdil, J. Esparza, and A. Kucera. Analysis and prediction of the long-run behavior of probabilistic sequential programs with recursion. In IEEE [IEE05], pages 521–

530. ISBN 0-7695-2468-0, 0-7695-2362-5. ISSN 0272-5428. LCCN QA76 .S979 2005. IEEE Computer Society order number P2468.

Bernstein:2009:FDA

[Ber09] A. Bernstein. Fully dynamic (2+ ϵ) approximate all-pairs shortest paths with fast query and close to linear update time. In IEEE [IEE09], pages 693–702. ISBN 0-7695-3850-9. ISSN 0272-5428. LCCN QA76 .S95 2009. IEEE Computer Society order number P3850.

Bohman:2001:ADP

[BF01] T. Bohman and A. Frieze. Arcdisjoint paths in expander digraphs. In IEEE [IEE01], pages 558–567. CODEN ASF-PDV. ISBN 0-7695-1390-5, 0-7695-1391-3 (case), 0-7695-1392-1 (microfiche). ISSN 0272-5428. LCCN ????

Batu:2001:TRV

[BFF^{+}01] T. Batu, E. Fischer, L. Fortnow, R. Kumar, R. Rubinfeld, and P. White. Testing random variables for independence and identity. In IEEE [IEE01], pages 442–451. CO-DEN ASFPDV. ISBN 0-7695-1390-5, 0-7695-1391-3 (case), 0-7695-1392-1 (microfiche). ISSN 0272-5428. LCCN ????

Berenbrink:2001:NWS

[BFG01] P. Berenbrink, T. Friedetzky, and L. A. Goldberg. The natural work-stealing algorithm is stable. In IEEE [IEE01], pages 178–187. CODEN ASF-PDV. ISBN 0-7695-1390-5, 0-7695-1391-3 (case), 0-7695-1392-1 (microfiche). ISSN 0272-5428. LCCN ????

Broadbent:2009:UBQ

[BFK09] A. Broadbent, J. Fitzsimons, and E. Kashefi. Universal blind quantum computation. In IEEE [IEE09], pages 517–526. ISBN 0-7695-3850-9. ISSN 0272-5428. LCCN QA76 .S95 2009. IEEE Computer Society order number P3850.

Bodlaender:2009:MK

[BFL^{+}09] H. L. Bodlaender, F. V. Fomin, D. Lokshtanov, E. Penninkx, S. Saurabh, and D. M. Thilikos. (meta) kernelization. In IEEE [IEE09], pages 629–638. ISBN 0-7695-3850-9. ISSN 0272-5428. LCCN QA76 .S95 2009. IEEE Computer Society order number P3850.

Batu:2000:TDC

[BFR^{+}00] T. Batu, L. Fortnow, R. Rubinfeld, W. D. Smith, and P. White. Testing that distributions are close. In IEEE [IEE00], pages 259–269. CO-DEN ASFPDV. ISBN 0-7695-0850-2, 0-7695-0851-0 (case), 0-7695-0852-9 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 2000. IEEE Computer Society Order Number PR00850.

Bshouty:2002:PPO

[BG02] N. H. Bshouty and D. Gavinsky. PAC = PAExact and other equivalent models in learning. In IEEE [IEE02], pages 167–176.

CODEN ASFPDV. ISBN 0-7695-1822-2. ISSN 0272-5428. LCCN QA267. URL http:/

/ieeexplore.ieee.org/iel5/
8411/26517/01181893.pdf?isnumber=
26517&prod=CNF&arnumber=1181893&
arSt=+167&ared=+176&arAuthor=
Bshouty%2C+N.H.%3B+Gavinsky%
2C+D.;   http://ieeexplore.
ieee.org/xpls/abs_all.jsp?
isnumber=26517&arnumber=1181893&
count=82&index=18.  IEEE

Computer Society Order Number PR01822.

Bhattacharjee:2003:IFA

[BG03] R. Bhattacharjee and A. Goel.

Instability of FIFO at arbitrarily low rates in the adversarial queuing model. In IEEE [IEE03], pages 160–167. CO-DEN ASFPDV. ISBN 0-7695-2040-5. ISSN 0272-5428. LCCN QA76 .S979 2003. URL http:/

/ieeexplore.ieee.org/iel5/
8767/27770/01238190.pdf?isnumber=
27770&prod=CNF&arnumber=1238190&
arSt=+160&ared=+167&arAuthor=
Bhattacharjee%2C+R.%3B+Goel%
2C+A.;   http://ieeexplore.
ieee.org/xpls/abs_all.jsp?
isnumber=27770&arnumber=1238190&
count=66&index=17.  IEEE

Computer Society Order Number PR02040.

Blelloch:2007:SHI

[BG07] G. E. Blelloch and D. Golovin.

Strongly history-independent hashing with applications. In IEEE [IEE07], pages 272–

282. ISBN 0-7695-3010-9. ISSN 0272-5428. LCCN QA76 .S974 2007. URL http://

                                        ieeexplore.ieee.org/servlet/
                                        opac?punumber=4389466. IEEE

Computer Society order number

P3010.

Barak:2001:RSZ

[BGGL01] B. Barak, O. Goldreich, S. Goldwasser, and Y. Lindell. Resettablysound zero-knowledge and its applications. In IEEE [IEE01], pages 116–125. CODEN ASF-PDV. ISBN 0-7695-1390-5, 0-7695-1391-3 (case), 0-7695-1392-1 (microfiche). ISSN 0272-5428. LCCN ????

Boneh:2007:SEI

[BGH07] D. Boneh, C. Gentry, and M. Hamburg. Space-efficient identity based encryption without pairings. In IEEE [IEE07], pages 647–657. ISBN 0-7695-3010-9. ISSN 0272-5428. LCCN QA76 .S974 2007. URL http://

                                        ieeexplore.ieee.org/servlet/
                                        opac?punumber=4389466. IEEE

Computer Society order number

P3010.

Bhatnagar:2003:SPZ

[BGL03] N. Bhatnagar, P. Gopalan, and R. J. Lipton. Symmetric polynomials over Z_{m} and simultaneous communication protocols. In IEEE [IEE03], pages 450–459. CODEN ASF-PDV. ISBN 0-7695-2040-5. ISSN 0272-5428. LCCN QA76 .S979 2003. URL http:/

                                        /ieeexplore.ieee.org/iel5/
                                        8767/27770/01238218.pdf?isnumber=
                                        27770&prod=CNF&arnumber=1238218&
                                        arSt=+450&ared=+459&arAuthor=
                                        Bhatnagar%2C+N.%3B+Gopalan%

2C+P.%3B+Lipton%2C+R.J.;
http://ieeexplore.ieee.org/
xpls/abs_all.jsp?isnumber=
27770&arnumber=1238218&count=
66&index=45.  IEEE Com-

puter Society Order Number PR02040.

Bjorklund:2006:IEA

[BH06] Andreas Bjorklund and Thore Husfeldt. Inclusion-exclusion algorithms for counting set partitions. In IEEE [IEE06], pages 575–582. ISBN 0-7695-2720-5, 0-7695-2362-5. ISSN 0272-5428. LCCN QA76 .S974 2006. URL http://

ieeexplore.ieee.org/servlet/
opac?punumber=4031329. IEEE

Computer Society Order Number P2720.

Barak:2008:RPR

[BHH^{+}08] B. Barak, M. Hardt, I. Haviv, A. Rao, O. Regev, and D. Steurer. Rounding parallel repetitions of unique games. In IEEE [IEE08], pages 374– 383. ISBN 0-7695-3436-8. ISSN 0272-5428. LCCN QA76 .S95 2008. URL http://

ieeexplore.ieee.org/servlet/
opac?punumber=4690923. IEEE

Computer Society order number P3436.

Bjorklund:2008:CTP

[BHKK08] A. Bjorklund, T. Husfeldt, P. Kaski, and M. Koivisto. Computing the Tutte polynomial in vertex-exponential time. In IEEE [IEE08], pages 677–686. ISBN 0-7695-3436-8. ISSN 0272-5428. LCCN QA76 .S95 2008. URL http://

                                        ieeexplore.ieee.org/servlet/
                                        opac?punumber=4690923. IEEE

Computer Society order number

P3436.

Beame:2008:VMR

[BHN08] P. Beame and Dang-Trinh Huynh-Ngoc. On the value of multiple read/write streams for approximating frequency moments. In IEEE [IEE08], pages 499–508. ISBN 0-7695-3436-

8. ISSN 0272-5428. LCCN QA76 .S95 2008. URL http://

                                        ieeexplore.ieee.org/servlet/
                                        opac?punumber=4690923. IEEE

Computer Society order number

P3436.

Beame:2009:MCC

[BHN09] P. Beame and Dang-Trinh Huynh-Ngoc. Multiparty communication complexity and threshold circuit size of AC^{0}. In IEEE [IEE09], pages 53–62. ISBN 0-7695-3850-9. ISSN 0272-5428. LCCN QA76 .S95 2009. IEEE Computer Society order number P3850.

Beimel:2002:BSB

[BIKR02] A. Beimel, Y. Ishai, E. Kushilevitz, and J.-F. Raymond. Breaking the O(n^{1/(2k−1)}) barrier for information-theoretic Private Information Retrieval. In IEEE [IEE02], pages 261–270. CODEN ASFPDV. ISBN 0-7695-1822-2. ISSN 0272-5428. LCCN QA267. URL http:/

                                        /ieeexplore.ieee.org/iel5/
                                        8411/26517/01181949.pdf?isnumber=
                                        26517&prod=CNF&arnumber=1181949&

arSt=+261&ared=+270&arAuthor=
Beimel%2C+A.%3B+Ishai%2C+
Y.%3B+Kushilevitz%2C+E.%3B+
Raymond%2C+J.-F.;   http:/
/ieeexplore.ieee.org/xpls/
abs_all.jsp?isnumber=26517&
arnumber=1181949&count=82&

index=27. IEEE Computer Society Order Number PR01822.

Barak:2004:ERU

[BIW04] B. Barak, R. Impagliazzo, and A. Wigderson. Extracting randomness using few independent sources. In IEEE [IEE04], pages 384–393. CO-DEN ASFPDV. ISBN 0-7695-2228-9. ISSN 0272-5428. LCCN QA276. URL http:/

/ieeexplore.ieee.org/iel5/
9430/29918/01366258.pdf?isnumber=
29918&prod=CNF&arnumber=1366258&
arSt=+384&ared=+393&arAuthor=
Barak%2C+B.%3B+Impagliazzo%
2C+R.%3B+Wigderson%2C+A.;
 http://ieeexplore.ieee.org/
xpls/abs_all.jsp?isnumber=
29918&arnumber=1366258&count=

64&index=39. IEEE Computer Society Order Number P2228.

Brodal:2002:DPC

[BJ02] G. S. Brodal and R. Jacob. Dynamic planar convex hull. In IEEE [IEE02], pages 617–626. CODEN ASFPDV. ISBN 0-7695-1822-2. ISSN 0272-5428. LCCN QA267. URL http:/

/ieeexplore.ieee.org/iel5/
8411/26517/01181985.pdf?isnumber=
26517&prod=CNF&arnumber=1181985&
arSt=+617&ared=+626&arAuthor=
Brodal%2C+G.S.%3B+Jacob%2C+

                                        R.;   http://ieeexplore.ieee.
                                        org/xpls/abs_all.jsp?isnumber=
                                        26517&arnumber=1181985&count=
                                        82&index=61.  IEEE Com-

puter Society Order Number PR01822.

Baswana:2006:FAA

[BK06] Surender Baswana and Telikepalli Kavitha. Faster algorithms for approximate distance oracles and all-pairs small stretch paths. In IEEE [IEE06], pages 591–602. ISBN 0-7695-2720-5, 0-7695-2362-5. ISSN 0272-5428. LCCN QA76 .S974 2006. URL http://

                                        ieeexplore.ieee.org/servlet/
                                        opac?punumber=4031329. IEEE

Computer Society Order Number P2720.

Bansal:2009:OLC

[BK09a] N. Bansal and S. Khot. Optimal long code test with one free bit. In IEEE [IEE09], pages 453– 462. ISBN 0-7695-3850-9. ISSN 0272-5428. LCCN QA76 .S95 2009. IEEE Computer Society order number P3850.

Barto:2009:CSP

[BK09b] L. Barto and M. Kozik. Constraint satisfaction problems of bounded width. In IEEE [IEE09], pages 595–603. ISBN 0-7695-3850-9. ISSN 0272-5428. LCCN QA76 .S95 2009. IEEE Computer Society order number P3850.

Borradaile:2008:PTA

[BKM08] G. Borradaile, P. N. Klein, and C. Mathieu. A polynomialtime approximation scheme for Euclidean Steiner forest. In IEEE [IEE08], pages 115–

124. ISBN 0-7695-3436-8. ISSN 0272-5428. LCCN QA76 .S95 2008. URL http://

ieeexplore.ieee.org/servlet/
opac?punumber=4690923. IEEE

Computer Society order number P3436.

Bansal:2004:DSS

[BKP04] N. Bansal, T. Kimbrel, and K. Pruhs. Dynamic speed scaling to manage energy and temperature. In IEEE [IEE04], pages 520–529. CO-DEN ASFPDV. ISBN 0-7695-2228-9. ISSN 0272-5428. LCCN QA276. URL http:/

/ieeexplore.ieee.org/iel5/
9430/29918/01366272.pdf?isnumber=
29918&prod=CNF&arnumber=1366272&
arSt=+520&ared=+529&arAuthor=
Bansal%2C+N.%3B+Kimbrel%2C+
T.%3B+Pruhs%2C+K.;   http:
//ieeexplore.ieee.org/xpls/
abs_all.jsp?isnumber=29918&
arnumber=1366272&count=64&

index=53. IEEE Computer Society Order Number P2228.

Benjamini:2002:GIL

[BL02a] I. Benjamini and L. Lovasz.

Global information from local observation. In IEEE [IEE02], pages 701–710. CO-DEN ASFPDV. ISBN 0-7695-1822-2. ISSN 0272-5428. LCCN QA267. URL http:/

/ieeexplore.ieee.org/iel5/
8411/26517/01181995.pdf?isnumber=
26517&prod=CNF&arnumber=1181995&
arSt=+701&ared=+710&arAuthor=
Benjamini%2C+I.%3B+Lovasz%

                                        2C+L.;   http://ieeexplore.
                                        ieee.org/xpls/abs_all.jsp?
                                        isnumber=26517&arnumber=1181995&
                                        count=82&index=69.  IEEE

Computer Society Order Num-

ber PR01822.

Burgisser:2002:LBB

[BL02b] P. Burgisser and M. Lotz. Lower bounds on the bounded coefficient complexity of bilinear maps. In IEEE [IEE02], pages 659–668. CODEN ASFPDV. ISBN 0-7695-1822-2. ISSN 0272-5428. LCCN QA267. URL http:/

                                        /ieeexplore.ieee.org/iel5/
                                        8411/26517/01181991.pdf?isnumber=
                                        26517&prod=CNF&arnumber=1181991&
                                        arSt=+659&ared=+668&arAuthor=
                                        Burgisser%2C+P.%3B+Lotz%2C+
                                        M.;   http://ieeexplore.ieee.
                                        org/xpls/abs_all.jsp?isnumber=
                                        26517&arnumber=1181991&count=
                                        82&index=65.  IEEE Com-

puter Society Order Number

PR01822.

Bilu:2004:CEG

[BL04] Y. Bilu and N. Linial. Constructing expander graphs by 2-lifts and discrepancy vs. spectral gap. In IEEE [IEE04], pages 404–412. CO-DEN ASFPDV. ISBN 0-7695-2228-9. ISSN 0272-5428. LCCN QA276. URL http:/

                                        /ieeexplore.ieee.org/iel5/
                                        9430/29918/01366260.pdf?isnumber=
                                        29918&prod=CNF&arnumber=1366260&
                                        arSt=+404&ared=+412&arAuthor=
                                        Bilu%2C+Y.%3B+Linial%2C+N.
                                        ;   http://ieeexplore.ieee.
                                        org/xpls/abs_all.jsp?isnumber=

29918&arnumber=1366260&count=

64&index=41. IEEE Computer Society Order Number P2228.

Becchetti:2003:ACS

[BLMS^{+}03] L. Becchetti, S. Leonardi, A. Marchetti-Spaccamela, G. Schafer, and T. Vredeveld. Average case and smoothed competitive analysis of the multi-level feedback algorithm. In IEEE [IEE03], pages 462–471. CO-DEN ASFPDV. ISBN 0-7695-2040-5. ISSN 0272-5428. LCCN QA76 .S979 2003. URL http:/

/ieeexplore.ieee.org/iel5/
8767/27770/01238219.pdf?isnumber=
27770&prod=CNF&arnumber=1238219&
arSt=+462&ared=+471&arAuthor=
Becchetti%2C+L.%3B+Leonardi%
2C+S.%3B+Marchetti-Spaccamela%
2C+A.%3B+Schafer%2C+G.%3B+
Vredeveld%2C+T.;   http://
ieeexplore.ieee.org/xpls/
abs_all.jsp?isnumber=27770&
arnumber=1238219&count=66&

index=46. IEEE Computer Society Order Number PR02040.

Biswal:2008:EBS

[BLR08] P. Biswal, J. R. Lee, and S. Rao. Eigenvalue bounds, spectral partitioning, and metrical deformations via flows. In IEEE [IEE08], pages 751–

760. ISBN 0-7695-3436-8. ISSN 0272-5428. LCCN QA76 .S95 2008. URL http://

ieeexplore.ieee.org/servlet/
opac?punumber=4690923. IEEE

Computer Society order number P3436.

Bansal:2005:TTD

[BLS05] N. Bansal, A. Lodi, and M. Sviridenko. A tale of two dimensional bin packing. In IEEE [IEE05], pages 657–666. ISBN 0-7695-2468-0, 0-7695-2362-5. ISSN 0272-5428. LCCN QA76 .S979 2005. IEEE Computer Society order number P2468.

Blum:2003:MLM

[Blu03] A. Blum. Machine learning: my favorite results, directions, and open problems. In IEEE [IEE03], page 2. CODEN ASFPDV. ISBN 0-7695-2040-

5. ISSN 0272-5428. LCCN QA76 .S979 2003. URL http:/

                                        /ieeexplore.ieee.org/iel5/
                                        8767/27770/01238174.pdf?isnumber=
                                        27770&prod=CNF&arnumber=1238174&
                                        arSt=+2&ared=+2&arAuthor=
                                        Blum%2C+A.;   http://ieeexplore.
                                        ieee.org/xpls/abs_all.jsp?
                                        isnumber=27770&arnumber=1238174&
                                        count=66&index=1. IEEE Com-

puter Society Order Number

PR02040.

Barak:2003:LBN

[BLV03] B. Barak, Y. Lindell, and S. Vadhan. Lower bounds for non-black-box zero knowledge. In IEEE [IEE03], pages 384–393. CODEN ASF-PDV. ISBN 0-7695-2040-5. ISSN 0272-5428. LCCN QA76 .S979 2003. URL http:/

                                        /ieeexplore.ieee.org/iel5/
                                        8767/27770/01238212.pdf?isnumber=
                                        27770&prod=CNF&arnumber=1238212&
                                        arSt=+384&ared=+393&arAuthor=
                                        Barak%2C+B.%3B+Lindell%2C+

Y.%3B+Vadhan%2C+S.;   http:
//ieeexplore.ieee.org/xpls/
abs_all.jsp?isnumber=27770&
arnumber=1238212&count=66&

index=39. IEEE Computer Society Order Number PR02040.

Buchin:2009:DTS

[BM09] K. Buchin and W. Mulzer.

Delaunay triangulations in O(sort(n)) time and more. In IEEE [IEE09], pages 139–148. ISBN 0-7695-3850-9. ISSN 0272-5428. LCCN QA76 .S95 2009. IEEE Computer Society order number P3850.

Barak:2007:LBS

[BMG07] B. Barak and M. Mahmoody-Ghidary. Lower bounds on signatures from symmetric primitives. In IEEE [IEE07], pages 680–688. ISBN 0-7695-3010-

9. ISSN 0272-5428. LCCN QA76 .S974 2007. URL http://

ieeexplore.ieee.org/servlet/
opac?punumber=4389466. IEEE

Computer Society order number P3010.

Bshouty:2003:LDR

[BMOS03] N. Bshouty, E. Mossel, R. O’Donnell, and R. A. Servedio. Learning DNF from random walks. In IEEE [IEE03], pages 189–

198. CODEN ASFPDV. ISBN 0-7695-2040-5. ISSN 0272-5428. LCCN QA76 .S979 2003. URL http:/

/ieeexplore.ieee.org/iel5/
8767/27770/01238193.pdf?isnumber=
27770&prod=CNF&arnumber=1238193&
arSt=+189&ared=+198&arAuthor=
Bshouty%2C+N.%3B+Mossel%2C+

                                        E.%3B+O%27Donnell%2C+R.%3B+
                                        Servedio%2C+R.A.;   http:/
                                        /ieeexplore.ieee.org/xpls/
                                        abs_all.jsp?isnumber=27770&
                                        arnumber=1238193&count=66&

index=20. IEEE Computer Society Order Number PR02040.

Blumrosen:2002:ASB

[BN02] L. Blumrosen and N. Nisan. Auctions with severely bounded communication. In IEEE [IEE02], pages 406–415. CO-DEN ASFPDV. ISBN 0-7695-1822-2. ISSN 0272-5428. LCCN QA267. URL http:/

                                        /ieeexplore.ieee.org/iel5/
                                        8411/26517/01181965.pdf?isnumber=
                                        26517&prod=CNF&arnumber=1181965&
                                        arSt=+406&ared=+415&arAuthor=
                                        Blumrosen%2C+L.%3B+Nisan%
                                        2C+N.;   http://ieeexplore.
                                        ieee.org/xpls/abs_all.jsp?
                                        isnumber=26517&arnumber=1181965&
                                        count=82&index=41.  IEEE

Computer Society Order Num-

ber PR01822.

Buchbinder:2006:IBO

[BN06] Niv Buchbinder and Joseph (Seffi) Naor. Improved bounds for online routing and packing via a primal-dual approach. In IEEE [IEE06], pages 293–304. ISBN 0-7695-2720-5, 0-7695-2362-5. ISSN 0272-5428. LCCN QA76 .S974 2006. URL http://

                                        ieeexplore.ieee.org/servlet/
                                        opac?punumber=4031329. IEEE

Computer Society Order Num-

ber P2720.

Belkin:2006:HFF

[BNN06] Mikhail Belkin, Hariharan Narayanan, and Partha Niyogi. Heat flow and a faster algorithm to compute the surface area of a convex body. In IEEE [IEE06], pages 47–56. ISBN 0-7695-2720-5, 0-7695-2362-5. ISSN 0272-5428. LCCN QA76 .S974 2006. URL http://

ieeexplore.ieee.org/servlet/
opac?punumber=4031329. IEEE

Computer Society Order Number P2720.

Braverman:2007:SHS

[BO07] V. Braverman and R. Ostrovsky.

Smooth histograms for sliding windows. In IEEE [IEE07], pages 283–293. ISBN 0-7695-3010-9. ISSN 0272-5428. LCCN QA76 .S974 2007. URL http://

ieeexplore.ieee.org/servlet/
opac?punumber=4389466. IEEE

Computer Society order number P3010.

Buresh-Oppenheim:2002:BDF

[BOBP^{+}02] J. Buresh-Oppenheim, P. Beame, T. Pitassi, R. Raz, and A. Sabharwal. Bounded-depth Frege lower bounds for weaker pigeonhole principles. In IEEE [IEE02], pages 583–592. CO-DEN ASFPDV. ISBN 0-7695-1822-2. ISSN 0272-5428. LCCN QA267. URL http:/

/ieeexplore.ieee.org/iel5/
8411/26517/01181982.pdf?isnumber=
26517&prod=CNF&arnumber=1181982&
arSt=+583&ared=+592&arAuthor=
Buresh-Oppenheim%2C+J.%3B+
Beame%2C+P.%3B+Pitassi%2C+
T.%3B+Raz%2C+R.%3B+Sabharwal%
2C+A.;   http://ieeexplore.
ieee.org/xpls/abs_all.jsp?

                                        isnumber=26517&arnumber=1181982&
                                        count=82&index=58.  IEEE

Computer Society Order Num-

ber PR01822.

Ben-Or:2006:SMQ

[BOCG^{+}06] Michael Ben-Or, Claude Crepeau, Daniel Gottesman, Avinatan Hassidim, and Adam Smith. Secure multiparty quantum computation with (only) a strict honest majority. In IEEE [IEE06], pages 249–260. ISBN 0-7695-2720-5, 0-7695-2362-5. ISSN 0272-5428. LCCN QA76 .S974 2006. URL http://

                                        ieeexplore.ieee.org/servlet/
                                        opac?punumber=4031329. IEEE

Computer Society Order Num-

ber P2720.

Buresh-Oppenheim:2003:RBI

[BOGH^{+}03] J. Buresh-Oppenheim, N. Galesi, S. Hoory, A. Magen, and T. Pitassi. Rank bounds and integrality gaps for cutting planes procedures. In IEEE [IEE03], pages 318–327. CODEN ASF-PDV. ISBN 0-7695-2040-5. ISSN 0272-5428. LCCN QA76 .S979 2003. URL http:/

                                        /ieeexplore.ieee.org/iel5/
                                        8767/27770/01238206.pdf?isnumber=
                                        27770&prod=CNF&arnumber=1238206&
                                        arSt=+318&ared=+327&arAuthor=
                                        Buresh-Oppenheim%2C+J.%3B+
                                        Galesi%2C+N.%3B+Hoory%2C+
                                        S.%3B+Magen%2C+A.%3B+Pitassi%
                                        2C+T.;   http://ieeexplore.
                                        ieee.org/xpls/abs_all.jsp?
                                        isnumber=27770&arnumber=1238206&
                                        count=66&index=33.  IEEE

Computer Society Order Num-

ber PR02040.

Boneh:2007:BLP

[Bon07]  D. Boneh.  A brief look

at pairings based cryptography. In IEEE [IEE07], pages 19–26. ISBN 0-7695-3010-9. ISSN 0272-5428. LCCN QA76 .S974 2007. URL http://

ieeexplore.ieee.org/servlet/
opac?punumber=4389466. IEEE

Computer Society order number P3010.

Bogdanov:2002:LBT

[BOT02] A. Bogdanov, K. Obata, and L. Trevisan. A lower bound for testing 3-colorability in bounded-degree graphs. In IEEE [IEE02], pages 93–102. CODEN ASFPDV. ISBN 0-7695-1822-2. ISSN 0272-5428. LCCN QA267. URL http:/

/ieeexplore.ieee.org/iel5/
8411/26517/01181886.pdf?isnumber=
26517&prod=CNF&arnumber=1181886&
arSt=+93&ared=+102&arAuthor=
Bogdanov%2C+A.%3B+Obata%2C+
K.%3B+Trevisan%2C+L.;   http:
//ieeexplore.ieee.org/xpls/
abs_all.jsp?isnumber=26517&
arnumber=1181886&count=82&

index=11. IEEE Computer Society Order Number PR01822.

Boneh:2008:IBI

[BPR^{+}08] D. Boneh, P. Papakonstantinou, C. Rackoff, Y. Vahlis, and B. Waters. On the impossibility of basing identity based encryption on trapdoor permutations. In IEEE [IEE08], pages 283–292. ISBN 0-7695-3436-

8. ISSN 0272-5428. LCCN QA76 .S95 2008. URL http://

                                        ieeexplore.ieee.org/servlet/
                                        opac?punumber=4690923. IEEE

Computer Society order number P3436.

Barak:2006:CNM

[BPS06] Boaz Barak, Manoj Prabhakaran, and Amit Sahai. Concurrent non-malleable zero knowledge. In IEEE [IEE06], pages 345–354. ISBN 0-7695-2720-5, 0-7695-2362-5. ISSN 0272-5428. LCCN QA76 .S974 2006. URL http://

                                        ieeexplore.ieee.org/servlet/
                                        opac?punumber=4031329. IEEE

Computer Society Order Number P2720.

Braverman:2005:CRF

[Bra05] M. Braverman. On the complexity of real functions. In IEEE [IEE05], pages 155–164. ISBN 0-7695-2468-0, 0-7695-2362-5. ISSN 0272-5428. LCCN QA76 .S979 2005. IEEE Computer Society order number P2468.

Barak:2005:HPA

[BS05] B. Barak and A. Sahai. How to play almost any mental game over the net — concurrent composition via super-polynomial simulation. In IEEE [IEE05], pages 543–552. ISBN 0-7695-2468-0, 0-7695-2362-5. ISSN 0272-5428. LCCN QA76 .S979 2005. IEEE Computer Society order number P2468.

Baumann:2006:SEP

[BS06] Nadine Baumann and Martin Skutella. Solving evacuation problems efficiently — earliest arrival flows with multiple sources. In IEEE [IEE06], pages 399–410. ISBN 0-7695-2720-5, 0-7695-2362-5. ISSN 0272-5428. LCCN QA76 .S974 2006. URL http://

ieeexplore.ieee.org/servlet/
opac?punumber=4031329. IEEE

Computer Society Order Number P2720.

Bogdanov:2007:HAE

[BS07] A. Bogdanov and M. Safra.

Hardness amplification for errorless heuristics. In IEEE [IEE07], pages 418–426. ISBN 0-7695-3010-9. ISSN 0272-5428. LCCN QA76 .S974 2007. URL http://

ieeexplore.ieee.org/servlet/
opac?punumber=4389466. IEEE

Computer Society order number P3010.

Ben-Sasson:2006:SPL

[BSKR06] Eli Ben-Sasson, Swastik Kopparty, and Jaikumar Radhakrishnan. Subspace polynomials and list decoding of Reed– Solomon codes. In IEEE [IEE06], pages 207–216. ISBN 0-7695-2720-5, 0-7695-2362-5. ISSN 0272-5428. LCCN QA76 .S974 2006. URL http://

ieeexplore.ieee.org/servlet/
opac?punumber=4031329. IEEE

Computer Society Order Number P2720.

Ben-Sasson:2008:SPM

[BSN08] E. Ben-Sasson and J. Nordstrom. Short proofs may be spacious: An optimal separation of space and length in resolution. In IEEE [IEE08], pages 709–718. ISBN 0-7695-3436-

8. ISSN 0272-5428. LCCN QA76 .S95 2008. URL http://

ieeexplore.ieee.org/servlet/
opac?punumber=4690923. IEEE

Computer Society order number P3436.

Babai:2003:LTC

[BSS03] L. Babai, A. Shpilka, and D. Stefankovic. Locally testable cyclic codes. In IEEE [IEE03], pages 116–125. CODEN ASF-PDV. ISBN 0-7695-2040-5. ISSN 0272-5428. LCCN QA76 .S979 2003. URL http:/

/ieeexplore.ieee.org/iel5/
8767/27770/01238186.pdf?isnumber=
27770&prod=CNF&arnumber=1238186&
arSt=+116&ared=+125&arAuthor=
Babai%2C+L.%3B+Shpilka%2C+
A.%3B+Stefankovic%2C+D.;
http://ieeexplore.ieee.org/
xpls/abs_all.jsp?isnumber=
27770&arnumber=1238186&count=
66&index=13.  IEEE Com-

puter Society Order Number PR02040.

Beame:2000:SLT

[BSSV00] P. Beame, M. Saks, Xiaodong Sun, and E. Vee. Superlinear time-space tradeoff lower bounds for randomized computation. In IEEE [IEE00], pages 169–179. CODEN ASF-PDV. ISBN 0-7695-0850-2, 0-7695-0851-0 (case), 0-7695-0852-9 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 2000. IEEE Computer Society Order Number PR00850.

Bogdanov:2003:WCA

[BT03] A. Bogdanov and L. Trevisan. On worst-case to average-case reductions for NP problems. In IEEE [IEE03], pages 308–317. CODEN ASF-PDV. ISBN 0-7695-2040-5. ISSN 0272-5428. LCCN QA76 .S979 2003. URL http:/

/ieeexplore.ieee.org/iel5/
8767/27770/01238205.pdf?isnumber=
27770&prod=CNF&arnumber=1238205&
arSt=+308&ared=+317&arAuthor=
Bogdanov%2C+A.%3B+Trevisan%
2C+L.;   http://ieeexplore.
ieee.org/xpls/abs_all.jsp?
isnumber=27770&arnumber=1238205&
count=66&index=32.  IEEE

Computer Society Order Number PR02040.

Bulatov:2002:DTC

[Bul02] A. A. Bulatov. A dichotomy theorem for constraints on a three-element set. In IEEE [IEE02], pages 649–658. CO-DEN ASFPDV. ISBN 0-7695-1822-2. ISSN 0272-5428. LCCN QA267. URL http:/

/ieeexplore.ieee.org/iel5/
8411/26517/01181990.pdf?isnumber=
26517&prod=CNF&arnumber=1181990&
arSt=+649&ared=+658&arAuthor=
Bulatov%2C+A.A.;   http://
ieeexplore.ieee.org/xpls/
abs_all.jsp?isnumber=26517&
arnumber=1181990&count=82&

index=64. IEEE Computer Society Order Number PR01822.

Burgisser:2001:CFM

[Bur01] P. Burgisser. The complexity of factors of multivariate polynomials. In IEEE [IEE01], pages 378–385. CODEN ASF-PDV. ISBN 0-7695-1390-5, 0-7695-1391-3 (case), 0-7695-1392-1 (microfiche). ISSN 0272-5428. LCCN ????

Bogdanov:2007:PBP

[BV07] A. Bogdanov and E. Viola. Pseudorandom bits for polynomials. In IEEE [IEE07], pages 41–51. ISBN 0-7695-3010-9. ISSN 0272-5428. LCCN QA76 .S974 2007. URL http://

                                        ieeexplore.ieee.org/servlet/
                                        opac?punumber=4389466. IEEE

Computer Society order number P3010.

Brubaker:2008:IPA

[BV08] S. C. Brubaker and S. Vempala. Isotropic PCA and affine-invariant clustering. In IEEE [IEE08], pages 551– 560. ISBN 0-7695-3436-8. ISSN 0272-5428. LCCN QA76 .S95 2008. URL http://

                                        ieeexplore.ieee.org/servlet/
                                        opac?punumber=4690923. IEEE

Computer Society order number P3436.

Barany:2005:NER

[BVV05] I. Barany, S. Vempala, and A. Vetta. Nash equilibria in random games. In IEEE [IEE05], pages 123–131. ISBN 0-7695-2468-0, 0-7695-2362-5. ISSN 0272-5428. LCCN QA76 .S979 2005. IEEE Computer Society order number P2468.

Beimel:2003:SPM

[BW03] A. Beimel and E. Weinreb. Separating the power of monotone span programs over different fields. In IEEE [IEE03], pages 428–501. CODEN ASF-PDV. ISBN 0-7695-2040-5. ISSN 0272-5428. LCCN QA76 .S979 2003. URL http:/

/ieeexplore.ieee.org/iel5/
8767/27770/01238216.pdf?isnumber=
27770&prod=CNF&arnumber=1238216&
arSt=+428&ared=+501&arAuthor=
Beimel%2C+A.%3B+Weinreb%2C+
E.;   http://ieeexplore.ieee.
org/xpls/abs_all.jsp?isnumber=
27770&arnumber=1238216&count=
66&index=43.  IEEE Com-

puter Society Order Number PR02040.

Bansal:2009:RLC

[BW09] N. Bansal and R. Williams.

Regularity lemmas and combinatorial algorithms. In IEEE [IEE09], pages 745–754. ISBN 0-7695-3850-9. ISSN 0272-5428. LCCN QA76 .S95 2009. IEEE Computer Society order number P3850.

Bar-Yossef:2006:ICS

[BYBJK06] Ziv Bar-Yossef, Yitzhak Birk, T. S. Jayram, and Tomer Kol. Index coding with side information. In IEEE [IEE06], pages 197–206. ISBN 0-7695-2720-5, 0-7695-2362-5. ISSN 0272-5428. LCCN QA76 .S974 2006. URL http://

ieeexplore.ieee.org/servlet/
opac?punumber=4031329. IEEE

Computer Society Order Number P2720.

Bar-Yossef:2004:AED

[BYJKK04] Z. Bar-Yossef, T. S. Jayram, R. Krauthgamer, and R. Kumar. Approximating edit distance efficiently. In IEEE [IEE04], pages 550–559. CO-DEN ASFPDV. ISBN 0-7695-2228-9. ISSN 0272-5428. LCCN QA276. URL http:/

/ieeexplore.ieee.org/iel5/
9430/29918/01366275.pdf?isnumber=
29918&prod=CNF&arnumber=1366275&
arSt=+550&ared=+559&arAuthor=
Bar-Yossef%2C+Z.%3B+Jayram%
2C+T.S.%3B+Krauthgamer%2C+
R.%3B+Kumar%2C+R.;   http:
//ieeexplore.ieee.org/xpls/
abs_all.jsp?isnumber=29918&
arnumber=1366275&count=64&

index=56. IEEE Computer Society Order Number P2228.

Bar-Yossef:2002:ISA

[BYJKS02] Z. Bar-Yossef, T. S. Jayram, R. Kumar, and D. Sivakumar. An information statistics approach to data stream and communication complexity. In IEEE [IEE02], pages 209–218. CODEN ASFPDV. ISBN 0-7695-1822-2. ISSN 0272-5428. LCCN QA267. URL http:/

/ieeexplore.ieee.org/iel5/
8411/26517/01181944.pdf?isnumber=
26517&prod=CNF&arnumber=1181944&
arSt=+209&ared=+218&arAuthor=
Bar-Yossef%2C+Z.%3B+Jayram%
2C+T.S.%3B+Kumar%2C+R.%3B+
Sivakumar%2C+D.;   http://
ieeexplore.ieee.org/xpls/
abs_all.jsp?isnumber=26517&

arnumber=1181944&count=82&

index=22. IEEE Computer Society Order Number PR01822.

Basu:2009:PHB

[BZ09] S. Basu and T. Zell. Polynomial hierarchy, Betti numbers and a real analogue of Toda’s theorem. In IEEE [IEE09], pages 73–82. ISBN 0-7695-3850-9. ISSN 0272-5428. LCCN QA76 .S95 2009. IEEE Computer Society order number P3850.

Cai:2001:Z

[Cai01] Jin-Yi Cai. S_{2} ZPP^{NP}. In IEEE [IEE01], pages 620–628. CO-DEN ASFPDV. ISBN 0-7695-1390-5, 0-7695-1391-3 (case), 0-7695-1392-1 (microfiche). ISSN 0272-5428. LCCN ????

Canetti:2001:UCS

[Can01] R. Canetti. Universally composable security: a new paradigm for cryptographic protocols. In IEEE [IEE01], pages 136–

145. CODEN ASFPDV. ISBN 0-7695-1390-5, 0-7695-1391-3 (case), 0-7695-1392-1 (microfiche). ISSN 0272-5428. LCCN ????

Caprara:2002:PDB

[Cap02]  A. Caprara.  Packing 2-

dimensional bins in harmony. In IEEE [IEE02], pages 490–499. CODEN ASFPDV. ISBN 0-7695-1822-2. ISSN 0272-5428. LCCN QA267. URL http:/

/ieeexplore.ieee.org/iel5/
8411/26517/01181973.pdf?isnumber=
26517&prod=CNF&arnumber=1181973&
arSt=+490&ared=+499&arAuthor=

                                        Caprara%2C+A.;   http://ieeexplore.
                                        ieee.org/xpls/abs_all.jsp?
                                        isnumber=26517&arnumber=1181973&
                                        count=82&index=49.  IEEE

Computer Society Order Num-

ber PR01822.

Chakrabarty:2009:AGM

[CCK09] D. Chakrabarty, J. Chuzhoy, and S. Khanna. On allocating goods to maximize fairness. In IEEE [IEE09], pages 107–116. ISBN 0-7695-3850-9. ISSN 0272-5428. LCCN QA76 .S95 2009. IEEE Computer Society order number P3850.

Chen:2006:SCT

[CD06] Xi Chen and Xiaotie Deng. Settling the complexity of twoplayer Nash equilibrium. In IEEE [IEE06], pages 261–272. ISBN 0-7695-2720-5, 0-7695-2362-5. ISSN 0272-5428. LCCN QA76 .S974 2006. URL http://

                                        ieeexplore.ieee.org/servlet/
                                        opac?punumber=4031329. IEEE

Computer Society Order Num-

ber P2720.

Chen:2009:SCA

[CDDT09] Xi Chen, Decheng Dai, Ye Du, and Shang-Hua Teng. Settling the complexity of arrow–Debreu equilibria in markets with additively separable utilities. In IEEE [IEE09], pages 273–282. ISBN 0-7695-3850-9. ISSN 0272-5428. LCCN QA76 .S95 2009. IEEE Computer Society order number P3850.

Cryan:2002:RMM

[CDG^{+}02] M. Cryan, M. Dyer, L. A. Goldberg, M. Jerrum, and R. Martin. Rapidly mixing Markov chains for sampling contingency tables with a constant number of rows. In IEEE [IEE02], pages 711–

720. CODEN ASFPDV. ISBN 0-7695-1822-2. ISSN 0272-5428. LCCN QA267. URL http:/

/ieeexplore.ieee.org/iel5/
8411/26517/01181996.pdf?isnumber=
26517&prod=CNF&arnumber=1181996&
arSt=+711&ared=+720&arAuthor=
Cryan%2C+M.%3B+Dyer%2C+M.%
3B+Goldberg%2C+L.A.%3B+Jerrum%
2C+M.%3B+Martin%2C+R.;   http:
//ieeexplore.ieee.org/xpls/
abs_all.jsp?isnumber=26517&
arnumber=1181996&count=82&

index=70. IEEE Computer Society Order Number PR01822.

Connelly:2000:SPA

[CDR00] R. Connelly, E. D. Demaine, and G. Rote. Straightening polygonal arcs and convexifying polygonal cycles. In IEEE [IEE00], pages 432–442. CO-DEN ASFPDV. ISBN 0-7695-0850-2, 0-7695-0851-0 (case), 0-7695-0852-9 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 2000. IEEE Computer Society Order Number PR00850.

Chen:2006:CNE

[CDT06] Xi Chen, Xiaotie Deng, and Shang-Hua Teng. Computing Nash equilibria: Approximation and smoothed complexity. In IEEE [IEE06], pages 603–612.

ISBN 0-7695-2720-5, 0-7695-2362-5. ISSN 0272-5428. LCCN QA76 .S974 2006. URL http://

ieeexplore.ieee.org/servlet/
opac?punumber=4031329. IEEE

Computer Society Order Number P2720.

Ciriani:2002:SOT

[CFLM02] V. Ciriani, P. Ferragina, F. Luccio, and S. Muthukrishnan. Static optimality theorem for external memory string access. In IEEE [IEE02], pages 219–

227. CODEN ASFPDV. ISBN 0-7695-1822-2. ISSN 0272-5428. LCCN QA267. URL http:/

/ieeexplore.ieee.org/iel5/
8411/26517/01181945.pdf?isnumber=
26517&prod=CNF&arnumber=1181945&
arSt=+219&ared=+227&arAuthor=
Ciriani%2C+V.%3B+Ferragina%
2C+P.%3B+Luccio%2C+F.%3B+
Muthukrishnan%2C+S.;   http:
//ieeexplore.ieee.org/xpls/
abs_all.jsp?isnumber=26517&
arnumber=1181945&count=82&

index=23. IEEE Computer Society Order Number PR01822.

Chan:2008:DBN

[CFLY08] Y. H. Chan, W. S. Fung, L. C. Lau, and C. K. Yung. Degree bounded network design with metric costs. In IEEE [IEE08], pages 125–134. ISBN 0-7695-3436-8. ISSN 0272-5428. LCCN QA76 .S95 2008. URL http://

ieeexplore.ieee.org/servlet/
opac?punumber=4690923. IEEE

Computer Society order number P3436.

Chakrabarty:2008:ABA

[CG08] D. Chakrabarty and G. Goel.

On the approximability of budgeted allocations and improved lower bounds for submodular welfare maximization and GAP. In IEEE [IEE08], pages 687–696. ISBN 0-7695-3436-

8. ISSN 0272-5428. LCCN QA76 .S95 2008. URL http://

ieeexplore.ieee.org/servlet/
opac?punumber=4690923. IEEE

Computer Society order number P3436.

Charikar:2000:CFS

[CGK^{+}00] M. Charikar, V. Guruswami, R. Kumar, S. Rajagopalan, and A. Sahai. Combinatorial feature selection problems. In IEEE [IEE00], pages 631–640. CO-DEN ASFPDV. ISBN 0-7695-0850-2, 0-7695-0851-0 (case), 0-7695-0852-9 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 2000. IEEE Computer Society Order Number PR00850.

Charikar:2004:IRA

[CGK04] M. Charikar, M. X. Goemans, and H. Karloff. On the integrality ratio for asymmetric TSP. In IEEE [IEE04], pages 101–

107. CODEN ASFPDV. ISBN 0-7695-2228-9. ISSN 0272-5428. LCCN QA276. URL http:/

/ieeexplore.ieee.org/iel5/
9430/29918/01366229.pdf?isnumber=
29918&prod=CNF&arnumber=1366229&
arSt=+101&ared=+107&arAuthor=
Charikar%2C+M.%3B+Goemans%
2C+M.X.%3B+Karloff%2C+H.;
 http://ieeexplore.ieee.org/

                                        xpls/abs_all.jsp?isnumber=
                                        29918&arnumber=1366229&count=

64&index=10. IEEE Computer Society Order Number P2228.

Chuzhoy:2004:MMS

[CGKN04] J. Chuzhoy, S. Guha, S. Khanna, and J. S. Naor. Machine minimization for scheduling jobs with interval constraints. In IEEE [IEE04], pages 81–90. CODEN ASFPDV. ISBN 0-7695-2228-9. ISSN 0272-5428. LCCN QA276. URL http:/

                                        /ieeexplore.ieee.org/iel5/
                                        9430/29918/01366227.pdf?isnumber=
                                        29918&prod=CNF&arnumber=1366227&
                                        arSt=+81&ared=+90&arAuthor=
                                        Chuzhoy%2C+J.%3B+Guha%2C+
                                        S.%3B+Khanna%2C+S.%3B+Naor%
                                        2C+J.S.;   http://ieeexplore.
                                        ieee.org/xpls/abs_all.jsp?
                                        isnumber=29918&arnumber=1366227&
                                        count=64&index=8. IEEE Com-

puter Society Order Number

P2228.

Comon:2001:CGT

[CGN01] H. Comon, G. Godoy, and R. Nieuwenhuis. The confluence of ground term rewrite systems is decidable in polynomial time. In IEEE [IEE01], pages 298–307. CODEN ASF-PDV. ISBN 0-7695-1390-5, 0-7695-1391-3 (case), 0-7695-1392-1 (microfiche). ISSN 0272-5428. LCCN ????

Chandran:2007:CMP

[CGOS07] N. Chandran, V. Goyal, R. Ostrovsky, and A. Sahai. Covert multi-party computation. In IEEE [IEE07], pages 238– 248. ISBN 0-7695-3010-9. ISSN 0272-5428. LCCN QA76 .S974 2007. URL http://

ieeexplore.ieee.org/servlet/
opac?punumber=4389466. IEEE

Computer Society order number P3010.

Chattopadhyay:2006:LBC

[CGPT06] Arkadev Chattopadhyay, Navin Goyal, Pavel Pudlak, and Denis Therien. Lower bounds for circuits with MOD_{m} gates. In IEEE [IEE06], pages 709–718. ISBN 0-7695-2720-5, 0-7695-2362-5. ISSN 0272-5428. LCCN QA76 .S974 2006. URL http://

ieeexplore.ieee.org/servlet/
opac?punumber=4031329. IEEE

Computer Society Order Number P2720.

Chrobak:2000:FBG

[CGR00] M. Chrobak, L. Gasieniec, and W. Rytter. Fast broadcasting and gossiping in radio networks. In IEEE [IEE00], pages 575–581. CODEN ASF-PDV. ISBN 0-7695-0850-2, 0-7695-0851-0 (case), 0-7695-0852-9 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 2000. IEEE Computer Society Order Number PR00850.

Chaudhuri:2003:PTM

[CGRT03] K. Chaudhuri, B. Godfrey, S. Rao, and K. Talwar. Paths, trees, and minimum latency tours. In IEEE [IEE03], pages 36–45. CODEN ASF-PDV. ISBN 0-7695-2040-5. ISSN 0272-5428. LCCN QA76 .S979 2003. URL http:/

/ieeexplore.ieee.org/iel5/
8767/27770/01238179.pdf?isnumber=
27770&prod=CNF&arnumber=1238179&
arSt=+36&ared=+45&arAuthor=
Chaudhuri%2C+K.%3B+Godfrey%
2C+B.%3B+Rao%2C+S.%3B+Talwar%
2C+K.;   http://ieeexplore.
ieee.org/xpls/abs_all.jsp?
isnumber=27770&arnumber=1238179&
count=66&index=6. IEEE Com-

puter Society Order Number PR02040.

Charikar:2003:CQI

[CGW03] M. Charikar, V. Guruswami, and A. Wirth. Clustering with qualitative information. In IEEE [IEE03], pages 524–533. CODEN ASFPDV. ISBN 0-7695-2040-5. ISSN 0272-5428. LCCN QA76 .S979 2003. URL http:/

/ieeexplore.ieee.org/iel5/
8767/27770/01238225.pdf?isnumber=
27770&prod=CNF&arnumber=1238225&
arSt=+524&ared=+533&arAuthor=
Charikar%2C+M.%3B+Guruswami%
2C+V.%3B+Wirth%2C+A.;   http:
//ieeexplore.ieee.org/xpls/
abs_all.jsp?isnumber=27770&
arnumber=1238225&count=66&

index=52. IEEE Computer Society Order Number PR02040.

Chan:2000:LAC

[Cha00] T. M. Chan. On levels in arrangements of curves. In IEEE [IEE00], pages 219–227. CO-DEN ASFPDV. ISBN 0-7695-0850-2, 0-7695-0851-0 (case), 0-7695-0852-9 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 2000. IEEE Computer Society Order Number PR00850.

Chan:2002:LDL

[Cha02] T. M. Chan. Low-dimensional linear programming with violations. In IEEE [IEE02], pages 570–579. CODEN ASFPDV. ISBN 0-7695-1822-2. ISSN 0272-5428. LCCN QA267. URL http:/

/ieeexplore.ieee.org/iel5/
8411/26517/01181981.pdf?isnumber=
26517&prod=CNF&arnumber=1181981&
arSt=+570&ared=+579&arAuthor=
Chan%2C+T.M.;   http://ieeexplore.
ieee.org/xpls/abs_all.jsp?
isnumber=26517&arnumber=1181981&
count=82&index=57.  IEEE

Computer Society Order Number PR01822.

Chan:2003:LAC

[Cha03] T. M. Chan. On levels in arrangements of curves. II. A simple inequality and its consequences. In IEEE [IEE03], pages 544–550. CODEN ASF-PDV. ISBN 0-7695-2040-5. ISSN 0272-5428. LCCN QA76 .S979 2003. URL http:/

/ieeexplore.ieee.org/iel5/
8767/27770/01238227.pdf?isnumber=
27770&prod=CNF&arnumber=1238227&
arSt=+544&ared=+550&arAuthor=
Chan%2C+T.M.;   http://ieeexplore.
ieee.org/xpls/abs_all.jsp?
isnumber=27770&arnumber=1238227&
count=66&index=54.  IEEE

Computer Society Order Number PR02040.

Chazelle:2005:ATT

[Cha05] B. Chazelle. Algorithmic techniques and tools from computational geometry. In IEEE [IEE05], page 7. ISBN 0-7695-2468-0, 0-7695-2362-5. ISSN 0272-5428. LCCN QA76 .S979 2005. IEEE Computer Society order number P2468.

Chan:2006:PLL

[Cha06] Timothy M. Chan. Point location in o(log n) time, Voronoi diagrams in o(n log n) time, and other transdichotomous results in computational geometry. In IEEE [IEE06], pages 333–344. ISBN 0-7695-2720-5, 0-7695-2362-5. ISSN 0272-5428. LCCN QA76 .S974 2006. URL http://

                                        ieeexplore.ieee.org/servlet/
                                        opac?punumber=4031329. IEEE

Computer Society Order Number P2720.

Chattopadhyay:2007:DPB

[Cha07] A. Chattopadhyay. Discrepancy and the power of bottom fan-in in depth-three circuits. In IEEE [IEE07], pages 449–458. ISBN 0-7695-3010-9. ISSN 0272-5428. LCCN QA76 .S974 2007. URL http://

                                        ieeexplore.ieee.org/servlet/
                                        opac?punumber=4389466. IEEE

Computer Society order number P3010.

Cheng:2007:DSC

[Che07] Qi Cheng. Derandomization of sparse cyclotomic integer zero testing. In IEEE [IEE07], pages 74–80. ISBN 0-7695-3010-9. ISSN 0272-5428. LCCN QA76 .S974 2007. URL http://

                                        ieeexplore.ieee.org/servlet/

opac?punumber=4389466. IEEE

Computer Society order number P3010.

Chakraborty:2009:DNU

[CHK09] T. Chakraborty, Zhiyi Huang, and S. Khanna. Dynamic and non-uniform pricing strategies for revenue maximization. In IEEE [IEE09], pages 495–504. ISBN 0-7695-3850-9. ISSN 0272-5428. LCCN QA76 .S95 2009. IEEE Computer Society order number P3850.

Chekuri:2006:AAN

[CHKS06] C. Chekuri, M. T. Hajiaghayi, G. Kortsarz, and M. R. Salavatipour. Approximation algorithms for non-uniform buyat-bulk network design. In IEEE [IEE06], pages 677–686. ISBN 0-7695-2720-5, 0-7695-2362-5. ISSN 0272-5428. LCCN QA76 .S974 2006. URL http://

ieeexplore.ieee.org/servlet/
opac?punumber=4031329. IEEE

Computer Society Order Number P2720.

Chlamtac:2007:AAU

[Chl07] E. Chlamtac. Approximation algorithms using hierarchies of semidefinite programming relaxations. In IEEE [IEE07], pages 691–701. ISBN 0-7695-3010-9. ISSN 0272-5428. LCCN QA76 .S974 2007. URL http://

ieeexplore.ieee.org/servlet/
opac?punumber=4389466. IEEE

Computer Society order number P3010.

Chow:2008:ANP

[Cho08] T. Y. Chow. Almost-natural proofs. In IEEE [IEE08], pages 86–91. ISBN 0-7695-3436-8. ISSN 0272-5428. LCCN QA76 .S95 2008. URL http://

                                        ieeexplore.ieee.org/servlet/
                                        opac?punumber=4690923. IEEE

Computer Society order number P3436.

Chakrabarti:2008:ETG

[CJLV08] A. Chakrabarti, A. Jaffe, J. R. Lee, and J. Vincent. Embeddings of topological graphs: Lossy invariants, linearization, and 2-sums. In IEEE [IEE08], pages 761–770. ISBN 0-7695-3436-8. ISSN 0272-5428. LCCN QA76 .S95 2008. URL http://

                                        ieeexplore.ieee.org/servlet/
                                        opac?punumber=4690923. IEEE

Computer Society order number P3436.

Chuzhoy:2008:ASS

[CK08] J. Chuzhoy and S. Khanna. Algorithms for single-source vertex connectivity. In IEEE [IEE08], pages 105–114. ISBN 0-7695-3436-8. ISSN 0272-5428. LCCN QA76 .S95 2008. URL http://

                                        ieeexplore.ieee.org/servlet/
                                        opac?punumber=4690923. IEEE

Computer Society order number P3436.

Chailloux:2009:OQS

[CK09a] A. Chailloux and I. Kerenidis. Optimal quantum strong coin flipping. In IEEE [IEE09], pages 527–533. ISBN 0-7695-3850-9. ISSN 0272-5428. LCCN QA76 .S95 2009. IEEE Computer Society order number P3850.

Chuzhoy:2009:KLA

[CK09b] J. Chuzhoy and S. Khanna. An O(k^{3} log n)-approximation algorithm for vertex-connectivity survivable network design. In IEEE [IEE09], pages 437–441. ISBN 0-7695-3850-9. ISSN 0272-5428. LCCN QA76 .S95 2009. IEEE Computer Society order number P3850.

Chierichetti:2009:MCW

[CKL^{+}09] F. Chierichetti, R. Kumar, S. Lattanzi, A. Panconesi, and P. Raghavan. Models for the compressible Web. In IEEE [IEE09], pages 331–340. ISBN 0-7695-3850-9. ISSN 0272-5428. LCCN QA76 .S95 2009. IEEE Computer Society order number P3850.

Cheeger:2009:IGS

[CKN09] J. Cheeger, B. Kleiner, and A. Naor. A (log n)^{Omega(1)} integrality gap for the sparsest cut SDP. In IEEE [IEE09], pages 555–564. ISBN 0-7695-3850-9. ISSN 0272-5428. LCCN QA76 .S95 2009. IEEE Computer Society order number P3850.

Cheriyan:2001:ADM

[CKR01] J. Cheriyan, H. Karloff, and Y. Rabani. Approximating directed multicuts. In IEEE [IEE01], pages 320–328. CO-DEN ASFPDV. ISBN 0-7695-1390-5, 0-7695-1391-3 (case), 0-7695-1392-1 (microfiche). ISSN 0272-5428. LCCN ????

Chekuri:2004:EDP

[CKS04] C. Chekuri, S. Khanna, and F. B. Shepherd. Edge-disjoint paths in planar graphs. In IEEE [IEE04], pages 71–80. CODEN ASFPDV. ISBN 0-7695-2228-9. ISSN 0272-5428. LCCN QA276. URL http:/

/ieeexplore.ieee.org/iel5/
9430/29918/01366226.pdf?isnumber=
29918&prod=CNF&arnumber=1366226&
arSt=+71&ared=+80&arAuthor=
Chekuri%2C+C.%3B+Khanna%2C+
S.%3B+Shepherd%2C+F.B.;
http://ieeexplore.ieee.org/
xpls/abs_all.jsp?isnumber=
29918&arnumber=1366226&count=

64&index=7. IEEE Computer Society Order Number P2228.

Cohn:2005:GTA

[CKSU05] H. Cohn, R. Kleinberg, B. Szegedy, and C. Umans. Group-theoretic algorithms for matrix multiplication. In IEEE [IEE05], pages 379–388. ISBN 0-7695-2468-0, 0-7695-2362-5. ISSN 0272-5428. LCCN QA76 .S979 2005. IEEE Computer Society order number P2468.

Cornuejols:2003:PAR

[CLV03] G. Cornuejols, Xinming Liu, and K. Vuskovic. A polynomial algorithm for recognizing perfect graphs. In IEEE [IEE03], pages 20–27. CODEN ASFPDV. ISBN 0-7695-2040-

5. ISSN 0272-5428. LCCN QA76 .S979 2003. URL http:/

/ieeexplore.ieee.org/iel5/
8767/27770/01238177.pdf?isnumber=
27770&prod=CNF&arnumber=1238177&

arSt=+20&ared=+27&arAuthor=
Cornuejols%2C+G.%3B+Xinming+
Liu%3B+Vuskovic%2C+K.;   http:
//ieeexplore.ieee.org/xpls/
abs_all.jsp?isnumber=27770&
arnumber=1238177&count=66&

index=4. IEEE Computer Society Order Number PR02040.

Cai:2008:HAF

[CLX08] Jin-Yi Cai, Pinyan Lu, and Mingji Xia. Holographic algorithms by Fibonacci gates and holographic reductions for hardness. In IEEE [IEE08], pages 644–653. ISBN 0-7695-3436-

8. ISSN 0272-5428. LCCN QA76 .S95 2008. URL http://

ieeexplore.ieee.org/servlet/
opac?punumber=4690923. IEEE

Computer Society order number P3436.

Connamacher:2004:EST

[CM04] H. Connamacher and M. Molloy. The exact satisfiability threshold for a potentially intractable random constraint satisfaction problem. In IEEE [IEE04], pages 590–599. CO-DEN ASFPDV. ISBN 0-7695-2228-9. ISSN 0272-5428. LCCN QA276. URL http:/

/ieeexplore.ieee.org/iel5/
9430/29918/01366279.pdf?isnumber=
29918&prod=CNF&arnumber=1366279&
arSt=+590&ared=+599&arAuthor=
Connamacher%2C+H.%3B+Molloy%
2C+M.;   http://ieeexplore.
ieee.org/xpls/abs_all.jsp?
isnumber=29918&arnumber=1366279&
count=64&index=60.  IEEE

Computer Society Order Number P2228.

Chan:2008:DTR

[CM08] Siu On Chan and M. Molloy. A dichotomy theorem for the resolution complexity of random constraint satisfaction problems. In IEEE [IEE08], pages 634–643. ISBN 0-7695-3436-8. ISSN 0272-5428. LCCN QA76 .S95 2008. URL http://

                                        ieeexplore.ieee.org/servlet/
                                        opac?punumber=4690923. IEEE

Computer Society order number P3436.

Chlamtac:2006:HPU

[CMM06] Eden Chlamtac, Konstantin Makarychev, and Yury Makarychev. How to play unique games using embeddings. In IEEE [IEE06], pages 687–696. ISBN 0-7695-2720-5, 0-7695-2362-5. ISSN 0272-5428. LCCN QA76 .S974 2006. URL http://

                                        ieeexplore.ieee.org/servlet/
                                        opac?punumber=4031329. IEEE

Computer Society Order Number P2720.

Charikar:2007:LGT

[CMM07a] M. Charikar, K. Makarychev, and Y. Makarychev. Local global tradeoffs in metric embeddings. In IEEE [IEE07], pages 713–723. ISBN 0-7695-3010-9. ISSN 0272-5428. LCCN QA76 .S974 2007. URL http://

                                        ieeexplore.ieee.org/servlet/
                                        opac?punumber=4389466. IEEE

Computer Society order number P3010.

Charikar:2007:ARM

[CMM07b] M. Charikar, K. Makarychev, and Y. Makarychev. On the advantage over random for maximum acyclic subgraph. In IEEE [IEE07], pages 625–

633. ISBN 0-7695-3010-9. ISSN 0272-5428. LCCN QA76 .S974 2007. URL http://

ieeexplore.ieee.org/servlet/
opac?punumber=4389466. IEEE

Computer Society order number P3010.

Chuzhoy:2002:CPH

[CN02] J. Chuzhoy and J. S. Naor.

Covering problems with hard capacities. In IEEE [IEE02], pages 481–489. CODEN ASFPDV. ISBN 0-7695-1822-2. ISSN 0272-5428. LCCN QA267. URL http:/

/ieeexplore.ieee.org/iel5/
8411/26517/01181972.pdf?isnumber=
26517&prod=CNF&arnumber=1181972&
arSt=+481&ared=+489&arAuthor=
Chuzhoy%2C+J.%3B+Naor%2C+
J.S.;   http://ieeexplore.
ieee.org/xpls/abs_all.jsp?
isnumber=26517&arnumber=1181972&
count=82&index=48.  IEEE

Computer Society Order Number PR01822.

Chuzhoy:2004:HML

[CN04] J. Chuzhoy and J. S. Naor. The hardness of metric labeling. In IEEE [IEE04], pages 108–114. CODEN ASFPDV. ISBN 0-7695-2228-9. ISSN 0272-5428. LCCN QA276. URL http:/

/ieeexplore.ieee.org/iel5/
9430/29918/01366230.pdf?isnumber=
29918&prod=CNF&arnumber=1366230&
arSt=+108&ared=+114&arAuthor=
Chuzhoy%2C+J.%3B+Naor%2C+

                                        J.S.;   http://ieeexplore.
                                        ieee.org/xpls/abs_all.jsp?
                                        isnumber=29918&arnumber=1366230&
                                        count=64&index=11.  IEEE

Computer Society Order Num-

ber P2228.

Chuzhoy:2001:AAJ

[COR01] J. Chuzhoy, R. Ostrovsky, and Y. Rabani. Approximation algorithms for the job interval selection problem and related scheduling problems. In IEEE [IEE01], pages 348–356. CO-DEN ASFPDV. ISBN 0-7695-1390-5, 0-7695-1391-3 (case), 0-7695-1392-1 (microfiche). ISSN 0272-5428. LCCN ????

Chekuri:2005:RGA

[CP05] Chandra Chekuri and M. Pal. A recursive greedy algorithm for walks in directed graphs. In IEEE [IEE05], pages 245–

253. ISBN 0-7695-2468-0, 0-7695-2362-5. ISSN 0272-5428. LCCN QA76 .S979 2005. IEEE Computer Society order number P2468.

Chan:2008:DCC

[CPR08] T. M. Chan, M. Patrascu, and L. Roditty. Dynamic connectivity: Connecting to networks and geometry. In IEEE [IEE08], pages 95–104. ISBN 0-7695-3436-8. ISSN 0272-5428. LCCN QA76 .S95 2008. URL http://

                                        ieeexplore.ieee.org/servlet/
                                        opac?punumber=4690923. IEEE

Computer Society order number

P3436.

Canetti:2007:CSH

[CPS07] R. Canetti, R. Pass, and A. Shelat. Cryptography from sunspots: How to use an imperfect reference string. In IEEE [IEE07], pages 249–

259. ISBN 0-7695-3010-9. ISSN 0272-5428. LCCN QA76 .S974 2007. URL http://

ieeexplore.ieee.org/servlet/
opac?punumber=4389466. IEEE

Computer Society order number P3010.

Czumaj:2003:BAR

[CR03] A. Czumaj and W. Rytter.

Broadcasting algorithms in radio networks with unknown topology. In IEEE [IEE03], pages 492–501. CODEN ASF-PDV. ISBN 0-7695-2040-5. ISSN 0272-5428. LCCN QA76 .S979 2003. URL http:/

/ieeexplore.ieee.org/iel5/
8767/27770/01238222.pdf?isnumber=
27770&prod=CNF&arnumber=1238222&
arSt=+492&ared=+501&arAuthor=
Czumaj%2C+A.%3B+Rytter%2C+
W.;   http://ieeexplore.ieee.
org/xpls/abs_all.jsp?isnumber=
27770&arnumber=1238222&count=
66&index=49.  IEEE Com-

puter Society Order Number PR02040.

Chakrabarti:2004:ORC

[CR04] A. Chakrabarti and O. Regev.

An optimal randomised cell probe lower bound for approximate nearest neighbour searching. In IEEE [IEE04], pages 473–482. CODEN ASFPDV. ISBN 0-7695-2228-9. ISSN 0272-5428. LCCN QA276. URL http:/

/ieeexplore.ieee.org/iel5/
9430/29918/01366267.pdf?isnumber=
29918&prod=CNF&arnumber=1366267&
arSt=+473&ared=+482&arAuthor=
Chakrabarti%2C+A.%3B+Regev%
2C+O.;   http://ieeexplore.
ieee.org/xpls/abs_all.jsp?
isnumber=29918&arnumber=1366267&
count=64&index=48.  IEEE

Computer Society Order Number P2228.

Charikar:2002:DRS

[CS02a] M. Charikar and A. Sahai. Dimension reduction in the ℓ_{1} norm. In IEEE [IEE02], pages 551–560. CO-DEN ASFPDV. ISBN 0-7695-1822-2. ISSN 0272-5428. LCCN QA267. URL http:/

/ieeexplore.ieee.org/iel5/
8411/26517/01181979.pdf?isnumber=
26517&prod=CNF&arnumber=1181979&
arSt=+551&ared=+560&arAuthor=
Charikar%2C+M.%3B+Sahai%2C+
A.;   http://ieeexplore.ieee.
org/xpls/abs_all.jsp?isnumber=
26517&arnumber=1181979&count=
82&index=55.  IEEE Com-

puter Society Order Number PR01822.

Czumaj:2002:ACP

[CS02b] A. Czumaj and C. Sohler. Abstract combinatorial programs and efficient property testers. In IEEE [IEE02], pages 83–92. CODEN ASFPDV. ISBN 0-7695-1822-2. ISSN 0272-5428. LCCN QA267. URL http:/

/ieeexplore.ieee.org/iel5/
8411/26517/01181885.pdf?isnumber=

26517&prod=CNF&arnumber=1181885&
arSt=+83&ared=+92&arAuthor=
Czumaj%2C+A.%3B+Sohler%2C+
C.;   http://ieeexplore.ieee.
org/xpls/abs_all.jsp?isnumber=
26517&arnumber=1181885&count=
82&index=10.  IEEE Com-

puter Society Order Number PR01822.

Chien:2004:API

[CS04] S. Chien and A. Sinclair. Algebras with polynomial identities and computing the determinant. In IEEE [IEE04], pages 352–

361. CODEN ASFPDV. ISBN 0-7695-2228-9. ISSN 0272-5428. LCCN QA276. URL http:/

/ieeexplore.ieee.org/iel5/
9430/29918/01366255.pdf?isnumber=
29918&prod=CNF&arnumber=1366255&
arSt=+352&ared=+361&arAuthor=
Chien%2C+S.%3B+Sinclair%2C+
A.;   http://ieeexplore.ieee.
org/xpls/abs_all.jsp?isnumber=
29918&arnumber=1366255&count=

64&index=36. IEEE Computer Society Order Number P2228.

Czumaj:2007:TEB

[CS07] A. Czumaj and C. Sohler.

Testing expansion in boundeddegree graphs. In IEEE [IEE07], pages 570–578. ISBN 0-7695-3010-9. ISSN 0272-5428. LCCN QA76 .S974 2007. URL http://

ieeexplore.ieee.org/servlet/
opac?punumber=4389466. IEEE

Computer Society order number P3010.

Cheung:2008:AAS

[CS08] M. Cheung and C. Swamy.

Approximation algorithms for single-minded envy-free profitmaximization problems with limited supply. In IEEE [IEE08], pages 35–44. ISBN 0-7695-3436-8. ISSN 0272-5428. LCCN QA76 .S95 2008. URL http://

                                        ieeexplore.ieee.org/servlet/
                                        opac?punumber=4690923. IEEE

Computer Society order number P3436.

Childs:2007:QAH

[CSV07] A. M. Childs, L. J. Schulman, and U. V. Vazirani. Quantum algorithms for hidden nonlinear structures. In IEEE [IEE07], pages 395–404. ISBN 0-7695-3010-9. ISSN 0272-5428. LCCN QA76 .S974 2007. URL http://

                                        ieeexplore.ieee.org/servlet/
                                        opac?punumber=4389466. IEEE

Computer Society order number P3010.

Chakrabarti:2001:ICD

[CSWY01] A. Chakrabarti, Yaoyun Shi, A. Wirth, and A. Yao. Informational complexity and the direct sum problem for simultaneous message complexity. In IEEE [IEE01], pages 270–278. CO-DEN ASFPDV. ISBN 0-7695-1390-5, 0-7695-1391-3 (case), 0-7695-1392-1 (microfiche). ISSN 0272-5428. LCCN ????

Cohn:2003:GTA

[CU03] H. Cohn and C. Umans. A group-theoretic approach to fast matrix multiplication. In IEEE [IEE03], pages 438–449. CO-DEN ASFPDV. ISBN 0-7695-2040-5. ISSN 0272-5428. LCCN QA76 .S979 2003. URL http:/

/ieeexplore.ieee.org/iel5/
8767/27770/01238217.pdf?isnumber=
27770&prod=CNF&arnumber=1238217&
arSt=+438&ared=+449&arAuthor=
Cohn%2C+H.%3B+Umans%2C+C.;
 http://ieeexplore.ieee.org/
xpls/abs_all.jsp?isnumber=
27770&arnumber=1238217&count=
66&index=44.  IEEE Com-

puter Society Order Number PR02040.

Cleve:2000:FPC

[CW00] R. Cleve and J. Watrous. Fast parallel circuits for the quantum Fourier transform. In IEEE [IEE00], pages 526–536. CO-DEN ASFPDV. ISBN 0-7695-0850-2, 0-7695-0851-0 (case), 0-7695-0852-9 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 2000. IEEE Computer Society Order Number PR00850.

Charikar:2004:MQP

[CW04a] M. Charikar and A. Wirth.

Maximizing quadratic programs: extending Grothendieck’s inequality. In IEEE [IEE04], pages 54–60. CODEN ASF-PDV. ISBN 0-7695-2228-9. ISSN 0272-5428. LCCN QA276.

URL http://ieeexplore.ieee.
org/iel5/9430/29918/01366224.
pdf?isnumber=29918&prod=CNF&
arnumber=1366224&arSt=+54&
ared=+60&arAuthor=Charikar%
2C+M.%3B+Wirth%2C+A.;   http:
//ieeexplore.ieee.org/xpls/
abs_all.jsp?isnumber=29918&
arnumber=1366224&count=64&

index=5. IEEE Computer Society Order Number P2228.

Cheng:2004:LBD

[CW04b] Qi Cheng and Daqing Wan. On the list and bounded distance decodibility of Reed– Solomon codes. In IEEE [IEE04], pages 335–341. CO-DEN ASFPDV. ISBN 0-7695-2228-9. ISSN 0272-5428. LCCN QA276. URL http:/

                                        /ieeexplore.ieee.org/iel5/
                                        9430/29918/01366253.pdf?isnumber=
                                        29918&prod=CNF&arnumber=1366253&
                                        arSt=+335&ared=+341&arAuthor=
                                        Qi+Cheng%3B+Daqing+Wan;
                                        http://ieeexplore.ieee.org/
                                        xpls/abs_all.jsp?isnumber=
                                        29918&arnumber=1366253&count=

64&index=34. IEEE Computer Society Order Number P2228.

Chattopadhyay:2009:LSC

[CW09] A. Chattopadhyay and A. Wigderson. Linear systems over composite moduli. In IEEE [IEE09], pages 43–52. ISBN 0-7695-3850-

9. ISSN 0272-5428. LCCN QA76 .S95 2009. IEEE Computer Society order number P3850.

Dobzinski:2009:PRA

[DD09] S. Dobzinski and S. Dughmi. On the power of randomization in algorithmic mechanism design. In IEEE [IEE09], pages 505–514. ISBN 0-7695-3850-9. ISSN 0272-5428. LCCN QA76 .S95 2009. IEEE Computer Society order number P3850.

Dhangwatnotai:2008:TAS

[DDDR08] P. Dhangwatnotai, S. Dobzinski, S. Dughmi, and T. Roughgarden. Truthful approximation schemes for single-parameter agents. In IEEE [IEE08], pages 15–24. ISBN 0-7695-3436-8. ISSN 0272-5428. LCCN QA76 .S95 2008. URL http://

ieeexplore.ieee.org/servlet/
opac?punumber=4690923. IEEE

Computer Society order number P3436.

Dinitz:2008:SLL

[DES08] Y. Dinitz, M. Elkin, and S. Solomon. Shallow-low-light trees, and tight lower bounds for Euclidean spanners. In IEEE [IEE08], pages 519–528. ISBN 0-7695-3436-8. ISSN 0272-5428. LCCN QA76 .S95 2008. URL http://

ieeexplore.ieee.org/servlet/
opac?punumber=4690923. IEEE

Computer Society order number P3436.

Dyer:2001:RCG

[DF01] M. Dyer and A. Frieze. Randomly colouring graphs with lower bounds on girth and maximum degree. In IEEE [IEE01], pages 579–587. CO-DEN ASFPDV. ISBN 0-7695-1390-5, 0-7695-1391-3 (case), 0-7695-1392-1 (microfiche). ISSN 0272-5428. LCCN ????

Dyer:2004:RCC

[DFHV04] M. Dyer, A. Frieze, T. P. Hayes, and E. Vigoda. Randomly coloring constant degree graphs. In IEEE [IEE04], pages 582–589. CODEN ASFPDV. ISBN 0-7695-2228-9. ISSN 0272-5428. LCCN QA276. URL http:/

/ieeexplore.ieee.org/iel5/

                                        9430/29918/01366278.pdf?isnumber=
                                        29918&prod=CNF&arnumber=1366278&
                                        arSt=+582&ared=+589&arAuthor=
                                        Dyer%2C+M.%3B+Frieze%2C+A.
                                        %3B+Hayes%2C+T.P.%3B+Vigoda%
                                        2C+E.;   http://ieeexplore.
                                        ieee.org/xpls/abs_all.jsp?
                                        isnumber=29918&arnumber=1366278&
                                        count=64&index=59.  IEEE

Computer Society Order Num-

ber P2228.

Damgard:2005:CBQ

[DFSS05] I. B. Damgard, S. Fehr, L. Salvail, and C. Schaffner. Cryptography in the bounded quantumstorage model. In IEEE [IEE05], pages 449–458. ISBN 0-7695-2468-0, 0-7695-2362-5. ISSN 0272-5428. LCCN QA76 .S979 2005. IEEE Computer Society order number P2468.

Dinur:2008:LTD

[DG08] I. Dinur and E. Goldenberg. Locally testing direct product in the low error range. In IEEE [IEE08], pages 613–

622. ISBN 0-7695-3436-8. ISSN 0272-5428. LCCN QA76 .S95 2008. URL http://

                                        ieeexplore.ieee.org/servlet/
                                        opac?punumber=4690923. IEEE

Computer Society order number

P3436.

Diakonikolas:2009:BIF

[DGJ^{+}09] I. Diakonikolas, P. Gopalan, R. Jaiswal, R. A. Servedio, and E. Viola. Bounded independence fools halfspaces. In IEEE [IEE09], pages 171–180. ISBN 0-7695-3850-9. ISSN 0272-5428. LCCN QA76 .S95 2009. IEEE Computer Society order number P3850.

Dhamdhere:2005:HPC

[DGRS05] K. Dhamdhere, V. Goyal, R. Ravi, and M. Singh. How to pay, come what may: approximation algorithms for demandrobust covering problems. In IEEE [IEE05], pages 367–376. ISBN 0-7695-2468-0, 0-7695-2362-5. ISSN 0272-5428. LCCN QA76 .S979 2005. IEEE Computer Society order number P2468.

Deng:2009:RSR

[DGS09] Yi Deng, V. Goyal, and A. Sahai. Resolving the simultaneous resettability conjecture and a new non-black-box simulation strategy. In IEEE [IEE09], pages 251–260. ISBN 0-7695-3850-9. ISSN 0272-5428. LCCN QA76 .S95 2009. IEEE Computer Society order number P3850.

Dean:2004:ASK

[DGV04] B. C. Dean, M. X. Goemans, and J. Vondrdk. Approximating the stochastic knapsack problem: the benefit of adaptivity. In IEEE [IEE04], pages 208–

217. CODEN ASFPDV. ISBN 0-7695-2228-9. ISSN 0272-5428. LCCN QA276. URL http:/

/ieeexplore.ieee.org/iel5/
9430/29918/01366240.pdf?isnumber=
29918&prod=CNF&arnumber=1366240&
arSt=+208&ared=+217&arAuthor=
Dean%2C+B.C.%3B+Goemans%2C+
M.X.%3B+Vondrdk%2C+J.;   http:
//ieeexplore.ieee.org/xpls/
abs_all.jsp?isnumber=29918&

                                        arnumber=1366240&count=64&

index=21. IEEE Computer Society Order Number P2228.

Dvir:2007:ERE

[DGW07] Z. Dvir, A. Gabizon, and A. Wigderson. Extractors and rank extractors for polynomial sources. In IEEE [IEE07], pages 52–62. ISBN 0-7695-3010-9. ISSN 0272-5428. LCCN QA76 .S974 2007. URL http://

                                        ieeexplore.ieee.org/servlet/
                                        opac?punumber=4389466. IEEE

Computer Society order number

P3010.

Dinur:2009:CLE

[DH09] I. Dinur and P. Harsha. Composition of low-error 2-query PCPs using decodable PCPs. In IEEE [IEE09], pages 472–481. ISBN 0-7695-3850-9. ISSN 0272-5428. LCCN QA76 .S95 2009. IEEE Computer Society order number P3850.

Demaine:2004:DOA

[DHIP04] E. D. Demaine, D. Harmon, J. Iacono, and M. Patrascu. Dynamic optimality — almost [competitive online binary search tree]. In IEEE [IEE04], pages 484–490. CO-DEN ASFPDV. ISBN 0-7695-2228-9. ISSN 0272-5428. LCCN QA276. URL http:/

                                        /ieeexplore.ieee.org/iel5/
                                        9430/29918/01366268.pdf?isnumber=
                                        29918&prod=CNF&arnumber=1366268&
                                        arSt=+484&ared=+490&arAuthor=
                                        Demaine%2C+E.D.%3B+Harmon%
                                        2C+D.%3B+Iacono%2C+J.%3B+
                                        Patrascu%2C+M.;   http://

ieeexplore.ieee.org/xpls/
abs_all.jsp?isnumber=29918&
arnumber=1366268&count=64&

index=49. IEEE Computer Society Order Number P2228.

Demaine:2005:AGM

[DHK05] E. D. Demaine, M. T. Hajiaghayi, and K. Kawarabayashi. Algorithmic graph minor theory: Decomposition, approximation, and coloring. In IEEE [IEE05], pages 637–646. ISBN 0-7695-2468-0, 0-7695-2362-5. ISSN 0272-5428. LCCN QA76 .S979 2005. IEEE Computer Society order number P2468.

Dasgupta:2005:LMH

[DHKS05] A. Dasgupta, J. Hopcroft, J. Kleinberg, and M. Sandler. On learning mixtures of heavytailed distributions. In IEEE [IEE05], pages 491–500. ISBN 0-7695-2468-0, 0-7695-2362-5. ISSN 0272-5428. LCCN QA76 .S979 2005. IEEE Computer Society order number P2468.

Dasgupta:2004:SAR

[DHM04] A. Dasgupta, J. E. Hopcroft, and F. McSherry. Spectral analysis of random graphs with skewed degree distributions. In IEEE [IEE04], pages 602–610. CODEN ASFPDV. ISBN 0-7695-2228-9. ISSN 0272-5428. LCCN QA276. URL http:/

/ieeexplore.ieee.org/iel5/
9430/29918/01366280.pdf?isnumber=
29918&prod=CNF&arnumber=1366280&
arSt=+602&ared=+610&arAuthor=
Dasgupta%2C+A.%3B+Hopcroft%

                                        2C+J.E.%3B+McSherry%2C+F.;
                                         http://ieeexplore.ieee.org/
                                        xpls/abs_all.jsp?isnumber=
                                        29918&arnumber=1366280&count=

64&index=61. IEEE Computer Society Order Number P2228.

Demetrescu:2000:FDT

[DI00] C. Demetrescu and G. F. Italiano. Fully dynamic transitive closure: breaking through the O(n^{2}) barrier. In IEEE [IEE00], pages 381–389. CO-DEN ASFPDV. ISBN 0-7695-0850-2, 0-7695-0851-0 (case), 0-7695-0852-9 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 2000. IEEE Computer Society Order Number PR00850.

Demetrescu:2001:FDA

[DI01] C. Demetrescu and G. F. Italiano. Fully dynamic all pairs shortest paths with real edge weights. In IEEE [IEE01], pages 260–267. CODEN ASF-PDV. ISBN 0-7695-1390-5, 0-7695-1391-3 (case), 0-7695-1392-1 (microfiche). ISSN 0272-5428. LCCN ????

Devanur:2008:MEP

[DK08] N. R. Devanur and R. Kannan. Market equilibria in polynomial time for fixed number of goods or agents. In IEEE [IEE08], pages 45–53. ISBN 0-7695-3436-8. ISSN 0272-5428. LCCN QA76 .S95 2008. URL http://

                                        ieeexplore.ieee.org/servlet/
                                        opac?punumber=4690923. IEEE

Computer Society order number P3436.

Dvir:2009:EMM

[DKSS09] Z. Dvir, S. Kopparty, S. Saraf, and M. Sudan. Extensions to the method of multiplicities, with applications to Kakeya sets and mergers. In IEEE [IEE09], pages 181–190. ISBN 0-7695-3850-9. ISSN 0272-5428. LCCN QA76 .S95 2009. IEEE Computer Society order number P3850.

Diakonikolas:2007:TCR

[DLM^{+}07] I. Diakonikolas, H. K. Lee, K. Matule, K. Onak, R. Rubinfeld, R. A. Servedio, and A. Wan. Testing for concise representations. In IEEE [IEE07], pages 549–558. ISBN 0-7695-3010-9. ISSN 0272-5428. LCCN QA76 .S974 2007. URL http://

ieeexplore.ieee.org/servlet/
opac?punumber=4389466. IEEE

Computer Society order number P3010.

Dobzinski:2008:MUA

[DLN08] S. Dobzinski, R. Lavi, and N. Nisan. Multi-unit auctions with budget limits. In IEEE [IEE08], pages 260– 269. ISBN 0-7695-3436-8. ISSN 0272-5428. LCCN QA76 .S95 2008. URL http://

ieeexplore.ieee.org/servlet/
opac?punumber=4690923. IEEE

Computer Society order number P3436.

Dubois:2002:XT

[DM02] O. Dubois and J. Mandler.

The 3-XORSAT threshold. In IEEE [IEE02], pages 769–778.

CODEN ASFPDV. ISBN 0-7695-1822-2. ISSN 0272-5428. LCCN QA267. URL http:/

/ieeexplore.ieee.org/iel5/
8411/26517/01182002.pdf?isnumber=
26517&prod=CNF&arnumber=1182002&
arSt=+769&ared=+778&arAuthor=
Dubois%2C+O.%3B+Mandler%2C+
J.;   http://ieeexplore.ieee.
org/xpls/abs_all.jsp?isnumber=
26517&arnumber=1182002&count=
82&index=76.  IEEE Com-

puter Society Order Number PR01822.

Dantchev:2007:PPC

[DMS07] S. Dantchev, B. Martin, and S. Szeider. Parameterized proof complexity. In IEEE [IEE07], pages 150–160. ISBN 0-7695-3010-9. ISSN 0272-5428. LCCN QA76 .S974 2007. URL http://

ieeexplore.ieee.org/servlet/
opac?punumber=4389466. IEEE

Computer Society order number P3010.

Dwork:2000:ZTA

[DN00] C. Dwork and M. Naor. Zaps and their applications. In IEEE [IEE00], pages 283–293. CO-DEN ASFPDV. ISBN 0-7695-0850-2, 0-7695-0851-0 (case), 0-7695-0852-9 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 2000. IEEE Computer Society Order Number PR00850.

Dodis:2004:IPC

[DOPS04] Y. Dodis, Shien Jin Ong, M. Prabhakaran, and A. Sahai. On the (im)possibility of cryptography with imperfect randomness. In IEEE [IEE04], pages 196–205. CO-DEN ASFPDV. ISBN 0-7695-2228-9. ISSN 0272-5428. LCCN QA276. URL http:/

/ieeexplore.ieee.org/iel5/
9430/29918/01366239.pdf?isnumber=
29918&prod=CNF&arnumber=1366239&
arSt=+196&ared=+205&arAuthor=
Dodis%2C+Y.%3B+Shien+Jin+
Ong%3B+Prabhakaran%2C+M.%
3B+Sahai%2C+A.;   http://
ieeexplore.ieee.org/xpls/
abs_all.jsp?isnumber=29918&
arnumber=1366239&count=64&

index=20. IEEE Computer Society Order Number P2228.

Doty:2009:RSA

[Dot09] D. Doty. Randomized selfassembly for exact shapes. In IEEE [IEE09], pages 85–94. ISBN 0-7695-3850-9. ISSN 0272-5428. LCCN QA76 .S95 2009. IEEE Computer Society order number P3850.

Daskalakis:2007:CEA

[DP07a] C. Daskalakis and C. Papadimitriou. Computing equilibria in anonymous games. In IEEE [IEE07], pages 83–93. ISBN 0-7695-3010-9. ISSN 0272-5428. LCCN QA76 .S974 2007. URL http://

ieeexplore.ieee.org/servlet/
opac?punumber=4389466. IEEE

Computer Society order number P3010.

Dziembowski:2007:IRS

[DP07b] S. Dziembowski and K. Pietrzak.

Intrusion-resilient secret sharing. In IEEE [IEE07], pages 227–237. ISBN 0-7695-3010-9. ISSN 0272-5428. LCCN QA76 .S974 2007. URL http://

ieeexplore.ieee.org/servlet/
opac?punumber=4389466. IEEE

Computer Society order number P3010.

Daskalakis:2008:DMD

[DP08a] C. Daskalakis and C. H. Papadimitriou. Discretized multinomial distributions and Nash equilibria in anonymous games. In IEEE [IEE08], pages 25– 34. ISBN 0-7695-3436-8. ISSN 0272-5428. LCCN QA76 .S95 2008. URL http://

ieeexplore.ieee.org/servlet/
opac?punumber=4690923. IEEE

Computer Society order number P3436.

Dziembowski:2008:LRC

[DP08b] S. Dziembowski and K. Pietrzak. Leakage-resilient cryptography. In IEEE [IEE08], pages 293– 302. ISBN 0-7695-3436-8. ISSN 0272-5428. LCCN QA76 .S95 2008. URL http://

ieeexplore.ieee.org/servlet/
opac?punumber=4690923. IEEE

Computer Society order number P3436.

Devanur:2002:MEP

[DPSV02] N. R. Devanur, C. H. Papadimitriou, A. Saberi, and V. V. Vazirani. Market equilibrium via a primal-dual-type algorithm. In IEEE [IEE02], pages 389–395. CO-DEN ASFPDV. ISBN 0-7695-1822-2. ISSN 0272-5428. LCCN QA267. URL http:/

/ieeexplore.ieee.org/iel5/
8411/26517/01181963.pdf?isnumber=
26517&prod=CNF&arnumber=1181963&
arSt=+389&ared=+395&arAuthor=
Devanur%2C+N.R.%3B+Papadimitriou%
2C+C.H.%3B+Saberi%2C+A.%3B+
Vazirani%2C+V.V.;   http:/
/ieeexplore.ieee.org/xpls/
abs_all.jsp?isnumber=26517&
arnumber=1181963&count=82&

index=39. IEEE Computer Society Order Number PR01822.

Dantchev:2001:STH

[DR01] S. Dantchev and S. Riis. “planar” tautologies hard for resolution. In IEEE [IEE01], pages 220–229. CODEN ASF-PDV. ISBN 0-7695-1390-5, 0-7695-1391-3 (case), 0-7695-1392-1 (microfiche). ISSN 0272-5428. LCCN ????

Dinur:2004:ATT

[DR04] I. Dinur and O. Reingold.

Assignment testers: towards a combinatorial proof of the PCP-theorem. In IEEE [IEE04], pages 155–164. CO-DEN ASFPDV. ISBN 0-7695-2228-9. ISSN 0272-5428. LCCN QA276. URL http:/

/ieeexplore.ieee.org/iel5/
9430/29918/01366235.pdf?isnumber=
29918&prod=CNF&arnumber=1366235&
arSt=+155&ared=+164&arAuthor=
Dinur%2C+I.%3B+Reingold%2C+
O.;   http://ieeexplore.ieee.
org/xpls/abs_all.jsp?isnumber=
29918&arnumber=1366235&count=

64&index=16. IEEE Computer Society Order Number P2228.

Dutta:2008:LBN

[DR08] Chinmoy Dutta and J. Radhakrishnan. Lower bounds for noisy wireless networks using sampling algorithms. In IEEE [IEE08], pages 394–

402. ISBN 0-7695-3436-8. ISSN 0272-5428. LCCN QA76 .S95 2008. URL http://

                                        ieeexplore.ieee.org/servlet/
                                        opac?punumber=4690923. IEEE

Computer Society order number

P3436.

Dinur:2002:HUH

[DRS02] I. Dinur, O. Regev, and C. Smyth. The hardness of 3-uniform hypergraph coloring. In IEEE [IEE02], pages 33–40. CODEN ASFPDV. ISBN 0-7695-1822-2. ISSN 0272-5428. LCCN QA267. URL http:/

                                        /ieeexplore.ieee.org/iel5/
                                        8411/26517/01181880.pdf?isnumber=
                                        26517&prod=CNF&arnumber=1181880&
                                        arSt=+33&ared=+40&arAuthor=
                                        Dinur%2C+I.%3B+Regev%2C+O.
                                        %3B+Smyth%2C+C.;   http://
                                        ieeexplore.ieee.org/xpls/
                                        abs_all.jsp?isnumber=26517&
                                        arnumber=1181880&count=82&

index=5. IEEE Computer Society Order Number PR01822.

Dodis:2002:NUO

[DS02] Y. Dodis and J. Spencer. On the (non)universality of the one-time pad. In IEEE [IEE02], pages 376–385. CO-DEN ASFPDV. ISBN 0-7695-1822-2. ISSN 0272-5428. LCCN QA267. URL http:/

                                        /ieeexplore.ieee.org/iel5/

8411/26517/01181962.pdf?isnumber=
26517&prod=CNF&arnumber=1181962&
arSt=+376&ared=+385&arAuthor=
Dodis%2C+Y.%3B+Spencer%2C+
J.;   http://ieeexplore.ieee.
org/xpls/abs_all.jsp?isnumber=
26517&arnumber=1181962&count=
82&index=38.  IEEE Com-

puter Society Order Number PR01822.

Deshpande:2005:ISA

[DS05] A. Deshpande and D. A. Spielman. Improved smoothed analysis of the shadow vertex simplex method. In IEEE [IEE05], pages 349–356. ISBN 0-7695-2468-0, 0-7695-2362-5. ISSN 0272-5428. LCCN QA76 .S979 2005. IEEE Computer Society order number P2468.

deVerdiere:2002:OSL

[dVL02] E. C. de Verdiere and F. Lazarus.

Optimal system of loops on an orientable surface. In IEEE [IEE02], pages 627–636. CO-DEN ASFPDV. ISBN 0-7695-1822-2. ISSN 0272-5428. LCCN QA267. URL http:/

/ieeexplore.ieee.org/iel5/
8411/26517/01181986.pdf?isnumber=
26517&prod=CNF&arnumber=1181986&
arSt=+627&ared=+636&arAuthor=
de+Verdiere%2C+E.C.%3B+Lazarus%
2C+F.;   http://ieeexplore.
ieee.org/xpls/abs_all.jsp?
isnumber=26517&arnumber=1181986&
count=82&index=62.  IEEE

Computer Society Order Number PR01822.

Dvir:2008:KSN

[DW08] Z. Dvir and A. Wigderson.

Kakeya sets, new mergers and old extractors. In IEEE [IEE08], pages 625–633. ISBN 0-7695-3436-8. ISSN 0272-5428. LCCN QA76 .S95 2008. URL http://

ieeexplore.ieee.org/servlet/
opac?punumber=4690923. IEEE

Computer Society order number P3436.

Even:2002:CFC

[ELRS02] G. Even, Z. Lotker, D. Ron, and S. Smorodinsky. Conflictfree colorings of simple geometric regions with applications to frequency assignment in cellular networks. In IEEE [IEE02], pages 691–700. CO-DEN ASFPDV. ISBN 0-7695-1822-2. ISSN 0272-5428. LCCN QA267. URL http:/

/ieeexplore.ieee.org/iel5/
8411/26517/01181994.pdf?isnumber=
26517&prod=CNF&arnumber=1181994&
arSt=+691&ared=+700&arAuthor=
Even%2C+G.%3B+Lotker%2C+Z.
%3B+Ron%2C+D.%3B+Smorodinsky%
2C+S.;   http://ieeexplore.
ieee.org/xpls/abs_all.jsp?
isnumber=26517&arnumber=1181994&
count=82&index=68.  IEEE

Computer Society Order Number PR01822.

Edelsbrunner:2000:TPS

[ELZ00] H. Edelsbrunner, D. Letscher, and A. Zomorodian. Topological persistence and simplification. In IEEE [IEE00], pages 454–463. CODEN ASF-PDV. ISBN 0-7695-0850-2, 0-7695-0851-0 (case), 0-7695-0852-9 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 2000. IEEE Computer Society Order Number PR00850.

Englert:2008:PRO

[EOW08] M. Englert, D. Ozmen, and M. Westermann. The power of reordering for online minimum makespan scheduling. In IEEE [IEE08], pages 603– 612. ISBN 0-7695-3436-8. ISSN 0272-5428. LCCN QA76 .S95 2008. URL http://

ieeexplore.ieee.org/servlet/
opac?punumber=4690923. IEEE

Computer Society order number P3436.

Edmonds:2006:BAC

[EP06] Jeff Edmonds and Kirk Pruhs.

Balanced allocations of cake. In IEEE [IEE06], pages 623–634. ISBN 0-7695-2720-5, 0-7695-2362-5. ISSN 0272-5428. LCCN QA76 .S974 2006. URL http://

ieeexplore.ieee.org/servlet/
opac?punumber=4031329. IEEE

Computer Society Order Number P2720.

Englert:2009:ORL

[ER09] M. Englert and H. Racke. Oblivious routing for the L_{p}-norm. In IEEE [IEE09], pages 32–40. ISBN 0-7695-3850-9. ISSN 0272-5428. LCCN QA76 .S95 2009. IEEE Computer Society order number P3850.

Ezra:2007:ATB

[ES07] E. Ezra and M. E. Sharir. Almost tight bound for the union of fat tetrahedra in three dimensions. In IEEE [IEE07], pages 525–535. ISBN 0-7695-3010-9. ISSN 0272-5428. LCCN QA76 .S974 2007. URL http://

ieeexplore.ieee.org/servlet/
opac?punumber=4389466. IEEE

Computer Society order number P3010.

Etessami:2007:CNE

[EY07] K. Etessami and M. Yannakakis. On the complexity of Nash equilibria and other fixed points (extended abstract). In IEEE [IEE07], pages 113– 123. ISBN 0-7695-3010-9. ISSN 0272-5428. LCCN QA76 .S974 2007. URL http://

ieeexplore.ieee.org/servlet/
opac?punumber=4389466. IEEE

Computer Society order number P3010.

Ezra:2008:UCT

[Ezr08] E. Ezra. On the union of cylinders in three dimensions. In IEEE [IEE08], pages 179– 188. ISBN 0-7695-3436-8. ISSN 0272-5428. LCCN QA76 .S95 2008. URL http://

ieeexplore.ieee.org/servlet/
opac?punumber=4690923. IEEE

Computer Society order number P3436.

Feige:2007:RSF

[Fei07] U. Feige. Refuting smoothed 3CNF formulas. In IEEE [IEE07], pages 407–417. ISBN 0-7695-3010-9. ISSN 0272-5428. LCCN QA76 .S974 2007. URL

http://ieeexplore.ieee.org/
servlet/opac?punumber=4389466.

IEEE Computer Society order number P3010.

Feldman:2009:CCS

[Fel09] V. Feldman. A complete characterization of statistical query learning with applications to evolvability. In IEEE [IEE09], pages 375–384. ISBN 0-7695-3850-9. ISSN 0272-5428. LCCN QA76 .S95 2009. IEEE Computer Society order number P3850.

Feldman:2006:CWF

[FFS06] Dan Feldman, Amos Fiat, and Micha Sharir. Coresets for weighted facilities and their applications. In IEEE [IEE06], pages 315–324. ISBN 0-7695-2720-5, 0-7695-2362-5. ISSN 0272-5428. LCCN QA76 .S974 2006. URL http://

ieeexplore.ieee.org/servlet/
opac?punumber=4031329. IEEE

Computer Society Order Number P2720.

Flum:2002:PCC

[FG02] J. Flum and M. Grohe. The parameterized complexity of counting problems. In IEEE [IEE02], pages 538–547. CO-DEN ASFPDV. ISBN 0-7695-1822-2. ISSN 0272-5428. LCCN QA267. URL http:/

/ieeexplore.ieee.org/iel5/
8411/26517/01181978.pdf?isnumber=
26517&prod=CNF&arnumber=1181978&
arSt=+538&ared=+547&arAuthor=
Flum%2C+J.%3B+Grohe%2C+M.;
http://ieeexplore.ieee.org/
xpls/abs_all.jsp?isnumber=
26517&arnumber=1181978&count=
82&index=54.  IEEE Com-

puter Society Order Number PR01822.

Franceschini:2003:PSC

[FG03] G. Franceschini and V. Geffert. An in-place sorting with O(n log n) comparisons and O(n) moves. In IEEE [IEE03], pages 242–250. CO-DEN ASFPDV. ISBN 0-7695-2040-5. ISSN 0272-5428. LCCN QA76 .S979 2003. URL http:/

/ieeexplore.ieee.org/iel5/
8767/27770/01238198.pdf?isnumber=
27770&prod=CNF&arnumber=1238198&
arSt=+242&ared=+250&arAuthor=
Franceschini%2C+G.%3B+Geffert%
2C+V.;   http://ieeexplore.
ieee.org/xpls/abs_all.jsp?
isnumber=27770&arnumber=1238198&
count=66&index=25.  IEEE

Computer Society Order Number PR02040.

Franceschini:2004:NSB

[FG04] G. Franceschini and R. Grossi. No sorting? Better searching! [optimal array organization]. In IEEE [IEE04], pages 491–498. CODEN ASFPDV. ISBN 0-7695-2228-9. ISSN 0272-5428. LCCN QA276. URL http:/

/ieeexplore.ieee.org/iel5/
9430/29918/01366269.pdf?isnumber=
29918&prod=CNF&arnumber=1366269&
arSt=+491&ared=+498&arAuthor=
Franceschini%2C+G.%3B+Grossi%
2C+R.;   http://ieeexplore.
ieee.org/xpls/abs_all.jsp?
isnumber=29918&arnumber=1366269&
count=64&index=50.  IEEE

Computer Society Order Number P2228.

Feldman:2006:NRL

[FGKP06] Vitaly Feldman, Parikshit Gopalan, Subhash Khot, and Ashok Kumar Ponnuswami. New results for learning noisy parities and halfspaces. In IEEE [IEE06], pages 563–574. ISBN 0-7695-2720-5, 0-7695-2362-5. ISSN 0272-5428. LCCN QA76 .S974 2006. URL http://

ieeexplore.ieee.org/servlet/
opac?punumber=4031329. IEEE

Computer Society Order Number P2720.

Franceschini:2002:IBT

[FGMP02] G. Franceschini, R. Grossi, J. I.

Munro, and L. Pagli. Implicit B-trees: New results for the dictionary problem. In IEEE [IEE02], pages 145–154. CODEN ASFPDV. ISBN 0-7695-1822-2. ISSN 0272-5428. LCCN QA267. URL http:/

/ieeexplore.ieee.org/iel5/
8411/26517/01181891.pdf?isnumber=
26517&prod=CNF&arnumber=1181891&
arSt=+145&ared=+154&arAuthor=
Franceschini%2C+G.%3B+Grossi%
2C+R.%3B+Munro%2C+J.I.%3B+
Pagli%2C+L.;   http://ieeexplore.
ieee.org/xpls/abs_all.jsp?
isnumber=26517&arnumber=1181891&
count=82&index=16.  IEEE

Computer Society Order Number PR01822.

Feder:2006:LSM

[FGMS06] Tomas Feder, Adam Guetz, Milena Mihail, and Amin Saberi. A local switch Markov chain on given degree graphs with application in connectivity of peer-to-peer networks. In IEEE [IEE06], pages 69– 76. ISBN 0-7695-2720-5, 0-7695-2362-5. ISSN 0272-5428. LCCN QA76 .S974 2006. URL http://

                                        ieeexplore.ieee.org/servlet/
                                        opac?punumber=4031329. IEEE

Computer Society Order Number P2720.

Feldman:2009:ALM

[FGRW09] V. Feldman, V. Guruswami, P. Raghavendra, and Yi Wu. Agnostic learning of monomials by halfspaces is hard. In IEEE [IEE09], pages 385–394. ISBN 0-7695-3850-9. ISSN 0272-5428. LCCN QA76 .S95 2009. IEEE Computer Society order number P3850.

Fill:2000:RRN

[FH00] J. A. Fill and M. Hubert. The randomness recycler: a new technique for perfect sampling. In IEEE [IEE00], pages 503–511. CODEN ASFPDV. ISBN 0-7695-0850-2, 0-7695-0851-0 (case), 0-7695-0852-9 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 2000. IEEE Computer Society Order Number PR00850.

Fich:2005:LLB

[FHS05] F. E. Fich, D. Hendler, and N. Shavit. Linear lower bounds on real-world implementations of concurrent objects. In IEEE [IEE05], pages 165–173. ISBN 0-7695-2468-0, 0-7695-2362-5. ISSN 0272-5428. LCCN QA76 .S979 2005. IEEE Computer Society order number P2468.

Fleischer:2004:THS

[FJM04] L. Fleischer, K. Jain, and M. Mahdian. Tolls for heterogeneous selfish users in multicommodity networks and generalized congestion games. In IEEE [IEE04], pages 277–285. CODEN ASFPDV. ISBN 0-7695-2228-9. ISSN 0272-5428. LCCN QA276. URL http:/

/ieeexplore.ieee.org/iel5/
9430/29918/01366247.pdf?isnumber=
29918&prod=CNF&arnumber=1366247&
arSt=+277&ared=+285&arAuthor=
Fleischer%2C+L.%3B+Jain%2C+
K.%3B+Mahdian%2C+M.;   http:
//ieeexplore.ieee.org/xpls/
abs_all.jsp?isnumber=29918&
arnumber=1366247&count=64&

index=28. IEEE Computer Society Order Number P2228.

Fleischer:2001:IRA

[FJW01] L. Fleischer, K. Jain, and D. P. Williamson. An iterative rounding 2-approximation algorithm for the element connectivity problem. In IEEE [IEE01], pages 339–347. CO-DEN ASFPDV. ISBN 0-7695-1390-5, 0-7695-1391-3 (case), 0-7695-1392-1 (microfiche). ISSN 0272-5428. LCCN ????

Feige:2000:PAM

[FK00] U. Feige and R. Krauthgamer.

A polylogarithmic approximation of the minimum bisection. In IEEE [IEE00], pages 105–115. CODEN ASFPDV. ISBN 0-7695-0850-2, 0-7695-0851-0 (case), 0-7695-0852-9 (microfiche). ISSN 0272-5428.

LCCN TK7885.A1 S92 2000. IEEE Computer Society Order Number PR00850.

Feldman:2002:DTL

[FK02] J. Feldman and D. R. Karger. Decoding turbo-like codes via linear programming. In IEEE [IEE02], pages 251–260. CO-DEN ASFPDV. ISBN 0-7695-1822-2. ISSN 0272-5428. LCCN QA267. URL http:/

                                        /ieeexplore.ieee.org/iel5/
                                        8411/26517/01181948.pdf?isnumber=
                                        26517&prod=CNF&arnumber=1181948&
                                        arSt=+251&ared=+260&arAuthor=
                                        Feldman%2C+J.%3B+Karger%2C+
                                        D.R.;   http://ieeexplore.
                                        ieee.org/xpls/abs_all.jsp?
                                        isnumber=26517&arnumber=1181948&
                                        count=82&index=26.  IEEE

Computer Society Order Num-

ber PR01822.

Friedgut:2008:ECM

[FKN08] E. Friedgut, G. Kalai, and N. Nisan. Elections can be manipulated often. In IEEE [IEE08], pages 243–

249. ISBN 0-7695-3436-8. ISSN 0272-5428. LCCN QA76 .S95 2008. URL http://

                                        ieeexplore.ieee.org/servlet/
                                        opac?punumber=4690923. IEEE

Computer Society order number

P3436.

Feige:2006:WNS

[FKO06] Uriel Feige, Jeong Han Kim, and Eran Ofek. Witnesses for non-satisfiability of dense random 3CNF formulas. In IEEE [IEE06], pages 497–508. ISBN 0-7695-2720-5, 0-7695-2362-5.

ISSN 0272-5428. LCCN QA76 .S974 2006. URL http://

ieeexplore.ieee.org/servlet/
opac?punumber=4031329. IEEE

Computer Society Order Number P2720.

Fischer:2002:TJC

[FKR^{+}02] E. Fischer, G. Kindler, D. Ron, S. Safra, and A. Samorodnitsky. Testing juntas [combinatorial property testing]. In IEEE [IEE02], pages 103–112. CODEN ASFPDV. ISBN 0-7695-1822-2. ISSN 0272-5428. LCCN QA267. URL http:/

/ieeexplore.ieee.org/iel5/
8411/26517/01181887.pdf?isnumber=
26517&prod=CNF&arnumber=1181887&
arSt=+103&ared=+112&arAuthor=
Fischer%2C+E.%3B+Kindler%
2C+G.%3B+Ron%2C+D.%3B+Safra%
2C+S.%3B+Samorodnitsky%2C+
A.;   http://ieeexplore.ieee.
org/xpls/abs_all.jsp?isnumber=
26517&arnumber=1181887&count=
82&index=12.  IEEE Com-

puter Society Order Number PR01822.

Ferragina:2005:SLT

[FLMM05] P. Ferragina, F. Luccio, G. Manzini, and S. Muthukrishnan. Structuring labeled trees for optimal succinctness, and beyond. In IEEE [IEE05], pages 184–

193. ISBN 0-7695-2468-0, 0-7695-2362-5. ISSN 0272-5428. LCCN QA76 .S979 2005. IEEE Computer Society order number P2468.

Feige:2002:GTV

[FLS02] U. Feige, M. Langberg, and G. Schechtman. Graphs with tiny vector chromatic numbers and huge chromatic numbers. In IEEE [IEE02], pages 283–292. CODEN ASFPDV. ISBN 0-7695-1822-2. ISSN 0272-5428. LCCN QA267. URL http:/

/ieeexplore.ieee.org/iel5/
8411/26517/01181951.pdf?isnumber=
26517&prod=CNF&arnumber=1181951&
arSt=+283&ared=+292&arAuthor=
Feige%2C+U.%3B+Langberg%2C+
M.%3B+Schechtman%2C+G.;
http://ieeexplore.ieee.org/
xpls/abs_all.jsp?isnumber=
26517&arnumber=1181951&count=
82&index=29.  IEEE Com-

puter Society Order Number PR01822.

Ferragina:2000:ODS

[FM00] P. Ferragina and G. Manzini. Opportunistic data structures with applications. In IEEE [IEE00], pages 390–398. CO-DEN ASFPDV. ISBN 0-7695-0850-2, 0-7695-0851-0 (case), 0-7695-0852-9 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 2000. IEEE Computer Society Order Number PR00850.

Feldman:2009:OSM

[FMMM09] J. Feldman, A. Mehta, V. Mirrokni, and S. Muthukrishnan. Online stochastic matching: Beating 1 − 1/e. In IEEE [IEE09], pages 117–126. ISBN 0-7695-3850-9. ISSN 0272-5428. LCCN QA76 .S95 2009. IEEE Computer Society order number P3850.

Fischer:2007:AHP

[FMS07] E. Fischer, A. Matsliah, and A. Shapira. Approximate hypergraph partitioning and applications. In IEEE [IEE07], pages 579–589. ISBN 0-7695-3010-9. ISSN 0272-5428. LCCN QA76 .S974 2007. URL http://

ieeexplore.ieee.org/servlet/
opac?punumber=4389466. IEEE

Computer Society order number P3010.

Feigc:2007:MNM

[FMV07] U. Feigc, V. S. Mirrokni, and J. Vondrak. Maximizing nonmonotone submodular functions. In IEEE [IEE07], pages 461–471. ISBN 0-7695-3010-9. ISSN 0272-5428. LCCN QA76 .S974 2007. URL http://

ieeexplore.ieee.org/servlet/
opac?punumber=4389466. IEEE

Computer Society order number P3010.

Feldman:2005:LMP

[FOS05] J. Feldman, R. O’Donnell, and R. A. Servedio. Learning mixtures of product distributions over discrete domains. In IEEE [IEE05], pages 501– 510. ISBN 0-7695-2468-0, 0-7695-2362-5. ISSN 0272-5428. LCCN QA76 .S979 2005. IEEE Computer Society order number P2468.

Fakcharoenphol:2001:PGN

[FR01] J. Fakcharoenphol and S. Rao.

Planar graphs, negative weight edges, shortest paths, and near linear time. In IEEE [IEE01], pages 232–241. CODEN ASF-PDV. ISBN 0-7695-1390-5, 0-7695-1391-3 (case), 0-7695-1392-1 (microfiche). ISSN 0272-5428. LCCN ????

Frieze:2002:RST

[Fri02] A. Frieze. On random symmetric travelling salesman problems. In IEEE [IEE02], pages 789–798. CODEN ASFPDV. ISBN 0-7695-1822-2. ISSN 0272-5428. LCCN QA267. URL http:/

/ieeexplore.ieee.org/iel5/
8411/26517/01182004.pdf?isnumber=
26517&prod=CNF&arnumber=1182004&
arSt=+789&ared=+798&arAuthor=
Frieze%2C+A.;   http://ieeexplore.
ieee.org/xpls/abs_all.jsp?
isnumber=26517&arnumber=1182004&
count=82&index=78.  IEEE

Computer Society Order Number PR01822.

Fortnow:2004:HTP

[FS04] L. Fortnow and R. Santhanam. Hierarchy theorems for probabilistic polynomial time. In IEEE [IEE04], pages 316–324. CODEN ASFPDV. ISBN 0-7695-2228-9. ISSN 0272-5428. LCCN QA276. URL http:/

/ieeexplore.ieee.org/iel5/
9430/29918/01366251.pdf?isnumber=
29918&prod=CNF&arnumber=1366251&
arSt=+316&ared=+324&arAuthor=
Fortnow%2C+L.%3B+Santhanam%
2C+R.;   http://ieeexplore.
ieee.org/xpls/abs_all.jsp?
isnumber=29918&arnumber=1366251&
count=64&index=32.  IEEE

Computer Society Order Number P2228.

Friggstad:2008:MMM

[FS08] Z. Friggstad and M. R. Salavatipour.

Minimizing movement in mobile facility location problems. In IEEE [IEE08], pages 357– 366. ISBN 0-7695-3436-8. ISSN 0272-5428. LCCN QA76 .S95 2008. URL http://

ieeexplore.ieee.org/servlet/
opac?punumber=4690923. IEEE

Computer Society order number P3436.

Feige:2006:AAA

[FV06] Uriel Feige and Jan Vondrak.

Approximation algorithms for allocation problems: Improving the factor of 1 − 1/e. In IEEE [IEE06], pages 667–676. ISBN 0-7695-2720-5, 0-7695-2362-5. ISSN 0272-5428. LCCN QA76 .S974 2006. URL http://

ieeexplore.ieee.org/servlet/
opac?punumber=4031329. IEEE

Computer Society Order Number P2720.

Gabow:2000:UEG

[Gab00]  H. N. Gabow.  Using ex-

pander graphs to find vertex connectivity. In IEEE [IEE00], pages 410–420. CODEN ASF-PDV. ISBN 0-7695-0850-2, 0-7695-0851-0 (case), 0-7695-0852-9 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 2000. IEEE Computer Society Order Number PR00850.

Gafni:2002:SAC

[Gaf02] E. Gafni. A simple algorithmic characterization of uniform solvability. In IEEE [IEE02], pages 228–237. CO-DEN ASFPDV. ISBN 0-7695-1822-2. ISSN 0272-5428. LCCN QA267. URL http:/

/ieeexplore.ieee.org/iel5/
8411/26517/01181946.pdf?isnumber=
26517&prod=CNF&arnumber=1181946&
arSt=+228&ared=+237&arAuthor=
Gafni%2C+E.;   http://ieeexplore.
ieee.org/xpls/abs_all.jsp?
isnumber=26517&arnumber=1181946&
count=82&index=24.  IEEE

Computer Society Order Number PR01822.

Gal:2007:LBS

[GG07] A. Gal and P. Gopalan. Lower bounds on streaming algorithms for approximating the length of the longest increasing subsequence. In IEEE [IEE07], pages 294–304. ISBN 0-7695-3010-9. ISSN 0272-5428. LCCN QA76 .S974 2007. URL http://

ieeexplore.ieee.org/servlet/
opac?punumber=4389466. IEEE

Computer Society order number P3010.

Grandoni:2008:SCO

[GGL^{+}08] F. Grandoni, A. Gupta, S. Leonardi, P. Miettinen, P. Sankowski, and M. Singh. Set covering with our eyes closed. In IEEE [IEE08], pages 347–356. ISBN 0-7695-3436-8. ISSN 0272-5428. LCCN QA76 .S95 2008. URL http://

ieeexplore.ieee.org/servlet/
opac?punumber=4690923. IEEE

Computer Society order number P3436.

Goldreich:2003:IHR

[GGN03] O. Goldreich, S. Goldwasser, and A. Nussboim. On the implementation of huge random objects. In IEEE [IEE03], pages 68–79. CODEN ASF-PDV. ISBN 0-7695-2040-5. ISSN 0272-5428. LCCN QA76 .S979 2003. URL http:/

/ieeexplore.ieee.org/iel5/
8767/27770/01238182.pdf?isnumber=
27770&prod=CNF&arnumber=1238182&
arSt=+68&ared=+79&arAuthor=
Goldreich%2C+O.%3B+Goldwasser%
2C+S.%3B+Nussboim%2C+A.;
http://ieeexplore.ieee.org/
xpls/abs_all.jsp?isnumber=
27770&arnumber=1238182&count=
66&index=9.  IEEE Com-

puter Society Order Number PR02040.

Guruswami:2000:HAH

[GHS00] V. Guruswami, J. Håstad, and M. Sudan. Hardness of approximate hypergraph coloring. In IEEE [IEE00], pages 149–158. CODEN ASFPDV. ISBN 0-7695-0850-2, 0-7695-0851-0 (case), 0-7695-0852-9 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 2000. IEEE Computer Society Order Number PR00850.

Guruswami:2001:EBC

[GI01] V. Guruswami and P. Indyk. Expander-based constructions of efficiently decodable codes. In IEEE [IEE01], pages 658–667. CODEN ASFPDV. ISBN 0-7695-1390-5, 0-7695-1391-3 (case), 0-7695-1392-1 (microfiche). ISSN 0272-5428. LCCN ????

Gottesman:2009:QCC

[GI09] D. Gottesman and S. Irani. The quantum and classical complexity of translationally invariant tiling and Hamiltonian problems. In IEEE [IEE09], pages 95–104. ISBN 0-7695-3850-9. ISSN 0272-5428. LCCN QA76 .S95 2009. IEEE Computer Society order number P3850.

Gupta:2001:SSS

[GK01] A. Gupta and A. Kumar. Sorting and selection with structured costs. In IEEE [IEE01], pages 416–425. CODEN ASF-PDV. ISBN 0-7695-1390-5, 0-7695-1391-3 (case), 0-7695-1392-1 (microfiche). ISSN 0272-5428. LCCN ????

Garg:2002:FAA

[GK02] N. Garg and R. Khandekar. Fast approximation algorithms for fractional Steiner forest and related problems. In IEEE [IEE02], pages 500–509. CO-DEN ASFPDV. ISBN 0-7695-1822-2. ISSN 0272-5428. LCCN QA267. URL http:/

                                        /ieeexplore.ieee.org/iel5/
                                        8411/26517/01181974.pdf?isnumber=
                                        26517&prod=CNF&arnumber=1181974&
                                        arSt=+500&ared=+509&arAuthor=
                                        Garg%2C+N.%3B+Khandekar%2C+
                                        R.;   http://ieeexplore.ieee.
                                        org/xpls/abs_all.jsp?isnumber=
                                        26517&arnumber=1181974&count=
                                        82&index=50.  IEEE Com-

puter Society Order Number

PR01822.

Goldwasser:2003:SFS

[GK03] S. Goldwasser and Y. T. Kalai.

On the (in)security of the Fiat– Shamir paradigm. In IEEE [IEE03], pages 102–113. CO-DEN ASFPDV. ISBN 0-7695-2040-5. ISSN 0272-5428. LCCN QA76 .S979 2003. URL http:/

/ieeexplore.ieee.org/iel5/
8767/27770/01238185.pdf?isnumber=
27770&prod=CNF&arnumber=1238185&
arSt=+102&ared=+113&arAuthor=
Goldwasser%2C+S.%3B+Kalai%
2C+Y.T.;   http://ieeexplore.
ieee.org/xpls/abs_all.jsp?
isnumber=27770&arnumber=1238185&
count=66&index=12.  IEEE

Computer Society Order Number PR02040.

Goldwasser:2005:IOA

[GK05] S. Goldwasser and Yael Tauman Kalai. On the impossibility of obfuscation with auxiliary input. In IEEE [IEE05], pages 553–562. ISBN 0-7695-2468-0, 0-7695-2362-5. ISSN 0272-5428. LCCN QA76 .S979 2005. IEEE Computer Society order number P2468.

Garg:2007:MAF

[GK07] N. Garg and A. Kumar. Minimizing average flow-time : Upper and lower bounds. In IEEE [IEE07], pages 603–

613. ISBN 0-7695-3010-9. ISSN 0272-5428. LCCN QA76 .S974 2007. URL http://

ieeexplore.ieee.org/servlet/
opac?punumber=4389466. IEEE

Computer Society order number P3010.

Garay:2007:RCA

[GKKO07] J. A. Garay, J. Katz, Chiu-Yuen Koo, and R. Ostrovsky. Round complexity of authenticated broadcast with a dishonest majority. In IEEE [IEE07], pages 658–668. ISBN 0-7695-3010-9. ISSN 0272-5428. LCCN QA76 .S974 2007. URL http://

                                        ieeexplore.ieee.org/servlet/
                                        opac?punumber=4389466. IEEE

Computer Society order number

P3010.

Gupta:2003:BGF

[GKL03] A. Gupta, R. Krauthgamer, and J. R. Lee. Bounded geometries, fractals, and lowdistortion embeddings. In IEEE [IEE03], pages 534–543. CO-DEN ASFPDV. ISBN 0-7695-2040-5. ISSN 0272-5428. LCCN QA76 .S979 2003. URL http:/

                                        /ieeexplore.ieee.org/iel5/
                                        8767/27770/01238226.pdf?isnumber=
                                        27770&prod=CNF&arnumber=1238226&
                                        arSt=+534&ared=+543&arAuthor=
                                        Gupta%2C+A.%3B+Krauthgamer%
                                        2C+R.%3B+Lee%2C+J.R.;   http:
                                        //ieeexplore.ieee.org/xpls/
                                        abs_all.jsp?isnumber=27770&
                                        arnumber=1238226&count=66&

index=53. IEEE Computer Society Order Number PR02040.

Gertner:2000:RBP

[GKM^{+}00] Y. Gertner, S. Kannan, T. Malkin, O. Reingold, and M. Viswanathan. The relationship between public key encryption and oblivious transfer. In IEEE [IEE00], pages 325–335. CODEN ASF-PDV. ISBN 0-7695-0850-2, 0-7695-0851-0 (case), 0-7695-0852-9 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 2000. IEEE Computer Society Order Number PR00850.

Gupta:2003:ACS

[GKPR03] A. Gupta, A. Kumar, M. Pal, and T. Roughgarden. Approximation via cost-sharing: a simple approximation algorithm for the multicommodity rent-or-buy problem. In IEEE [IEE03], pages 606–615. CO-DEN ASFPDV. ISBN 0-7695-2040-5. ISSN 0272-5428. LCCN QA76 .S979 2003. URL http:/

/ieeexplore.ieee.org/iel5/
8767/27770/01238233.pdf?isnumber=
27770&prod=CNF&arnumber=1238233&
arSt=+606&ared=+615&arAuthor=
Gupta%2C+A.%3B+Kumar%2C+A.
%3B+Pal%2C+M.%3B+Roughgarden%
2C+T.;   http://ieeexplore.
ieee.org/xpls/abs_all.jsp?
isnumber=27770&arnumber=1238233&
count=66&index=60.  IEEE

Computer Society Order Number PR02040.

Gandhi:2002:DRB

[GKPS02] R. Gandhi, S. Khuller, S. Parthasarathy, and A. Srinivasan. Dependent rounding in bipartite graphs. In IEEE [IEE02], pages 323–332. CODEN ASFPDV. ISBN 0-7695-1822-2. ISSN 0272-5428. LCCN QA267. URL http:/

/ieeexplore.ieee.org/iel5/
8411/26517/01181955.pdf?isnumber=
26517&prod=CNF&arnumber=1181955&
arSt=+323&ared=+332&arAuthor=
Gandhi%2C+R.%3B+Khuller%2C+
S.%3B+Parthasarathy%2C+S.%

                                        3B+Srinivasan%2C+A.;   http:
                                        //ieeexplore.ieee.org/xpls/
                                        abs_all.jsp?isnumber=26517&
                                        arnumber=1181955&count=82&

index=33. IEEE Computer Society Order Number PR01822.

Gupta:2001:TPD

[GKR01] A. Gupta, A. Kumar, and R. Rastogi. Traveling with a PEZ dispenser (or, routing issues in MPLS). In IEEE [IEE01], pages 148–157. CO-DEN ASFPDV. ISBN 0-7695-1390-5, 0-7695-1391-3 (case), 0-7695-1392-1 (microfiche). ISSN 0272-5428. LCCN ????

Gottlob:2000:ESO

[GKS00] G. Gottlob, P. G. Kolaitis, and T. Schwentick. Existential second-order logic over graphs: charting the tractability frontier. In IEEE [IEE00], pages 664–674. CODEN ASF-PDV. ISBN 0-7695-0850-2, 0-7695-0851-0 (case), 0-7695-0852-9 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 2000. IEEE Computer Society Order Number PR00850.

Goyal:2005:LBN

[GKS05] N. Goyal, G. Kindler, and M. Saks. Lower bounds for the noisy broadcast problem. In IEEE [IEE05], pages 40– 49. ISBN 0-7695-2468-0, 0-7695-2362-5. ISSN 0272-5428. LCCN QA76 .S979 2005. IEEE Computer Society order number P2468.

Gopalan:2007:HRM

[GKS07] P. Gopalan, S. Khot, and R. Saket. Hardness of reconstructing multivariate polynomials over finite fields. In IEEE [IEE07], pages 349– 359. ISBN 0-7695-3010-9. ISSN 0272-5428. LCCN QA76 .S974 2007. URL http://

ieeexplore.ieee.org/servlet/
opac?punumber=4389466. IEEE

Computer Society order number P3010.

Goel:2009:ACP

[GKTW09] G. Goel, C. Karande, P. Tripathi, and Lei Wang. Approximability of combinatorial problems with multi-agent submodular cost functions. In IEEE [IEE09], pages 755–764. ISBN 0-7695-3850-9. ISSN 0272-5428. LCCN QA76 .S95 2009. IEEE Computer Society order number P3850.

Gu:2006:PCC

[GLM06] Xiaoyang Gu, Jack H. Lutz, and Elvira Mayordomo. Points on computable curves. In IEEE [IEE06], pages 469–474. ISBN 0-7695-2720-5, 0-7695-2362-5. ISSN 0272-5428. LCCN QA76 .S974 2006. URL http://

ieeexplore.ieee.org/servlet/
opac?punumber=4031329. IEEE

Computer Society Order Number P2720.

Garay:2000:COT

[GM00] J. A. Garay and P. MacKenzie. Concurrent oblivious transfer. In IEEE [IEE00], pages 314–324. CODEN ASFPDV. ISBN 0-7695-0850-2, 0-7695-0851-0 (case), 0-7695-0852-9 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 2000. IEEE Computer Society Order Number PR00850.

Gerschenfeld:2007:RMR

[GM07a] A. Gerschenfeld and A. Monianari. Reconstruction for models on random graphs. In IEEE [IEE07], pages 194– 204. ISBN 0-7695-3010-9. ISSN 0272-5428. LCCN QA76 .S974 2007. URL http://

ieeexplore.ieee.org/servlet/
opac?punumber=4389466. IEEE

Computer Society order number P3010.

Guha:2007:AAP

[GM07b] S. Guha and K. Munagala. Approximation algorithms for partial-information based stochastic control with Markovian rewards. In IEEE [IEE07], pages 483–493. ISBN 0-7695-3010-9. ISSN 0272-5428. LCCN QA76 .S974 2007. URL http://

ieeexplore.ieee.org/servlet/
opac?punumber=4389466. IEEE

Computer Society order number P3010.

Guha:2000:HPN

[GMM00] S. Guha, A. Meyerson, and K. Munagala. Hierarchical placement and network design problems. In IEEE [IEE00], pages 603–612. CODEN ASF-PDV. ISBN 0-7695-0850-2, 0-7695-0851-0 (case), 0-7695-0852-9 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 2000. IEEE Computer Society Order Number PR00850.

Guha:2000:CDS

[GMMO00] S. Guha, N. Mishra, R. Motwani, and L. O’Callaghan. Clustering data streams. In IEEE [IEE00], pages 359–366. CO-DEN ASFPDV. ISBN 0-7695-0850-2, 0-7695-0851-0 (case), 0-7695-0852-9 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 2000. IEEE Computer Society Order Number PR00850.

Goldberg:2004:SSM

[GMP04] L. A. Goldberg, R. Martin, and M. Paterson. Strong spatial mixing for lattice graphs with fewer colours. In IEEE [IEE04], pages 562–571. CO-DEN ASFPDV. ISBN 0-7695-2228-9. ISSN 0272-5428. LCCN QA276. URL http:/

/ieeexplore.ieee.org/iel5/
9430/29918/01366276.pdf?isnumber=
29918&prod=CNF&arnumber=1366276&
arSt=+562&ared=+571&arAuthor=
Goldberg%2C+L.A.%3B+Martin%
2C+R.%3B+Paterson%2C+M.;
http://ieeexplore.ieee.org/
xpls/abs_all.jsp?isnumber=
29918&arnumber=1366276&count=

64&index=57. IEEE Computer Society Order Number P2228.

Georgiou:2007:IGV

[GMPT07] K. Georgiou, A. Magen, T. Pitassi, and I. Tourlakis. Integrality gaps of 2 − o(1) for vertex cover SDPs in the Lovász–Schrijver hierarchy. In IEEE [IEE07], pages 702–712. ISBN 0-7695-3010-9. ISSN 0272-5428. LCCN QA76 .S974 2007. URL http://

                                        ieeexplore.ieee.org/servlet/
                                        opac?punumber=4389466. IEEE

Computer Society order number P3010.

Gertner:2001:IBT

[GMR01] Y. Gertner, T. Malkin, and O. Reingold. On the impossibility of basing trapdoor functions on trapdoor predicates. In IEEE [IEE01], pages 126–135. CO-DEN ASFPDV. ISBN 0-7695-1390-5, 0-7695-1391-3 (case), 0-7695-1392-1 (microfiche). ISSN 0272-5428. LCCN ????

Guruswami:2008:BRO

[GMR08] V. Guruswami, R. Manokaran, and P. Raghavendra. Beating the random ordering is hard: Inapproximability of maximum acyclic subgraph. In IEEE [IEE08], pages 573– 582. ISBN 0-7695-3436-8. ISSN 0272-5428. LCCN QA76 .S95 2008. URL http://

                                        ieeexplore.ieee.org/servlet/
                                        opac?punumber=4690923. IEEE

Computer Society order number P3436.

Goemans:2005:SEC

[GMV05] M. Goemans, Vahab Mirrokni, and A. Vetta. Sink equilibria and convergence. In IEEE [IEE05], pages 142–151. ISBN 0-7695-2468-0, 0-7695-2362-5. ISSN 0272-5428. LCCN QA76 .S979 2005. IEEE Computer Society order number P2468.

Goemans:2006:MBD

[Goe06] Michel X. Goemans. Minimum bounded degree spanning trees. In IEEE [IEE06], pages 273–282. ISBN 0-7695-2720-5, 0-7695-2362-5. ISSN 0272-5428. LCCN QA76 .S974 2006. URL http://

ieeexplore.ieee.org/servlet/
opac?punumber=4031329. IEEE

Computer Society Order Number P2720.

Goldreich:2002:ZKA

[Gol02] O. Goldreich. Zero-knowledge:

abstract of a tutorial. In IEEE [IEE02], page 3. CO-DEN ASFPDV. ISBN 0-7695-1822-2. ISSN 0272-5428. LCCN QA267. URL http:/

/ieeexplore.ieee.org/iel5/
8411/26517/01181876.pdf?isnumber=
26517&prod=CNF&arnumber=1181876&
arSt=+3&ared=&arAuthor=Goldreich%
2C+O.;   http://ieeexplore.
ieee.org/xpls/abs_all.jsp?
isnumber=26517&arnumber=1181876&
count=82&index=1. IEEE Com-

puter Society Order Number PR01822.

Guruswami:2006:CAG

[GP06] Venkatesan Guruswami and Anindya C. Patthak. Correlated algebraic-geometric codes: Improved list decoding over bounded alphabets. In IEEE [IEE06], pages 227–238. ISBN 0-7695-2720-5, 0-7695-2362-5. ISSN 0272-5428. LCCN QA76 .S974 2006. URL http://

ieeexplore.ieee.org/servlet/
opac?punumber=4031329. IEEE

Computer Society Order Number P2720.

Goel:2009:OAS

[GP09] A. Goel and I. Post. An oblivious O(1)-approximation for single source buy-at-bulk. In IEEE [IEE09], pages 442–450. ISBN 0-7695-3850-9. ISSN 0272-5428. LCCN QA76 .S95 2009. IEEE Computer Society order number P3850.

Goldwasser:2006:FTD

[GPV06] Shafi Goldwasser, Elan Pavlov, and Vinod Vaikuntanathan. Fault-tolerant distributed computing in full-information networks. In IEEE [IEE06], pages 15–26. ISBN 0-7695-2720-5, 0-7695-2362-5. ISSN 0272-5428. LCCN QA76 .S974 2006. URL http://

                                        ieeexplore.ieee.org/servlet/
                                        opac?punumber=4031329. IEEE

Computer Society Order Number P2720.

Gabizon:2005:DEA

[GR05] A. Gabizon and R. Raz. Deterministic extractors for affine sources over large fields. In IEEE [IEE05], pages 407–416. ISBN 0-7695-2468-0, 0-7695-2362-5. ISSN 0272-5428. LCCN QA76 .S979 2005. IEEE Computer Society order number P2468.

Guruswami:2006:HLH

[GR06] Venkatesan Guruswami and Prasad Raghavendra. Hardness of learning halfspaces with noise. In IEEE [IEE06], pages 543–552.

ISBN 0-7695-2720-5, 0-7695-2362-5. ISSN 0272-5428. LCCN QA76 .S974 2006. URL http://

ieeexplore.ieee.org/servlet/
opac?punumber=4031329. IEEE

Computer Society Order Number P2720.

Grohe:2003:CHC

[Gro03] M. Grohe. The complexity of homomorphism and constraint satisfaction problems seen from the other side. In IEEE [IEE03], pages 552–561. CODEN ASF-PDV. ISBN 0-7695-2040-5. ISSN 0272-5428. LCCN QA76 .S979 2003. URL http:/

/ieeexplore.ieee.org/iel5/
8767/27770/01238228.pdf?isnumber=
27770&prod=CNF&arnumber=1238228&
arSt=+552&ared=+561&arAuthor=
Grohe%2C+M.;   http://ieeexplore.
ieee.org/xpls/abs_all.jsp?
isnumber=27770&arnumber=1238228&
count=66&index=55.  IEEE

Computer Society Order Number PR02040.

Gabizon:2004:DEB

[GRS04a] A. Gabizon, R. Raz, and R. Shaltiel. Deterministic extractors for bit-fixing sources by obtaining an independent seed. In IEEE [IEE04], pages 394–

403. CODEN ASFPDV. ISBN 0-7695-2228-9. ISSN 0272-5428. LCCN QA276. URL http:/

/ieeexplore.ieee.org/iel5/
9430/29918/01366259.pdf?isnumber=
29918&prod=CNF&arnumber=1366259&
arSt=+394&ared=+403&arAuthor=
Gabizon%2C+A.%3B+Raz%2C+R.
%3B+Shaltiel%2C+R.;   http:
//ieeexplore.ieee.org/xpls/

                                        abs_all.jsp?isnumber=29918&
                                        arnumber=1366259&count=64&

index=40. IEEE Computer Society Order Number P2228.

Gupta:2004:ETS

[GRS04b] A. Gupta, R. Ravi, and A. Sinha. An edge in time saves nine: LP rounding approximation algorithms for stochastic network design. In IEEE [IEE04], pages 218–227. CO-DEN ASFPDV. ISBN 0-7695-2228-9. ISSN 0272-5428. LCCN QA276. URL http:/

                                        /ieeexplore.ieee.org/iel5/
                                        9430/29918/01366241.pdf?isnumber=
                                        29918&prod=CNF&arnumber=1366241&
                                        arSt=+218&ared=+227&arAuthor=
                                        Gupta%2C+A.%3B+Ravi%2C+R.%
                                        3B+Sinha%2C+A.;   http://
                                        ieeexplore.ieee.org/xpls/
                                        abs_all.jsp?isnumber=29918&
                                        arnumber=1366241&count=64&

index=22. IEEE Computer Society Order Number P2228.

Gentry:2006:EES

[GRW06] Craig Gentry, Zulfikar Ramzan, and David P. Woodruff. Explicit exclusive set systems with applications to broadcast encryption. In IEEE [IEE06], pages 27–

38. ISBN 0-7695-2720-5, 0-7695-2362-5. ISSN 0272-5428. LCCN QA76 .S974 2006. URL http://

                                        ieeexplore.ieee.org/servlet/
                                        opac?punumber=4031329. IEEE

Computer Society Order Num-

ber P2720.

Goldreich:2002:LTC

[GS02] O. Goldreich and M. Sudan. Locally testable codes and PCPs of almost-linear length. In IEEE [IEE02], pages 13–22. CODEN ASFPDV. ISBN 0-7695-1822-2. ISSN 0272-5428. LCCN QA267. URL http:/

/ieeexplore.ieee.org/iel5/
8411/26517/01181878.pdf?isnumber=
26517&prod=CNF&arnumber=1181878&
arSt=+13&ared=+22&arAuthor=
Goldreich%2C+O.%3B+Sudan%
2C+M.;   http://ieeexplore.
ieee.org/xpls/abs_all.jsp?
isnumber=26517&arnumber=1181878&
count=82&index=3. IEEE Com-

puter Society Order Number PR01822.

Guruswami:2000:SDD

[GSS00] V. Guruswami, A. Sahai, and M. Sudan. “soft-decision” decoding of Chinese remainder codes. In IEEE [IEE00], pages 159–168. CODEN ASFPDV. ISBN 0-7695-0850-2, 0-7695-0851-0 (case), 0-7695-0852-9 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 2000. IEEE Computer Society Order Number PR00850.

Gennaro:2000:LBE

[GT00] R. Gennaro and L. Trevisan.

Lower bounds on the efficiency of generic cryptographic constructions. In IEEE [IEE00], pages 305–313. CODEN ASF-PDV. ISBN 0-7695-0850-2, 0-7695-0851-0 (case), 0-7695-0852-9 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 2000. IEEE Computer Society Order Number PR00850.

Goldreich:2001:TTR

[GT01] O. Goldreich and L. Trevisan. Three theorems regarding testing graph properties. In IEEE [IEE01], pages 460–469. CO-DEN ASFPDV. ISBN 0-7695-1390-5, 0-7695-1391-3 (case), 0-7695-1392-1 (microfiche). ISSN 0272-5428. LCCN ????

Ghen:2007:PBL

[GT07] Xi Ghen and Shang-Hua Teng. Paths beyond local search: a tight bound for randomized fixed-point computation. In IEEE [IEE07], pages 124– 134. ISBN 0-7695-3010-9. ISSN 0272-5428. LCCN QA76 .S974 2007. URL http://

                                        ieeexplore.ieee.org/servlet/
                                        opac?punumber=4389466. IEEE

Computer Society order number P3010.

Guha:2000:NGD

[Guh00] S. Guha. Nested graph dissection and approximation algorithms. In IEEE [IEE00], pages 126–135. CODEN ASF-PDV. ISBN 0-7695-0850-2, 0-7695-0851-0 (case), 0-7695-0852-9 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 2000. IEEE Computer Society Order Number PR00850.

Gibson:2009:DCP

[GV09] M. Gibson and K. Varadarajan. Decomposing coverings and the planar sensor cover problem. In IEEE [IEE09], pages 159–168. ISBN 0-7695-3850-9. ISSN 0272-5428. LCCN QA76 .S95 2009.

IEEE Computer Society order number P3850.

Garg:2002:LEE

[GY02] N. Garg and N. E. Young.

On-line end-to-end congestion control. In IEEE [IEE02], pages 303–310. CODEN ASFPDV. ISBN 0-7695-1822-2. ISSN 0272-5428. LCCN QA267. URL http:/

/ieeexplore.ieee.org/iel5/
8411/26517/01181953.pdf?isnumber=
26517&prod=CNF&arnumber=1181953&
arSt=+303&ared=+310&arAuthor=
Garg%2C+N.%3B+Young%2C+N.
E.;   http://ieeexplore.ieee.
org/xpls/abs_all.jsp?isnumber=
26517&arnumber=1181953&count=
82&index=31.  IEEE Com-

puter Society Order Number PR01822.

Haitner:2009:PRT

[Hai09] I. Haitner. A parallel repetition theorem for any interactive argument. In IEEE [IEE09], pages 241–250. ISBN 0-7695-3850-9. ISSN 0272-5428. LCCN QA76 .S95 2009. IEEE Computer Society order number P3850.

Halman:2004:PDL

[Hal04] N. Halman. On the power of discrete and of lexicographic Helly-type theorems. In IEEE [IEE04], pages 463–472. CO-DEN ASFPDV. ISBN 0-7695-2228-9. ISSN 0272-5428. LCCN QA276. URL http:/

/ieeexplore.ieee.org/iel5/
9430/29918/01366266.pdf?isnumber=
29918&prod=CNF&arnumber=1366266&
arSt=+463&ared=+472&arAuthor=

                                        Halman%2C+N.;   http://ieeexplore.
                                        ieee.org/xpls/abs_all.jsp?
                                        isnumber=29918&arnumber=1366266&
                                        count=64&index=47.  IEEE

Computer Society Order Num-

ber P2228.

Harvey:2006:ASA

[Har06] Nicholas J. A. Harvey. Algebraic structures and algorithms for matching and matroid problems. In IEEE [IEE06], pages 531–542. ISBN 0-7695-2720-5, 0-7695-2362-5. ISSN 0272-5428. LCCN QA76 .S974 2006. URL http://

                                        ieeexplore.ieee.org/servlet/
                                        opac?punumber=4031329. IEEE

Computer Society Order Num-

ber P2720.

Hayes:2006:SCI

[Hay06] Thomas P. Hayes. A simple condition implying rapid mixing of single-site dynamics on spin systems. In IEEE [IEE06], pages 39–46. ISBN 0-7695-2720-5, 0-7695-2362-5. ISSN 0272-5428. LCCN QA76 .S974 2006. URL http://

                                        ieeexplore.ieee.org/servlet/
                                        opac?punumber=4031329. IEEE

Computer Society Order Num-

ber P2720.

Hales:2000:IQF

[HH00] L. Hales and S. Hallgren. An improved quantum Fourier transform algorithm and applications. In IEEE [IEE00], pages 515–525. CODEN ASFPDV. ISBN 0-7695-0850-2, 0-7695-0851-0 (case), 0-7695-0852-9 (microfiche). ISSN 0272-5428.

LCCN TK7885.A1 S92 2000. IEEE Computer Society Order Number PR00850.

Haitner:2007:FCI

[HHRS07] I. Haitner, J. J. Hoch, O. Reingold, and G. Segev. Finding collisions in interactive protocols — a tight lower bound on the round complexity of statistically-hiding commitments. In IEEE [IEE07], pages 669–679. ISBN 0-7695-3010-9. ISSN 0272-5428. LCCN QA76 .S974 2007. URL http://

ieeexplore.ieee.org/servlet/
opac?punumber=4389466. IEEE

Computer Society order number P3010.

Haastad:2001:QEP

[HK01] J. Håstad and S. Khot. Query efficient PCPs with perfect completeness. In IEEE [IEE01], pages 610–619. CODEN ASF-PDV. ISBN 0-7695-1390-5, 0-7695-1391-3 (case), 0-7695-1392-1 (microfiche). ISSN 0272-5428. LCCN ????

Hassidim:2009:LGP

[HKNO09] A. Hassidim, J. A. Kelner, H. N.

Nguyen, and K. Onak. Local graph partitions for approximation and testing. In IEEE [IEE09], pages 22–31. ISBN 0-7695-3850-9. ISSN 0272-5428. LCCN QA76 .S95 2009. IEEE Computer Society order number P3850.

Harnik:2006:CNI

[HN06] Danny Harnik and Moni Naor.

On the compressibility of NP instances and cryptographic applications. In IEEE [IEE06], pages 719–728. ISBN 0-7695-2720-5, 0-7695-2362-5. ISSN 0272-5428. LCCN QA76 .S974 2006. URL http://

ieeexplore.ieee.org/servlet/
opac?punumber=4031329. IEEE

Computer Society Order Number P2720.

Harvey:2008:SSE

[HNO08] N. J. A. Harvey, J. Nelson, and K. Onak. Sketching and streaming entropy via approximation theory. In IEEE [IEE08], pages 489–498. ISBN 0-7695-3436-8. ISSN 0272-5428. LCCN QA76 .S95 2008. URL http://

ieeexplore.ieee.org/servlet/
opac?punumber=4690923. IEEE

Computer Society order number P3436.

Haxell:2005:AVH

[HNR05] P. E. Haxell, B. Nagle, and V. Rodl. An algorithmic version of the hypergraph regularity method. In IEEE [IEE05], pages 439–446. ISBN 0-7695-2468-0, 0-7695-2362-5. ISSN 0272-5428. LCCN QA76 .S979 2005. IEEE Computer Society order number P2468.

Har-Peled:2001:CM

[HP01a] S. Har-Peled. Clustering motion. In IEEE [IEE01], pages 84–93. CODEN ASFPDV. ISBN 0-7695-1390-5, 0-7695-1391-3 (case), 0-7695-1392-1 (microfiche). ISSN 0272-5428. LCCN ????

Har-Peled:2001:RVD

[HP01b] S. Har-Peled. A replacement for Voronoi diagrams of near linear size. In IEEE [IEE01], pages 94–103. CODEN ASF-PDV. ISBN 0-7695-1390-5, 0-7695-1391-3 (case), 0-7695-1392-1 (microfiche). ISSN 0272-5428. LCCN ????

Herte:2007:ETS

[HP07] P. Herte and T. Pitassi. Exponential time/space speedups for resolution and the PSPACEcompleteness of black-white pebbling. In IEEE [IEE07], pages 137–149. ISBN 0-7695-3010-9. ISSN 0272-5428. LCCN QA76 .S974 2007. URL http://

ieeexplore.ieee.org/servlet/
opac?punumber=4389466. IEEE

Computer Society order number P3010.

Har-Peled:2001:ASF

[HPV01] S. Har-Peled and K. R.

Varadarajan. Approximate shape fitting via linearization. In IEEE [IEE01], pages 66–73. CO-DEN ASFPDV. ISBN 0-7695-1390-5, 0-7695-1391-3 (case), 0-7695-1392-1 (microfiche). ISSN 0272-5428. LCCN ????

Herlihy:2000:EBT

[HR00] M. Herlihy and E. Ruppert.

On the existence of booster types. In IEEE [IEE00], pages 653–663. CODEN ASFPDV. ISBN 0-7695-0850-2, 0-7695-0851-0 (case), 0-7695-0852-9 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 2000.

IEEE Computer Society Order Number PR00850.

Hershberger:2001:VPS

[HS01] J. Hershberger and Subhash Suri. Vickrey prices and shortest paths: what is an edge worth? In IEEE [IEE01], pages 252–259. CODEN ASF-PDV. ISBN 0-7695-1390-5, 0-7695-1391-3 (case), 0-7695-1392-1 (microfiche). ISSN 0272-5428. LCCN ???? See erratum [HS02].

Hershberger:2002:ESP

[HS02] J. Hershberger and Subhash Suri. Erratum to “Vickrey pricing and shortest paths: what is an edge worth?”. In IEEE [IEE02], page 809. CO-DEN ASFPDV. ISBN 0-7695-1822-2. ISSN 0272-5428. LCCN QA267. URL http:/

/ieeexplore.ieee.org/iel5/
8411/26517/01182006.pdf?isnumber=
26517&prod=CNF&arnumber=1182006&
arSt=+809&ared=+809&arAuthor=
Hershberger%2C+J.%3B+Subhash+
Suri;   http://ieeexplore.
ieee.org/xpls/abs_all.jsp?
isnumber=26517&arnumber=1182006&
count=82&index=80.  See

[HS01].

Hayes:2005:GLB

[HS05] T. P. Hayes and A. Sinclair. A general lower bound for mixing of single-site dynamics on graphs. In IEEE [IEE05], pages 511–520. ISBN 0-7695-2468-0, 0-7695-2362-5. ISSN 0272-5428. LCCN QA76 .S979 2005. IEEE Computer Society order number P2468.

Hon:2003:BTS

[HSS03] Wing-Kai Hon, K. Sadakane, and Wing-Kin Sung. Breaking a time-and-space barrier in constructing full-text indices. In IEEE [IEE03], pages 251–260. CODEN ASF-PDV. ISBN 0-7695-2040-5. ISSN 0272-5428. LCCN QA76 .S979 2003. URL http:/

/ieeexplore.ieee.org/iel5/
8767/27770/01238199.pdf?isnumber=
27770&prod=CNF&arnumber=1238199&
arSt=+251&ared=+260&arAuthor=
Wing-Kai+Hon%3B+Sadakane%
2C+K.%3B+Wing-Kin+Sung;
http://ieeexplore.ieee.org/
xpls/abs_all.jsp?isnumber=
27770&arnumber=1238199&count=
66&index=26.  IEEE Com-

puter Society Order Number PR02040.

Hon:2009:SEF

[HSV09] Wing-Kai Hon, R. Shah, and J. S. Vitter. Space-efficient framework for top-k string retrieval problems. In IEEE [IEE09], pages 713–722. ISBN 0-7695-3850-9. ISSN 0272-5428. LCCN QA76 .S95 2009. IEEE Computer Society order number P3850.

Han:2002:ISS

[HT02] Yijie Han and M. Thorup. Integer sorting in O(n\sqrt{(} log log n)) expected time and linear space. In IEEE [IEE02], pages 135–

144. CODEN ASFPDV. ISBN 0-7695-1822-2. ISSN 0272-5428.

LCCN QA267. URL http:/

/ieeexplore.ieee.org/iel5/
8411/26517/01181890.pdf?isnumber=
26517&prod=CNF&arnumber=1181890&
arSt=+135&ared=+144&arAuthor=
Yijie+Han%3B+Thorup%2C+M.;
 http://ieeexplore.ieee.org/
xpls/abs_all.jsp?isnumber=
26517&arnumber=1181890&count=
82&index=15.  IEEE Com-

puter Society Order Number PR01822.

Hayes:2003:NMC

[HV03] T. P. Hayes and E. Vigoda. A non-Markovian coupling for randomly sampling colorings. In IEEE [IEE03], pages 618–627. CODEN ASF-PDV. ISBN 0-7695-2040-5. ISSN 0272-5428. LCCN QA76 .S979 2003. URL http:/

/ieeexplore.ieee.org/iel5/
8767/27770/01238234.pdf?isnumber=
27770&prod=CNF&arnumber=1238234&
arSt=+618&ared=+627&arAuthor=
Hayes%2C+T.P.%3B+Vigoda%2C+
E.;   http://ieeexplore.ieee.
org/xpls/abs_all.jsp?isnumber=
27770&arnumber=1238234&count=
66&index=61.  IEEE Com-

puter Society Order Number PR02040.

IEEE:2000:ASF

[IEE00] IEEE, editor. 41st Annual Symposium on Foundations of Computer Science: proceedings: 12– 14 November, 2000, Redondo Beach, California. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2000. CODEN ASFPDV. ISBN 0-7695-0850-2, 0-7695-0851-0 (case), 0-7695-0852-9 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 2000. IEEE Computer Society Order Number PR00850.

IEEE:2001:ISF

[IEE01] IEEE, editor. 42nd IEEE Symposium on Foundations of Computer Science: proceedings: October 14–17, 2001, Las Vegas, Nevada, USA. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2001. CODEN ASFPDV. ISBN 0-7695-1390-5, 0-7695-1391-3 (case), 0-7695-1392-1 (microfiche). ISSN 0272-5428. LCCN ????

IEEE:2002:PAI

[IEE02] IEEE, editor. Proceedings of the 43rd Annual IEEE Symposium on Foundations of Computer Science, FOCS 2002, Vancouver, BC, Canada, 16–19 November 2002. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2002. CO-DEN ASFPDV. ISBN 0-7695-1822-2. ISSN 0272-5428. LCCN QA267. URL http:/

/ieeexplore.ieee.org/iel5/
8411/26517/01181875.pdf. IEEE

Computer Society Order Number PR01822.

IEEE:2003:PAI

[IEE03] IEEE, editor. Proceedings: 44th Annual IEEE Symposium on Foundations of Computer Science, FOCS 2003, 11–14 October 2003, Cambridge, Massachusetts. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2003. CODEN ASFPDV. ISBN 0-7695-2040-5. ISSN 0272-5428. LCCN QA76 .S979 2003. URL http:/

/ieeexplore.ieee.org/iel5/
8767/27770/01238173.pdf;
http://ieeexplore.ieee.org/
xpl/RecentCon.jsp?punumber=
8767&conhome=1000292. IEEE

Computer Society Order Number PR02040.

IEEE:2004:PAI

[IEE04] IEEE, editor. Proceedings: 45th Annual IEEE Symposium on Foundations of Computer Science: FOCS 2004, 17–19 October, 2004, Rome, Italy. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2004. CODEN ASFPDV. ISBN 0-7695-2228-9. ISSN 0272-5428. LCCN QA276. URL http:/

/ieeexplore.ieee.org/iel5/
9430/29918/01366212.pdf;
http://ieeexplore.ieee.org/
servlet/opac?punumber=9430.

IEEE Computer Society Order Number P2228.

IEEE:2005:AIS

[IEE05] IEEE, editor. 46th Annual IEEE Symposium on Foundations of Computer Science: FOCS 2005: 23–25 October, 2005, Pittsburgh, Pennsylvania, USA. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2005. ISBN 0-7695-2468-0, 0-7695-2362-5. ISSN 0272-5428. LCCN QA76 .S979 2005. IEEE Computer Society order number P2468.

IEEE:2006:AIS

[IEE06] IEEE, editor. 47th Annual IEEE Symposium on Foundations of Computer Science: FOCS 2006: 21–24 October, 2006, Berkeley, California. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2006. ISBN 0-7695-2720-5, 0-7695-2362-5. ISSN 0272-5428. LCCN QA76 .S974 2006. URL http://

ieeexplore.ieee.org/servlet/
opac?punumber=4031329. IEEE

Computer Society Order Number P2720.

IEEE:2007:PAI

[IEE07] IEEE, editor. Proceedings of the 48th Annual IEEE Symposium on Foundations of Computer Science: [FOCS 2007]: October 20–23, 2007, Providence, Rhode Island. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2007. ISBN 0-7695-3010-9. ISSN 0272-5428. LCCN QA76 .S974 2007. URL http://

ieeexplore.ieee.org/servlet/
opac?punumber=4389466. IEEE

Computer Society order number P3010.

IEEE:2008:PAI

[IEE08] IEEE, editor. Proceedings of the 49th Annual IEEE Symposium on Foundations of Computer Science: October 25–23, 2008, Philadelphia, Pennsylvania, USA. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2008. ISBN 0-7695-3436-8. ISSN 0272-5428. LCCN QA76 .S95 2008. URL http://

                                        ieeexplore.ieee.org/servlet/
                                        opac?punumber=4690923. IEEE

Computer Society order number P3436.

IEEE:2009:PAI

[IEE09] IEEE, editor. Proceedings of the 50th Annual IEEE Symposium on Foundations of Computer Science: October 25–27, 2009, Atlanta, Georgia, USA. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2009. ISBN 0-7695-3850-9. ISSN 0272-5428. LCCN QA76 .S95 2009. IEEE Computer Society order number P3850.

Impagliazzo:2006:ALD

[IJK06] Russell Impagliazzo, Ragesh Jaiswal, and Valentine Kabanets. Approximately listdecoding direct product codes and uniform hardness amplification. In IEEE [IEE06], pages 187–196. ISBN 0-7695-2720-5, 0-7695-2362-5. ISSN 0272-5428. LCCN QA76 .S974 2006. URL http://

                                        ieeexplore.ieee.org/servlet/
                                        opac?punumber=4031329. IEEE

Computer Society Order Number P2720.

Ishai:2000:RPN

[IK00] Y. Ishai and E. Kushilevitz.

Randomizing polynomials: A new representation with applications to round-efficient secure computation. In IEEE [IEE00], pages 294–304. CO-DEN ASFPDV. ISBN 0-7695-0850-2, 0-7695-0851-0 (case), 0-7695-0852-9 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 2000. IEEE Computer Society Order Number PR00850.

Impagliazzo:2003:LRA

[IK03] R. Impagliazzo and B. M.

Kapron. Logics for reasoning about cryptographic constructions. In IEEE [IEE03], pages 372–383. CODEN ASF-PDV. ISBN 0-7695-2040-5. ISSN 0272-5428. LCCN QA76 .S979 2003. URL http:/

/ieeexplore.ieee.org/iel5/
8767/27770/01238211.pdf?isnumber=
27770&prod=CNF&arnumber=1238211&
arSt=+372&ared=+383&arAuthor=
Impagliazzo%2C+R.%3B+Kapron%
2C+B.M.;   http://ieeexplore.
ieee.org/xpls/abs_all.jsp?
isnumber=27770&arnumber=1238211&
count=66&index=38.  IEEE

Computer Society Order Number PR02040.

Kawarabayashi:2009:PAF

[iK09] K. i. Kawarabayashi. Planarity allowing few error vertices in linear time. In IEEE [IEE09], pages 639–648. ISBN 0-7695-3850-9. ISSN 0272-5428. LCCN QA76 .S95 2009. IEEE Computer Society order number P3850.

Kawarabayashi:2008:SLT

[iKMR08] K. i. Kawarabayashi, B. Mohar, and B. Reed. A simpler linear time algorithm for embedding graphs into an arbitrary surface and the genus of graphs of bounded treewidth. In IEEE [IEE08], pages 771–780. ISBN 0-7695-3436-8. ISSN 0272-5428. LCCN QA76 .S95 2008. URL http://

                                        ieeexplore.ieee.org/servlet/
                                        opac?punumber=4690923. IEEE

Computer Society order number P3436.

Immorlica:2007:BPA

[IKMT07] N. Immorlica, A. R. Karlin, M. Mahdian, and K. Talvvar. Balloon popping with applications to ascending auctions. In IEEE [IEE07], pages 104–112. ISBN 0-7695-3010-9. ISSN 0272-5428. LCCN QA76 .S974 2007. URL http://

                                        ieeexplore.ieee.org/servlet/
                                        opac?punumber=4389466. IEEE

Computer Society order number P3010.

Ishai:2006:CA

[IKOS06] Yuval Ishai, Eyal Kushilevitz, Rafail Ostrovsky, and Amit Sahai. Cryptography from anonymity. In IEEE [IEE06], pages 239–248. ISBN 0-7695-2720-5, 0-7695-2362-5. ISSN 0272-5428. LCCN QA76 .S974 2006. URL http://

                                        ieeexplore.ieee.org/servlet/
                                        opac?punumber=4031329. IEEE

Computer Society Order Number P2720.

Ishai:2009:EC

[IKOS09] Y. Ishai, E. Kushilevitz, R. Ostrovsky, and A. Sahai. Extracting correlations. In IEEE [IEE09], pages 261–270. ISBN 0-7695-3850-9. ISSN 0272-5428. LCCN QA76 .S95 2009. IEEE Computer Society order number P3850.

Izmalkov:2005:RSC

[IML05] S. Izmalkov, S. Micali, and M. Lepinski. Rational secure computation and ideal mechanism design. In IEEE [IEE05], pages 585–594. ISBN 0-7695-2468-0, 0-7695-2362-5. ISSN 0272-5428. LCCN QA76 .S979 2005. IEEE Computer Society order number P2468.

Iwata:2009:SFM

[IN09] S. Iwata and K. Nagano. Submodular function minimization under covering constraints. In IEEE [IEE09], pages 671–680. ISBN 0-7695-3850-9. ISSN 0272-5428. LCCN QA76 .S95 2009. IEEE Computer Society order number P3850.

Indyk:2000:SDP

[Ind00] P. Indyk. Stable distributions, pseudorandom generators, embeddings and data stream computation. In IEEE [IEE00], pages 189–197. CODEN ASF-PDV. ISBN 0-7695-0850-2, 0-7695-0851-0 (case), 0-7695-0852-9 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 2000. IEEE Computer Society Order Number PR00850.

Indyk:2001:AAL

[Ind01] P. Indyk. Algorithmic applications of low-distortion geometric embeddings. In IEEE [IEE01], pages 10–33. CODEN ASFPDV. ISBN 0-7695-1390-5, 0-7695-1391-3 (case), 0-7695-1392-1 (microfiche). ISSN 0272-5428. LCCN ????

Indyk:2008:NOS

[IR08] P. Indyk and M. Ruzic. Nearoptimal sparse recovery in the L_{1} norm. In IEEE [IEE08], pages 199–207. ISBN 0-7695-3436-8. ISSN 0272-5428. LCCN QA76 .S95 2008. URL http://

                                        ieeexplore.ieee.org/servlet/
                                        opac?punumber=4690923. IEEE

Computer Society order number

P3436.

Impagliazzo:2001:CAD

[IS01] R. Impagliazzo and N. Segerlind. Counting axioms do not polynomially simulate counting gates. In IEEE [IEE01], pages 200–209. CODEN ASFPDV. ISBN 0-7695-1390-5, 0-7695-1391-3 (case), 0-7695-1392-1 (microfiche). ISSN 0272-5428. LCCN ????

Indyk:2003:TLB

[IW03] P. Indyk and D. Woodruff. Tight lower bounds for the distinct elements problem. In IEEE [IEE03], pages 283–288. CO-DEN ASFPDV. ISBN 0-7695-2040-5. ISSN 0272-5428. LCCN QA76 .S979 2003. URL http:/

                                        /ieeexplore.ieee.org/iel5/
                                        8767/27770/01238202.pdf?isnumber=

27770&prod=CNF&arnumber=1238202&
arSt=+283&ared=+288&arAuthor=
Indyk%2C+P.%3B+Woodruff%2C+
D.;   http://ieeexplore.ieee.
org/xpls/abs_all.jsp?isnumber=
27770&arnumber=1238202&count=
66&index=29.  IEEE Com-

puter Society Order Number PR02040.

Jain:2004:PTA

[Jai04] K. Jain. A polynomial time algorithm for computing an Arrow–Debreu market equilibrium for linear utilities. In IEEE [IEE04], pages 286–294. CODEN ASFPDV. ISBN 0-7695-2228-9. ISSN 0272-5428. LCCN QA276. URL http:/

/ieeexplore.ieee.org/iel5/
9430/29918/01366248.pdf?isnumber=
29918&prod=CNF&arnumber=1366248&
arSt=+286&ared=+294&arAuthor=
Jain%2C+K.;   http://ieeexplore.
ieee.org/xpls/abs_all.jsp?
isnumber=29918&arnumber=1366248&
count=64&index=29.  IEEE

Computer Society Order Number P2228.

Jutla:2004:TLD

[JPRZ04] C. S. Jutla, A. C. Patthak, A. Rudra, and D. Zuckerman. Testing low-degree polynomials over prime fields. In IEEE [IEE04], pages 423–432. CODEN ASFPDV. ISBN 0-7695-2228-9. ISSN 0272-5428. LCCN QA276. URL http:/

/ieeexplore.ieee.org/iel5/
9430/29918/01366262.pdf?isnumber=
29918&prod=CNF&arnumber=1366262&
arSt=+423&ared=+432&arAuthor=
Jutla%2C+C.S.%3B+Patthak%

                                        2C+A.C.%3B+Rudra%2C+A.%3B+
                                        Zuckerman%2C+D.;   http://
                                        ieeexplore.ieee.org/xpls/
                                        abs_all.jsp?isnumber=29918&
                                        arnumber=1366262&count=64&

index=43. IEEE Computer Society Order Number P2228.

Jain:2003:LBB

[JRP03] R. Jain, J. Radhakrishnan, and Sen P. A lower bound for the bounded round quantum communication complexity of set disjointness. In IEEE [IEE03], pages 220–229. CO-DEN ASFPDV. ISBN 0-7695-2040-5. ISSN 0272-5428. LCCN QA76 .S979 2003. URL http:/

                                        /ieeexplore.ieee.org/iel5/
                                        8767/27770/01238196.pdf?isnumber=
                                        27770&prod=CNF&arnumber=1238196&
                                        arSt=+220&ared=+229&arAuthor=
                                        Jain%2C+R.%3B+Radhakrishnan%
                                        2C+J.%3B+Sen+P;   http://
                                        ieeexplore.ieee.org/xpls/
                                        abs_all.jsp?isnumber=27770&
                                        arnumber=1238196&count=66&

index=23. IEEE Computer Society Order Number PR02040.

Jain:2002:PIQ

[JRS02] R. Jain, J. Radhakrishnan, and P. Sen. Privacy and interaction in quantum communication complexity and a theorem about the relative entropy of quantum states. In IEEE [IEE02], pages 429–438. CODEN ASFPDV. ISBN 0-7695-1822-2. ISSN 0272-5428. LCCN QA267. URL http:/

                                        /ieeexplore.ieee.org/iel5/
                                        8411/26517/01181967.pdf?isnumber=

26517&prod=CNF&arnumber=1181967&
arSt=+429&ared=+438&arAuthor=
Jain%2C+R.%3B+Radhakrishnan%
2C+J.%3B+Sen%2C+P.;   http:
//ieeexplore.ieee.org/xpls/
abs_all.jsp?isnumber=26517&
arnumber=1181967&count=82&

index=43. IEEE Computer Society Order Number PR01822.

Jerrum:2002:SGL

[JS02] Mark Jerrum and Jung-Bae Son.

Spectral gap and log-Sobolev constant for balanced matroids. In IEEE [IEE02], pages 721–

729. CODEN ASFPDV. ISBN 0-7695-1822-2. ISSN 0272-5428. LCCN QA267. URL http:/

/ieeexplore.ieee.org/iel5/
8411/26517/01181997.pdf?isnumber=
26517&prod=CNF&arnumber=1181997&
arSt=+721&ared=+729&arAuthor=
Mark+Jerrum%3B+Jung-Bae+Son;
http://ieeexplore.ieee.org/
xpls/abs_all.jsp?isnumber=
26517&arnumber=1181997&count=
82&index=71.  IEEE Com-

puter Society Order Number PR01822.

Jain:2009:TMQ

[JUW09] R. Jain, S. Upadhyay, and J. Watrous. Two-message quantum interactive proofs are in PSPACE. In IEEE [IEE09], pages 534–543. ISBN 0-7695-3850-9. ISSN 0272-5428. LCCN QA76 .S95 2009. IEEE Computer Society order number P3850.

Jayram:2009:DSS

[JW09] T. S. Jayram and D. P.

Woodruff. The data stream space complexity of cascaded norms. In IEEE [IEE09], pages 765–774. ISBN 0-7695-3850-9. ISSN 0272-5428. LCCN QA76 .S95 2009. IEEE Computer Society order number P3850.

Kannan:2009:NPI

[Kan09] R. Kannan. A new probability inequality using typical moments and concentration results. In IEEE [IEE09], pages 211–220. ISBN 0-7695-3850-9. ISSN 0272-5428. LCCN QA76 .S95 2009. IEEE Computer Society order number P3850.

Kempe:2003:GBC

[KDG03] D. Kempe, A. Dobra, and J. Gehrke. Gossip-based computation of aggregate information. In IEEE [IEE03], pages 482–491. CODEN ASF-PDV. ISBN 0-7695-2040-5. ISSN 0272-5428. LCCN QA76 .S979 2003. URL http:/

/ieeexplore.ieee.org/iel5/
8767/27770/01238221.pdf?isnumber=
27770&prod=CNF&arnumber=1238221&
arSt=+482&ared=+491&arAuthor=
Kempe%2C+D.%3B+Dobra%2C+A.
%3B+Gehrke%2C+J.;   http:/
/ieeexplore.ieee.org/xpls/
abs_all.jsp?isnumber=27770&
arnumber=1238221&count=66&

index=48. IEEE Computer Society Order Number PR02040.

Kumar:2002:CFA

[KGR02] A. Kumar, A. Gupta, and T. Roughgarden. A constantfactor approximation algorithm for the multicommodity rentor-buy problem. In IEEE [IEE02], pages 333–342. CO-DEN ASFPDV. ISBN 0-7695-1822-2. ISSN 0272-5428. LCCN QA267. URL http:/

/ieeexplore.ieee.org/iel5/
8411/26517/01181956.pdf?isnumber=
26517&prod=CNF&arnumber=1181956&
arSt=+333&ared=+342&arAuthor=
Kumar%2C+A.%3B+Gupta%2C+A.
%3B+Roughgarden%2C+T.;   http:
//ieeexplore.ieee.org/xpls/
abs_all.jsp?isnumber=26517&
arnumber=1181956&count=82&

index=34. IEEE Computer Society Order Number PR01822.

Khot:2002:HRC

[Kho02]  S. Khot.  Hardness results

for coloring 3-colorable 3uniform hypergraphs. In IEEE [IEE02], pages 23–32. CO-DEN ASFPDV. ISBN 0-7695-1822-2. ISSN 0272-5428. LCCN QA267. URL http:/

/ieeexplore.ieee.org/iel5/
8411/26517/01181879.pdf?isnumber=
26517&prod=CNF&arnumber=1181879&
arSt=+23&ared=+32&arAuthor=
Khot%2C+S.;   http://ieeexplore.
ieee.org/xpls/abs_all.jsp?
isnumber=26517&arnumber=1181879&
count=82&index=4. IEEE Com-

puter Society Order Number PR01822.

Khot:2003:HAS

[Kho03] S. Khot. Hardness of approximating the shortest vector problem in high L_{p} norms. In IEEE [IEE03], pages 290–297. CO-DEN ASFPDV. ISBN 0-7695-2040-5. ISSN 0272-5428. LCCN QA76 .S979 2003. URL http:/

/ieeexplore.ieee.org/iel5/

                                        8767/27770/01238203.pdf?isnumber=
                                        27770&prod=CNF&arnumber=1238203&
                                        arSt=+290&ared=+297&arAuthor=
                                        Khot%2C+S.;   http://ieeexplore.
                                        ieee.org/xpls/abs_all.jsp?
                                        isnumber=27770&arnumber=1238203&
                                        count=66&index=30.  IEEE

Computer Society Order Num-

ber PR02040.

Khot:2004:HAS

[Kho04a] S. Khot. Hardness of approximating the shortest vector problem in lattices. In IEEE [IEE04], pages 126–135. CODEN ASFPDV. ISBN 0-7695-2228-9. ISSN 0272-5428. LCCN QA276. URL http:/

                                        /ieeexplore.ieee.org/iel5/
                                        9430/29918/01366232.pdf?isnumber=
                                        29918&prod=CNF&arnumber=1366232&
                                        arSt=+126&ared=+135&arAuthor=
                                        Khot%2C+S.;   http://ieeexplore.
                                        ieee.org/xpls/abs_all.jsp?
                                        isnumber=29918&arnumber=1366232&
                                        count=64&index=13.  IEEE

Computer Society Order Num-

ber P2228.

Khot:2004:RPG

[Kho04b] S. Khot. Ruling out PTAS for graph min-bisection, densest subgraph and bipartite clique. In IEEE [IEE04], pages 136–

145. CODEN ASFPDV. ISBN 0-7695-2228-9. ISSN 0272-5428. LCCN QA276. URL http:/

                                        /ieeexplore.ieee.org/iel5/
                                        9430/29918/01366233.pdf?isnumber=
                                        29918&prod=CNF&arnumber=1366233&
                                        arSt=+136&ared=+145&arAuthor=
                                        Khot%2C+S.;   http://ieeexplore.
                                        ieee.org/xpls/abs_all.jsp?
                                        isnumber=29918&arnumber=1366233&

count=64&index=14.  IEEE

Computer Society Order Number P2228.

Khot:2005:UGCa

[Kho05] Subhash Khot. On the unique games conjecture. In IEEE [IEE05], page ?? ISBN 0-7695-2468-0, 0-7695-2362-5. ISSN 0272-5428. LCCN QA76 .S979 2005. IEEE Computer Society order number P2468.

Kumar:2000:FMR

[KK00] A. Kumar and J. Kleinberg.

Fairness measures for resource allocation. In IEEE [IEE00], pages 75–85. CODEN ASF-PDV. ISBN 0-7695-0850-2, 0-7695-0851-0 (case), 0-7695-0852-9 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 2000. IEEE Computer Society Order Number PR00850.

Kempe:2002:PIR

[KK02] D. Kempe and J. Kleinberg.

Protocols and impossibility results for gossip-based communication mechanisms. In IEEE [IEE02], pages 471–480. CO-DEN ASFPDV. ISBN 0-7695-1822-2. ISSN 0272-5428. LCCN QA267. URL http:/

/ieeexplore.ieee.org/iel5/
8411/26517/01181971.pdf?isnumber=
26517&prod=CNF&arnumber=1181971&
arSt=+471&ared=+480&arAuthor=
Kempe%2C+D.%3B+Kleinberg%
2C+J.;   http://ieeexplore.
ieee.org/xpls/abs_all.jsp?
isnumber=26517&arnumber=1181971&
count=82&index=47.  IEEE

Computer Society Order Number PR01822.

Karakostas:2004:EPM

[KK04] G. Karakostas and S. G. Kolliopoulos. Edge pricing of multicommodity networks for heterogeneous selfish users. In IEEE [IEE04], pages 268–276. CODEN ASFPDV. ISBN 0-7695-2228-9. ISSN 0272-5428. LCCN QA276. URL http:/

/ieeexplore.ieee.org/iel5/
9430/29918/01366246.pdf?isnumber=
29918&prod=CNF&arnumber=1366246&
arSt=+268&ared=+276&arAuthor=
Karakostas%2C+G.%3B+Kolliopoulos%
2C+S.G.;   http://ieeexplore.
ieee.org/xpls/abs_all.jsp?
isnumber=29918&arnumber=1366246&
count=64&index=27.  IEEE

Computer Society Order Number P2228.

Karlin:2005:BVF

[KK05] A. R. Karlin and D. Kempe. Beyond VCG: frugality of truthful mechanisms. In IEEE [IEE05], pages 615–624. ISBN 0-7695-2468-0, 0-7695-2362-5. ISSN 0272-5428. LCCN QA76 .S979 2005. IEEE Computer Society order number P2468.

Kahn:2000:CTB

[KKLV00] J. Kahn, J. H. Kim, L. Lovasz, and V. H. Vu. The cover time, the blanket time, and the Matthews bound. In IEEE [IEE00], pages 467–475. CO-DEN ASFPDV. ISBN 0-7695-0850-2, 0-7695-0851-0 (case), 0-7695-0852-9 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 2000. IEEE Computer Society Order Number PR00850.

Kempe:2008:EGH

[KKM^{+}08] J. Kempe, H. Kobayashi, K. Matsumoto, B. Toner, and T. Vidick. Entangled games are hard to approximate. In IEEE [IEE08], pages 447–

456. ISBN 0-7695-3436-8. ISSN 0272-5428. LCCN QA76 .S95 2008. URL http://

ieeexplore.ieee.org/servlet/
opac?punumber=4690923. IEEE

Computer Society order number P3436.

Khot:2004:OIR

[KKMO04] S. Khot, G. Kindler, E. Mossel, and R. O’Donnell. Optimal inapproximability results for MAX-CUT and other 2variable CSPs? In IEEE [IEE04], pages 146–154. CO-DEN ASFPDV. ISBN 0-7695-2228-9. ISSN 0272-5428. LCCN QA276. URL http:/

/ieeexplore.ieee.org/iel5/
9430/29918/01366234.pdf?isnumber=
29918&prod=CNF&arnumber=1366234&
arSt=+146&ared=+154&arAuthor=
Khot%2C+S.%3B+Kindler%2C+
G.%3B+Mossel%2C+E.%3B+O%27Donnell%
2C+R.;   http://ieeexplore.
ieee.org/xpls/abs_all.jsp?
isnumber=29918&arnumber=1366234&
count=64&index=15.  IEEE

Computer Society Order Number P2228.

Kalai:2005:ALH

[KKMS05] A. T. Kalai, A. R. Klivans, Yishay Mansour, and R. A. Servedio. Agnostically learning halfspaces. In IEEE [IEE05], pages 11–20. ISBN 0-7695-2468-0, 0-7695-2362-5. ISSN 0272-5428. LCCN QA76 .S979 2005. IEEE Computer Society order number P2468.

Karp:2000:OPC

[KKPS00] R. Karp, E. Koutsoupias, C. Papadimitriou, and S. Shenker. Optimization problems in congestion control. In IEEE [IEE00], pages 66–74. CODEN ASFPDV. ISBN 0-7695-0850-2, 0-7695-0851-0 (case), 0-7695-0852-9 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 2000. IEEE Computer Society Order Number PR00850.

Kleinberg:2003:VKD

[KL03] R. Kleinberg and T. Leighton. The value of knowing a demand curve: bounds on regret for online posted-price auctions. In IEEE [IEE03], pages 594–605. CODEN ASF-PDV. ISBN 0-7695-2040-5. ISSN 0272-5428. LCCN QA76 .S979 2003. URL http:/

                                        /ieeexplore.ieee.org/iel5/
                                        8767/27770/01238232.pdf?isnumber=
                                        27770&prod=CNF&arnumber=1238232&
                                        arSt=+594&ared=+605&arAuthor=
                                        Kleinberg%2C+R.%3B+Leighton%
                                        2C+T.;   http://ieeexplore.
                                        ieee.org/xpls/abs_all.jsp?
                                        isnumber=27770&arnumber=1238232&
                                        count=66&index=59.  IEEE

Computer Society Order Num-

ber PR02040.

Kaufman:2005:AOL

[KL05] T. Kaufman and S. Litsyn. Almost orthogonal linear codes are locally testable. In IEEE [IEE05], pages 317–326. ISBN 0-7695-2468-0, 0-7695-2362-5. ISSN 0272-5428. LCCN QA76 .S979 2005. IEEE Computer Society order number P2468.

Kiraly:2006:AMM

[KL06a] Tamas Kiraly and Lap Chi Lau. Approximate min-max theorems of Steiner rootedorientations of hypergraphs. In IEEE [IEE06], pages 283–292. ISBN 0-7695-2720-5, 0-7695-2362-5. ISSN 0272-5428. LCCN QA76 .S974 2006. URL http://

ieeexplore.ieee.org/servlet/
opac?punumber=4031329. IEEE

Computer Society Order Number P2720.

Krauthgamer:2006:ANC

[KL06b] Robert Krauthgamer and James R.

Lee. Algorithms on negatively curved spaces. In IEEE [IEE06], pages 119–132. ISBN 0-7695-2720-5, 0-7695-2362-5. ISSN 0272-5428. LCCN QA76 .S974 2006. URL http://

ieeexplore.ieee.org/servlet/
opac?punumber=4031329. IEEE

Computer Society Order Number P2720.

Kaufman:2008:WCA

[KL08] T. Kaufman and S. Lovett.

Worst case to average case reductions for polynomials. In IEEE [IEE08], pages 166– 175. ISBN 0-7695-3436-8. ISSN 0272-5428. LCCN QA76 .S95 2008. URL http://

ieeexplore.ieee.org/servlet/

                                        opac?punumber=4690923. IEEE

Computer Society order number P3436.

Klauck:2001:LBQ

[Kla01] H. Klauck. Lower bounds for quantum communication complexity. In IEEE [IEE01], pages 288–297. CODEN ASF-PDV. ISBN 0-7695-1390-5, 0-7695-1391-3 (case), 0-7695-1392-1 (microfiche). ISSN 0272-5428. LCCN ????

Kleinberg:2000:DNF

[Kle00] J. Kleinberg. Detecting a network failure. In IEEE [IEE00], pages 231–239. CODEN ASF-PDV. ISBN 0-7695-0850-2, 0-7695-0851-0 (case), 0-7695-0852-9 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 2000. IEEE Computer Society Order Number PR00850.

Klein:2005:LTA

[Kle05a] P. N. Klein. A linear-time approximation scheme for planar weighted TSP. In IEEE [IEE05], pages 647–656. ISBN 0-7695-2468-0, 0-7695-2362-5. ISSN 0272-5428. LCCN QA76 .S979 2005. IEEE Computer Society order number P2468.

Kleinberg:2005:AAD

[Kle05b] J. Kleinberg. An approximation algorithm for the disjoint paths problem in even-degree planar graphs. In IEEE [IEE05], pages 627–636. ISBN 0-7695-2468-0, 0-7695-2362-5. ISSN 0272-5428. LCCN QA76 .S979 2005. IEEE Computer Society order number P2468.

Krauthgamer:2004:MDN

[KLMN04] R. Krauthgamer, J. R. Lee, M. Mendel, and A. Naor. Measured descent: a new embedding method for finite metrics. In IEEE [IEE04], pages 434–443. CODEN ASFPDV. ISBN 0-7695-2228-9. ISSN 0272-5428. LCCN QA276. URL http:/

/ieeexplore.ieee.org/iel5/
9430/29918/01366263.pdf?isnumber=
29918&prod=CNF&arnumber=1366263&
arSt=+434&ared=+443&arAuthor=
Krauthgamer%2C+R.%3B+Lee%
2C+J.R.%3B+Mendel%2C+M.%3B+
Naor%2C+A.;   http://ieeexplore.
ieee.org/xpls/abs_all.jsp?
isnumber=29918&arnumber=1366263&
count=64&index=44.  IEEE

Computer Society Order Number P2228.

Kasiviswanathan:2008:WCW

[KLN^{+}08] S. P. Kasiviswanathan, H. K.

Lee, K. Nissim, S. Raskhodnikova, and A. Smith. What can we learn privately? In IEEE [IEE08], pages 531–

540. ISBN 0-7695-3436-8. ISSN 0272-5428. LCCN QA76 .S95 2008. URL http://

ieeexplore.ieee.org/servlet/
opac?punumber=4690923. IEEE

Computer Society order number P3436.

Kelner:2009:HEG

[KLPT09] J. A. Kelner, J. R. Lee, G. N.

Price, and Shang-Hua Teng. Higher eigenvalues of graphs. In IEEE [IEE09], pages 735–744.

ISBN 0-7695-3850-9. ISSN 0272-5428. LCCN QA76 .S95 2009. IEEE Computer Society order number P3850.

Kalai:2009:SEU

[KLR09] Y. T. Kalai, Xin Li, and A. Rao. 2-source extractors under computational assumptions and cryptography with defective randomness. In IEEE [IEE09], pages 617–626. ISBN 0-7695-3850-9. ISSN 0272-5428. LCCN QA76 .S95 2009. IEEE Computer Society order number P3850.

Kalai:2008:NEP

[KLRZ08] Y. T. Kalai, Xin Li, A. Rao, and D. Zuckerman. Network extractor protocols. In IEEE [IEE08], pages 654–663. ISBN 0-7695-3436-8. ISSN 0272-5428. LCCN QA76 .S95 2008. URL http://

                                        ieeexplore.ieee.org/servlet/
                                        opac?punumber=4690923. IEEE

Computer Society order number

P3436.

Kaplan:2003:AAA

[KLSS03] H. Kaplan, M. Lewenstein, N. Shafrir, and M. Sviridenko. Approximation algorithms for asymmetric TSP by decomposing directed regular multigraphs. In IEEE [IEE03], pages 56–65. CODEN ASF-PDV. ISBN 0-7695-2040-5. ISSN 0272-5428. LCCN QA76 .S979 2003. URL http:/

                                        /ieeexplore.ieee.org/iel5/
                                        8767/27770/01238181.pdf?isnumber=
                                        27770&prod=CNF&arnumber=1238181&
                                        arSt=+56&ared=+65&arAuthor=

Kaplan%2C+H.%3B+Lewenstein%
2C+M.%3B+Shafrir%2C+N.%3B+
Sviridenko%2C+M.;   http:/
/ieeexplore.ieee.org/xpls/
abs_all.jsp?isnumber=27770&
arnumber=1238181&count=66&

index=8. IEEE Computer Society Order Number PR02040.

Karger:2000:BST

[KM00a] D. R. Karger and M. Minkoff.

Building Steiner trees with incomplete global knowledge. In IEEE [IEE00], pages 613– 623. CODEN ASFPDV. ISBN 0-7695-0850-2, 0-7695-0851-0 (case), 0-7695-0852-9 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 2000. IEEE Computer Society Order Number PR00850.

Kenyon:2000:LWB

[KM00b] C. Kenyon and M. Mitzenmacher. Linear waste of best fit bin packing on skewed distributions. In IEEE [IEE00], pages 582–589. CODEN ASF-PDV. ISBN 0-7695-0850-2, 0-7695-0851-0 (case), 0-7695-0852-9 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 2000. IEEE Computer Society Order Number PR00850.

Kumar:2005:AAS

[KM05] V. S. Anil Kumar and M. V.

Marathe. Approximation algorithms for scheduling on multiple machines. In IEEE [IEE05], pages 254–263. ISBN 0-7695-2468-0, 0-7695-2362-5. ISSN 0272-5428. LCCN QA76 .S979 2005. IEEE Computer Society order number P2468.

Kelner:2009:FGR

[KM09] J. A. Kelner and A. Madry. Faster generation of random spanning trees. In IEEE [IEE09], pages 13–21. ISBN 0-7695-3850-9. ISSN 0272-5428. LCCN QA76 .S95 2009. IEEE Computer Society order number P3850.

Kenyon:2001:GDT

[KMP01] C. Kenyon, E. Mossel, and Y. Peres. Glauber dynamics on trees and hyperbolic graphs. In IEEE [IEE01], pages 568–578. CODEN ASF-PDV. ISBN 0-7695-1390-5, 0-7695-1391-3 (case), 0-7695-1392-1 (microfiche). ISSN 0272-5428. LCCN ????

Kim:2007:NOB

[KMT07] Jeong Han Kim, R. Montenegro, and P. Tetali. Near optimal bounds for collision in Pollard rho for discrete log. In IEEE [IEE07], pages 215– 223. ISBN 0-7695-3010-9. ISSN 0272-5428. LCCN QA76 .S974 2007. URL http://

                                        ieeexplore.ieee.org/servlet/
                                        opac?punumber=4389466. IEEE

Computer Society order number P3010.

Khot:2005:NTF

[KN05] Subhash Khot and Assaf Naor. Nonembeddability theorems via Fourier analysis. In IEEE [IEE05], pages 101–110. ISBN 0-7695-2468-0, 0-7695-2362-5.

ISSN 0272-5428. LCCN QA76 .S979 2005. IEEE Computer Society order number P2468.

Kelner:2007:HSC

[KN07a] J. A. Kelner and E. Nikolova.

On the hardness and smoothed complexity of quasi-concave minimization. In IEEE [IEE07], pages 472–482. ISBN 0-7695-3010-9. ISSN 0272-5428. LCCN QA76 .S974 2007. URL http://

ieeexplore.ieee.org/servlet/
opac?punumber=4389466. IEEE

Computer Society order number P3010.

Khot:2007:LEM

[KN07b] S. Khot and A. Naor. Linear equations modulo 2 and the L_{1} diameter of convex bodies. In IEEE [IEE07], pages 318–328. ISBN 0-7695-3010-9. ISSN 0272-5428. LCCN QA76 .S974 2007. URL http://

ieeexplore.ieee.org/servlet/
opac?punumber=4389466. IEEE

Computer Society order number P3010.

Khot:2008:AKC

[KN08] S. Khot and A. Naor. Approximate kernel clustering. In IEEE [IEE08], pages 561– 570. ISBN 0-7695-3436-8. ISSN 0272-5428. LCCN QA76 .S95 2008. URL http://

ieeexplore.ieee.org/servlet/
opac?punumber=4690923. IEEE

Computer Society order number P3436.

Klivans:2002:LIT

[KO02] A. R. Klivans and R. O’Donnell. Learning intersections and thresholds of halfspaces. In IEEE [IEE02], pages 177–186. CODEN ASFPDV. ISBN 0-7695-1822-2. ISSN 0272-5428. LCCN QA267. URL http:/

/ieeexplore.ieee.org/iel5/
8411/26517/01181894.pdf?isnumber=
26517&prod=CNF&arnumber=1181894&
arSt=+177&ared=+186&arAuthor=
Klivans%2C+A.R.%3B+O%27Donnell%
2C+R.;   http://ieeexplore.
ieee.org/xpls/abs_all.jsp?
isnumber=26517&arnumber=1181894&
count=82&index=19.  IEEE

Computer Society Order Number PR01822.

Khot:2006:SGU

[KO06] Subhash Khot and Ryan O’Donnell. SDP gaps and UGC-hardness for MAXCUTGAIN. In IEEE [IEE06], pages 217–226. ISBN 0-7695-2720-5, 0-7695-2362-5. ISSN 0272-5428. LCCN QA76 .S974 2006. URL http://

ieeexplore.ieee.org/servlet/
opac?punumber=4031329. IEEE

Computer Society Order Number P2720.

Koivisto:2006:AGC

[Koi06] Mikko Koivisto. An O ∗ (2^{n}) algorithm for graph coloring and other partitioning problems via inclusion-exclusion. In IEEE [IEE06], pages 583–590. ISBN 0-7695-2720-5, 0-7695-2362-5. ISSN 0272-5428. LCCN QA76 .S974 2006. URL http://

ieeexplore.ieee.org/servlet/

opac?punumber=4031329. IEEE

Computer Society Order Number P2720.

Koltun:2001:ATU

[Kol01] V. Koltun. Almost tight upper bounds for vertical decompositions in four dimensions. In IEEE [IEE01], pages 56–65. CO-DEN ASFPDV. ISBN 0-7695-1390-5, 0-7695-1391-3 (case), 0-7695-1392-1 (microfiche). ISSN 0272-5428. LCCN ????

Kindler:2008:SCR

[KORW08] G. Kindler, R. O’Donnell, A. Rao, and A. Wigderson. Spherical cubes and rounding in high dimensions. In IEEE [IEE08], pages 189– 198. ISBN 0-7695-3436-8. ISSN 0272-5428. LCCN QA76 .S95 2008. URL http://

ieeexplore.ieee.org/servlet/
opac?punumber=4690923. IEEE

Computer Society order number P3436.

Klivans:2008:LGC

[KOS08] A. R. Klivans, R. O’Donnell, and R. A. Servedio. Learning geometric concepts via Gaussian surface area. In IEEE [IEE08], pages 541–550. ISBN 0-7695-3436-8. ISSN 0272-5428. LCCN QA76 .S95 2008. URL http://

ieeexplore.ieee.org/servlet/
opac?punumber=4690923. IEEE

Computer Society order number P3436.

Kowalski:2002:DBT

[KP02] D. R. Kowalski and A. Pelc.

Deterministic broadcasting time in radio networks of unknown topology. In IEEE [IEE02], pages 63–72. CO-DEN ASFPDV. ISBN 0-7695-1822-2. ISSN 0272-5428. LCCN QA267. URL http:/

/ieeexplore.ieee.org/iel5/
8411/26517/01181883.pdf?isnumber=
26517&prod=CNF&arnumber=1181883&
arSt=+63&ared=+72&arAuthor=
Kowalski%2C+D.R.%3B+Pelc%
2C+A.;   http://ieeexplore.
ieee.org/xpls/abs_all.jsp?
isnumber=26517&arnumber=1181883&
count=82&index=8. IEEE Com-

puter Society Order Number PR01822.

Kintali:2009:RAF

[KPR^{+}09] S. Kintali, L. J. Poplawski, R. Rajaraman, R. Sundaram, and Shang-Hua Teng. Reducibility among fractional stability problems. In IEEE [IEE09], pages 283–292. ISBN 0-7695-3850-9. ISSN 0272-5428. LCCN QA76 .S95 2009. IEEE Computer Society order number P3850.

Kale:2008:NTE

[KPS08] S. Kale, Y. Peres, and C. Seshadhri. Noise tolerance of expanders and sublinear expander reconstruction. In IEEE [IEE08], pages 719–

728. ISBN 0-7695-3436-8. ISSN 0272-5428. LCCN QA76 .S95 2008. URL http://

ieeexplore.ieee.org/servlet/
opac?punumber=4690923. IEEE

Computer Society order number P3436.

Kaufman:2004:TPG

[KR04] T. Kaufman and D. Ron.

Testing polynomials over general fields. In IEEE [IEE04], pages 413–422. CODEN ASFPDV. ISBN 0-7695-2228-9. ISSN 0272-5428. LCCN QA276. URL http:/

/ieeexplore.ieee.org/iel5/
9430/29918/01366261.pdf?isnumber=
29918&prod=CNF&arnumber=1366261&
arSt=+413&ared=+422&arAuthor=
Kaufman%2C+T.%3B+Ron%2C+D.
;   http://ieeexplore.ieee.
org/xpls/abs_all.jsp?isnumber=
29918&arnumber=1366261&count=

64&index=42. IEEE Computer Society Order Number P2228.

Kleinberg:2005:QIN

[KR05] J. Kleinberg and Prabhakar Raghavan. Query incentive networks. In IEEE [IEE05], pages 132–141. ISBN 0-7695-2468-0, 0-7695-2362-5. ISSN 0272-5428. LCCN QA76 .S979 2005. IEEE Computer Society order number P2468.

Kalai:2006:SNI

[KR06] Yael Tauman Kalai and Ran Raz. Succinct non-interactive zero-knowledge proofs with preprocessing for LOGSNP. In IEEE [IEE06], pages 355–366. ISBN 0-7695-2720-5, 0-7695-2362-5. ISSN 0272-5428. LCCN QA76 .S974 2006. URL http://

ieeexplore.ieee.org/servlet/
opac?punumber=4031329. IEEE

Computer Society Order Number P2720.

Kumar:2000:SMW

[KRR^{+}00] R. Kumar, P. Raghavan, S. Rajagopalan, D. Sivakumar, A. Tomkins, and E. Upfal. Stochastic models for the Web graph. In IEEE [IEE00], pages 57–65. CODEN ASF-PDV. ISBN 0-7695-0850-2, 0-7695-0851-0 (case), 0-7695-0852-9 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 2000. IEEE Computer Society Order Number PR00850.

Kempe:2008:UGE

[KRT08] J. Kempe, O. Regev, and B. Toner. Unique games with entangled provers are easy. In IEEE [IEE08], pages 457–466. ISBN 0-7695-3436-

8. ISSN 0272-5428. LCCN QA76 .S95 2008. URL http://

ieeexplore.ieee.org/servlet/
opac?punumber=4690923. IEEE

Computer Society order number P3436.

Koltun:2002:PTO

[KS02] V. Koltun and M. Sharir. The partition technique for overlays of envelopes. In IEEE [IEE02], pages 637–646. CO-DEN ASFPDV. ISBN 0-7695-1822-2. ISSN 0272-5428. LCCN QA267. URL http:/

/ieeexplore.ieee.org/iel5/
8411/26517/01181989.pdf?isnumber=
26517&prod=CNF&arnumber=1181989&
arSt=+637&ared=+646&arAuthor=
Koltun%2C+V.%3B+Sharir%2C+
M.;   http://ieeexplore.ieee.
org/xpls/abs_all.jsp?isnumber=
26517&arnumber=1181989&count=

82&index=63.  IEEE Com-

puter Society Order Number PR01822.

Klivans:2006:CHL

[KS06] Adam R. Klivans and Alexander A. Sherstov. Cryptographic hardness for learning intersections of halfspaces. In IEEE [IEE06], pages 553–562. ISBN 0-7695-2720-5, 0-7695-2362-5. ISSN 0272-5428. LCCN QA76 .S974 2006. URL http://

ieeexplore.ieee.org/servlet/
opac?punumber=4031329. IEEE

Computer Society Order Number P2720.

Kaufman:2007:SRL

[KS07] T. Kaufman and M. Sudan.

Sparse random linear codes are locally decodable and testable. In IEEE [IEE07], pages 590– 600. ISBN 0-7695-3010-9. ISSN 0272-5428. LCCN QA76 .S974 2007. URL http://

ieeexplore.ieee.org/servlet/
opac?punumber=4389466. IEEE

Computer Society order number P3010.

Khot:2008:HML

[KS08] S. Khot and R. Saket. Hardness of minimizing and learning DNF expressions. In IEEE [IEE08], pages 231–240. ISBN 0-7695-3436-8. ISSN 0272-5428. LCCN QA76 .S95 2008. URL http://

ieeexplore.ieee.org/servlet/
opac?punumber=4690923. IEEE

Computer Society order number P3436.

Kayal:2009:BPI

[KS09a] N. Kayal and S. Saraf. Blackbox polynomial identity testing for depth 3 circuits. In IEEE [IEE09], pages 198–207. ISBN 0-7695-3850-9. ISSN 0272-5428. LCCN QA76 .S95 2009. IEEE Computer Society order number P3850.

Khot:2009:SIG

[KS09b] S. Khot and R. Saket. SDP integrality gaps with local ℓ_{1}- embeddability. In IEEE [IEE09], pages 565–574. ISBN 0-7695-3850-9. ISSN 0272-5428. LCCN QA76 .S95 2009. IEEE Computer Society order number P3850.

Klauck:2004:QCS

[KSdW04] H. Klauck, R. Spalek, and R. de Wolf. Quantum and classical strong direct product theorems and optimal timespace tradeoffs. In IEEE [IEE04], pages 12–21. CO-DEN ASFPDV. ISBN 0-7695-2228-9. ISSN 0272-5428. LCCN QA276. URL http:/

                                        /ieeexplore.ieee.org/iel5/
                                        9430/29918/01366220.pdf?isnumber=
                                        29918&prod=CNF&arnumber=1366220&
                                        arSt=+12&ared=+21&arAuthor=
                                        Klauck%2C+H.%3B+Spalek%2C+
                                        R.%3B+de+Wolf%2C+R.;   http:
                                        //ieeexplore.ieee.org/xpls/
                                        abs_all.jsp?isnumber=29918&
                                        arnumber=1366220&count=64&

index=1. IEEE Computer Society Order Number P2228.

Kumar:2004:SLT

[KSS04] A. Kumar, Y. Sabharwal, and S. Sen. A simple linear time (1 + ϵ)-approximation algorithm for k-means clustering in any dimensions. In IEEE [IEE04], pages 454–462. CODEN ASFPDV. ISBN 0-7695-2228-9. ISSN 0272-5428. LCCN QA276. URL http:/

/ieeexplore.ieee.org/iel5/
9430/29918/01366265.pdf?isnumber=
29918&prod=CNF&arnumber=1366265&
arSt=+454&ared=+462&arAuthor=
Kumar%2C+A.%3B+Sabharwal%
2C+Y.%3B+Sen%2C+S.;   http:
//ieeexplore.ieee.org/xpls/
abs_all.jsp?isnumber=29918&
arnumber=1366265&count=64&

index=46. IEEE Computer Society Order Number P2228.

Karp:2000:RRS

[KSSV00] R. Karp, C. Schindelhauert, S. Shenkert, and B. Vocking. Randomized rumor spreading. In IEEE [IEE00], pages 565–574. CODEN ASFPDV. ISBN 0-7695-0850-2, 0-7695-0851-0 (case), 0-7695-0852-9 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 2000. IEEE Computer Society Order Number PR00850.

King:2006:TSS

[KSSV06] Valerie King, Jared Saia, Vishal Sanwalani, and Erik Vee. Towards secure and scalable computation in peer-to-peer networks. In IEEE [IEE06], pages 87–98. ISBN 0-7695-2720-5, 0-7695-2362-5. ISSN 0272-5428. LCCN QA76 .S974 2006. URL http://

                                        ieeexplore.ieee.org/servlet/
                                        opac?punumber=4031329. IEEE

Computer Society Order Num-

ber P2720.

Kalai:2009:LSA

[KST09] A. T. Kalai, A. Samorodnitsky, and Shang-Hua Teng. Learning and smoothed analysis. In IEEE [IEE09], pages 395–404. ISBN 0-7695-3850-9. ISSN 0272-5428. LCCN QA76 .S95 2009. IEEE Computer Society order number P3850.

Kleinberg:2004:TEU

[KSW04] J. Kleinberg, A. Slivkins, and T. Wexler. Triangulation and embedding using small sets of beacons. In IEEE [IEE04], pages 444–453. CO-DEN ASFPDV. ISBN 0-7695-2228-9. ISSN 0272-5428. LCCN QA276. URL http:/

                                        /ieeexplore.ieee.org/iel5/
                                        9430/29918/01366264.pdf?isnumber=
                                        29918&prod=CNF&arnumber=1366264&
                                        arSt=+444&ared=+453&arAuthor=
                                        Kleinberg%2C+J.%3B+Slivkins%
                                        2C+A.%3B+Wexler%2C+T.;   http:
                                        //ieeexplore.ieee.org/xpls/
                                        abs_all.jsp?isnumber=29918&
                                        arnumber=1366264&count=64&

index=45. IEEE Computer Society Order Number P2228.

Kedlaya:2008:FMC

[KU08] K. S. Kedlaya and C. Umans. Fast modular composition in any characteristic. In IEEE [IEE08], pages 146–155. ISBN 0-7695-3436-8. ISSN 0272-5428.

LCCN QA76 .S95 2008. URL

http://ieeexplore.ieee.org/
servlet/opac?punumber=4690923.

IEEE Computer Society order number P3436.

Kalyanaraman:2009:CRN

[KU09] S. Kalyanaraman and C. Umans.

The complexity of rationalizing network formation. In IEEE [IEE09], pages 485–494. ISBN 0-7695-3850-9. ISSN 0272-5428. LCCN QA76 .S95 2009. IEEE Computer Society order number P3850.

Kalai:2000:EAU

[KV00] A. Kalai and S. Vempala. Efficient algorithms for universal portfolios. In IEEE [IEE00], pages 486–491. CODEN ASF-PDV. ISBN 0-7695-0850-2, 0-7695-0851-0 (case), 0-7695-0852-9 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 2000. IEEE Computer Society Order Number PR00850.

Khot:2005:UGCb

[KV05a] S. A. Khot and N. K. Vishnoi. The unique games conjecture, integrality gap for cut problems and embeddability of negative type metrics into l_{1}. In IEEE [IEE05], pages 53– 62. ISBN 0-7695-2468-0, 0-7695-2362-5. ISSN 0272-5428. LCCN QA76 .S979 2005. IEEE Computer Society order number P2468.

Kupferman:2005:SDP

[KV05b] O. Kupferman and M. Y. Vardi.

Safraless decision procedures.

In IEEE [IEE05], pages 531– 540. ISBN 0-7695-2468-0, 0-7695-2362-5. ISSN 0272-5428. LCCN QA76 .S979 2005. IEEE Computer Society order number P2468.

Kannan:2000:CGB

[KVV00] R. Kannan, S. Vempala, and A. Vetta. On clusterings — good, bad and spectral. In IEEE [IEE00], pages 367–377. CO-DEN ASFPDV. ISBN 0-7695-0850-2, 0-7695-0851-0 (case), 0-7695-0852-9 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 2000. IEEE Computer Society Order Number PR00850.

Kushilevitz:2009:CCS

[KW09] E. Kushilevitz and E. Weinreb. The communication complexity of set-disjointness with small sets and 0–1 intersection. In IEEE [IEE09], pages 63–72. ISBN 0-7695-3850-9. ISSN 0272-5428. LCCN QA76 .S95 2009. IEEE Computer Society order number P3850.

Kolliopoulos:2001:TAR

[KY01] S. G. Kolliopoulos and N. E. Young. Tight approximation results for general covering integer programs. In IEEE [IEE01], pages 522–528. CO-DEN ASFPDV. ISBN 0-7695-1390-5, 0-7695-1391-3 (case), 0-7695-1392-1 (microfiche). ISSN 0272-5428. LCCN ????

Koufogiannakis:2007:BSF

[KY07] C. Koufogiannakis and N. E. Young. Beating simplex for fractional packing and covering linear programs. In IEEE [IEE07], pages 494–504. ISBN 0-7695-3010-9. ISSN 0272-5428. LCCN QA76 .S974 2007. URL http://

ieeexplore.ieee.org/servlet/
opac?punumber=4389466. IEEE

Computer Society order number P3010.

Kamp:2003:DEB

[KZ03] J. Kamp and D. Zuckerman. Deterministic extractors for bitfixing sources and exposureresilient cryptography. In IEEE [IEE03], pages 92–101. CO-DEN ASFPDV. ISBN 0-7695-2040-5. ISSN 0272-5428. LCCN QA76 .S979 2003. URL http:/

/ieeexplore.ieee.org/iel5/
8767/27770/01238184.pdf?isnumber=
27770&prod=CNF&arnumber=1238184&
arSt=+92&ared=+101&arAuthor=
Kamp%2C+J.%3B+Zuckerman%2C+
D.;   http://ieeexplore.ieee.
org/xpls/abs_all.jsp?isnumber=
27770&arnumber=1238184&count=
66&index=11.  IEEE Com-

puter Society Order Number PR02040.

Langberg:2004:PCS

[Lan04] M. Langberg. Private codes or succinct random codes that are (almost) perfect. In IEEE [IEE04], pages 325–334. CO-DEN ASFPDV. ISBN 0-7695-2228-9. ISSN 0272-5428. LCCN QA276. URL http:/

/ieeexplore.ieee.org/iel5/
9430/29918/01366252.pdf?isnumber=
29918&prod=CNF&arnumber=1366252&
arSt=+325&ared=+334&arAuthor=
Langberg%2C+M.;   http://

                                        ieeexplore.ieee.org/xpls/
                                        abs_all.jsp?isnumber=29918&
                                        arnumber=1366252&count=64&

index=33. IEEE Computer Society Order Number P2228.

Lau:2004:AMS

[Lau04] Lap Chi Lau. An approximate max-Steiner-tree-packing min-Steiner-cut theorem. In IEEE [IEE04], pages 61–70. CODEN ASFPDV. ISBN 0-7695-2228-9. ISSN 0272-5428. LCCN QA276. URL http:/

                                        /ieeexplore.ieee.org/iel5/
                                        9430/29918/01366225.pdf?isnumber=
                                        29918&prod=CNF&arnumber=1366225&
                                        arSt=+61&ared=+70&arAuthor=
                                        Lap+Chi+Lau;   http://ieeexplore.
                                        ieee.org/xpls/abs_all.jsp?
                                        isnumber=29918&arnumber=1366225&
                                        count=64&index=6. IEEE Com-

puter Society Order Number

P2228.

Levin:2002:FI

[Lev02] L. A. Levin. Forbidden information. In IEEE [IEE02], pages 761–765. CODEN ASFPDV. ISBN 0-7695-1822-2. ISSN 0272-5428. LCCN QA267. URL http:/

                                        /ieeexplore.ieee.org/iel5/
                                        8411/26517/01182001.pdf?isnumber=
                                        26517&prod=CNF&arnumber=1182001&
                                        arSt=+761&ared=+765&arAuthor=
                                        Levin%2C+L.A.;   http://ieeexplore.
                                        ieee.org/xpls/abs_all.jsp?
                                        isnumber=26517&arnumber=1182001&
                                        count=82&index=75.  IEEE

Computer Society Order Num-

ber PR01822.

Lindell:2003:GCU

[Lin03] Y. Lindell. General composition and universal composability in secure multi-party computation. In IEEE [IEE03], pages 394–403. CODEN ASF-PDV. ISBN 0-7695-2040-5. ISSN 0272-5428. LCCN QA76 .S979 2003. URL http:/

/ieeexplore.ieee.org/iel5/
8767/27770/01238213.pdf?isnumber=
27770&prod=CNF&arnumber=1238213&
arSt=+394&ared=+403&arAuthor=
Lindell%2C+Y.;   http://ieeexplore.
ieee.org/xpls/abs_all.jsp?
isnumber=27770&arnumber=1238213&
count=66&index=40.  IEEE

Computer Society Order Number PR02040.

Lenzen:2008:CSB

[LLW08] C. Lenzen, T. Locher, and R. Wattenhofer. Clock synchronization with bounded global and local skew. In IEEE [IEE08], pages 509–

518. ISBN 0-7695-3436-8. ISSN 0272-5428. LCCN QA76 .S95 2008. URL http://

ieeexplore.ieee.org/servlet/
opac?punumber=4690923. IEEE

Computer Society order number P3436.

Lavi:2003:TCT

[LMN03] R. Lavi, Ahuva Mu’alem, and N. Nisan. Towards a characterization of truthful combinatorial auctions. In IEEE [IEE03], pages 574–583. CO-DEN ASFPDV. ISBN 0-7695-2040-5. ISSN 0272-5428. LCCN QA76 .S979 2003. URL http:/

                                        /ieeexplore.ieee.org/iel5/
                                        8767/27770/01238230.pdf?isnumber=
                                        27770&prod=CNF&arnumber=1238230&
                                        arSt=+574&ared=+583&arAuthor=
                                        Lavi%2C+R.%3B+Ahuva+Mu%27alem%
                                        3B+Nisan%2C+N.;   http://
                                        ieeexplore.ieee.org/xpls/
                                        abs_all.jsp?isnumber=27770&
                                        arnumber=1238230&count=66&

index=57. IEEE Computer Society Order Number PR02040.

Lee:2006:MHG

[LN06] James R. Lee and Assaf Naor. L_{p} metrics on the Heisenberg group and the Goemans– Linial conjecture. In IEEE [IEE06], pages 99–108. ISBN 0-7695-2720-5, 0-7695-2362-5. ISSN 0272-5428. LCCN QA76 .S974 2006. URL http://

                                        ieeexplore.ieee.org/servlet/
                                        opac?punumber=4031329. IEEE

Computer Society Order Num-

ber P2720.

Long:2007:MTP

[LNP07] Yun Long, A. Nachmias, and Y. Peres. Mixing time power laws at criticality. In IEEE [IEE07], pages 205–

214. ISBN 0-7695-3010-9. ISSN 0272-5428. LCCN QA76 .S974 2007. URL http://

                                        ieeexplore.ieee.org/servlet/
                                        opac?punumber=4389466. IEEE

Computer Society order number

P3010.

Lavi:2005:TNO

[LS05] R. Lavi and C. Swamy. Truthful and near-optimal mechanism design via linear programming.

In IEEE [IEE05], pages 595–

604. ISBN 0-7695-2468-0, 0-7695-2362-5. ISSN 0272-5428. LCCN QA76 .S979 2005. IEEE Computer Society order number P2468.

Lubetzky:2007:NLI

[LS07] E. Lubetzky and U. Stav. Nonlinear index coding outperforming the linear optimum. In IEEE [IEE07], pages 161–

168. ISBN 0-7695-3010-9. ISSN 0272-5428. LCCN QA76 .S974 2007. URL http://

ieeexplore.ieee.org/servlet/
opac?punumber=4389466. IEEE

Computer Society order number P3010.

Luby:2002:LC

[Lub02] M. Luby. LT codes. In IEEE [IEE02], pages 271–280. CODEN ASFPDV. ISBN 0-7695-1822-2. ISSN 0272-5428. LCCN QA267. URL http:/

/ieeexplore.ieee.org/iel5/
8411/26517/01181950.pdf?isnumber=
26517&prod=CNF&arnumber=1181950&
arSt=+271&ared=+280&arAuthor=
Luby%2C+M.;   http://ieeexplore.
ieee.org/xpls/abs_all.jsp?
isnumber=26517&arnumber=1181950&
count=82&index=28.  IEEE

Computer Society Order Number PR01822.

Lovasz:2003:LFG

[LV03a] L. Lovasz and S. Vempala.

Logconcave functions: geometry and efficient sampling algorithms. In IEEE [IEE03], pages 640–649. CODEN ASF-PDV. ISBN 0-7695-2040-5.

ISSN 0272-5428. LCCN QA76 .S979 2003. URL http:/

/ieeexplore.ieee.org/iel5/
8767/27770/01238236.pdf?isnumber=
27770&prod=CNF&arnumber=1238236&
arSt=+640&ared=+649&arAuthor=
Lovasz%2C+L.%3B+Vempala%2C+
S.;   http://ieeexplore.ieee.
org/xpls/abs_all.jsp?isnumber=
27770&arnumber=1238236&count=
66&index=63.  IEEE Com-

puter Society Order Number PR02040.

Lovasz:2003:SAC

[LV03b] L. Lovasz and S. Vempala. Simulated annealing in convex bodies and an O ∗ (n^{4}) volume algorithm. In IEEE [IEE03], pages 650–659. CO-DEN ASFPDV. ISBN 0-7695-2040-5. ISSN 0272-5428. LCCN QA76 .S979 2003. URL http:/

/ieeexplore.ieee.org/iel5/
8767/27770/01238237.pdf?isnumber=
27770&prod=CNF&arnumber=1238237&
arSt=+650&ared=+659&arAuthor=
Lovasz%2C+L.%3B+Vempala%2C+
S.;   http://ieeexplore.ieee.
org/xpls/abs_all.jsp?isnumber=
27770&arnumber=1238237&count=
66&index=64.  IEEE Com-

puter Society Order Number PR02040.

Lovasz:2006:FAL

[LV06] Laszlo Lovasz and Santosh Vempala. Fast algorithms for logconcave functions: Sampling, rounding, integration and optimization. In IEEE [IEE06], pages 57–68. ISBN 0-7695-2720-5, 0-7695-2362-5. ISSN 0272-5428. LCCN QA76 .S974 2006. URL http://

ieeexplore.ieee.org/servlet/
opac?punumber=4031329. IEEE

Computer Society Order Number P2720.

Maidl:2000:CFC

[Mai00] M. Maidl. The common fragment of CTL and LTL. In IEEE [IEE00], pages 643–652. CO-DEN ASFPDV. ISBN 0-7695-0850-2, 0-7695-0851-0 (case), 0-7695-0852-9 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 2000. IEEE Computer Society Order Number PR00850.

Marx:2005:CSP

[Mar05] D. Marx. The closest substring problem with small distances. In IEEE [IEE05], pages 63–72. ISBN 0-7695-2468-0, 0-7695-2362-5. ISSN 0272-5428. LCCN QA76 .S979 2005. IEEE Computer Society order number P2468.

Marx:2007:CYB

[Mar07a]  D. Marx.  Can you beat

treewidth? In IEEE [IEE07], pages 169–179. ISBN 0-7695-3010-9. ISSN 0272-5428. LCCN QA76 .S974 2007. URL http://

ieeexplore.ieee.org/servlet/
opac?punumber=4389466. IEEE

Computer Society order number P3010.

Marx:2007:OPG

[Mar07b] D. Marx. On the optimality of planar and geometric approximation schemes. In IEEE [IEE07], pages 338– 348. ISBN 0-7695-3010-9.

ISSN 0272-5428. LCCN QA76 .S974 2007. URL http://

                                        ieeexplore.ieee.org/servlet/
                                        opac?punumber=4389466. IEEE

Computer Society order number P3010.

McConnell:2001:LTR

[McC01] R. M. McConnell. Lineartime recognition of circular-arc graphs. In IEEE [IEE01], pages 386–394. CODEN ASF-PDV. ISBN 0-7695-1390-5, 0-7695-1391-3 (case), 0-7695-1392-1 (microfiche). ISSN 0272-5428. LCCN ????

McSherry:2001:SPR

[McS01] F. McSherry. Spectral partitioning of random graphs. In IEEE [IEE01], pages 529–537. CO-DEN ASFPDV. ISBN 0-7695-1390-5, 0-7695-1391-3 (case), 0-7695-1392-1 (microfiche). ISSN 0272-5428. LCCN ????

Meir:2009:CPE

[Mei09] O. Meir. Combinatorial PCPs with efficient verifiers. In IEEE [IEE09], pages 463–471. ISBN 0-7695-3850-9. ISSN 0272-5428. LCCN QA76 .S95 2009. IEEE Computer Society order number P3850.

Meyerson:2001:OFL

[Mey01] A. Meyerson. Online facility location. In IEEE [IEE01], pages 426–431. CODEN ASF-PDV. ISBN 0-7695-1390-5, 0-7695-1391-3 (case), 0-7695-1392-1 (microfiche). ISSN 0272-5428. LCCN ????

Meyerson:2005:PPP

[Mey05] A. Meyerson. The parking permit problem. In IEEE [IEE05], pages 274–282. ISBN 0-7695-2468-0, 0-7695-2362-5. ISSN 0272-5428. LCCN QA76 .S979 2005. IEEE Computer Society order number P2468.

Micciancio:2002:GCK

[Mic02]  D. Micciancio.  Generalized

compact knapsacks, cyclic lattices, and efficient one-way functions from worst-case complexity assumptions. In IEEE [IEE02], pages 356–365. CO-DEN ASFPDV. ISBN 0-7695-1822-2. ISSN 0272-5428. LCCN QA267. URL http:/

/ieeexplore.ieee.org/iel5/
8411/26517/01181960.pdf?isnumber=
26517&prod=CNF&arnumber=1181960&
arSt=+356&ared=+365&arAuthor=
Micciancio%2C+D.;   http:/
/ieeexplore.ieee.org/xpls/
abs_all.jsp?isnumber=26517&
arnumber=1181960&count=82&

index=36. IEEE Computer Society Order Number PR01822.

Moitra:2008:SRG

[ML08] A. Moitra and T. Leighton.

Some results on greedy embeddings in metric spaces. In IEEE [IEE08], pages 337–

346. ISBN 0-7695-3436-8. ISSN 0272-5428. LCCN QA76 .S95 2008. URL http://

ieeexplore.ieee.org/servlet/
opac?punumber=4690923. IEEE

Computer Society order number P3436.

Meyerson:2000:CDT

[MMP00] A. Meyerson, K. Munagala, and S. Plotkin. COST DIS-TANCE: two metric network design. In IEEE [IEE00], pages 624–630. CODEN ASF-PDV. ISBN 0-7695-0850-2, 0-7695-0851-0 (case), 0-7695-0852-9 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 2000. IEEE Computer Society Order Number PR00850.

Meyerson:2001:DNI

[MMP01] A. Meyerson, K. Munagala, and S. Plotkin. Designing networks incrementally. In IEEE [IEE01], pages 406–415. CO-DEN ASFPDV. ISBN 0-7695-1390-5, 0-7695-1391-3 (case), 0-7695-1392-1 (microfiche). ISSN 0272-5428. LCCN ????

Mendel:2006:RPP

[MN06] Manor Mendel and Assaf Naor. Ramsey partitions and proximity data structures. In IEEE [IEE06], pages 109–118. ISBN 0-7695-2720-5, 0-7695-2362-5. ISSN 0272-5428. LCCN QA76 .S974 2006. URL http://

                                        ieeexplore.ieee.org/servlet/
                                        opac?punumber=4031329. IEEE

Computer Society Order Number P2720.

Mochon:2004:QWC

[Moc04] C. Mochon. Quantum weak coin-flipping with bias of 0.192. In IEEE [IEE04], pages 2–11. CODEN ASFPDV. ISBN 0-7695-2228-9. ISSN 0272-5428. LCCN QA276. URL http:/

/ieeexplore.ieee.org/iel5/
9430/29918/01366219.pdf?isnumber=
29918&prod=CNF&arnumber=1366219&
arSt=+2&ared=+11&arAuthor=
Mochon%2C+C.;   http://ieeexplore.
ieee.org/xpls/abs_all.jsp?
isnumber=29918&arnumber=1366219&
count=64&index=0. IEEE Com-

puter Society Order Number P2228.

Moitra:2009:AAM

[Moi09] A. Moitra. Approximation algorithms for multicommodity-type problems with guarantees independent of the graph size. In IEEE [IEE09], pages 3–12. ISBN 0-7695-3850-9. ISSN 0272-5428. LCCN QA76 .S95 2009. IEEE Computer Society order number P3850.

Mossel:2005:NSF

[MOO05] E. Mossel, R. O’Donnell, and K. Oleszkiewicz. Noise stability of functions with low influences: Invariance and optimality. In IEEE [IEE05], pages 21–30. ISBN 0-7695-2468-0, 0-7695-2362-5. ISSN 0272-5428. LCCN QA76 .S979 2005. IEEE Computer Society order number P2468.

Mossel:2008:GBN

[Mos08] E. Mossel. Gaussian bounds for noise correlation of functions and tight analysis of long codes. In IEEE [IEE08], pages 156–165. ISBN 0-7695-3436-

8. ISSN 0272-5428. LCCN QA76 .S95 2008. URL http://

ieeexplore.ieee.org/servlet/
opac?punumber=4690923. IEEE

Computer Society order number

P3436.

Mettu:2000:OMP

[MP00] R. R. Mettu and C. G. Plaxton. The online median problem. In IEEE [IEE00], pages 339–348. CODEN ASFPDV. ISBN 0-7695-0850-2, 0-7695-0851-0 (case), 0-7695-0852-9 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 2000. IEEE Computer Society Order Number PR00850.

Micali:2006:IIC

[MPR06] Silvio Micali, Rafael Pass, and Alon Rosen. Inputindistinguishable computation. In IEEE [IEE06], pages 367–378. ISBN 0-7695-2720-5, 0-7695-2362-5. ISSN 0272-5428. LCCN QA76 .S974 2006. URL http://

                                        ieeexplore.ieee.org/servlet/
                                        opac?punumber=4031329. IEEE

Computer Society Order Num-

ber P2720.

Mitzenmacher:2002:LBM

[MPS02] M. Mitzenmacher, B. Prabhakar, and D. Shah. Load balancing with memory. In IEEE [IEE02], pages 799–808. CODEN ASFPDV. ISBN 0-7695-1822-2. ISSN 0272-5428. LCCN QA267. URL http:/

                                        /ieeexplore.ieee.org/iel5/
                                        8411/26517/01182005.pdf?isnumber=
                                        26517&prod=CNF&arnumber=1182005&
                                        arSt=+799&ared=+808&arAuthor=
                                        Mitzenmacher%2C+M.%3B+Prabhakar%
                                        2C+B.%3B+Shah%2C+D.;   http:
                                        //ieeexplore.ieee.org/xpls/
                                        abs_all.jsp?isnumber=26517&

arnumber=1182005&count=82&

index=79. IEEE Computer Society Order Number PR01822.

Mihail:2003:CCP

[MPS03] M. Mihail, C. Papadimitriou, and A. Saberi. On certain connectivity properties of the Internet topology. In IEEE [IEE03], pages 28–35. CODEN ASFPDV. ISBN 0-7695-2040-

5. ISSN 0272-5428. LCCN QA76 .S979 2003. URL http:/

/ieeexplore.ieee.org/iel5/
8767/27770/01238178.pdf?isnumber=
27770&prod=CNF&arnumber=1238178&
arSt=+28&ared=+35&arAuthor=
Mihail%2C+M.%3B+Papadimitriou%
2C+C.%3B+Saberi%2C+A.;   http:
//ieeexplore.ieee.org/xpls/
abs_all.jsp?isnumber=27770&
arnumber=1238178&count=66&

index=5. IEEE Computer Society Order Number PR02040.

Mossel:2004:SSR

[MPS04] E. Mossel, Y. Peres, and A. Sinclair. Shuffling by semirandom transpositions. In IEEE [IEE04], pages 572–581. CODEN ASFPDV. ISBN 0-7695-2228-9. ISSN 0272-5428. LCCN QA276. URL http:/

/ieeexplore.ieee.org/iel5/
9430/29918/01366277.pdf?isnumber=
29918&prod=CNF&arnumber=1366277&
arSt=+572&ared=+581&arAuthor=
Mossel%2C+E.%3B+Peres%2C+
Y.%3B+Sinclair%2C+A.;   http:
//ieeexplore.ieee.org/xpls/
abs_all.jsp?isnumber=29918&
arnumber=1366277&count=64&

index=58. IEEE Computer Society Order Number P2228.

Martin:2000:SAS

[MR00] R. A. Martin and D. Randall. Sampling adsorbing staircase walks using a new Markov chain decomposition method. In IEEE [IEE00], pages 492–502. CO-DEN ASFPDV. ISBN 0-7695-0850-2, 0-7695-0851-0 (case), 0-7695-0852-9 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 2000. IEEE Computer Society Order Number PR00850.

Micciancio:2004:WCA

[MR04] D. Micciancio and O. Regev. Worst-case to average-case reductions based on Gaussian measures. In IEEE [IEE04], pages 372–381. CO-DEN ASFPDV. ISBN 0-7695-2228-9. ISSN 0272-5428. LCCN QA276. URL http:/

                                        /ieeexplore.ieee.org/iel5/
                                        9430/29918/01366257.pdf?isnumber=
                                        29918&prod=CNF&arnumber=1366257&
                                        arSt=+372&ared=+381&arAuthor=
                                        Micciancio%2C+D.%3B+Regev%
                                        2C+O.;   http://ieeexplore.
                                        ieee.org/xpls/abs_all.jsp?
                                        isnumber=29918&arnumber=1366257&
                                        count=64&index=38.  IEEE

Computer Society Order Num-

ber P2228.

Moshkovitz:2008:TQP

[MR08] D. Moshkovitz and R. Raz. Two query PCP with sub-constant error. In IEEE [IEE08], pages 314–323. ISBN 0-7695-3436-

8. ISSN 0272-5428. LCCN QA76 .S95 2008. URL http://

ieeexplore.ieee.org/servlet/
opac?punumber=4690923. IEEE

Computer Society order number P3436.

Micali:2003:ZKS

[MRK03] S. Micali, M. Rabin, and J. Kilian. Zero-knowledge sets. In IEEE [IEE03], pages 80–91. CO-DEN ASFPDV. ISBN 0-7695-2040-5. ISSN 0272-5428. LCCN QA76 .S979 2003. URL http:/

/ieeexplore.ieee.org/iel5/
8767/27770/01238183.pdf?isnumber=
27770&prod=CNF&arnumber=1238183&
arSt=+80&ared=+91&arAuthor=
Micali%2C+S.%3B+Rabin%2C+
M.%3B+Kilian%2C+J.;   http:
//ieeexplore.ieee.org/xpls/
abs_all.jsp?isnumber=27770&
arnumber=1238183&count=66&

index=10. IEEE Computer Society Order Number PR02040.

Moore:2005:SGD

[MRS05] C. Moore, A. Russell, and L. J.

Schulman. The symmetric group defies strong Fourier sampling. In IEEE [IEE05], pages 479–

488. ISBN 0-7695-2468-0, 0-7695-2362-5. ISSN 0272-5428. LCCN QA76 .S979 2005. IEEE Computer Society order number P2468.

Molloy:2003:RCR

[MS03] M. Molloy and M. Salavatipour.

The resolution complexity of random constraint satisfaction problems. In IEEE [IEE03], pages 330–339. CODEN ASF-PDV. ISBN 0-7695-2040-5. ISSN 0272-5428. LCCN QA76 .S979 2003. URL http:/

/ieeexplore.ieee.org/iel5/
8767/27770/01238207.pdf?isnumber=
27770&prod=CNF&arnumber=1238207&
arSt=+330&ared=+339&arAuthor=
Molloy%2C+M.%3B+Salavatipour%
2C+M.;   http://ieeexplore.
ieee.org/xpls/abs_all.jsp?
isnumber=27770&arnumber=1238207&
count=66&index=34.  IEEE

Computer Society Order Number PR02040.

Matousek:2004:REC

[MS04a] J. Matousek and T. Szabo. Random edge can be exponential on abstract cubes. In IEEE [IEE04], pages 92–100. CODEN ASFPDV. ISBN 0-7695-2228-9. ISSN 0272-5428. LCCN QA276. URL http:/

/ieeexplore.ieee.org/iel5/
9430/29918/01366228.pdf?isnumber=
29918&prod=CNF&arnumber=1366228&
arSt=+92&ared=+100&arAuthor=
Matousek%2C+J.%3B+Szabo%2C+
T.;   http://ieeexplore.ieee.
org/xpls/abs_all.jsp?isnumber=
29918&arnumber=1366228&count=

64&index=9. IEEE Computer Society Order Number P2228.

Mucha:2004:MMG

[MS04b] M. Mucha and P. Sankowski. Maximum matchings via Gaussian elimination. In IEEE [IEE04], pages 248–255. CO-DEN ASFPDV. ISBN 0-7695-2228-9. ISSN 0272-5428. LCCN QA276. URL http:/

/ieeexplore.ieee.org/iel5/
9430/29918/01366244.pdf?isnumber=
29918&prod=CNF&arnumber=1366244&
arSt=+248&ared=+255&arAuthor=

Mucha%2C+M.%3B+Sankowski%
2C+P.;   http://ieeexplore.
ieee.org/xpls/abs_all.jsp?
isnumber=29918&arnumber=1366244&
count=64&index=25.  IEEE

Computer Society Order Number P2228.

Mastrolilli:2008:AJS

[MS08a] M. Mastrolilli and O. Svensson.

(acyclic) job shops are hard to approximate. In IEEE [IEE08], pages 583–592. ISBN 0-7695-3436-8. ISSN 0272-5428. LCCN QA76 .S95 2008. URL http://

ieeexplore.ieee.org/servlet/
opac?punumber=4690923. IEEE

Computer Society order number P3436.

Matousek:2008:IME

[MS08b] J. Matousek and A. Sidiropoulos. Inapproximability for metric embeddings into R^{d}. In IEEE [IEE08], pages 405–

413. ISBN 0-7695-3436-8. ISSN 0272-5428. LCCN QA76 .S95 2008. URL http://

ieeexplore.ieee.org/servlet/
opac?punumber=4690923. IEEE

Computer Society order number P3436.

Montanari:2009:CEL

[MS09a] A. Montanari and A. Saberi.

Convergence to equilibrium in local interaction games. In IEEE [IEE09], pages 303–312. ISBN 0-7695-3850-9. ISSN 0272-5428. LCCN QA76 .S95 2009. IEEE Computer Society order number P3850.

Myers:2009:BEC

[MS09b] S. Myers and A. Shelat. Bit encryption is complete. In IEEE [IEE09], pages 607–616. ISBN 0-7695-3850-9. ISSN 0272-5428. LCCN QA76 .S95 2009. IEEE Computer Society order number P3850.

Mossel:2003:SEB

[MST03] E. Mossel, A. Shpilka, and L. Trevisan. On ϵ-biased generators in NC^{0}. In IEEE [IEE03], pages 136–145. CO-DEN ASFPDV. ISBN 0-7695-2040-5. ISSN 0272-5428. LCCN QA76 .S979 2003. URL http:/

                                        /ieeexplore.ieee.org/iel5/
                                        8767/27770/01238188.pdf?isnumber=
                                        27770&prod=CNF&arnumber=1238188&
                                        arSt=+136&ared=+145&arAuthor=
                                        Mossel%2C+E.%3B+Shpilka%2C+
                                        A.%3B+Trevisan%2C+L.;   http:
                                        //ieeexplore.ieee.org/xpls/
                                        abs_all.jsp?isnumber=27770&
                                        arnumber=1238188&count=66&

index=15. IEEE Computer Society Order Number PR02040.

Mehta:2005:AGL

[MSVV05] A. Mehta, A. Saberi, U. Vazirani, and V. Vazirani. AdWords and generalized on-line matching. In IEEE [IEE05], pages 264–273. ISBN 0-7695-2468-0, 0-7695-2362-5. ISSN 0272-5428. LCCN QA76 .S979 2005. IEEE Computer Society order number P2468.

Martinelli:2003:IMT

[MSW03] F. Martinelli, A. Sinclair, and D. Weitz. The Ising model on trees: boundary conditions and mixing time. In IEEE [IEE03], pages 628–639. CO-DEN ASFPDV. ISBN 0-7695-2040-5. ISSN 0272-5428. LCCN QA76 .S979 2003. URL http:/

/ieeexplore.ieee.org/iel5/
8767/27770/01238235.pdf?isnumber=
27770&prod=CNF&arnumber=1238235&
arSt=+628&ared=+639&arAuthor=
Martinelli%2C+F.%3B+Sinclair%
2C+A.%3B+Weitz%2C+D.;   http:
//ieeexplore.ieee.org/xpls/
abs_all.jsp?isnumber=27770&
arnumber=1238235&count=66&

index=62. IEEE Computer Society Order Number PR02040.

McSherry:2007:MDD

[MT07] F. McSherry and K. Talwar.

Mechanism design via differential privacy. In IEEE [IEE07], pages 94–103. ISBN 0-7695-3010-9. ISSN 0272-5428. LCCN QA76 .S974 2007. URL http://

ieeexplore.ieee.org/servlet/
opac?punumber=4389466. IEEE

Computer Society order number P3010.

Newman:2000:TFS

[New00] I. Newman. Testing of function that have small width branching programs. In IEEE [IEE00], pages 251–258. CO-DEN ASFPDV. ISBN 0-7695-0850-2, 0-7695-0851-0 (case), 0-7695-0852-9 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 2000. IEEE Computer Society Order Number PR00850.

Nguyen:2008:CTA

[NO08] H. N. Nguyen and K. Onak.

Constant-time approximation algorithms via local improvements. In IEEE [IEE08], pages 327–336. ISBN 0-7695-3436-

8. ISSN 0272-5428. LCCN QA76 .S95 2008. URL http://

ieeexplore.ieee.org/servlet/
opac?punumber=4690923. IEEE

Computer Society order number P3436.

Nguyen:2006:SZK

[NOV06] Minh-Huyen Nguyen, Shien Jin Ong, and Salil Vadhan. Statistical zero-knowledge arguments for NP from any one-way function. In IEEE [IEE06], pages 3–

14. ISBN 0-7695-2720-5, 0-7695-2362-5. ISSN 0272-5428. LCCN QA76 .S974 2006. URL http://

ieeexplore.ieee.org/servlet/
opac?punumber=4031329. IEEE

Computer Society Order Number P2720.

Nair:2003:PPC

[NPS03] C. Nair, B. Prabhakar, and M. Sharma. Proofs of the Parisi and Coppersmith–Sorkin conjectures for the finite random assignment problem. In IEEE [IEE03], pages 168–178. CO-DEN ASFPDV. ISBN 0-7695-2040-5. ISSN 0272-5428. LCCN QA76 .S979 2003. URL http:/

/ieeexplore.ieee.org/iel5/
8767/27770/01238191.pdf?isnumber=
27770&prod=CNF&arnumber=1238191&
arSt=+168&ared=+178&arAuthor=
Nair%2C+C.%3B+Prabhakar%2C+
B.%3B+Sharma%2C+M.;   http:
//ieeexplore.ieee.org/xpls/
abs_all.jsp?isnumber=27770&
arnumber=1238191&count=66&

index=18. IEEE Computer Society Order Number PR02040.

Naor:2005:COM

[NR05] M. Naor and G. N. Rothblum.

The complexity of online memory checking. In IEEE [IEE05], pages 573–582. ISBN 0-7695-2468-0, 0-7695-2362-5. ISSN 0272-5428. LCCN QA76 .S979 2005. IEEE Computer Society order number P2468.

Navon:2005:DLP

[NS05] M. Navon and A. Samorodnitsky. On Delsarte’s linear programming bounds for binary codes. In IEEE [IEE05], pages 327–336. ISBN 0-7695-2468-0, 0-7695-2362-5. ISSN 0272-5428. LCCN QA76 .S979 2005. IEEE Computer Society order number P2468.

Naor:2006:PEL

[NS06] Assaf Naor and Gideon Schechtman. Planar Earthmover is not in L_{1}. In IEEE [IEE06], pages 655–666. ISBN 0-7695-2720-5, 0-7695-2362-5. ISSN 0272-5428. LCCN QA76 .S974 2006. URL http://

ieeexplore.ieee.org/servlet/
opac?punumber=4031329. IEEE

Computer Society Order Number P2720.

Nutov:2009:AMC

[Nut09] Z. Nutov. Approximating minimum cost connectivity problems via uncrossable bifamilies and spider-cover decompositions. In IEEE [IEE09], pages 417–426.

ISBN 0-7695-3850-9. ISSN 0272-5428. LCCN QA76 .S95 2009. IEEE Computer Society order number P3850.

Or:2008:BLO

[OH08] M. Ben Or and A. Hassidim. The Bayesian learner is optimal for noisy binary search (and pretty good for quantum as well). In IEEE [IEE08], pages 221–230. ISBN 0-7695-3436-8. ISSN 0272-5428. LCCN QA76 .S95 2008. URL http://

ieeexplore.ieee.org/servlet/
opac?punumber=4690923. IEEE

Computer Society order number P3436.

Or:2008:QMP

[OHP08] M. B. Or, A. Hassidim, and H. Pilpel. Quantum multi prover interactive proofs with communicating provers. In IEEE [IEE08], pages 467– 476. ISBN 0-7695-3436-8. ISSN 0272-5428. LCCN QA76 .S95 2008. URL http://

ieeexplore.ieee.org/servlet/
opac?punumber=4690923. IEEE

Computer Society order number P3436.

Onak:2006:GBS

[OP06] Krzysztof Onak and Pawel Parys. Generalization of binary search: Searching in trees and forest-like partial orders. In IEEE [IEE06], pages 379–388. ISBN 0-7695-2720-5, 0-7695-2362-5. ISSN 0272-5428. LCCN QA76 .S974 2006. URL http://

ieeexplore.ieee.org/servlet/
opac?punumber=4031329. IEEE

Computer Society Order Number P2720.

Ostrovsky:2000:PTA

[OR00] R. Ostrovsky and Y. Rabani.

Polynomial time approximation schemes for geometric kclustering. In IEEE [IEE00], pages 349–358. CODEN ASF-PDV. ISBN 0-7695-0850-2, 0-7695-0851-0 (case), 0-7695-0852-9 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 2000. IEEE Computer Society Order Number PR00850.

Ostrovsky:2005:ECC

[ORS05] R. Ostrovsky, Y. Rabani, and L. J. Schulman. Error-correcting codes for automatic control. In IEEE [IEE05], pages 309– 316. ISBN 0-7695-2468-0, 0-7695-2362-5. ISSN 0272-5428. LCCN QA76 .S979 2005. IEEE Computer Society order number P2468.

Ostrovsky:2006:ELT

[ORSS06] Rafail Ostrovsky, Yuval Rabani, Leonard J. Schulman, and Chaitanya Swamy. The effectiveness of Lloyd-type methods for the k-means problem. In IEEE [IEE06], pages 165–176. ISBN 0-7695-2720-5, 0-7695-2362-5. ISSN 0272-5428. LCCN QA76 .S974 2006. URL http://

ieeexplore.ieee.org/servlet/
opac?punumber=4031329. IEEE

Computer Society Order Number P2720.

ODonnell:2005:EDT

[OSSS05] R. O’Donnell, M. Saks, O. Schramm, and R. A. Servedio. Every decision tree has an influential variable. In IEEE [IEE05], pages 31–39. ISBN 0-7695-2468-0, 0-7695-2362-5. ISSN 0272-5428. LCCN QA76 .S979 2005. IEEE Computer Society order number P2468.

Orlitsky:2003:AGT

[OSZ03] A. Orlitsky, N. P. Santhanam, and J. Zhang. Always Good Turing: asymptotically optimal probability estimation. In IEEE [IEE03], pages 179–188. CO-DEN ASFPDV. ISBN 0-7695-2040-5. ISSN 0272-5428. LCCN QA76 .S979 2003. URL http:/

/ieeexplore.ieee.org/iel5/
8767/27770/01238192.pdf?isnumber=
27770&prod=CNF&arnumber=1238192&
arSt=+179&ared=+188&arAuthor=
Orlitsky%2C+A.%3B+Santhanam%
2C+N.P.%3B+Zhang%2C+J.;
http://ieeexplore.ieee.org/
xpls/abs_all.jsp?isnumber=
27770&arnumber=1238192&count=
66&index=19.  IEEE Com-

puter Society Order Number PR02040.

ODonnell:2009:KKK

[OW09] R. O’Donnell and K. Wimmer. KKL, Kruskal–Katona, and monotone nets. In IEEE [IEE09], pages 725–734. ISBN 0-7695-3850-9. ISSN 0272-5428. LCCN QA76 .S95 2009. IEEE Computer Society order number P3850.

Pak:2000:PRA

[Pak00]  I. Pak.  The product re-

placement algorithm is polynomial. In IEEE [IEE00], pages 476–485. CODEN ASF-PDV. ISBN 0-7695-0850-2, 0-7695-0851-0 (case), 0-7695-0852-9 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 2000. IEEE Computer Society Order Number PR00850.

Papadimitriou:2001:GTM

[Pap01] C. H. Papadimitriou. Game theory and mathematical economics: a theoretical computer scientist’s introduction. In IEEE [IEE01], pages 4–8. CODEN ASFPDV. ISBN 0-7695-1390-5, 0-7695-1391-3 (case), 0-7695-1392-1 (microfiche). ISSN 0272-5428. LCCN ????

Patrascu:2006:PPL

[Pat06] Mihai Patrascu. Planar point location in sublogarithmic time. In IEEE [IEE06], pages 325–332. ISBN 0-7695-2720-5, 0-7695-2362-5. ISSN 0272-5428. LCCN QA76 .S974 2006. URL http://

ieeexplore.ieee.org/servlet/
opac?punumber=4031329. IEEE

Computer Society Order Number P2720.

Patrascu:2008:DS

[Pat08a] M. Patrascu. (data) STRUC-TURES. In IEEE [IEE08], pages 434–443. ISBN 0-7695-3436-8. ISSN 0272-5428. LCCN QA76 .S95 2008. URL http://

ieeexplore.ieee.org/servlet/
opac?punumber=4690923. IEEE

Computer Society order number P3436.

Patrascu:2008:S

[Pat08b] M. Patrascu. Succincter. In IEEE [IEE08], pages 305–

313. ISBN 0-7695-3436-8. ISSN 0272-5428. LCCN QA76 .S95 2008. URL http://

ieeexplore.ieee.org/servlet/
opac?punumber=4690923. IEEE

Computer Society order number P3436.

Pettie:2002:IAS

[Pet02] S. Pettie. An inverse-Ackermann style lower bound for the online minimum spanning tree verification problem. In IEEE [IEE02], pages 155–163. CO-DEN ASFPDV. ISBN 0-7695-1822-2. ISSN 0272-5428. LCCN QA267. URL http:/

/ieeexplore.ieee.org/iel5/
8411/26517/01181892.pdf?isnumber=
26517&prod=CNF&arnumber=1181892&
arSt=+155&ared=+163&arAuthor=
Pettie%2C+S.;   http://ieeexplore.
ieee.org/xpls/abs_all.jsp?
isnumber=26517&arnumber=1181892&
count=82&index=17.  IEEE

Computer Society Order Number PR01822.

Pettie:2005:TFA

[Pet05]  S. Pettie.  Towards a fi-

nal analysis of pairing heaps. In IEEE [IEE05], pages 174–

183. ISBN 0-7695-2468-0, 0-7695-2362-5. ISSN 0272-5428. LCCN QA76 .S979 2005. IEEE Computer Society order number P2468.

Porat:2009:EAP

[PP09] B. Porat and E. Porat. Exact and approximate pattern matching in the streaming model. In IEEE [IEE09], pages 315–323. ISBN 0-7695-3850-9. ISSN 0272-5428. LCCN QA76 .S95 2009. IEEE Computer Society order number P3850.

Pass:2003:BCS

[PR03] R. Pass and A. Rosen. Boundedconcurrent secure two-party computation in a constant number of rounds. In IEEE [IEE03], pages 404–413. CODEN ASF-PDV. ISBN 0-7695-2040-5. ISSN 0272-5428. LCCN QA76 .S979 2003. URL http:/

/ieeexplore.ieee.org/iel5/
8767/27770/01238214.pdf?isnumber=
27770&prod=CNF&arnumber=1238214&
arSt=+404&ared=+413&arAuthor=
Pass%2C+R.%3B+Rosen%2C+A.;
http://ieeexplore.ieee.org/
xpls/abs_all.jsp?isnumber=
27770&arnumber=1238214&count=
66&index=41.  IEEE Com-

puter Society Order Number PR02040.

Pass:2005:CNM

[PR05] R. Pass and A. Rosen. Concurrent non-malleable commitments. In IEEE [IEE05], pages 563–572. ISBN 0-7695-2468-0, 0-7695-2362-5. ISSN 0272-5428. LCCN QA76 .S979 2005. IEEE Computer Society order number P2468.

Prabhakaran:2002:CZK

[PRS02] M. Prabhakaran, A. Rosen, and A. Sahai. Concurrent zero knowledge with logarithmic round-complexity. In IEEE [IEE02], pages 366–375. CO-DEN ASFPDV. ISBN 0-7695-1822-2. ISSN 0272-5428. LCCN QA267. URL http:/

/ieeexplore.ieee.org/iel5/
8411/26517/01181961.pdf?isnumber=
26517&prod=CNF&arnumber=1181961&
arSt=+366&ared=+375&arAuthor=
Prabhakaran%2C+M.%3B+Rosen%
2C+A.%3B+Sahai%2C+A.;   http:
//ieeexplore.ieee.org/xpls/
abs_all.jsp?isnumber=26517&
arnumber=1181961&count=82&

index=37. IEEE Computer Society Order Number PR01822.

Pandurangan:2001:BLD

[PRU01] G. Pandurangan, P. Raghavan, and E. Upfal. Building low-diameter P2P networks. In IEEE [IEE01], pages 492–499. CODEN ASFPDV. ISBN 0-7695-1390-5, 0-7695-1391-3 (case), 0-7695-1392-1 (microfiche). ISSN 0272-5428. LCCN ????

Papadimitriou:2008:HBT

[PSS08] C. Papadimitriou, M. Schapira, and Y. Singer. On the hardness of being truthful. In IEEE [IEE08], pages 250–

259. ISBN 0-7695-3436-8. ISSN 0272-5428. LCCN QA76 .S95 2008. URL http://

ieeexplore.ieee.org/servlet/
opac?punumber=4690923. IEEE

Computer Society order number P3436.

Pach:2000:BCU

[PT00] J. Pach and G. Tardos. On the boundary complexity of the union of fat triangles. In IEEE [IEE00], pages 423–431. CO-DEN ASFPDV. ISBN 0-7695-0850-2, 0-7695-0851-0 (case), 0-7695-0852-9 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 2000. IEEE Computer Society Order Number PR00850.

Pal:2003:GSP

[PT03] M. Pal and E. Tardos. Group strategy proof mechanisms via primal-dual algorithms. In IEEE [IEE03], pages 584–

593. CODEN ASFPDV. ISBN 0-7695-2040-5. ISSN 0272-5428. LCCN QA76 .S979 2003. URL http:/

/ieeexplore.ieee.org/iel5/
8767/27770/01238231.pdf?isnumber=
27770&prod=CNF&arnumber=1238231&
arSt=+584&ared=+593&arAuthor=
Pal%2C+M.%3B+Tardos%2C+E.;
 http://ieeexplore.ieee.org/
xpls/abs_all.jsp?isnumber=
27770&arnumber=1238231&count=
66&index=58.  IEEE Com-

puter Society Order Number PR02040.

Patrascu:2006:HLB

[PT06] Mihai Patrascu and Mikkel Thorup. Higher lower bounds for near-neighbor and further rich problems. In IEEE [IEE06], pages 646–654. ISBN 0-7695-2720-5, 0-7695-2362-5. ISSN 0272-5428. LCCN QA76 .S974 2006. URL http://

ieeexplore.ieee.org/servlet/

                                        opac?punumber=4031329. IEEE

Computer Society Order Number P2720.

Patrascu:2007:PFC

[PT07] M. Patrascu and M. Thorup. Planning for fast connectivity updates. In IEEE [IEE07], pages 263–271. ISBN 0-7695-3010-9. ISSN 0272-5428. LCCN QA76 .S974 2007. URL http://

                                        ieeexplore.ieee.org/servlet/
                                        opac?punumber=4389466. IEEE

Computer Society order number P3010.

Pal:2001:FLN

[PTW01] M. Pal, E. Tardos, and T. Wexler. Facility location with nonuniform hard capacities. In IEEE [IEE01], pages 329–338. CODEN ASFPDV. ISBN 0-7695-1390-5, 0-7695-1391-3 (case), 0-7695-1392-1 (microfiche). ISSN 0272-5428. LCCN ????

Panigrahy:2008:GAL

[PTW08] R. Panigrahy, K. Talwar, and U. Wieder. A geometric approach to lower bounds for approximate near-neighbor search and partial match. In IEEE [IEE08], pages 414– 423. ISBN 0-7695-3436-8. ISSN 0272-5428. LCCN QA76 .S95 2008. URL http://

                                        ieeexplore.ieee.org/servlet/
                                        opac?punumber=4690923. IEEE

Computer Society order number P3436.

Parvaresh:2005:CEB

[PV05] F. Parvaresh and A. Vardy.

Correcting errors beyond the Guruswami–Sudan radius in polynomial time. In IEEE [IEE05], pages 285–294. ISBN 0-7695-2468-0, 0-7695-2362-5. ISSN 0272-5428. LCCN QA76 .S979 2005. IEEE Computer Society order number P2468.

Papadimitriou:2000:ATO

[PY00] C. H. Papadimitriou and M. Yannakakis. On the approximability of trade-offs and optimal access of Web sources. In IEEE [IEE00], pages 86–92. CO-DEN ASFPDV. ISBN 0-7695-0850-2, 0-7695-0851-0 (case), 0-7695-0852-9 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 2000. IEEE Computer Society Order Number PR00850.

Racke:2002:MCG

[Rac02] H. Racke. Minimizing congestion in general networks. In IEEE [IEE02], pages 43–52. CODEN ASFPDV. ISBN 0-7695-1822-2. ISSN 0272-5428. LCCN QA267. URL http:/

/ieeexplore.ieee.org/iel5/
8411/26517/01181881.pdf?isnumber=
26517&prod=CNF&arnumber=1181881&
arSt=+43&ared=+52&arAuthor=
Racke%2C+H.;   http://ieeexplore.
ieee.org/xpls/abs_all.jsp?
isnumber=26517&arnumber=1181881&
count=82&index=6. IEEE Com-

puter Society Order Number PR01822.

Randall:2003:MMC

[Ran03] D. Randall. Mixing [markov chain]. In IEEE [IEE03], pages 4–15. CODEN ASF-PDV. ISBN 0-7695-2040-5. ISSN 0272-5428. LCCN QA76 .S979 2003. URL http:/

/ieeexplore.ieee.org/iel5/
8767/27770/01238175.pdf?isnumber=
27770&prod=CNF&arnumber=1238175&
arSt=+4&ared=+15&arAuthor=
Randall%2C+D.;   http://ieeexplore.
ieee.org/xpls/abs_all.jsp?
isnumber=27770&arnumber=1238175&
count=66&index=2. IEEE Com-

puter Society Order Number PR02040.

Raz:2004:MNS

[Raz04] R. Raz. Multilinear-NC_{1} ≠ multilinear-NC_{2}. In IEEE [IEE04], pages 344–351. CO-DEN ASFPDV. ISBN 0-7695-2228-9. ISSN 0272-5428. LCCN QA276. URL http:/

/ieeexplore.ieee.org/iel5/
9430/29918/01366254.pdf?isnumber=
29918&prod=CNF&arnumber=1366254&
arSt=+344&ared=+351&arAuthor=
Raz%2C+R.;   http://ieeexplore.
ieee.org/xpls/abs_all.jsp?
isnumber=29918&arnumber=1366254&
count=64&index=35.  IEEE

Computer Society Order Number P2228.

Raz:2005:QIP

[Raz05] R. Raz. Quantum information and the PCP theorem. In IEEE [IEE05], pages 459–

468. ISBN 0-7695-2468-0, 0-7695-2362-5. ISSN 0272-5428. LCCN QA76 .S979 2005. IEEE Computer Society order number P2468.

Raz:2008:CSP

[Raz08]  R. Raz.  A counterexam-

ple to strong parallel repetition. In IEEE [IEE08], pages 369–373. ISBN 0-7695-3436-

8. ISSN 0272-5428. LCCN QA76 .S95 2008. URL http://

ieeexplore.ieee.org/servlet/
opac?punumber=4690923. IEEE

Computer Society order number P3436.

Regev:2002:QCL

[Reg02] O. Regev. Quantum computation and lattice problems. In IEEE [IEE02], pages 520–529. CODEN ASFPDV. ISBN 0-7695-1822-2. ISSN 0272-5428. LCCN QA267. URL http:/

/ieeexplore.ieee.org/iel5/
8411/26517/01181976.pdf?isnumber=
26517&prod=CNF&arnumber=1181976&
arSt=+520&ared=+529&arAuthor=
Regev%2C+O.;   http://ieeexplore.
ieee.org/xpls/abs_all.jsp?
isnumber=26517&arnumber=1181976&
count=82&index=52.  IEEE

Computer Society Order Number PR01822.

Reichardt:2006:PTA

[Rei06] Ben W. Reichardt. Postselection threshold against biased noise. In IEEE [IEE06], pages 420–428. ISBN 0-7695-2720-5, 0-7695-2362-5. ISSN 0272-5428. LCCN QA76 .S974 2006. URL http://

ieeexplore.ieee.org/servlet/
opac?punumber=4031329. IEEE

Computer Society Order Number P2720.

Reichardt:2009:SPQ

[Rei09] B. W. Reichardt. Span programs and quantum query complexity: The general adversary bound is nearly tight for every Boolean function. In IEEE [IEE09], pages 544–551. ISBN 0-7695-3850-9. ISSN 0272-5428. LCCN QA76 .S95 2009. IEEE Computer Society order number P3850.

Roch:2008:SLR

[Roc08]  S. Roch.  Sequence length

requirement of distance-based phylogeny reconstruction: Breaking the polynomial barrier. In IEEE [IEE08], pages 729– 738. ISBN 0-7695-3436-8. ISSN 0272-5428. LCCN QA76 .S95 2008. URL http://

                                        ieeexplore.ieee.org/servlet/
                                        opac?punumber=4690923. IEEE

Computer Society order number P3436.

Roughgarden:2001:DNS

[Rou01] T. Roughgarden. Designing networks for selfish users is hard. In IEEE [IEE01], pages 472–481. CODEN ASFPDV. ISBN 0-7695-1390-5, 0-7695-1391-3 (case), 0-7695-1392-1 (microfiche). ISSN 0272-5428. LCCN ????

Raskhodnikova:2007:SLB

[RRSS07] S. Raskhodnikova, D. Ron, A. Shpilka, and A. Smith. Strong lower bounds for approximating distribution support size and the distinct elements problem. In IEEE [IEE07], pages 559–569. ISBN 0-7695-3010-9. ISSN 0272-5428. LCCN QA76 .S974 2007. URL http://

ieeexplore.ieee.org/servlet/
opac?punumber=4389466. IEEE

Computer Society order number P3010.

Ronen:2002:HOA

[RS02] A. Ronen and A. Saberi. On the hardness of optimal auctions. In IEEE [IEE02], pages 396–405. CODEN ASFPDV. ISBN 0-7695-1822-2. ISSN 0272-5428. LCCN QA267. URL http:/

/ieeexplore.ieee.org/iel5/
8411/26517/01181964.pdf?isnumber=
26517&prod=CNF&arnumber=1181964&
arSt=+396&ared=+405&arAuthor=
Ronen%2C+A.%3B+Saberi%2C+
A.;   http://ieeexplore.ieee.
org/xpls/abs_all.jsp?isnumber=
26517&arnumber=1181964&count=
82&index=40.  IEEE Com-

puter Society Order Number PR01822.

Razborov:2008:SRA

[RS08] A. A. Razborov and A. A.

Sherstov. The sign-rank of AC^{0}. In IEEE [IEE08], pages 57–66. ISBN 0-7695-3436-

8. ISSN 0272-5428. LCCN QA76 .S95 2008. URL http://

ieeexplore.ieee.org/servlet/
opac?punumber=4690923. IEEE

Computer Society order number P3436.

Raghavendra:2009:HRC

[RS09a] P. Raghavendra and D. Steurer.

How to round any CSP. In IEEE [IEE09], pages 586–594. ISBN 0-7695-3850-9. ISSN 0272-5428.

LCCN QA76 .S95 2009. IEEE Computer Society order number P3850.

Raghavendra:2009:IGS

[RS09b] P. Raghavendra and D. Steurer. Integrality gaps for strong SDP relaxations of UNIQUE GAMES. In IEEE [IEE09], pages 575–585. ISBN 0-7695-3850-9. ISSN 0272-5428. LCCN QA76 .S95 2009. IEEE Computer Society order number P3850.

Radhakrishnan:2000:QCS

[RSV00] J. Radhakrishnan, Pranab Sen, and S. Venkatesh. The quantum complexity of set membership. In IEEE [IEE00], pages 554–562. CODEN ASF-PDV. ISBN 0-7695-0850-2, 0-7695-0851-0 (case), 0-7695-0852-9 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 2000. IEEE Computer Society Order Number PR00850.

Reingold:2000:ERR

[RSW00] O. Reingold, R. Shaltiel, and A. Wigderson. Extracting randomness via repeated condensing. In IEEE [IEE00], pages 22–31. CODEN ASF-PDV. ISBN 0-7695-0850-2, 0-7695-0851-0 (case), 0-7695-0852-9 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 2000. IEEE Computer Society Order Number PR00850.

Raz:2007:LBS

[RSY07] R. Raz, A. Shpilka, and A. Yehudayoff. A lower bound for the size of syntactically multilinear arithmetic circuits. In IEEE [IEE07], pages 438– 448. ISBN 0-7695-3010-9. ISSN 0272-5428. LCCN QA76 .S974 2007. URL http://

ieeexplore.ieee.org/servlet/
opac?punumber=4389466. IEEE

Computer Society order number P3010.

Roughgarden:2000:HBS

[RT00] T. Roughgarden and E. Tardos. How bad is selfish routing? In IEEE [IEE00], pages 93–102. CODEN ASF-PDV. ISBN 0-7695-0850-2, 0-7695-0851-0 (case), 0-7695-0852-9 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 2000. IEEE Computer Society Order Number PR00850.

Regev:2007:SQC

[RT07] O. Regev and B. Toner. Simulating quantum correlations with finite communication. In IEEE [IEE07], pages 384– 394. ISBN 0-7695-3010-9. ISSN 0272-5428. LCCN QA76 .S974 2007. URL http://

ieeexplore.ieee.org/servlet/
opac?punumber=4389466. IEEE

Computer Society order number P3010.

Roglin:2009:SAM

[RT09] H. Roglin and Shang-Hua Teng.

Smoothed analysis of multiobjective optimization. In IEEE [IEE09], pages 681–690. ISBN 0-7695-3850-9. ISSN 0272-5428. LCCN QA76 .S95 2009. IEEE Computer Society order number P3850.

Reingold:2008:DSP

[RTTV08] O. Reingold, L. Trevisan, M. Tulsiani, and S. Vadhan. Dense subsets of pseudorandom sets. In IEEE [IEE08], pages 76–85. ISBN 0-7695-3436-8. ISSN 0272-5428. LCCN QA76 .S95 2008. URL http://

                                        ieeexplore.ieee.org/servlet/
                                        opac?punumber=4690923. IEEE

Computer Society order number P3436.

Rudelson:2006:NIR

[Rud06] Mark Rudelson. Norm of the inverse of a random matrix. In IEEE [IEE06], pages 487–496. ISBN 0-7695-2720-5, 0-7695-2362-5. ISSN 0272-5428. LCCN QA76 .S974 2006. URL http://

                                        ieeexplore.ieee.org/servlet/
                                        opac?punumber=4031329. IEEE

Computer Society Order Number P2720.

Rademacher:2006:DMC

[RV06] Luis Rademacher and Santosh Vempala. Dispersion of mass and the complexity of randomized geometric algorithms. In IEEE [IEE06], pages 729–738. ISBN 0-7695-2720-5, 0-7695-2362-5. ISSN 0272-5428. LCCN QA76 .S974 2006. URL http://

                                        ieeexplore.ieee.org/servlet/
                                        opac?punumber=4031329. IEEE

Computer Society Order Number P2720.

Reingold:2000:EWZ

[RVW00] O. Reingold, S. Vadhan, and A. Wigderson. Entropy waves, the zig-zag graph product, and new constant-degree expanders and extractors. In IEEE [IEE00], pages 3–13. CODEN ASFPDV. ISBN 0-7695-0850-2, 0-7695-0851-0 (case), 0-7695-0852-9 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 2000. IEEE Computer Society Order Number PR00850.

Razborov:2006:LBB

[RY06] Alexander A. Razborov and Sergey Yekhanin. An Ω(n^{1/3}) lower bound for bilinear group based private information retrieval. In IEEE [IEE06], pages 739–748. ISBN 0-7695-2720-5, 0-7695-2362-5. ISSN 0272-5428. LCCN QA76 .S974 2006. URL http://

ieeexplore.ieee.org/servlet/
opac?punumber=4031329. IEEE

Computer Society Order Number P2720.

Raz:2008:MFM

[RY08] R. Raz and A. Yehudayoff. Multilinear formulas, maximal-partition discrepancy and mixed-sources extractors. In IEEE [IEE08], pages 273– 282. ISBN 0-7695-3436-8. ISSN 0272-5428. LCCN QA76 .S95 2008. URL http://

ieeexplore.ieee.org/servlet/
opac?punumber=4690923. IEEE

Computer Society order number P3436.

Roditty:2002:IDR

[RZ02] L. Roditty and U. Zwick. Improved dynamic reachability algorithms for directed graphs. In IEEE [IEE02], pages 679–688. CODEN ASFPDV. ISBN 0-7695-1822-2. ISSN 0272-5428. LCCN QA267. URL http:/

/ieeexplore.ieee.org/iel5/
8411/26517/01181993.pdf?isnumber=
26517&prod=CNF&arnumber=1181993&
arSt=+679&ared=+688&arAuthor=
Roditty%2C+L.%3B+Zwick%2C+
U.;   http://ieeexplore.ieee.
org/xpls/abs_all.jsp?isnumber=
26517&arnumber=1181993&count=
82&index=67.  IEEE Com-

puter Society Order Number PR01822.

Roditty:2004:DAA

[RZ04] L. Roditty and U. Zwick. Dynamic approximate all-pairs shortest paths in undirected graphs. In IEEE [IEE04], pages 499–508. CODEN ASFPDV. ISBN 0-7695-2228-9. ISSN 0272-5428. LCCN QA276. URL http:/

/ieeexplore.ieee.org/iel5/
9430/29918/01366270.pdf?isnumber=
29918&prod=CNF&arnumber=1366270&
arSt=+499&ared=+508&arAuthor=
Roditty%2C+L.%3B+Zwick%2C+
U.;   http://ieeexplore.ieee.
org/xpls/abs_all.jsp?isnumber=
29918&arnumber=1366270&count=

64&index=51. IEEE Computer Society Order Number P2228.

Sankowski:2004:DTC

[San04] P. Sankowski. Dynamic transitive closure via dynamic matrix inverse: extended abstract. In IEEE [IEE04], pages 509–517. CODEN ASFPDV. ISBN 0-7695-2228-9. ISSN 0272-5428. LCCN QA276. URL http:/

/ieeexplore.ieee.org/iel5/
9430/29918/01366271.pdf?isnumber=
29918&prod=CNF&arnumber=1366271&
arSt=+509&ared=+517&arAuthor=
Sankowski%2C+P.;   http://
ieeexplore.ieee.org/xpls/
abs_all.jsp?isnumber=29918&
arnumber=1366271&count=64&

index=52. IEEE Computer Society Order Number P2228.

Sarlos:2006:IAA

[Sar06] Tamas Sarlos. Improved approximation algorithms for large matrices via random projections. In IEEE [IEE06], pages 143–152. ISBN 0-7695-2720-5, 0-7695-2362-5. ISSN 0272-5428. LCCN QA76 .S974 2006. URL http://

ieeexplore.ieee.org/servlet/
opac?punumber=4031329. IEEE

Computer Society Order Number P2720.

Segerlind:2002:SLS

[SBI02] N. Segerlind, S. Buss, and R. Impagliazzo. Switching lemma for small restrictions and lower bounds for k-DNF resolution. In IEEE [IEE02], pages 604–613. CO-DEN ASFPDV. ISBN 0-7695-1822-2. ISSN 0272-5428. LCCN QA267. URL http:/

/ieeexplore.ieee.org/iel5/
8411/26517/01181984.pdf?isnumber=
26517&prod=CNF&arnumber=1181984&
arSt=+604&ared=+613&arAuthor=
Segerlind%2C+N.%3B+Buss%2C+

                                        S.%3B+Impagliazzo%2C+R.;
                                        http://ieeexplore.ieee.org/
                                        xpls/abs_all.jsp?isnumber=
                                        26517&arnumber=1181984&count=
                                        82&index=60.  IEEE Com-

puter Society Order Number PR01822.

Schoenebeck:2008:LLL

[Sch08] G. Schoenebeck. Linear level Lasserre lower bounds for certain k-CSPs. In IEEE [IEE08], pages 593–602. ISBN 0-7695-3436-8. ISSN 0272-5428. LCCN QA76 .S95 2008. URL http://

                                        ieeexplore.ieee.org/servlet/
                                        opac?punumber=4690923. IEEE

Computer Society order number P3436.

Svitkina:2008:SAS

[SF08] Z. Svitkina and L. Fleischer. Submodular approximation: Sampling-based algorithms and lower bounds. In IEEE [IEE08], pages 697–706. ISBN 0-7695-3436-8. ISSN 0272-5428. LCCN QA76 .S95 2008. URL http://

                                        ieeexplore.ieee.org/servlet/
                                        opac?punumber=4690923. IEEE

Computer Society order number P3436.

Sherstov:2008:UEC

[She08] A. A. Sherstov. The unboundederror communication complexity of symmetric functions. In IEEE [IEE08], pages 384– 393. ISBN 0-7695-3436-8. ISSN 0272-5428. LCCN QA76 .S95 2008. URL http://

                                        ieeexplore.ieee.org/servlet/
                                        opac?punumber=4690923. IEEE

Computer Society order number P3436.

Sherman:2009:BMF

[She09a]  J. Sherman.  Breaking the

multicommodity flow barrier for O(v log n)-approximations to sparsest cut. In IEEE [IEE09], pages 363–372. ISBN 0-7695-3850-9. ISSN 0272-5428. LCCN QA76 .S95 2009. IEEE Computer Society order number P3850.

Sherstov:2009:ITH

[She09b] A. A. Sherstov. The intersection of two halfspaces has high threshold degree. In IEEE [IEE09], pages 343–362. ISBN 0-7695-3850-9. ISSN 0272-5428. LCCN QA76 .S95 2009. IEEE Computer Society order number P3850.

Shi:2002:QLB

[Shi02] Yaoyun Shi. Quantum lower bounds for the collision and the element distinctness problems. In IEEE [IEE02], pages 513–

519. CODEN ASFPDV. ISBN 0-7695-1822-2. ISSN 0272-5428. LCCN QA267. URL http:/

/ieeexplore.ieee.org/iel5/
8411/26517/01181975.pdf?isnumber=
26517&prod=CNF&arnumber=1181975&
arSt=+513&ared=+519&arAuthor=
Yaoyun+Shi;   http://ieeexplore.
ieee.org/xpls/abs_all.jsp?
isnumber=26517&arnumber=1181975&
count=82&index=51.  IEEE

Computer Society Order Number PR01822.

Shpilka:2001:LBM

[Shp01] A. Shpilka. Lower bounds for matrix product. In IEEE [IEE01], pages 358–367. CO-DEN ASFPDV. ISBN 0-7695-1390-5, 0-7695-1391-3 (case), 0-7695-1392-1 (microfiche). ISSN 0272-5428. LCCN ????

Skutella:2000:ASS

[Sku00] M. Skutella. Approximating the single source unsplittable mincost flow problem. In IEEE [IEE00], pages 136–145. CO-DEN ASFPDV. ISBN 0-7695-0850-2, 0-7695-0851-0 (case), 0-7695-0852-9 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 2000. IEEE Computer Society Order Number PR00850.

Spielman:2007:SGT

[Spi07]  D. A. Spielman.  Spectral

graph theory and its applications. In IEEE [IEE07], pages 29–38. ISBN 0-7695-3010-9. ISSN 0272-5428. LCCN QA76 .S974 2007. URL http://

                                        ieeexplore.ieee.org/servlet/
                                        opac?punumber=4389466. IEEE

Computer Society order number P3010.

Srinivasan:2001:DLS

[Sri01] A. Srinivasan. Distributions on level-sets with applications to approximation algorithms. In IEEE [IEE01], pages 588–597. CODEN ASFPDV. ISBN 0-7695-1390-5, 0-7695-1391-3 (case), 0-7695-1392-1 (microfiche). ISSN 0272-5428. LCCN ????

Shmoys:2004:SOA

[SS04] D. B. Shmoys and C. Swamy.

Stochastic optimization is (almost) as easy as deterministic optimization. In IEEE [IEE04], pages 228–237. CO-DEN ASFPDV. ISBN 0-7695-2228-9. ISSN 0272-5428. LCCN QA276. URL http:/

/ieeexplore.ieee.org/iel5/
9430/29918/01366242.pdf?isnumber=
29918&prod=CNF&arnumber=1366242&
arSt=+228&ared=+237&arAuthor=
Shmoys%2C+D.B.%3B+Swamy%2C+
C.;   http://ieeexplore.ieee.
org/xpls/abs_all.jsp?isnumber=
29918&arnumber=1366242&count=

64&index=23. IEEE Computer Society Order Number P2228.

Swamy:2005:SBA

[SS05] C. Swamy and D. B. Shmoys.

Sampling-based approximation algorithms for multi-stage stochastic optimization. In IEEE [IEE05], pages 357–366. ISBN 0-7695-2468-0, 0-7695-2362-5. ISSN 0272-5428. LCCN QA76 .S979 2005. IEEE Computer Society order number P2468.

Spielman:2003:SSS

[ST03] D. A. Spielman and Shang-Hua Teng. Solving sparse, symmetric, diagonally-dominant linear systems in time O(m^{1.31}). In IEEE [IEE03], pages 416–

427. CODEN ASFPDV. ISBN 0-7695-2040-5. ISSN 0272-5428. LCCN QA76 .S979 2003. URL http:/

/ieeexplore.ieee.org/iel5/
8767/27770/01238215.pdf?isnumber=

                                        27770&prod=CNF&arnumber=1238215&
                                        arSt=+416&ared=+427&arAuthor=
                                        Spielman%2C+D.A.%3B+Shang-
                                        Hua+Teng;   http://ieeexplore.
                                        ieee.org/xpls/abs_all.jsp?
                                        isnumber=27770&arnumber=1238215&
                                        count=66&index=42.  IEEE

Computer Society Order Num-

ber PR02040.

Storjohann:2001:DCF

[Sto01] A. Storjohann. Deterministic computation of the Frobenius form. In IEEE [IEE01], pages 368–377. CODEN ASF-PDV. ISBN 0-7695-1390-5, 0-7695-1391-3 (case), 0-7695-1392-1 (microfiche). ISSN 0272-5428. LCCN ????

Streinu:2000:CAP

[Str00]  I. Streinu.  A combinato-

rial approach to planar noncolliding robot arm motion planning. In IEEE [IEE00], pages 443–453. CODEN ASFPDV. ISBN 0-7695-0850-2, 0-7695-0851-0 (case), 0-7695-0852-9 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 2000. IEEE Computer Society Order Number PR00850.

Shaltiel:2001:SEA

[SU01] R. Shaltiel and C. Umans. Simple extractors for all minentropies and a new pseudorandom generator. In IEEE [IEE01], pages 648–657. CO-DEN ASFPDV. ISBN 0-7695-1390-5, 0-7695-1391-3 (case), 0-7695-1392-1 (microfiche). ISSN 0272-5428. LCCN ????

Sudan:2001:CTT

[Sud01] M. Sudan. Coding theory:

tutorial & survey. In IEEE [IEE01], pages 36–53. CODEN ASFPDV. ISBN 0-7695-1390-5, 0-7695-1391-3 (case), 0-7695-1392-1 (microfiche). ISSN 0272-5428. LCCN ????

Savani:2004:EMS

[SvS04] R. Savani and B. von Stengel. Exponentially many steps for finding a Nash equilibrium in a bimatrix game. In IEEE [IEE04], pages 258–267. CODEN ASFPDV. ISBN 0-7695-2228-9. ISSN 0272-5428. LCCN QA276. URL http:/

/ieeexplore.ieee.org/iel5/
9430/29918/01366245.pdf?isnumber=
29918&prod=CNF&arnumber=1366245&
arSt=+258&ared=+267&arAuthor=
Savani%2C+R.%3B+von+Stengel%
2C+B.;   http://ieeexplore.
ieee.org/xpls/abs_all.jsp?
isnumber=29918&arnumber=1366245&
count=64&index=26.  IEEE

Computer Society Order Number P2228.

Stefankovic:2007:ASA

[SVV07] D. Stefankovic, S. Vempala, and E. Vigoda. Adaptive simulated annealing: a near-optimal connection between sampling and counting. In IEEE [IEE07], pages 183–193. ISBN 0-7695-3010-9. ISSN 0272-5428. LCCN QA76 .S974 2007. URL http://

ieeexplore.ieee.org/servlet/
opac?punumber=4389466. IEEE

Computer Society order number P3010.

Sommer:2009:DOS

[SVY09] C. Sommer, E. Verbin, and Wei Yu. Distance oracles for sparse graphs. In IEEE [IEE09], pages 703–712. ISBN 0-7695-3850-9. ISSN 0272-5428. LCCN QA76 .S95 2009. IEEE Computer Society order number P3850.

Sviridenko:2000:AAR

[SW00] M. Sviridenko and G. J. Woeginger. Approximability and inapproximability results for nowait shop scheduling. In IEEE [IEE00], pages 116–125. CO-DEN ASFPDV. ISBN 0-7695-0850-2, 0-7695-0851-0 (case), 0-7695-0852-9 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 2000. IEEE Computer Society Order Number PR00850.

Szabo:2001:USO

[SW01] T. Szabo and E. Welzl. Unique sink orientations of cubes. In IEEE [IEE01], pages 547– 555. CODEN ASFPDV. ISBN 0-7695-1390-5, 0-7695-1391-3 (case), 0-7695-1392-1 (microfiche). ISSN 0272-5428. LCCN ????

Sun:2006:QQC

[SY06] Xiaoming Sun and Andrew C. Yao. On the quantum query complexity of local search in two and three dimensions. In IEEE [IEE06], pages 429–438. ISBN 0-7695-2720-5, 0-7695-2362-5. ISSN 0272-5428. LCCN QA76 .S974 2006. URL http://

                                        ieeexplore.ieee.org/servlet/
                                        opac?punumber=4031329. IEEE

Computer Society Order Number P2720.

Szegedy:2004:QSM

[Sze04] M. Szegedy. Quantum speedup of Markov chain based algorithms. In IEEE [IEE04], pages 32–41. CODEN ASFPDV. ISBN 0-7695-2228-9. ISSN 0272-5428. LCCN QA276. URL

http://ieeexplore.ieee.org/
iel5/9430/29918/01366222.
pdf?isnumber=29918&prod=CNF&
arnumber=1366222&arSt=+32&
ared=+41&arAuthor=Szegedy%
2C+M.;   http://ieeexplore.
ieee.org/xpls/abs_all.jsp?
isnumber=29918&arnumber=1366222&
count=64&index=3. IEEE Com-

puter Society Order Number P2228.

Tao:2007:SRC

[Tao07] T. Tao. Structure and randomness in combinatorics. In IEEE [IEE07], pages 3–15. ISBN 0-7695-3010-9. ISSN 0272-5428. LCCN QA76 .S974 2007. URL http://

ieeexplore.ieee.org/servlet/
opac?punumber=4389466. IEEE

Computer Society order number P3010.

Thorup:2001:COR

[Tho01] M. Thorup. Compact oracles for reachability and approximate distances in planar digraphs. In IEEE [IEE01], pages 242–251. CODEN ASF-PDV. ISBN 0-7695-1390-5, 0-7695-1391-3 (case), 0-7695-1392-1 (microfiche). ISSN 0272-5428. LCCN ????

Thorup:2002:EBP

[Tho02] M. Thorup. Equivalence between priority queues and sorting. In IEEE [IEE02], pages 125–134. CODEN ASFPDV. ISBN 0-7695-1822-2. ISSN 0272-5428. LCCN QA267. URL http:/

                                        /ieeexplore.ieee.org/iel5/
                                        8411/26517/01181889.pdf?isnumber=
                                        26517&prod=CNF&arnumber=1181889&
                                        arSt=+125&ared=+134&arAuthor=
                                        Thorup%2C+M.;   http://ieeexplore.
                                        ieee.org/xpls/abs_all.jsp?
                                        isnumber=26517&arnumber=1181889&
                                        count=82&index=14.  IEEE

Computer Society Order Num-

ber PR01822.

Toran:2000:HGI

[Tor00] J. Toran. On the hardness of graph isomorphism. In IEEE [IEE00], pages 180–186. CO-DEN ASFPDV. ISBN 0-7695-0850-2, 0-7695-0851-0 (case), 0-7695-0852-9 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 2000. IEEE Computer Society Order Number PR00850.

Trevisan:2003:LDU

[Tre03] L. Trevisan. List-decoding using the XOR lemma. In IEEE [IEE03], pages 126–135. CO-DEN ASFPDV. ISBN 0-7695-2040-5. ISSN 0272-5428. LCCN QA76 .S979 2003. URL http:/

                                        /ieeexplore.ieee.org/iel5/
                                        8767/27770/01238187.pdf?isnumber=
                                        27770&prod=CNF&arnumber=1238187&
                                        arSt=+126&ared=+135&arAuthor=
                                        Trevisan%2C+L.;   http://
                                        ieeexplore.ieee.org/xpls/

abs_all.jsp?isnumber=27770&
arnumber=1238187&count=66&

index=14. IEEE Computer Society Order Number PR02040.

Trevisan:2005:AAU

[Tre05] L. Trevisan. Approximation algorithms for unique games. In IEEE [IEE05], pages 197– 205. ISBN 0-7695-2468-0, 0-7695-2362-5. ISSN 0272-5428. LCCN QA76 .S979 2005. IEEE Computer Society order number P2468.

Trevisan:2008:ACC

[Tre08] L. Trevisan. Average-case complexity. In IEEE [IEE08], page 11. ISBN 0-7695-3436-8. ISSN 0272-5428. LCCN QA76 .S95 2008. URL http://

ieeexplore.ieee.org/servlet/
opac?punumber=4690923. IEEE

Computer Society order number P3436.

Ta-Shma:2006:BLC

[TSU06] Amnon Ta-Shma and Christopher Umans. Better lossless condensers through derandomized curve samplers. In IEEE [IEE06], pages 177–186. ISBN 0-7695-2720-5, 0-7695-2362-5. ISSN 0272-5428. LCCN QA76 .S974 2006. URL http://

ieeexplore.ieee.org/servlet/
opac?punumber=4031329. IEEE

Computer Society Order Number P2720.

Ta-Shma:2001:ERM

[TSZS01] A. Ta-Shma, D. Zuckerman, and S. Safra. Extractors from Reed–Muller codes. In IEEE [IEE01], pages 638–647. CO-DEN ASFPDV. ISBN 0-7695-1390-5, 0-7695-1391-3 (case), 0-7695-1392-1 (microfiche). ISSN 0272-5428. LCCN ????

Trevisan:2000:ERS

[TV00] L. Trevisan and S. Vadhan. Extracting randomness from samplable distributions. In IEEE [IEE00], pages 32–42. CODEN ASFPDV. ISBN 0-7695-0850-2, 0-7695-0851-0 (case), 0-7695-0852-9 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 2000. IEEE Computer Society Order Number PR00850.

Unger:2009:PIA

[Ung09] F. Unger. A probabilistic inequality with applications to threshold direct-product theorems. In IEEE [IEE09], pages 221–229. ISBN 0-7695-3850-9. ISSN 0272-5428. LCCN QA76 .S95 2009. IEEE Computer Society order number P3850.

Upfal:2003:PAD

[Upf03] E. Upfal. Performance analysis of dynamic processes. In IEEE [IEE03], page 18. CO-DEN ASFPDV. ISBN 0-7695-2040-5. ISSN 0272-5428. LCCN QA76 .S979 2003. URL http:/

/ieeexplore.ieee.org/iel5/
8767/27770/01238176.pdf?isnumber=
27770&prod=CNF&arnumber=1238176&
arSt=+18&ared=&arAuthor=Upfal%
2C+E.;   http://ieeexplore.
ieee.org/xpls/abs_all.jsp?
isnumber=27770&arnumber=1238176&
count=66&index=3. IEEE Com-

puter Society Order Number PR02040.

Vadhan:2002:RET

[Vad02] S. P. Vadhan. Randomness extractors and their many guises. In IEEE [IEE02], page 9. CO-DEN ASFPDV. ISBN 0-7695-1822-2. ISSN 0272-5428. LCCN QA267. URL http:/

/ieeexplore.ieee.org/iel5/
8411/26517/01181877.pdf?isnumber=
26517&prod=CNF&arnumber=1181877&
arSt=+9&ared=&arAuthor=Vadhan%
2C+S.P.;   http://ieeexplore.
ieee.org/xpls/abs_all.jsp?
isnumber=26517&arnumber=1181877&
count=82&index=2. IEEE Com-

puter Society Order Number PR01822.

Vadhan:2004:USC

[Vad04] S. P. Vadhan. An unconditional study of computational zero knowledge. In IEEE [IEE04], pages 176–185. CO-DEN ASFPDV. ISBN 0-7695-2228-9. ISSN 0272-5428. LCCN QA276. URL http:/

/ieeexplore.ieee.org/iel5/
9430/29918/01366237.pdf?isnumber=
29918&prod=CNF&arnumber=1366237&
arSt=+176&ared=+185&arAuthor=
Vadhan%2C+S.P.;   http://
ieeexplore.ieee.org/xpls/
abs_all.jsp?isnumber=29918&
arnumber=1366237&count=64&

index=18. IEEE Computer Society Order Number P2228.

Valiant:2004:HA

[Val04] L. G. Valiant. Holographic algorithms. In IEEE [IEE04], pages 306–315. CODEN ASFPDV. ISBN 0-7695-2228-9. ISSN 0272-5428. LCCN QA276. URL http:/

                                        /ieeexplore.ieee.org/iel5/
                                        9430/29918/01366250.pdf?isnumber=
                                        29918&prod=CNF&arnumber=1366250&
                                        arSt=+306&ared=+315&arAuthor=
                                        Valiant%2C+L.G.;   http://
                                        ieeexplore.ieee.org/xpls/
                                        abs_all.jsp?isnumber=29918&
                                        arnumber=1366250&count=64&

index=31. IEEE Computer Society Order Number P2228.

Valiant:2006:AA

[Val06] Leslie G. Valiant. Accidental algorithms. In IEEE [IEE06], pages 509–517. ISBN 0-7695-2720-5, 0-7695-2362-5. ISSN 0272-5428. LCCN QA76 .S974 2006. URL http://

                                        ieeexplore.ieee.org/servlet/
                                        opac?punumber=4031329. IEEE

Computer Society Order Num-

ber P2720.

vanDam:2001:HPA

[vDMV01] W. van Dam, M. Mosca, and U. Vazirani. How powerful is adiabatic quantum computation? In IEEE [IEE01], pages 279–287. CODEN ASF-PDV. ISBN 0-7695-1390-5, 0-7695-1391-3 (case), 0-7695-1392-1 (microfiche). ISSN 0272-5428. LCCN ????

Vershynin:2006:BHC

[Ver06]  Roman Vershynin.  Beyond

Hirsch conjecture: Walks on random polytopes and smoothed complexity of the simplex method. In IEEE [IEE06], pages 133–142. ISBN 0-7695-2720-5, 0-7695-2362-5. ISSN 0272-5428. LCCN QA76 .S974 2006. URL http://

ieeexplore.ieee.org/servlet/
opac?punumber=4031329. IEEE

Computer Society Order Number P2720.

Vetta:2002:NEC

[Vet02] A. Vetta. Nash equilibria in competitive societies, with applications to facility location, traffic routing and auctions. In IEEE [IEE02], pages 416–425. CODEN ASFPDV. ISBN 0-7695-1822-2. ISSN 0272-5428. LCCN QA267. URL http:/

/ieeexplore.ieee.org/iel5/
8411/26517/01181966.pdf?isnumber=
26517&prod=CNF&arnumber=1181966&
arSt=+416&ared=+425&arAuthor=
Vetta%2C+A.;   http://ieeexplore.
ieee.org/xpls/abs_all.jsp?
isnumber=26517&arnumber=1181966&
count=82&index=42.  IEEE

Computer Society Order Number PR01822.

Vondrak:2009:SAS

[Von09] J. Vondrak. Symmetry and approximability of submodular maximization problems. In IEEE [IEE09], pages 651–670. ISBN 0-7695-3850-9. ISSN 0272-5428. LCCN QA76 .S95 2009. IEEE Computer Society order number P3850.

Vovk:2002:LCM

[Vov02] V. Vovk. On-line confidence machines are well-calibrated. In IEEE [IEE02], pages 187–196. CODEN ASFPDV. ISBN 0-7695-1822-2. ISSN 0272-5428.

LCCN QA267. URL http:/

/ieeexplore.ieee.org/iel5/
8411/26517/01181895.pdf?isnumber=
26517&prod=CNF&arnumber=1181895&
arSt=+187&ared=+196&arAuthor=
Vovk%2C+V.;   http://ieeexplore.
ieee.org/xpls/abs_all.jsp?
isnumber=26517&arnumber=1181895&
count=82&index=20.  IEEE

Computer Society Order Number PR01822.

Vereshchagin:2002:KSF

[VV02] N. Vereshchagin and P. Vitanyi. Kolmogorov’s structure functions with an application to the foundations of model selection. In IEEE [IEE02], pages 751–

760. CODEN ASFPDV. ISBN 0-7695-1822-2. ISSN 0272-5428. LCCN QA267. URL http:/

/ieeexplore.ieee.org/iel5/
8411/26517/01182000.pdf?isnumber=
26517&prod=CNF&arnumber=1182000&
arSt=+751&ared=+760&arAuthor=
Vereshchagin%2C+N.%3B+Vitanyi%
2C+P.;   http://ieeexplore.
ieee.org/xpls/abs_all.jsp?
isnumber=26517&arnumber=1182000&
count=82&index=74.  IEEE

Computer Society Order Number PR01822.

Varadarajan:2002:ARP

[VVZ02] K. R. Varadarajan, S. Venkatesh, and Jiawei Zhang. On approximating the radii of point sets in high dimensions. In IEEE [IEE02], pages 561–569. CODEN ASFPDV. ISBN 0-7695-1822-2. ISSN 0272-5428. LCCN QA267. URL http:/

/ieeexplore.ieee.org/iel5/
8411/26517/01181980.pdf?isnumber=

26517&prod=CNF&arnumber=1181980&
arSt=+561&ared=+569&arAuthor=
Varadarajan%2C+K.R.%3B+Venkatesh%
2C+S.%3B+Jiawei+Zhang;   http:
//ieeexplore.ieee.org/xpls/
abs_all.jsp?isnumber=26517&
arnumber=1181980&count=82&

index=56. IEEE Computer Society Order Number PR01822.

Vempala:2002:SAL

[VW02] S. Vempala and G. Wang. A spectral algorithm for learning mixtures of distributions. In IEEE [IEE02], pages 113–122. CODEN ASFPDV. ISBN 0-7695-1822-2. ISSN 0272-5428. LCCN QA267. URL http:/

/ieeexplore.ieee.org/iel5/
8411/26517/01181888.pdf?isnumber=
26517&prod=CNF&arnumber=1181888&
arSt=+113&ared=+122&arAuthor=
Vempala%2C+S.%3B+Wang%2C+
G.;   http://ieeexplore.ieee.
org/xpls/abs_all.jsp?isnumber=
26517&arnumber=1181888&count=
82&index=13.  IEEE Com-

puter Society Order Number PR01822.

Viola:2007:OWM

[VW07] E. Viola and A. Wigderson. One-way multi-party communication lower bound for pointer jumping with applications. In IEEE [IEE07], pages 427–437. ISBN 0-7695-3010-

9. ISSN 0272-5428. LCCN QA76 .S974 2007. URL http://

ieeexplore.ieee.org/servlet/
opac?punumber=4389466. IEEE

Computer Society order number P3010.

Wagner:2006:GGU

[Wag06] Uli Wagner. On a geometric generalization of the upper bound theorem. In IEEE [IEE06], pages 635–645. ISBN 0-7695-2720-5, 0-7695-2362-5. ISSN 0272-5428. LCCN QA76 .S974 2006. URL http://

                                        ieeexplore.ieee.org/servlet/
                                        opac?punumber=4031329. IEEE

Computer Society Order Num-

ber P2720.

Watrous:2000:SQP

[Wat00] J. Watrous. Succinct quantum proofs for properties of finite groups. In IEEE [IEE00], pages 537–546. CODEN ASF-PDV. ISBN 0-7695-0850-2, 0-7695-0851-0 (case), 0-7695-0852-9 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 2000. IEEE Computer Society Order Number PR00850.

Watrous:2002:LPQ

[Wat02] J. Watrous. Limits on the power of quantum statistical zero-knowledge. In IEEE [IEE02], pages 459–468. CO-DEN ASFPDV. ISBN 0-7695-1822-2. ISSN 0272-5428. LCCN QA267. URL http:/

                                        /ieeexplore.ieee.org/iel5/
                                        8411/26517/01181970.pdf?isnumber=
                                        26517&prod=CNF&arnumber=1181970&
                                        arSt=+459&ared=+468&arAuthor=
                                        Watrous%2C+J.;   http://ieeexplore.
                                        ieee.org/xpls/abs_all.jsp?
                                        isnumber=26517&arnumber=1181970&
                                        count=82&index=46.  IEEE

Computer Society Order Num-

ber PR01822.

Woods:2006:TCT

[WN06] Damien Woods and Turlough Neary. On the time complexity of 2-tag systems and small universal Turing machines. In IEEE [IEE06], pages 439–448. ISBN 0-7695-2720-5, 0-7695-2362-5. ISSN 0272-5428. LCCN QA76 .S974 2006. URL http://

ieeexplore.ieee.org/servlet/
opac?punumber=4031329. IEEE

Computer Society Order Number P2720.

Woodruff:2006:LBA

[Woo06] David P. Woodruff. Lower bounds for additive spanners, emulators, and more. In IEEE [IEE06], pages 389–398. ISBN 0-7695-2720-5, 0-7695-2362-5. ISSN 0272-5428. LCCN QA76 .S974 2006. URL http://

ieeexplore.ieee.org/servlet/
opac?punumber=4031329. IEEE

Computer Society Order Number P2720.

Wigderson:2005:RES

[WX05] A. Wigderson and D. Xiao. A randomness-efficient sampler for matrix-valued functions and applications. In IEEE [IEE05], pages 397–406. ISBN 0-7695-2468-0, 0-7695-2362-5. ISSN 0272-5428. LCCN QA76 .S979 2005. IEEE Computer Society order number P2468.

Young:2001:SPA

[You01] N. E. Young. Sequential and parallel algorithms for mixed packing and covering. In IEEE [IEE01], pages 538–546. CO-DEN ASFPDV. ISBN 0-7695-1390-5, 0-7695-1391-3 (case), 0-7695-1392-1 (microfiche). ISSN 0272-5428. LCCN ????

Yuster:2008:MSR

[Yus08] R. Yuster. Matrix sparsification for rank and determinant computations via nested dissection. In IEEE [IEE08], pages 137–145. ISBN 0-7695-3436-8. ISSN 0272-5428. LCCN QA76 .S95 2008. URL http://

                                        ieeexplore.ieee.org/servlet/
                                        opac?punumber=4690923. IEEE

Computer Society order number P3436.

Yuster:2005:ADQ

[YZ05] R. Yuster and U. Zwick. Answering distance queries in directed graphs using fast matrix multiplication. In IEEE [IEE05], pages 389–396. ISBN 0-7695-2468-0, 0-7695-2362-5. ISSN 0272-5428. LCCN QA76 .S979 2005. IEEE Computer Society order number P2468.